A Complete Bibliography of Publications in Fishes

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

30 August 2024 Version 1.19

Title word cross-reference

−80^{∘} [LJK^{+}22]. ^{1} [WGRM^{+}19]. _{2} [Kim23a, Kim23b]. α [AASQPU^{+}23, PLJ^{+}23]. β [EAE^{+}23, JSJ^{+}24, NRÁGPM^{+}18, TU18, VMS^{+}23, WLC^{+}23]. δ^{13} [BFM23]. δ^{15} [BFM23]. γ [HGSE23]. × [FHHC23, HAC^{+}23, Kim23a, Kim23b, MMVVCJ^{+}23, PYJ^{+}23, SZZ^{+}23, WZL^{+}24, YGD^{+}23].

-1 [EAE^{+}23]. -Estradiol [JSJ^{+}24, VMS^{+}23]. -Glucans [NRÁGPM^{+}18]. -Irradiated [HGSE23]. -Limonene [dSJC^{+}21]. -Omics [EBCM24]. -Sitosterol [TU18]. -Stimulated [SOW^{+}23]. -Tocopherol [AASQPU^{+}23].

1 [AHJ^{+}23, Gae17, PLJ^{+}23, TBPJ23, ZSZ^{+}23]. 100-Year [APS^{+}21]. 113 [Kim23a]. 17 [VMS^{+}23, WSZ^{+}23].

2-methylisoborneol [LLKKV20]. 200 [ZLZ^{+}22]. 222 [BEMC23, CWMX21]. 24 [JML^{+}24].

3 [LL18a]. 3-Aminobenzoic [CWMX21]. 36 [LL18a].

41 [PKSN23a]. 411 [APD^{+}23a]. 4n

[ASM^{+}22]. 4n-6 [ASM^{+}22].

51 [Gae17]. 54 [XYC^{+}23]. 552 [MWPS24].

7 [XYC^{+}23]. 7-NAChR [ZAM^{+}23]. 77 [CPJK23a].

8 [APD^{+}23a, CPJK23a, Kim23a, MWPS24, PKSN23a]. 8-like [LYL^{+}24a].

90-Day [GLW^{+}22].

abdominalis [STZ^{+}23, ZQLW23]. Ability [JHNF24]. Abiotic [CH23]. Abnormal [DAF^{+}22, LHZ^{+}24]. Aboriginal [KSAB^{+}23]. Absorption [JCR^{+}22, RMSPMC^{+}22]. Abudefduf [LDBL19]. Abundance [AGE^{+}18, DMA22, HWZ21, LCZ^{+}23, NIN^{+}19, RCR^{+}23, SMM^{+}18, Tri23, TMPP23, WZG^{+}23]. Acanthocybium [GPD^{+}23]. Acanthopagrus [ZJM^{+}23]. Acceptability [RLAE23]. Acceptance [CWP^{+}21]. Access [OOG24]. Acclimated [RHUJ24]. Accompanied [DAHM19]. According [Fra23, YCR^{+}23]. Accumulation [LLKKV20]. Accuracy [CLZ^{+}24]. Accurate [CMP^{+}23, KLD^{+}23]. Acetyl [ZYT^{+}24]. Acetyl-CoA [ZYT^{+}24]. Acetylcholine [ZAM^{+}23]. Achievements [HGC^{+}23]. Achieving [YAS21]. Achilles [BAA^{+}23]. Acid [ASM^{+}22, BBCJ23, CWMX21, CHZ^{+}24, HZL^{+}22, HS18, LZM^{+}23, MLZ^{+}21, PSW^{+}23, TPC^{+}23, TMD^{+}19, VKP^{+}24, ZYT^{+}24]. Acid-Based [HS18]. Acidification [DLL^{+}23, KWK^{+}24, PFM^{+}20, ZSZSS^{+}22]. Acidifiers [ZWZ22]. Acids [BKBR^{+}23, MNP^{+}16, ZSZ^{+}23]. Acipenser [AVT18, LDW^{+}21, AVT18, CZJ^{+}24, HCZ^{+}23, JCR^{+}22, Kim23a, Kim23b, ZLX^{+}23]. AcIV [AVT18]. AcIV-E [AVT18]. Acknowledgement [Off18, Off19, Off20]. Acknowledgment [Off21, Off23]. Acoustic [BFLC19, CXT^{+}24, HBL^{+}22, VPPF^{+}19]. across [CMP^{+}20, JJB^{+}24, MWPS23, MWPS24, NHC^{+}23, RRG22, ZZY^{+}23]. Actinopterygii [KRAFO23]. Actions [LKU21]. Activates [PLJ^{+}23]. Activation [LWS^{+}23a]. Activators [SSK^{+}22]. Activities [CFCE20, DCL^{+}23a, DDN19, DZC^{+}22, ERE21, LYL^{+}24b, WZL^{+}24, YBL^{+}22, ZZL^{+}23]. Activity [BB23, BSH^{+}23, DRFCL23, EEdCSOPJ^{+}23, GMBR^{+}21, GMDMT^{+}23, JCL^{+}24, MMVVCJ^{+}23, NRÁGPM^{+}18, OGMG^{+}17, PJPMMV^{+}22, dCPRG^{+}21, SSK^{+}22, TCV^{+}19, WLW^{+}23, XYL^{+}23, YHZ^{+}23, dSJC^{+}21, dAdSCC^{+}23]. aculeatus [BEMC23, DLL^{+}23, GMMNRS18, JBK^{+}23, KAJ^{+}24, SAL18]. Acute [AMT^{+}24, LJK^{+}22, LWS^{+}23b, WLN^{+}23]. Adaptation [MSK^{+}22, MML22, WMZ^{+}22]. Adaptive [DMB^{+}20, FCT19]. Addendum [LL18a]. Addiction [LMLH22]. Addition [CCFP19, GZX^{+}22, GSH^{+}24]. Additional [KRAFO23, Näs18]. Additive [HMP^{+}24]. Additives [CHZ^{+}24, MBAM19, OO24, ZWD^{+}23]. Adjacent [AM24, LHL^{+}24]. Adjusting [MOW^{+}18]. Adjustments [dAPAA^{+}24, ZSZSS^{+}22]. Administration [GCFA^{+}22, HSD^{+}24, JE18, LCWH22, LDD^{+}22, SSSP21, TU18, YZH^{+}24]. Adriatic [LOBTL22, HB24, LKD22]. Adult [DOB^{+}17, DM24, DZC^{+}22, ZLZ^{+}24]. Advanced [DOB^{+}17, NLTL23]. Advances [EBCM24, LYY^{+}24]. Adverse [DK22, KJ22]. Aegean [KMB18, KKB20, TK24, BEG^{+}23]. Aequidens [LBH^{+}24]. Aequipecten [KŽM^{+}23]. aerea [GBT^{+}24]. Aerial [RGABD20]. Aerobic [LWS^{+}21]. Aeromonas [AEME^{+}23, GWT^{+}24, JLW^{+}24, LCWH22, MYW^{+}24, PYJ^{+}23, PSN18, dCPRG^{+}21, QXAY22, SSSP21, SMO^{+}22, SNK^{+}23, WCX^{+}24, XYF^{+}24, XXL^{+}22, ZLSB22]. aestivum [WLL^{+}22]. Affect [CCFP19, FÅG^{+}23, JCL^{+}24]. Affected [LUM18]. Affecting [WM22]. Affects [GRKC19, PLJ^{+}23, RSA17, RJTVC^{+}19, ZLZ^{+}22, ZHQ^{+}23]. Africa [MML22, OOGAS23]. African [AZY^{+}24, AAAF^{+}21, BC23, BHR^{+}23, Eny17, EPKV17, MSK^{+}22, SSSP21, SBB^{+}19, WSE^{+}21]. After [ŠT19, CHJ^{+}23, EPKV17, GWT^{+}24, GLY^{+}23, GSHGE18, HIO^{+}19, HSD^{+}24, Mil23, MBZ^{+}21, RJR^{+}22, RJFCJC^{+}21, SEA^{+}23, VKP^{+}24, WBK^{+}23, YLX^{+}22, ZZW^{+}24]. against [CCP^{+}24, DCL^{+}23b, JML^{+}24, LTZ^{+}22, LCWH22, PSN18, dCPRG^{+}21, SML^{+}23b, XYF^{+}24, dSJC^{+}21]. agalactiae [AMT^{+}24, GGL^{+}23, GCFA^{+}22, GYH^{+}23, HLZ^{+}22, LDD^{+}22, PSN18, RIF^{+}23]. Age [BFB^{+}23, BPOS19, BPO19, CLL^{+}22, CZJ^{+}24, FZZ^{+}24, FB22, HS24, KZC^{+}24, LLWW23, LOT^{+}22, MCSB^{+}19, MPK^{+}23, NZVB20, PSP^{+}22, RGTPSCCC24, SNSVFL23, SCT^{+}24, XSZ^{+}23, YCR^{+}23]. Age-Based [HS24]. Ageing [TMM^{+}18]. Agent [FCT19, QXAY22]. Agent-Based [FCT19]. Agents [LBH^{+}24, SML^{+}23b]. Ages [MOW^{+}18]. Aggression [WWW19]. Aging [MTM^{+}19]. Agonist [WSI^{+}19]. Agreement [QL22]. Agro [KYB^{+}23]. Agro-Ecological [KYB^{+}23]. AI [PSP^{+}22]. Air [BMSGS^{+}18, LZW^{+}24]. AIS [FYH^{+}23, LSJ24]. akashiwo [JXW^{+}23]. al [APD^{+}23a, CPJK23a, MWPS24, PKSN23a, XYC^{+}23]. Alarm [LYL^{+}23]. Alaska [DMB^{+}20]. albacares [LYL^{+}24b]. Albendazole [NDC^{+}23]. albicaudus [ARH^{+}23]. albomaculosus [KMS^{+}17]. Albula [KBK^{+}23]. Albulidae [KBK^{+}23]. album [BPO19]. Alburnus [HV24]. ALC [ZJC^{+}22b]. Alcichthys [PKY^{+}23]. alexandri [SdSdOSSL23]. Alfalfa [CWP^{+}21]. alfredi [LCW23a]. Alga [JXW^{+}23]. Algae [GBT^{+}24, RDG^{+}20, VRG^{+}24]. Algal [SAHS18]. Algorithm [NLTL23, WYL23]. Alien [SKD^{+}23]. Alimentary [BMMD22]. Alizarin [ZJC^{+}22b]. Alkalinity [LWZ^{+}22]. Alleviate [WHX^{+}23]. Alleviated [SPQ^{+}24]. Alleviates [MYW^{+}24]. Allocated [NNL^{+}23]. Allometric [STZ^{+}23]. Almost [YLW^{+}23a]. Alone [HYN^{+}24]. along [ARH^{+}23, EKL^{+}23, SML^{+}23a]. Alosa [SCT^{+}24]. Alpha [LHL^{+}24, ZAM^{+}23]. Alpine [APD^{+}23a, APD^{+}23b]. alpinus [WGRM^{+}19]. Alter [LLY^{+}24]. Altered [ZZW^{+}22]. Alternation [EAJ^{+}23]. Alternative [DBP^{+}20, LWT^{+}24]. Alternatives [HIO^{+}19]. Alters [BEMC23]. Amaranth [MBL20]. Amaranthus [MBL20]. Amarilladesma [GGP^{+}23]. Amaro [KRAFO23]. Amazon [dSCFQB^{+}23, GAR^{+}24, HVRCG18, PCH^{+}24, dCPFS^{+}24, VRKV24]. Amberjack [IKT19, JFP^{+}18, NRKT19]. Ambient [MAR^{+}18]. amblycephala [GWT^{+}24, MYW^{+}24]. Amblygobius [CHH^{+}23]. amboinensis [KSAB^{+}23]. Ameiurus [SMH^{+}22a]. America [WWD^{+}23]. American [dAMPS^{+}23, WHX^{+}23, ZWZ22]. americanus [BFM23, PPAB^{+}18, PLV^{+}19]. amh [ZLZ^{+}22, CMP^{+}23]. Amino [BBCJ23, MLZ^{+}21, TPC^{+}23]. Aminobenzoic [CWMX21]. Aminotransferase [GLY^{+}23]. Ammocetes [KSO^{+}23]. Ammonia [CSL24]. among [AWCS23, BBCJ23, CZJ^{+}24, VMBT24, ZLL^{+}23]. Amorgos [TK24]. Amphibians [OMC^{+}19]. Amphipoda [BSH^{+}23]. Amplification [LDZ^{+}24]. Amur [WLL^{+}23, ZLZ^{+}24]. Amyloodiniosis [MCSB^{+}19]. Anadromous [APS^{+}21, HCWH20]. Anaesthetic [BEMC23]. Analog [JML^{+}24]. Analyses [CZL23, EM23, HWX^{+}23, LOT^{+}22, SUL^{+}23, VV23, WGRM^{+}19, XYT23, ZZZ^{+}23]. Analysis [APAHBMAG23, APS^{+}21, BBF22, CYL^{+}23, CCCFE18, CVANRD^{+}21, DLL^{+}22, DCD^{+}24, DL23, DSC^{+}19, DDG^{+}22, FLX^{+}22, FHHC23, FM21a, GGL^{+}23, GLY^{+}23, GWH21, HHL^{+}20, HWW^{+}24b, HAC^{+}23, HHM^{+}24, HYXY23, IFA^{+}23, JJK21, JLW^{+}24, KTT24, KAB^{+}23, LPK^{+}23b, LSJ^{+}23, LFH^{+}23, LLY^{+}22, LYL^{+}24a, LKD22, LYL^{+}24b, LZW^{+}24, LYW^{+}24, MYY^{+}23, MPM^{+}18, MTM^{+}19, MLSC^{+}23, NTP^{+}21, PKY^{+}23, RWF^{+}23, SGANM^{+}24, SLY^{+}21, SJY^{+}22, TJW^{+}22, TMM^{+}18, USRDFO^{+}22, VMBT24, VKP^{+}24, WYG^{+}23, WWW^{+}24, WLN^{+}23, XWW^{+}24, XXL^{+}24, YMD^{+}21, YLL22, YZL^{+}23, YSG^{+}23, ZLW^{+}23, ZJJ^{+}22, ZLF23, ZQL^{+}23, ZZT^{+}23, ZCT^{+}23, ZGY^{+}23, ZSL^{+}23, ZJM^{+}23, dSSBdS23]. Analyze [GFDPSR22]. Anchovies [DQC^{+}23]. Anchovy [AAN22, FZZ^{+}24, LZZ^{+}22, ŠT19]. Andrias [CWW^{+}23, ZHF^{+}22b]. Anemone [HZS^{+}21]. Anemonefish [HZS^{+}21]. Anesthesia [BSH^{+}23]. Anesthetic [JBdFS^{+}22]. Anesthetics [LF24, SCCM23]. Angel [UGH^{+}24]. Angler [LLTM17, PB24]. Angola [BC23]. Anguilla [BGMM^{+}24, DMA22, DCL^{+}23b, EFG^{+}23, HMVRFD19, JKP^{+}23, LPB^{+}24, MLK^{+}19, SLYY23, WHX^{+}23, ZWZ22]. anguillicaudatus [ZSL^{+}23]. Animal [BCG^{+}23, SdSdOSSL23]. Animal-Based [BCG^{+}23]. Animals [KKPL23]. Anisakid [DMT^{+}19]. Anisakis [ŠT19]. Anisotremus [BPO19]. Annual [MKC^{+}22, PPAB^{+}18]. Anodonta [CLL^{+}23b]. Antarctic [MPM^{+}18]. Anthropogenic [MH23]. Anti [CMP^{+}23, TCV^{+}19]. Anti-Mullerian [CMP^{+}23]. Anti-Osteogenic [TCV^{+}19]. Antibacterial [EES^{+}23, JBdFS^{+}22, dCPRG^{+}21, WML^{+}21]. Antibiofilm [dSJC^{+}21]. Antibiotics [JSLE23, LXW^{+}24]. Antigens [ZLSB22]. Antimicrobial [CAC^{+}17, DO24, GMBR^{+}21, ZAM^{+}23, dSJC^{+}21]. Antimycin [BBCJ23]. Antinematode [NWN^{+}22]. Antioxidant [AZY^{+}24, AEME^{+}23, CXW^{+}23, ERE21, HGSE23, HFEH^{+}23, JKK24, JBdFS^{+}22, LPF^{+}23, LWZ^{+}22, LHX^{+}23, LCX^{+}23, MWZ^{+}23, RIF^{+}23, RHU^{+}23, RXL^{+}24, SGG^{+}21, TDN^{+}22, WSD^{+}23, WZL^{+}24, WLL^{+}24, WLZ^{+}22b, XYC^{+}22, XYC^{+}23, XCW^{+}23, XYL^{+}23, XWP^{+}24, YLW^{+}23b]. Antioxidation [HYN^{+}24, ZYG^{+}23]. Antipredator [AWCS23, LYL^{+}23]. Antithrombotic [MKN^{+}16]. Antiviral [JML^{+}24, SML^{+}23b]. Antognazza [APD^{+}23a]. Ao [TTT23]. Apex [DMB^{+}20]. Aphanomyces [PSS^{+}18]. Apoptosis [LYL^{+}24a]. Apostichopus [YBL^{+}22]. Appalachian [UH19]. Applicability [AGC23]. Application [APD24, BLA^{+}22, CHZ^{+}24, CZCW23, GHS20, JTS^{+}24, QPDGF^{+}23, SOW^{+}23, SMH^{+}22b, TDN^{+}22, YCL^{+}23]. Applications [DO24, Fer23, LJR^{+}24, LYY^{+}24, LW18, SDA23]. Applying [GL23]. Approach [CKMT23, CVANRD^{+}21, CBCL23, FM21a, MBP^{+}24, QL22, RDANPA^{+}24, SCSR22, ST17, Sha19, SS23, VPPF^{+}19, ZTWK24]. Approaches [Dul23]. Approaching [CFM^{+}23]. Aqua [DCL^{+}23a, SAB^{+}22]. Aqua-Ento-Ponics [SAB^{+}22]. Aquaculture [AJF23, BÁD^{+}22, CG20, CMB^{+}24, CLZ^{+}24, DCL^{+}23a, DÁ23, DO24, GHS20, GLH^{+}23, HMN^{+}22, HARB23a, HHP^{+}24, HGC^{+}23, HIO^{+}19, IKT19, IFA^{+}23, Ims23, IBN^{+}23, JE18, JBdFS^{+}22, KYB^{+}23, LCW^{+}23b, LZC^{+}23a, LWZ^{+}23, LYY^{+}24, MKH24, MIHH23, NNL^{+}23, OOTC24, OO24, PLV^{+}19, QPDGF^{+}23, RXL^{+}24, RE21, SGANM^{+}24, SIZ^{+}22, VMDV^{+}22, VRG^{+}24, WM22, WLZ^{+}23, WXL^{+}23, XLP23, YCL^{+}23, ZSH^{+}23, ZLF23, ZTWK24]. Aquaculture-Triggered [RE21]. Aquafeed [BF16]. Aquafeeds [FLB^{+}21]. Aquaponic [NSK^{+}23, PLC^{+}24]. Aquaponics [MWZ^{+}23, YCL^{+}23, ZWP^{+}23]. Aquaporin [ZLW^{+}23]. Aquarana [XWW^{+}24]. Aquarium [VM19]. Aquatic [CHZ^{+}24, Dul23, DD23, EM23, EBCM24, GMRJ22, HVRCG18, KKPL23, PSS^{+}18, SCHT23, dSSBdS23]. aquatica [RXL^{+}24]. aquimarina [YBL^{+}22]. Arachidonic [ASM^{+}22]. Aral [APD24]. ArcFace [LWZ^{+}23]. Archipelago [JJB^{+}24]. Archived [LJK^{+}22]. Arctic [HJE^{+}23, SBG^{+}24, WGRM^{+}19]. Area [CFM^{+}23]. Areas [CXL23, FYH^{+}23, FCF19, FJL^{+}23, JLT22, UGH^{+}24, VKP^{+}24]. areolata [DLL^{+}22]. argenteus [GP17, ZSZ^{+}23]. Arginine [HAdM^{+}24]. Arginine-Vasotocin [HAdM^{+}24]. Argopecten [BÁD^{+}22, CPVMA^{+}24, HHM^{+}24]. argus [JSJ^{+}24, LWL^{+}23]. Argyrosomus [DDN19, GSHGE18, RJMVC^{+}18, RJTVC^{+}19, TPC^{+}23, VMDV^{+}22]. ariakensis [GLY^{+}23, WLC^{+}23]. Ariopsis [OJC^{+}23]. Arrhenius [CTY^{+}21]. Arsenic [RKHAMM22]. Artemia [SdSdOSSL23]. Arthromitus [RCR^{+}23]. Artificial [BEF^{+}23, CZCW23, GSH^{+}24, HLK^{+}23, JKP^{+}23, JCL^{+}24, LPB^{+}24, LDX^{+}23, SCBSSMA24, ZLX^{+}23]. Artificially [HAdM^{+}24, MLK^{+}19]. Artisanal [MBC^{+}24]. ASC [YSG^{+}23]. Ascending [PRPW23]. Ascidian [TPA^{+}24]. Ascorbate [OMC^{+}19]. ASCT2 [MLZ^{+}21]. Ash [WZX^{+}24]. Asia [KZC^{+}24]. asiaeorientalis [CYL^{+}23, CLL^{+}24, CYL^{+}23, CLL^{+}24, FLX^{+}22]. Asian [LCZ^{+}23, NPT^{+}24]. Asiatic [JXW^{+}23]. Aspartate [GSHGE18]. Aspects [Mil23]. Aspergillus [EAJ^{+}23]. Assay [CMP^{+}23, DDN19, SdSdOSSL23, SOW^{+}23]. Assemblage [RJR^{+}22]. Assemblages [GTC^{+}17, MWPS23, MWPS24, PYP17, SMAR24]. Assess [WMZ^{+}22]. Assessing [BASBW24, CSL24, CBP^{+}24, EMFZ^{+}18, FGG^{+}22, HLK^{+}23, ISA^{+}22, LSL^{+}24, NPT^{+}24, SCBSSMA24, VM19, ZSH^{+}23, ZJC^{+}22b]. Assessment [AIŠRB22, BSR^{+}20, BLA^{+}22, BFB^{+}23, CKMT23, CCCFE18, CMB^{+}24, HCWH20, HKS^{+}18, KTT24, KJ22, KSWT22, LDBL19, MSK^{+}22, MKN^{+}16, MBL20, NHC^{+}23, SALC^{+}19, SLS^{+}22, WM22, WFZ^{+}23, WFZ^{+}24b]. Assessments [EBM24, SWS22]. Assisted [LDX^{+}23]. Asso [DDN19]. Associated [ATM^{+}24, AVT18, FCB^{+}21, GMCF^{+}22, JKP^{+}23, KJ22, LJK^{+}22, LCWH22, NHC^{+}23, RWF^{+}23, SLY^{+}21, SNZ^{+}23, UDG^{+}19, WGW^{+}23, WHY^{+}24, ZVRH23]. Association [HMP^{+}24, LZL^{+}23]. astaci [PSS^{+}18]. Astaxanthin [WLL^{+}24]. Asterias [WKB^{+}23]. Astragalus [PYJ^{+}23, XCW^{+}23]. Asymmetric [XLC^{+}24]. Asymmetry [MH23]. Asynchronous [GFDPSR22]. Atlantic [EKL^{+}23, GPD^{+}23, MSRGCG^{+}23, ANA^{+}23, AGC23, AHLC19, AKM23, BCG^{+}23, BMTR23, CG20, EKL^{+}23, FHF23, GP17, GLH^{+}23, IFA^{+}23, IR22, IM24, LOB^{+}23, MIHH23, MLSC^{+}23, MPK^{+}23, PSO^{+}19, PH23, SML^{+}23a, SCSS23, SGG^{+}21, Tri23, YHH^{+}20, vKRNL^{+}19]. Atoll [LL22]. Atomic [CCCFE18]. ATPases [LGZ^{+}23]. Atractosteus [FQÁGTR^{+}17, MPM^{+}21, MVPMAV^{+}22, NRÁGPM^{+}18, PJPMMV^{+}22, dlCRMB^{+}22]. Atrazine [MTHPJS^{+}23]. Attention [EFG^{+}23]. Attenuation [KCCK23]. Attractive [HLK^{+}23, SRL^{+}19]. Attractiveness [BHP^{+}24]. Attributes [LXT^{+}22, PYP17]. Auction [JBGCG^{+}24]. Augment [SOW^{+}23]. Aurantiochytrium [HMP^{+}24]. aurata [BSGMC^{+}22, CSB^{+}20, CCCFE18, CFCE20, ERE21, GMBR^{+}21, GMDMT^{+}23, TJTV^{+}23, TMD^{+}19]. auratus [ULR^{+}23, BF16, HGSE23, LHZ^{+}23, LCC^{+}24, MWZ^{+}23, XHC^{+}22, XLX^{+}22, XYC^{+}22, XYC^{+}23]. australasicus [KMSO18]. Australia [MS20, RKL24]. Australian [BF16]. australis [WLL^{+}22]. Australoheros [BSR^{+}20]. Autism [VLCCA^{+}23]. Automated [JBGCG^{+}24]. Automatic [KMSO18, EEE21, VPPF^{+}19]. Automating [PSP^{+}22]. Autophagy [MDVM^{+}23]. Auxis [ZCX^{+}22].

Availability [MPM^{+}21, SMO^{+}23]. Aveiro [XQLA18]. avenaceum [PSS^{+}18]. Avian [VF23]. avidus [CCP^{+}24]. Axis [CZW^{+}23, LPB^{+}24]. Azolla [DRH18].

B [RSA17, ZSL^{+}24]. Babitonga [CFM^{+}23]. Babylonia [DLL^{+}22]. Bacillus [HXS^{+}23, JKP^{+}23, SAC23, CBCL23, GCFA^{+}22, JCY^{+}23, LJR^{+}24, WFZ^{+}24a, XYC^{+}22, XYC^{+}23]. Bacillus-Supplemented [JKP^{+}23]. Back [QWR^{+}23]. Back-Reef [QWR^{+}23]. Bacteria [CSGE23, DO24, KSO^{+}23, dSJC^{+}21]. Bacterial [AML^{+}24, EES^{+}23, GLX^{+}23, GSH^{+}24, HPJ^{+}23, LS19, LJR^{+}24, SNZ^{+}23, ŢPN^{+}23]. Bad [SBG^{+}24]. baerii [AVT18]. Baghdad [MS18]. Bagre [OJC^{+}23]. Bagrid [HZZ^{+}23]. Bait [XQLA18]. Baits [DZC^{+}22]. Bala [PKV^{+}22]. Balance [LCX^{+}23]. Balancing [CL21]. Balantiocheilos [PKV^{+}22]. Balearic [CSSMV^{+}23, FPRAT^{+}23, TCB^{+}24]. Bali [ST17, SZSW21]. Balistes [HS24]. Balitoridae [GWH^{+}22, LZC^{+}23b]. Balkan [APS^{+}23]. Baltic [AIŠRB22, KAR^{+}23, Ols19, VKP^{+}24]. Bambaranut [EPKV17]. Ban [ZZW^{+}24]. Bangladesh [BLA^{+}22, FM21a, HHM^{+}18, HVRCG18, HKS^{+}18, RKHAMM22, SLS^{+}22]. Barb [DRH18, HKS^{+}18]. Barbel [SRBCGV^{+}21, SPQ^{+}24]. Barbonymus [DRH18]. Barcoding [BAA^{+}19, OJC^{+}23, SBS^{+}24]. Bark [PSS^{+}18]. Barracuda [FB22]. Barred [CHH^{+}23, NNN23]. Barrier [CL19, MVPMAV^{+}22, NRÁGPM^{+}18, PJPMMV^{+}22, WLW^{+}23]. Barriers [Sus20, Zac22]. bartramii [XLC^{+}24]. Based [BCG^{+}23, BLA^{+}22, CHW21, CMP^{+}23, CTY^{+}21, CSJ^{+}23, FYL^{+}23, FCT19, GP17, HS24, HS18, KBK^{+}23, KBB^{+}21, KSWT22, LSJ^{+}22, LJX^{+}21, LV23, LSJ24, MSK^{+}22, MSA24, MGMG24, NPT^{+}24, RJMVC^{+}18, SFK^{+}23, SZT^{+}23, SXQ^{+}22, UGH^{+}24, WGRM^{+}19, WLZ^{+}22a, WYL23, WSA^{+}23, WHM21, WFL^{+}23, XLC^{+}24, YYW^{+}23, YWL^{+}24, ZLF23, ZXH^{+}23, ZJM^{+}23, ZWD^{+}23, dSSBdS23, SLS^{+}22]. Basic [DAF^{+}22]. basilicum [VAT^{+}23]. Basin [BTSK24, BAA^{+}19, FWJ21, GAR^{+}24, SKD^{+}23, APD24, PCH^{+}24, RKL24]. Basis [MBPB24, NTP^{+}21]. Bass [AHLC19, BGMM^{+}24, CFCE20, CMC^{+}24, HCWH20, HBL^{+}22, KKPY22, LZH^{+}23, RSA17, SOW^{+}23, SAB^{+}22, SJY^{+}22, SHW^{+}23, WSZ^{+}23, XYL^{+}23, XXL^{+}22, YSG^{+}23, ZZY^{+}23]. Batch [DTS^{+}17]. Bathed [XXL^{+}24]. Bathing [XJC^{+}22]. Bay [HHAG24, HWW^{+}24a, HHM^{+}24, MMSK21, SXQ^{+}22, BLA^{+}22, KSAB^{+}23, SLS^{+}22]. Bayesian [SLS^{+}22]. BBNJ [QL22]. Be [NdNFK^{+}24b, RJMVC^{+}18, SB20, ZS21, SMM^{+}18]. Beaches [JSSD23, RGTPSCCC24]. Beak [CLL^{+}22]. beani [MCÁGHA^{+}17]. Bearded [CSJ^{+}23]. Bed [SLSC^{+}24]. before [ANA^{+}23]. Behavior [CZCW23, CHH^{+}23, FÅG^{+}23, dFBdSG^{+}19, HV24, ILA22, KTT24, LMC21, LYL^{+}23, LLS^{+}23, Mee24, SZF^{+}21, SZT^{+}23, VK23, VLCCA^{+}23, WTC^{+}22, WWD^{+}23, ZSH^{+}23]. Behavioral [AWCS23, CFLK21, CMP^{+}20, CXT^{+}24, MU21, SCSS23]. Behaviors [KKPY22, LHZ^{+}24]. Behaviour [DSC^{+}19, JBK^{+}23, LCW23a, LMB^{+}23]. Behavioural [AAB^{+}18, KBCM19, ZSZSS^{+}22]. Behaviours [EM23]. Beibu [LWD^{+}23, SXQ^{+}22]. Being [SEA^{+}23]. Bellied [STZ^{+}23, ZQLW23]. Benchmark [ANA^{+}23]. Beneficial [SNK^{+}23]. Benefits [SFK^{+}23]. Bengal [BLA^{+}22, KSAB^{+}23, SLS^{+}22]. Benin [AHK^{+}23]. Bennett [SCT^{+}24]. Benthic [KRAFO23]. Berg [MNP^{+}16]. bernacchii [GGF^{+}22]. Berryteuthis [LOT^{+}22]. Bertalanffy [BMTR23]. Berwickshire [EBRS23]. Best [CBANCM^{+}21]. Bester [Kim23a, Kim23b]. Beta [LHL^{+}24]. Betanodavirus [VPP^{+}22]. Better [AGC23]. between [AGE^{+}18, AJF^{+}22, BC23, CYL^{+}23, CP24, CFLK21, HZS^{+}21, HWZ21, KAB^{+}23, KKPY22, KMSO18, LCW^{+}23b, LYL^{+}23, LYL^{+}24b, NHR20, PKC^{+}19, SML^{+}23a, TRM^{+}23, WM22, WHL^{+}24, WWY^{+}23, XWW^{+}24, YJK23]. Bias [RH19, RRG22]. bicuspidata [LYW^{+}24]. bidens [YXS^{+}24]. Biferno [RDE^{+}23]. Big [Tay19]. Bigeye [PBS^{+}22, ST17, USRDFO^{+}22]. Bigheaded [BGT^{+}20, Sus20, ZS21]. biloba [CXW^{+}23]. bimaculatus [MSB^{+}23]. Binding [BBCJ23]. Bio [ZWD^{+}23]. Bio-Products [ZWD^{+}23]. Bioaccumulation [JKK24, WFZ^{+}23]. Bioacoustics [LL18a, LL18b]. Bioactive [DDG^{+}22]. Bioactivities [NWN^{+}21]. Biochemical [DCR^{+}23, FMXQ23, KŽM^{+}23, OSM23, PAVCCJVV24, RMA^{+}18, SAGG^{+}23, THS^{+}22, XLX^{+}22, YLH^{+}24, ZWZ22]. Biochemistry [BDŞ^{+}24, LHX^{+}23, SNK^{+}23, ZLX^{+}23, dAdSCC^{+}23]. Biocontrol [MS20]. Biodistribution [CFP^{+}23]. Biodiversity [GSK^{+}21]. Bioeconomic [FHHC23, HAC^{+}23]. Biofloc [CdOCH^{+}23, dOCCH^{+}23, GL23, LZC^{+}23a, MSB^{+}23, RHU^{+}23, YCL^{+}23]. Bioimpedance [KAB^{+}23]. Bioindicator [LL22]. Biological [BLA^{+}22, dOCCH^{+}23, EEdCSOPJ^{+}23, KŽM^{+}23, RTBL^{+}18, TJW^{+}22, TMPP23]. Biology [HHM^{+}18, HS24, Sor21, WLL^{+}23]. Biomarker [LMLH22]. Biomarkers [ADM^{+}24, CMC^{+}24, GRKC19, OOTC24]. Biomass [ATEfA^{+}21, JSSD23, SLS^{+}22, TMPP23]. Biomedical [LW18, SDA23]. Bionic [WMZ^{+}22]. Bioproduction [NWN^{+}21]. Bioremediation [AZY^{+}24, GLH^{+}23].

Biosecurity [MDF^{+}23]. Biosensors [DD23]. Biotechnology [Hal23]. Biotic [CH23, SXQ^{+}22]. Biotransformation [KSSI24]. Biozonation [APS^{+}23]. birdiae [HSAF^{+}23]. Bisphenol [vKRNL^{+}19]. Bites [CBP^{+}24]. Bivalent [XYF^{+}24]. Bivalvia [ZQL^{+}23]. Black [BAA^{+}23, BLA^{+}22, BPO19, CSR22, LWT^{+}24, LZW^{+}24, MMY^{+}17, NNL^{+}23, NYS^{+}23, RE21, SZF^{+}21, SMH^{+}22a, SSK^{+}22, SSSS23, ŢPN^{+}23, TMPP23, WWW^{+}24, YAEAB23, ZJM^{+}23, BAA^{+}23, DDD^{+}23, SSSS23]. Blackmouth [BMMD22, DAF^{+}22]. Blackspot [LDBL19, RJFCJC^{+}21]. Blastema [VCL20]. Bleaching [HZS^{+}21]. Bleak [HV24]. bleekeri [LDX^{+}23, WWY^{+}23]. Blend [SRL^{+}19]. Blends [GMFG^{+}24]. Blimp [RGABD20]. Bloch [GMA^{+}24]. Blocked [ZS21]. Blood [CYL^{+}23, CZJ^{+}24, EAJ^{+}23, FLX^{+}22, FMXQ23, GMDMT^{+}23, INCD23, SNK^{+}23, TMM^{+}18, ZLX^{+}23]. Blood-Chemistry [CZJ^{+}24]. Blooms [DÁ23, SAHS18]. Blotched [ZWZ^{+}24]. Blue [And23, DAHM19, DM24, FHHC23]. Bluegill [CH23]. Blunt [GWT^{+}24, MYW^{+}24]. BMP [ZYH^{+}24]. BMP2 [LCC^{+}24]. BMP4 [LCC^{+}24]. Bocachico [MMMNAG23]. Body [EAJ^{+}23, dSGBdF23, LFH^{+}23, SCSS23, ULR^{+}23, WSE^{+}21, YBL^{+}22, ZWZ22]. bogaraveo [RJFCJC^{+}21]. Bolbometopon [TCD^{+}21]. Boldness [AJF^{+}22]. bonariensis [BSH^{+}23, GRKC19]. Bonefish [KBK^{+}23]. Boone [HSAF^{+}23, NdNFK^{+}24b, NdNFK^{+}24a]. Boreogadus [KWK^{+}24, YHH^{+}20]. Borne [Sha19]. Both [Poi24, YXS^{+}24]. Bottle [SLSC^{+}24]. Bottom [CLL23a]. Bowdich [DTS^{+}17]. Box [LJX^{+}21, TOB^{+}23]. Bracciano [SAL18]. Brachymystax [PXH^{+}24]. brachysoma [JTS^{+}24]. Brachyura [KMB18, KKB20]. Brackish [EFG^{+}23, YJK23]. Brain [CBK^{+}21, DAHM19, DM24, LPB^{+}24, Näs18, PHB^{+}23]. Brain- [Näs18]. Brazil [CFM^{+}23, CASMK23, dSCFQB^{+}23, JSRE^{+}24, dAMPS^{+}23, dCPFS^{+}24, SNSVFL23]. Brazilian [ARH^{+}23]. Bream [AA23, CSB^{+}20, GWT^{+}24, GMFG^{+}24, HHL^{+}20, JJK21, KJK^{+}23, LSY^{+}17, MYW^{+}24, SNSG^{+}19, TMD^{+}19, ZMLFS^{+}20]. Breeders [GGP^{+}23]. Breeding [HZZ^{+}23, NTP^{+}21, SMH^{+}24, STZ^{+}23, WHY^{+}24, ZQLW23, ZHF^{+}22b]. Brine [WDL^{+}23]. Brine-Preserved [WDL^{+}23]. British [TBPJ23]. Bronze [ZHX^{+}24]. Broodstock [PSN18]. Brook [MH23]. Brotula [CSJ^{+}23, CSJ^{+}23]. Brotulella [CSJ^{+}23]. Brown [APD^{+}23a, APD^{+}23b, EAJG24, FÅG^{+}23, MSRGCG^{+}23, Näs18, PMFBI22, RDI^{+}21, RDE^{+}23, VMS^{+}23, WYG^{+}23]. Brown-Marbled [WYG^{+}23]. Bruguière [ADM^{+}24]. Brycon [DRFCL23, EMFZ^{+}18]. Bubbles [SIZ^{+}22]. Budgets [GZX^{+}22]. Building [HWW^{+}24a]. Bulatmai [SPQ^{+}24]. Bulgaria [APS^{+}23]. Bullfrog [XWW^{+}24]. Bullheads [SMH^{+}22a]. bungeanum [WZX^{+}24]. Burchell [BHR^{+}23]. Burden [EvSCB23]. Butachlor [WWZ^{+}24]. Butter [MSB^{+}23]. butyricum [WLW^{+}24]. By-Product [NWN^{+}22]. By-Products [AAAF^{+}21]. Bycatch [FM21a]. bZIP [SHW^{+}23].

C [BFM23, CSF^{+}23, LJK^{+}22, MYW^{+}24, WSZ^{+}23]. C-Type [CSF^{+}23]. Ca [JCR^{+}22, NZVB20]. Ca/P [JCR^{+}22, NZVB20]. cachuatsi [JML^{+}24]. Cádiz [MSRGCG^{+}23, RGTPSCCC24]. Cadmium [GWT^{+}24, JKK24, RLB^{+}23, SPQ^{+}24, TCV^{+}19, WCY^{+}24]. Cadmium-Induced [SPQ^{+}24]. Caeca [GMFG^{+}24]. Cage [NNL^{+}23, SNZ^{+}23, ZLF23]. Cages [AHK^{+}23, IBN^{+}23]. Calanus [BHP^{+}24]. Calcified [ZJC^{+}22b]. Caledonia [LCW23a].

Caliber [JBGCG^{+}24]. California [USRDFO^{+}22, AFTÁPA^{+}23, ASM^{+}22]. Caligus [PSO^{+}19]. Callinectes [KKB20]. Calmodulin [GWT^{+}24]. Cambeva [CFM^{+}23]. Caml [XSZ^{+}23, XSZ^{+}23]. Campeche [dRCCLRNRWK21]. Can [KSO^{+}23, NdNFK^{+}24b, NYS^{+}23, RJMVC^{+}18, SB20, VMS^{+}23, ZS21]. Canada [LOB^{+}23, TBPJ23]. canadum [MKC^{+}22]. Canal [BMMD22]. canaliculata [GMRJ22]. Canary [JFP^{+}18]. Cancer [EBRS23]. Cancridae [EBRS23]. Candidate [SLY^{+}21]. Candidates [WBK^{+}23]. Candidatus [RCR^{+}23]. canicula [BMSGS^{+}18]. Cannibalism [CZC23, CCFP19]. Cantharidin [CSGE23]. Capacities [LLS^{+}23, PKC^{+}19]. Capacity [CXW^{+}23, DK22, FCT19, GLH^{+}23, LWZ^{+}22, WZL^{+}24, WLL^{+}24, WLZ^{+}22b, XCW^{+}23, XYL^{+}23, XWP^{+}24, YLW^{+}23b, YGD^{+}23, ZLZ^{+}24]. capito [SPQ^{+}24]. Captive [BFM23, CYL^{+}23, CHH^{+}23, STZ^{+}23, VPPF^{+}19, ZQLW23]. Captivity [CLL^{+}24, GGP^{+}23, HAdM^{+}24, OGMG^{+}17, PPAB^{+}18]. Capture [RJFCJC^{+}21]. Captured [JTS^{+}24]. Car [KSAB^{+}23]. Carangidae [FGHYCA23, MBD^{+}23]. Carapace [XXL^{+}24]. Carassius [HGSE23, LHZ^{+}23, LCC^{+}24, MWZ^{+}23, ULR^{+}23, XHC^{+}22, XLX^{+}22, XYC^{+}22, XYC^{+}23, ZJC^{+}22a]. Carbohydrate [BSGMC^{+}22, YGD^{+}23]. Carbohydrate/Lipid [YGD^{+}23]. Carbon [DL23, GZX^{+}22, GSH^{+}24, INCD23, LZC^{+}23a, LCX^{+}23, RHU^{+}23, Sus20]. Carboxylase [ZYT^{+}24]. Carcharhinid [BSB^{+}23]. Carcharhiniformes [WSON24]. Carcharhinus [KSWT22]. carcharias [HL24]. Carcharodon [HL24]. Carcinus [YE20]. Cardiac [PLY^{+}24]. Cardiomyocytes [ZMA^{+}24]. CARES [VM19]. Caribbean [HS24, ZVRH23]. carinicauda [HWX^{+}23]. Carotenoid [GBT^{+}24]. Carp [BAP22, CXW^{+}23, CHJ^{+}23, GLW^{+}22, HHP^{+}24, HBG^{+}20, KBCM19, LHZ^{+}23, LCC^{+}24, LWS^{+}21, LWT^{+}24, LJR^{+}24, MLZ^{+}21, MS20, MGS^{+}23, NDC^{+}23, NPT^{+}24, RKHAMM22, RMA^{+}18, SRL^{+}19, SB20, Sus20, ULR^{+}23, WML^{+}21, WMZ^{+}22, WSD^{+}23, WZL^{+}24, WZX^{+}24, WWD^{+}23, XHC^{+}22, XLX^{+}22, XYC^{+}22, XYC^{+}23, XCW^{+}23, YWDP21, YXS^{+}24, YLX^{+}22, ZSQ^{+}21, ZLZL23, ZSL^{+}24, ZZC^{+}22, ZJC^{+}22a, ZS21]. carpio [CHJ^{+}23, LWT^{+}24, WSD^{+}23, BAP22, BDŞ^{+}24, CXW^{+}23, KBCM19, MS20, XCW^{+}23, YWDP21]. Carps [BGT^{+}20, PVY^{+}21]. Cas9 [KEA^{+}23]. Cas9-Mediated [XWR^{+}23]. Case [CMB^{+}24, CMC^{+}24, DNP^{+}23, GFDPSR22, Ho22, HAC^{+}23, HXY^{+}23, KAB^{+}20, LV23, MDF^{+}23, MFKS23, NNL^{+}23, NRKT19, PAMG19, RLAE23, RJFCJC^{+}21, SHT^{+}23, SB20, TVL21, TTT23, XQLA18, XLC^{+}24, XW24, XYT23, ZJC^{+}22a]. Caspase [LYL^{+}24a]. Caspase1 [YSG^{+}23]. Catabolism [TPC^{+}23]. Catch [CGY^{+}23, FGBA^{+}23, HWZ21, RH19, SHT^{+}23, WLM^{+}20]. Catch-per-Unit-Effort [SHT^{+}23]. Catches [LWD^{+}23, MGMG24, NGMR23, NNN23, PKSN23a, PKSN23b]. Catechol [KSSI24]. catesbeiana [XWW^{+}24]. Catfish [AZY^{+}24, AAAF^{+}21, BHR^{+}23, BBN^{+}24, Eny17, EPKV17, FHHC23, FGR19, HVRCG18, HWW^{+}24b, JLW^{+}24, KEA^{+}23, KBCM19, MSB^{+}23, OJC^{+}23, PYJ^{+}23, QXAY22, SBB^{+}19, SIZ^{+}22, VH20, WSE^{+}21, YLW^{+}23a, YAEAB23, ZLSB22]. Catfishes [CASMK23, HZZ^{+}23, PCO^{+}23]. Catla [PVY^{+}21]. Catostomus [RHUJ24]. Catshark [BMSGS^{+}18, BMMD22, DAF^{+}22]. Caudal [VCL20]. Caught [AJF^{+}22]. Caulerpa [LMNN21]. Causal [QXAY22]. Causative [LBH^{+}24]. Caused [GLX^{+}23, MYW^{+}24]. Causes [BKJ^{+}24, XLP23]. Causing [WCX^{+}24]. Cautious [SMM^{+}18]. Caves [KGP24]. caviae [XXL^{+}22]. Caviar [LVB^{+}20, RE21]. Cefotaxime [HSD^{+}24]. Cell [CSGE23, CLJ^{+}23, GHS20, QWY^{+}24, SMH^{+}24, ZQLW23, vKRNL^{+}19]. Cells [BMMD22, DCL^{+}23b, KKPY22, SMO^{+}22, VCL20, YXS^{+}24, ZMA^{+}24]. Cellular [LWS^{+}23b]. Central [BTSK24, CTTW23, DAF^{+}22, SS23, TK24, YYW^{+}23, BFB^{+}23, KBK^{+}23, KZC^{+}24, NHC^{+}23, VV23]. Cephalophilis [BPOS19]. cephalus [ATEfA^{+}21, EES^{+}23, GL23, GMDMT^{+}23, MABÁMSM22, MBAM19]. Cerastoderma [ADM^{+}24]. cereus [JCY^{+}23]. cernua [NHR20]. Certain [SSK^{+}22]. cettii [RDE^{+}23]. cf. [BMOH23]. Chad [OOGAS23]. Chaetomorpha [GBT^{+}24]. Chain [CMP^{+}23, EN22, KLD^{+}23, SLC^{+}22, TRM^{+}23, YLL22]. Challenge [APS^{+}23, CAC^{+}17, PYJ^{+}23, RIF^{+}23, SNSG^{+}19, ZSL^{+}24]. Challenged [EAJ^{+}23, GCFA^{+}22, HMP^{+}24, JLW^{+}24, JCY^{+}23, LDD^{+}22, NDC^{+}23, SSSP21]. Challenges [And23, BKJ^{+}24, DSC^{+}19, GWT^{+}24, SS23]. Chamber [CFLK21]. Champlain [YAS21]. Change [CS23, HB24, HHAG24, KKPL23, Kim23a, Kim23b, MSK^{+}22, MML22, NPT^{+}24, SMO^{+}23, SRHCO23, SXQ^{+}22, SBG^{+}24, VSH23, WSA^{+}23, dlCRMB^{+}22]. Changed [ANA^{+}23]. Changes [AGE^{+}18, CGSBGN24, CYL^{+}23, CBK^{+}21, DAHM19, ERE21, Fra23, GLX^{+}23, MJL^{+}24, MSB^{+}23, MAR^{+}18, MLK^{+}19, NGMR23, RJR^{+}22, RGVG19, SPJ^{+}24, TU18, THS^{+}22, TJW^{+}22, WKB^{+}23, WM22, WDL^{+}23, WSS^{+}19, YLH^{+}24, ZSZSS^{+}22, ZZW^{+}24, HMVRFD19]. Changing [CL19, FÅG^{+}23, HMX^{+}21, LOB^{+}23]. Channa [ZWZ^{+}24]. Channel [FHHC23, KEA^{+}23, KBCM19, QXAY22, SYL^{+}24, YAEAB23, ZLSB22, DMA22]. Characteristics [CLL^{+}24, DSC^{+}23, Gae16, Gae17, HHP^{+}24, HSZ^{+}22, HPJ^{+}23, KCCK23, LHG^{+}23, SMH^{+}22a, SJY^{+}22, WGW^{+}23, WXL^{+}23, WDL^{+}23, XZZ^{+}24, XXL^{+}22, YCR^{+}23, YYH^{+}24, ZQLW23, ZZW^{+}24, ZWZ^{+}24]. Characterization [APD24, AM24, BMMD22, CWW^{+}23, CLJ^{+}23, FCB^{+}21, GWT^{+}24, GMA^{+}24, KJK^{+}23, LMLH22, LLY^{+}22, LWS^{+}21, LYL^{+}24a, MdM^{+}23, MLZ^{+}21, MCÁGHA^{+}17, MPM^{+}18, NWN^{+}23, PBS^{+}22, WLC^{+}23, WHY^{+}24, WCX^{+}24, XHC^{+}22, YSG^{+}23, ZLSB22, ZSZ^{+}23, ZLZL23, ZSL^{+}24, ZZC^{+}22, ZHX^{+}24]. Characterized [BAA^{+}19]. Charr [SBG^{+}24, WGRM^{+}19]. Checklist [GSK^{+}21, GASS^{+}22]. Chelidonichthys [FDM^{+}23a]. Chelon [GMCF^{+}22]. Chemical [EEdCSOPJ^{+}23, LLY^{+}24, SGG^{+}21]. Chemical-Structural [EEdCSOPJ^{+}23]. Chemistry [CZJ^{+}24]. Chemokine [YLX^{+}22]. Chemotherapeutants [MKH24]. Chestnut [CSV^{+}19]. Chief [Hal23]. Chile [AAN22, BÁD^{+}22]. Chilean [AFTÁPA^{+}23, BS23, FGHYCA23]. chilensis [AFTÁPA^{+}23]. Chilled [WDL^{+}23]. Chimaeras [SDA23]. Chimaeridae [SQ23]. China [CLL^{+}22, HWZ21, HSZ^{+}22, LWD^{+}23, ZCX^{+}22, CHW21, FZZ^{+}24, GWH^{+}22, HXY^{+}23, LKJ22, LCZ^{+}23, LLWW23, LFH^{+}23, LHL^{+}24, LJP^{+}22, LXT^{+}22, LZL^{+}23, QXAY22, RJR^{+}22, STZ^{+}23, WM22, WLL^{+}23, WFL^{+}23, WWY^{+}23, WFZ^{+}23, WFZ^{+}24b, XW24, XLP23, XZZ^{+}24, XJC^{+}22, ZLF23, ZQLW23]. chinensis [GYH^{+}23, TJW^{+}22]. chinensis-Supplemented [GYH^{+}23]. Chinese [AML^{+}24, CZC^{+}22, CWW^{+}23, CZJ^{+}24, GLX^{+}23, HCZ^{+}23, JLT22, LWZ^{+}22, LXT^{+}22, LZM^{+}23, LLS^{+}23, PLY^{+}24, PNW^{+}22, WHL^{+}24, WWZ^{+}24, XZD^{+}24, XXL^{+}24, XJC^{+}22, YMD^{+}21, YXS^{+}24, ZHF^{+}22b, ZLX^{+}23, ZGY^{+}23]. Chionoecetes [BASBW24, FBSB24].

Chirolophis [LPK^{+}23a]. Chitinase

[HGC^{+}23]. Chitosan

[MMVVCJ^{+}23, RIF^{+}23]. Chlorella [AZY^{+}24, Eny17]. Chloride [BF16]. Chlorophyll [WLM^{+}20]. Choi [CPJK23a]. Cholinergic [ZAM^{+}23]. Chondrichthyans [GSK^{+}21, GASS^{+}22]. Chondrichthyes [ARH^{+}23, BMMD22, GAR^{+}24, VV23]. Chromatography [JSLE23]. Chromatography-Tandem [JSLE23]. Chromium [HYN^{+}24]. Chronic [CSL24]. Chrosomus [BMOH23]. chuatsi [DZC^{+}22, LMLH22, XZD^{+}24]. Chub [CKMT23, CPJK23a, CPJK23b]. Chum [WLL^{+}23]. Cichlasoma [MCÁGHA^{+}17]. Cichlid [JHNF24, LBH^{+}24, MCÁGHA^{+}17, SSSP21]. Cimei [SLL22]. Cinnamon [JBdFS^{+}22]. Circadian [WLN^{+}23]. Circle [KMS^{+}17]. Circularity [CMB^{+}24]. Cirrhinus [PVY^{+}21]. Cis [XWR^{+}23]. Cis-Regulatory [XWR^{+}23]. Citizen [GSK^{+}21]. Citizen-Science [GSK^{+}21]. Citral [MMM^{+}24]. Citral-Supplemented [MMM^{+}24]. Citronella [GMBR^{+}21]. Clam [FCB^{+}21, GGP^{+}23, GWH21, JXW^{+}23, LLY^{+}22, PSW^{+}23, SML^{+}23a]. Clarias [AZY^{+}24, AAAF^{+}21, BHR^{+}23, Eny17, EPKV17, SBB^{+}19, WSE^{+}21]. Clarifies [LBC^{+}24]. clarkae [CSJ^{+}23]. Class [GLY^{+}23, JSLE23, ZSL^{+}24]. Classe [RGVG19]. Classes [VKP^{+}24]. Classification [JBGCG^{+}24, JTS^{+}24]. clavata [FPRAT^{+}23, SSK^{+}22, SSSS23, TMPP23]. Cleaner [Ims23, IBN^{+}23]. Clearhead [TJW^{+}22]. Climate [ANA^{+}23, CS23, FÅG^{+}23, HB24, HMX^{+}21, Ho22, LUM18, LZZ^{+}22, LOB^{+}23, MSK^{+}22, MML22, NPT^{+}24, SMO^{+}23, SRHCO23, SBG^{+}24, VSH23, WZG^{+}23, WSA^{+}23, dlCRMB^{+}22]. Climate-Driven [LZZ^{+}22]. Climbing [PRPW23]. Clinch [BMOH23]. Clinical [AMT^{+}24]. Cloning [FLR^{+}22, LYL^{+}24a, LXZ^{+}22, MKC^{+}22, WPLK23, YLX^{+}22, ZHF^{+}22a, ZSL^{+}24]. Close [Tri23]. Close-Kin [Tri23]. Closed [KTT24]. Clostridium [WLW^{+}24]. Clupea [BMTR23, MNP^{+}16]. Clupeid [DTS^{+}17, VKP^{+}24]. Clupeidae [MNP^{+}16]. Clupeiformes [MNP^{+}16]. CMSY [AIŠRB22]. Co [CMC^{+}24, EES^{+}23, MTPK23, SAB^{+}22, WCY^{+}24, XYC^{+}22, XYC^{+}23]. Co-Culture [CMC^{+}24, SAB^{+}22]. Co-Exposure [WCY^{+}24]. Co-Fermented [XYC^{+}22, XYC^{+}23]. Co-Infection [EES^{+}23]. Co-Operation [MTPK23]. CoA [ZYT^{+}24]. Coast [FGG^{+}22, MPK^{+}23, ARH^{+}23, EKL^{+}23]. Coastal [AKM23, CLZ^{+}24, FZZ^{+}24, LKJ22, LCZ^{+}23, LZL^{+}23, MD21, NIN^{+}19, Ols19, RGABD20, XQLA18, XW24]. Cobia [MKC^{+}22, SCCM23]. Cobitidae [FŠS^{+}23]. Cocaine [CGSBGN24, RCL^{+}23]. Coccidian [DCL^{+}23b]. Cod [KWK^{+}24, MKN^{+}16, PH23, SCSS23, YHH^{+}20, vKRNL^{+}19]. Coelomic [WKB^{+}23]. Coercion [MTPK23, QAC23]. Coexistence [MD21]. Coexisting [RGTPSCCC24]. Cognition [QAC23]. Cognitive [JHNF24]. Coho [PSO^{+}19, ZYG^{+}23]. COI [WFL^{+}23]. Coilia [WTC^{+}22, WHY^{+}24]. Coinfected [XZD^{+}24]. coioides [ZHQ^{+}23]. Cold [LOB^{+}23, PBS^{+}22, ZYT^{+}24]. Cold-Water [LOB^{+}23]. Colder [SCSS23]. coli [KSO^{+}23]. Collagen [LWS^{+}21]. Collected [KYB^{+}23, MLH^{+}20]. Collection [RDI^{+}21]. Collichthys [XZZ^{+}24]. Colombian [SRHCO23]. Color [HHM^{+}24, PNW^{+}22]. Coloration [WLL^{+}24]. Colossoma [dAPAA^{+}24, SBS^{+}23]. Colours [FŠS^{+}23]. Columbia [TBPJ23]. Column [INCD23]. Comber [MSRGCG^{+}23]. Combination [HLK^{+}23, JSJ^{+}24, ZS21]. Combine [SMAR24]. Combined [GZX^{+}22, HYN^{+}24, PLC^{+}24, SSSP21].

Combining [AMK23]. Commercial [BHR^{+}23, BKBR^{+}23, IBN^{+}23, ISA^{+}22, KŽM^{+}23, MMAO22, TRM^{+}23, VMDV^{+}22, WZG^{+}23, WFL^{+}23]. Commercially [RKHAMM22]. commerson [NNN23]. Commission [SS23]. Common [BEMC23, KBCM19, MS20, RKL24, RGVG19, SB20, WKB^{+}23, YWDP21]. Commonalities [Mil23]. Communities [AWCS23, AGE^{+}18, ENO21, EFG^{+}23, GLX^{+}23, HXS^{+}23, HPJ^{+}23, LHL^{+}24]. Community [AML^{+}24, GSH^{+}24, HHAG24, HSZ^{+}22, KBB^{+}21, MBC^{+}24, YLH^{+}24, ZZW^{+}24]. Community-Based [KBB^{+}21]. Comparative [AEME^{+}23, FLX^{+}22, GWH21, IFA^{+}23, KSSI24, LYL^{+}24b, MSK^{+}21, MTM^{+}19, RWF^{+}23, SLY^{+}21, SJY^{+}22, TMM^{+}18, YZL^{+}23, ZZ24]. Compared [AJF23]. Comparing [EBRS23]. Comparison [AML^{+}24, CZJ^{+}24, HZS^{+}21, KZC^{+}24, LZZ^{+}22, MBC^{+}24, MMY^{+}17, SEA^{+}23, WHL^{+}24, YJZ^{+}24, YHZ^{+}23, ZLX^{+}23, ZZY^{+}23]. Compatibility [JLT22]. Compensation [SYL^{+}24]. Compensatory [Näs18]. Competition [AGE^{+}18, NHR20, YJK23]. Competitive [PB24]. Competitiveness [EvSCB23]. Complete [JJK21, LPK^{+}23a, LPK^{+}23b, PKY^{+}23]. Completely [YLW^{+}23a]. Complex [AHK^{+}23, FWJ21, HBL^{+}22, LHZ^{+}24, SSK^{+}22]. Complexity [DSC^{+}19, HWW^{+}24a]. Complexone [ZJC^{+}22b]. Compliance [EvSCB23]. Components [JKK24, MZA^{+}23]. Composition [AM24, AML^{+}24, BTSK24, BEG^{+}23, DZC^{+}22, EAJ^{+}23, FGBA^{+}23, GSH^{+}24, JSSD23, KŽM^{+}23, LZM^{+}23, PNW^{+}22, PAVCCJVV24, RGVG19, SGG^{+}21, TMD^{+}19, ULR^{+}23, WHL^{+}24, WSE^{+}21, XRX^{+}23, YJZ^{+}24, ZWZ22, ZHX^{+}24]. Compositions [WZL^{+}24, YYH^{+}24].

Compound [HXS^{+}23, ZWZ22]. Compounds [DDG^{+}22, WDL^{+}23]. Comprehensive [CS23, HGC^{+}23, MSRGCG^{+}23, PNW^{+}22]. Computer [JBGCG^{+}24]. Concentrate [CWP^{+}21, XWP^{+}24]. Concentration [dRCCLRNRWK21, PLC^{+}24, ZJC^{+}22b]. Concentrations [dOCCH^{+}23, KSO^{+}23, KJ22, MSK^{+}21]. Conch [SPJ^{+}24]. Condition [EHGS23, MH23, SGAC20]. Conditionally [CLJ^{+}23]. Conditioned [GGP^{+}23]. Conditioning [CPVMA^{+}24]. Conditions [CL19, CBCMRD^{+}23, EAJG24, EA18, GL23, GRO^{+}17, HYXY23, RCR^{+}23, STZ^{+}23, WMZ^{+}22, WXL^{+}23, ZQLW23]. Confronted [vSRB^{+}18]. Congruent [HBL^{+}22]. Connectivity [KBK^{+}23, QWR^{+}23]. Consequences [DCD^{+}24, Dul23, RCL^{+}23]. Conservation [APD24, ANA^{+}23, BMOH23, DNP^{+}23, EBM24, FPRAT^{+}23, JLT22, KSWT22, LKD22, MMAO22, PCO^{+}23, SKSL23, SCBSSMA24, VM19, VMBT24]. Conservation-Status [MMAO22]. Conservative [FM21a]. Consistent [AWCS23]. Conspecific [LYL^{+}23]. constricta [LLY^{+}22]. Constructed [KMS^{+}17]. Consumed [DMT^{+}19]. Consumer [RLAE23]. Consumption [AHLC19, GLW^{+}22, IFA^{+}23, LLY^{+}24, Poi24]. Containing [SBS^{+}23]. Contaminant [DCD^{+}24]. Contamination [LL22, ŢPN^{+}23]. Contemporary [RKL24]. Content [GRKC19, GBT^{+}24, MBAM19, RSA17]. Contents [OJC^{+}23]. Context [dlCRMB^{+}22]. Continental [dCPFS^{+}24]. Continuous [RGABD20]. Contributing [YZL^{+}23]. Contribution [LdRCV24]. Contributions [DDN19]. Control [CG20, GGP^{+}23, IR22, LXW^{+}24, Sor21, XLP23, YAS21]. Controls [WWD^{+}23]. Conventional [ZHHO23]. Conversion [NWN^{+}22, TDN^{+}22]. Convulsion [LKU21]. Coordinated [LJP^{+}22]. Copepods [OSM23]. Coping [GTC^{+}17]. Copper [CLL^{+}23b, WCY^{+}24]. Copper-Induced [CLL^{+}23b]. Coptis [GYH^{+}23]. Copy [EAJG24]. Coral [CZC23, CP24, LL22, SMAR24, ZVRH23]. Coral-Associated [ZVRH23]. Cord [LKU21]. Coregonus [BWKS20, LLKKV20]. Coreius [ZHX^{+}24]. Cornetfish [TK24]. Coronal [YCR^{+}23]. Correction [APD^{+}23a, CPJK23a, DL23, Gae17, Kim23a, MWPS24, PKSN23a, XYC^{+}23]. Correlation [CSB^{+}20, KKPY22, LPF^{+}23, LWS^{+}21, PKC^{+}19, XWW^{+}24]. Correlations [WWY^{+}23, YYH^{+}24]. Cortisol [WLT23]. Coryphaena [Poi24]. Cosmonauts [XSZ^{+}23]. Cost [APAHBMAG23, DRH18, MSB^{+}23, NYS^{+}23]. Cost-Effectively [NYS^{+}23]. Cottidae [PKY^{+}23]. Cottonseed [WLW^{+}23, XWP^{+}24]. Cottus [CTTW23]. Could [Tri23]. Counting [LHZ^{+}24]. Countries [BBF22, EN22]. County [XW24]. Cover [FÅG^{+}23, NGMR23]. CpG [CCP^{+}24]. Crab [BASBW24, EBRS23, FBSB24, GLX^{+}23, JJB^{+}24, KMB18, LWZ^{+}22, LXT^{+}22, LWS^{+}23b, LLS^{+}23, MBP^{+}24, PLY^{+}24, PNW^{+}22, RWF^{+}23, SYL^{+}24, WLZ^{+}22a, WHL^{+}24, WWZ^{+}24, YMD^{+}21, ZZL^{+}23, ZGY^{+}23]. Crabs [AML^{+}24, XXL^{+}24, XJC^{+}22, YE20]. crassicornis [WDL^{+}23]. Crassostrea [LYL^{+}24a, dOCCH^{+}23, GLY^{+}23, MSK^{+}22, QWY^{+}24, SLSC^{+}24, WLC^{+}23]. Creatine [BB23]. Crew [WYL23]. Crimson [LGZ^{+}23]. Crisis [BFP^{+}23]. CRISPR [KEA^{+}23, XWR^{+}23]. CRISPR-Cas9 [KEA^{+}23]. CRISPR/Cas9 [XWR^{+}23]. CRISPR/Cas9-Mediated [XWR^{+}23]. Critical [HVRCG18, MPM^{+}21, UGH^{+}24, VAT^{+}23]. Critically [HCZ^{+}23, LDX^{+}23, VV23]. Croaker [CTY^{+}21, CXT^{+}24, LTZ^{+}22, dAMPS^{+}23, USRDFO^{+}22]. crocea [CTY^{+}21]. Crop [HPJ^{+}23]. Cross [DPGFL^{+}22, HHAG24]. Cross-Decadal [HHAG24]. Cross-Generational [DPGFL^{+}22]. Crossing [EBM24]. Crowd [Tay19]. Crucian [LHZ^{+}23, LCC^{+}24, MLZ^{+}21, WML^{+}21, WMZ^{+}22, WZL^{+}24, XHC^{+}22, XLX^{+}22, XYC^{+}22, XYC^{+}23, ZZC^{+}22, ZJC^{+}22a]. Crude [EEdCSOPJ^{+}23, SdSdOSSL23]. cruentata [BPOS19]. Crustacea [BSH^{+}23, KMB18]. Crustacean [HHAG24, JSSD23, MKH24, SFK^{+}23]. Crustaceans [YJZ^{+}24]. Cryobank [RDI^{+}21]. Cryptobenthic [ZVRH23]. Ctenopharyngodon [HHP^{+}24, LWS^{+}21, WLL^{+}22, WZX^{+}24, ZSQ^{+}21, ZSL^{+}24]. Cube [XYT23]. Cues [LS19, LLY^{+}24]. Culter [XWP^{+}24]. Cultural [Tay19]. Culture [CMC^{+}24, GLX^{+}23, GL23, KŽM^{+}23, MJL^{+}24, Moy18, NYS^{+}23, PNW^{+}22, RJR^{+}22, SNZ^{+}23, SAB^{+}22, VCL20]. Cultured [CWMX21, DCL^{+}23a, HXY^{+}23, JJK21, LVB^{+}20, RHU^{+}23, RJMVC^{+}18, WHL^{+}24, XJC^{+}22]. cumingii [ZZT^{+}23]. Cuora [KSAB^{+}23]. Curcumin [EAJ^{+}23, JML^{+}24]. Cured [EA18]. Curonian [AIŠRB22, ISA^{+}22]. Current [APD^{+}23a, APD^{+}23b, Ols19, SKSL23, TK24, WXL^{+}23]. Curves [BMTR23]. Cutthroat [MD21]. Cuttlefish [OGMG^{+}17]. Cuvier [GRO^{+}17]. Cycle [CFCE20, dSCFQB^{+}23, PPAB^{+}18, PMFBI22, QWY^{+}24]. Cycles [DSC^{+}23]. Cyclin [QWY^{+}24]. Cyclopterus [BKJ^{+}24, CMP^{+}23, IR22]. Cymbopogon [GMBR^{+}21]. Cynoglossus [CL21, LZM^{+}23]. Cyprinid [PKV^{+}22]. Cyprinidae [CZL23, LBC^{+}24]. Cyprinids [PAMG19, VMBT24]. Cypriniformes [LMB^{+}23, LBC^{+}24, LZC^{+}23b]. Cyprinus [BAP22, BDŞ^{+}24, CXW^{+}23, CHJ^{+}23, KBCM19, LWT^{+}24, MS20, WSD^{+}23, XCW^{+}23, YWDP21]. Cyprus [GSK^{+}21]. CYS06 [LJR^{+}24]. Cytb [WFL^{+}23]. Cytokine [WPLK23]. Cytokines [BAP22]. Cytotoxic [MTHPJS^{+}23].

D [Kim23a]. dabryanus [LDW^{+}21, CSL24]. Dace [BMOH23]. dactylifera [HGSE23]. Daily [BBN^{+}24, GFDPSR22, OGMG^{+}17]. Dalatiid [DNP^{+}23]. Dam [APS^{+}21]. Damage [BB23, HWW^{+}24b, MTHPJS^{+}23]. Dams [ZS21]. damselae [MMD^{+}23]. Damselfish [LDBL19, LL22]. Danio [BGMM^{+}24, CAC^{+}17, HFEH^{+}23, LCWH22, MBPB24, Ord19, TU18]. Danube [BAA^{+}23, BAA^{+}23, SCT^{+}24]. Dar [MBL20]. Dark [CFCE20]. Darling [RKL24]. Darters [BAA^{+}19]. Data [CVANRD^{+}21, CHW21, DOB^{+}17, FGG^{+}22, FYH^{+}23, FB22, FWJ21, LSJ24, MBC^{+}24, MKH24, MSA24, PB24, SKT23, SGANM^{+}24, YYW^{+}23, ZCX^{+}22]. Data-Driven [SGANM^{+}24]. Data-Limited [SKT23]. Data-Poor [DOB^{+}17]. Databases [GSK^{+}21]. Dataset [LSJ^{+}22]. Datasets [APS^{+}21]. Date [HGSE23]. dauricus [JCR^{+}22]. davidianus [CWW^{+}23, ZHF^{+}22b]. Day [GLW^{+}22]. Dazl [YXS^{+}24]. DDEYOLOv9 [LHZ^{+}24]. ddRAD [KBK^{+}23]. ddRAD-Based [KBK^{+}23]. Decadal [HHAG24]. Decade [DMB^{+}20]. Decapoda [KMB18, KKB20]. Decapterus [MBD^{+}23]. Deck [WYL23]. Decoding [SDA23]. Decomposing [UH19]. Decrease [CXW^{+}23]. decussatus [SML^{+}23a]. Deep [HZS^{+}21, HARB23a, ILA22, JBGCG^{+}24, SZT^{+}23, XLC^{+}24, ZLF23]. Deep-Sea [ZLF23]. Deepest [KRAFO23]. DeepOtolith [PSP^{+}22]. Deepwater [SBP23, WSON24]. Defense [ADM^{+}24, DCL^{+}23b, LPF^{+}23, SSSP21]. Deficiency [VKP^{+}24]. Deficient [BF16]. Defogging [SYL^{+}24]. Deformity [dAMPS^{+}23]. Degradation [YLW^{+}23b].

Dehydrogenase [LTZ^{+}22]. Delayed [LKU21]. Deletion [LWS^{+}23a]. Delivery [JFP^{+}18]. Delsman [SNSVFL23]. Delta [BAA^{+}23, TVL21, BAA^{+}23]. Deltaic [dRCCLRNRWK21]. Demersal [HMX^{+}21, TMPP23]. Demographic [SMH^{+}22a, WTC^{+}22]. Demographics [CH23, HS24]. Demonstrate [SB20]. Demonstrates [ZS21]. Densities [HSAF^{+}23, LMNN21, NDC^{+}23]. Density [AHK^{+}23, BHR^{+}23, dSGBdF23, HZS^{+}21, JSSD23, LWG^{+}23, LSL^{+}24, LF24, MSB^{+}23, NYS^{+}23, SIZ^{+}22, SLSC^{+}24, TPA^{+}24, ZWP^{+}23, ZJM^{+}23]. dentex [MNO^{+}22]. Deoxygenation [KFS23]. Dependent [FDM^{+}23a, LYL^{+}23, dlCRMB^{+}22]. Deposit [CMC^{+}24]. Deprivation [LLS^{+}23, SZF^{+}21]. Depth [BSB^{+}23]. Derived [FYH^{+}23, MNO^{+}22]. Describe [RDANPA^{+}24]. Description [IFA^{+}23, PPAB^{+}18, WSON24]. Descriptive [TMD^{+}19]. Design [AHLC19, FLFF23, KBB^{+}21, LLL^{+}23, EEE21]. Designating [JLT22]. Designation [NNL^{+}23]. Designing [WMZ^{+}22]. Designs [EBM24]. Destructive [LOBTL22]. Detect [HBG^{+}20]. Detected [VKP^{+}24]. Detecting [LHZ^{+}24]. Detection [AVT18, BFLC19, ITG^{+}18, LSJ^{+}22, LDZ^{+}24, LJX^{+}21, NLTL23, PKSN23a, PKSN23b, SPEGMC24, SLC^{+}22, SYL^{+}24, WLZ^{+}23, WYL23]. Deter [Sus20]. Determinants [EN22]. Determination [Fed23, JSLE23, MS18]. Determine [VK23]. Determined [MLH^{+}20, SML^{+}23a]. Determines [dSCFQB^{+}23]. Determining [FBSB24, RGABD20]. Deterrents [ZS21]. Detour [BEF^{+}23, JHNF24]. Detrimental [WHX^{+}23]. Developed [VKP^{+}24]. Developing [BFP^{+}23, EN22, NRKT19, NIN^{+}19]. Development [APD24, AJF^{+}22, CAC^{+}17, CMP^{+}23, CBK^{+}21, CZC^{+}22, CZC23, CHH^{+}23, CFP^{+}23, CCFP19, HZG^{+}21, JHLW24, KMS^{+}17, KLD^{+}23, KAR^{+}23, LTZ^{+}22, LKJ22, LMLH22, LJP^{+}22, Liu24, LXZ^{+}22, MKC^{+}22, MMMNAG23, MPM^{+}21, NNL^{+}23, PHB^{+}23, RMSPMC^{+}22, SLC^{+}22, SAB^{+}22, Sus20, THS^{+}22, TJW^{+}22, UDG^{+}19, VMDV^{+}22, WHY^{+}24, WMD^{+}24, WWD^{+}23, WSS^{+}19, YYH^{+}24, ZLZ^{+}22, ZJJ^{+}22, ZQLW23, ZCT^{+}23, ZHX^{+}24, dlCRMB^{+}22]. Development-Related [ZJJ^{+}22]. Developmental [FYL^{+}23, Kim23a, Kim23b, MPM^{+}21, RLB^{+}23]. Deviation [CVANRD^{+}21]. Device [CLL23a, LLL^{+}23]. Diagnosis [EES^{+}23, ILA22, SKT23]. Diandric [MPK^{+}23]. Dicentrarchus [ALNVDG^{+}22, BGMM^{+}24, CFCE20, FDM^{+}23b, NZVB20, RGTPSCCC24, RSA17, SAB^{+}22, VPP^{+}22]. Dicentrarchuss [RRM^{+}20]. Dichichthyidae [WSON24]. dieffenbachii [MLK^{+}19]. Diel [BSB^{+}23]. Diet [BEG^{+}23, BHP^{+}24, CWP^{+}21, DZC^{+}22, EPKV17, FGR19, GYH^{+}23, MABÁMSM22, MMVVCJ^{+}23, MMM^{+}24, NdNFK^{+}24b, NdNFK^{+}24a, SFK^{+}23, SLSC^{+}24, WLL^{+}22, WLW^{+}24, WZX^{+}24, WWD^{+}23, YZH^{+}24]. Dietary [ASM^{+}22, AASQPU^{+}23, BSGMC^{+}22, BKBR^{+}23, CSR22, CSL24, EAJ^{+}23, GRKC19, GSHGE18, HLZ^{+}22, JKK24, LHX^{+}23, OSM23, OJC^{+}23, QXM^{+}24, SSSP21, SNSG^{+}19, SCHT23, TPC^{+}23, WGRM^{+}19, WHX^{+}23, WZL^{+}24, WFZ^{+}24a, XHC^{+}22, XLX^{+}22, XYC^{+}22, XYC^{+}23, XYL^{+}23, XWP^{+}24, YGD^{+}23, YAEAB23, ZWZ22, ZYG^{+}23, ZSZ^{+}23, dAdSCC^{+}23]. Diets [AAAF^{+}21, BCG^{+}23, CSR22, DCL^{+}23a, Eny17, FQÁGTR^{+}17, GMDMT^{+}23, GMFG^{+}24, GBT^{+}24, JKP^{+}23, MVPMAV^{+}22, NRÁGPM^{+}18, NYS^{+}23, PAVCCJVV24, PJPMMV^{+}22, PSW^{+}23, SBS^{+}23, TMD^{+}19, XYL^{+}23, ZWD^{+}23]. Differences [AJF^{+}22, BBCJ23, DM24, LZM^{+}23, Näs18, WYG^{+}23, ZZ24, ZJC^{+}22a]. Different [AGA^{+}21, CZJ^{+}24, CHH^{+}23, CCP^{+}24, dOCCH^{+}23, CPVMA^{+}24, DQC^{+}23, GMFG^{+}24, GGP^{+}23, HSAF^{+}23, IM24, JCR^{+}22, KYB^{+}23, LUM18, LDD^{+}22, LWL^{+}23, LVB^{+}20, LMNN21, MNP^{+}16, NdNFK^{+}24b, NZVB20, PAVCCJVV24, RHU^{+}23, RKS^{+}24, SdSdOSSL23, SPEGMC24, SBS^{+}23, SUL^{+}23, ŠT19, SJY^{+}22, WSL^{+}23, WML^{+}21, WYG^{+}23, XRX^{+}23, YYW^{+}23, YGD^{+}23, YLH^{+}24, YWL^{+}24, ZZT^{+}23, ZCT^{+}23, ZZY^{+}23, ZWD^{+}23, ZHX^{+}24]. Differential [GWT^{+}24, LMMC^{+}22]. Differentially [ZLSB22]. Differentiation [Fed23, MLSC^{+}23, SML^{+}23a, SLY^{+}21, VCL20, WLZ^{+}22a]. Diffusion [NRKT19]. Diformate [CHZ^{+}24]. Digestibility [BRB^{+}23, DBP^{+}20, FDM^{+}23b, MdM^{+}23, MLH^{+}20, SdSdOSSL23, TJTV^{+}23]. Digestive [DZC^{+}22, LGZ^{+}23, LLS^{+}23, MCÁGHA^{+}17, MVPMAV^{+}22, MMVVCJ^{+}23, MMM^{+}24, NRÁGPM^{+}18, PJPMMV^{+}22, SSSP21, TCB^{+}24, WZL^{+}24, XYL^{+}23, YBL^{+}22, YHZ^{+}23, dAdSCC^{+}23]. Digital [JHLW24]. Dihydrolipoamide [LTZ^{+}22]. Dimensional [ZTWK24]. Dimensions [KKFC^{+}23]. Dioxide [INCD23, Sus20]. Dipole [SZSW21]. Direct [LPB^{+}24, ŠT19]. Directions [MS20]. Discarded [MSRGCG^{+}23]. Discards [SWS22]. Discharge [WSA^{+}23]. Discordance [CFM^{+}23]. Discovery [KMS^{+}17]. Discrimination [BEF^{+}23, KMSO18, ZXH^{+}23]. Disease [JKK^{+}20, LJK^{+}22, LDD^{+}22, NHC^{+}23, SNK^{+}23, WCX^{+}24, WBK^{+}23, XLX^{+}22]. Diseased [LBH^{+}24, SNZ^{+}23, WLZ^{+}23, WFX^{+}23, XXL^{+}22]. Diseases [BAP22, LJR^{+}24, Sha19]. Disentangling [BÁD^{+}22, RRG22]. Disinfectant [HS18]. Dismissed [MSRGCG^{+}23]. Disparate [AWCS23]. Dispersal [EMFZ^{+}18]. Display [SFP17]. Disputed [CXL23]. Disruption [XWR^{+}23]. Disruptor [RLB^{+}23, SAGG^{+}23]. Dissolved [WSE^{+}21]. Distant [And23]. Distant-Water [And23]. Distiller [FDM^{+}23b]. Distinct [AML^{+}24, KAJ^{+}24, SFP17]. Distinction [MBD^{+}23]. Distinguish [FJL^{+}23]. Distributed [LZC^{+}23b]. Distribution [BSR^{+}20, BBN^{+}24, CASMK23, DCD^{+}24, Gae16, Gae17, GP17, Gre17, LCZ^{+}23, LWD^{+}23, LXW^{+}24, MSA24, MDLA22, RSA17, RKS^{+}24, SKD^{+}23, SKF^{+}23, TK24, WWY^{+}23, WWZ^{+}24, XZZ^{+}24, ZZL^{+}23, ZCX^{+}22]. Distributions [HBL^{+}22, LBC^{+}24]. District [MFKS23]. Disturbance [CLL^{+}23b]. Divergence [WTC^{+}22]. Diverse [ZAM^{+}23]. Diversity [ARH^{+}23, ANA^{+}23, ATM^{+}24, BAA^{+}23, CTTW23, GSH^{+}24, JZX22, JCL^{+}24, LHL^{+}24, LZC^{+}23b, RXL^{+}24, SWUH23, SLDC23, WLZ^{+}22a, XRX^{+}23, YJK23, ZLL^{+}23, ZGY^{+}23, ZZZ^{+}23]. DNA [AM24, AMK23, BAA^{+}19, EAJG24, HBG^{+}20, HZG^{+}21, JZX22, OJC^{+}23, SBS^{+}24, WFZ^{+}24b]. dnd [YXS^{+}24]. Do [FÅG^{+}23, IBN^{+}23, YGD^{+}23]. Docking [NWN^{+}23]. Does [BF16, CCFP19, JCL^{+}24]. Dogfish [ARH^{+}23]. dolomieu [SOW^{+}23]. Dolphin [JLT22]. Dolphinfish [Poi24]. domestica [HIO^{+}19]. Domesticated [AJF^{+}22, FHF23]. Domestication [LMLH22]. Dominant [LHZ^{+}23]. Door [VCC^{+}18]. Dopamine [CGSBGN24]. Dormitator [RMSPMC^{+}22]. Dorsal [HL24]. dorsalis [ASM^{+}22]. Dose [LYL^{+}23]. Dosidicus [HZL^{+}22, ZXY^{+}24]. Dottyback [CZC23]. Double [CCP^{+}24, RDANPA^{+}24]. Double-Stranded [CCP^{+}24]. Downstream [WLL^{+}23]. Downstream-Migrating [WLL^{+}23]. Drainage [CASMK23]. Dried [ATEfA^{+}21, FDM^{+}23b, HKS^{+}18, MAR^{+}18, RMA^{+}18]. Drift [BAA^{+}19]. Drifting [NIN^{+}19, WMD^{+}24]. Drinking [WFZ^{+}24b].

Driven [LZZ^{+}22, SGANM^{+}24]. Drivers [FCT19, RRG22]. Driving [BÁD^{+}22, XZZ^{+}24]. Drug [RCL^{+}23]. Drying [RMA^{+}18]. Due [BAA^{+}23]. dumerili [JFP^{+}18]. Duplication [MBZ^{+}21]. Duration [SZF^{+}21]. During [MAR^{+}18, BKBR^{+}23, BEF^{+}23, CAC^{+}17, CWMX21, CBK^{+}21, CTY^{+}21, EAJG24, Kim23a, Kim23b, LPB^{+}24, LGZ^{+}23, MKC^{+}22, MSB^{+}23, MLK^{+}19, NdNFK^{+}24b, NdNFK^{+}24a, PXH^{+}24, PSO^{+}19, RWF^{+}23, SMH^{+}24, SZSW21, THS^{+}22, TJW^{+}22, TRM^{+}23, UDG^{+}19, VCC^{+}18, WLL^{+}24, WHY^{+}24, WDL^{+}23, WSS^{+}19, ZHF^{+}22a, ZCX^{+}22]. Dusky [AGA^{+}21]. DUSP2 [LWS^{+}23a]. Dwarf [HSZ^{+}22]. Dwarf-Form [HSZ^{+}22]. Dynamic [EFQ23, HWZ21, SGANM^{+}24, TJW^{+}22, WLM^{+}20, ZZL^{+}23]. Dynamics [AGC23, APS^{+}21, BKBR^{+}23, FGG^{+}22, GFDPSR22, HWW^{+}24a, IFA^{+}23, SCBSSMA24, SB20, Tri23, YE20].

E-Nose [PBS^{+}22]. E. [KSO^{+}23, SZZ^{+}23, YGD^{+}23]. E2 [LPB^{+}24]. E20 [CBCL23]. Earliest [KMS^{+}17]. Earliest-Stage [KMS^{+}17]. Early [BKBR^{+}23, CAC^{+}17, CHH^{+}23, CBCMRD^{+}23, IM24, JKP^{+}23, Kim23a, Kim23b, MPM^{+}21, PH23, SMO^{+}22, SMO^{+}23, STZ^{+}23, TJW^{+}22, VPP^{+}22, VH20, ZSL^{+}23]. Early-Life [SMO^{+}23]. Early-Stage [JKP^{+}23]. East [HWZ21, LCZ^{+}23, YSY^{+}23]. Eastern [CASMK23, GSK^{+}21, HCWH20, UGH^{+}24, AGE^{+}18, DMA22, MMD^{+}23, MGMG24, MLSC^{+}23, RKS^{+}24, SMM^{+}18]. Eat [WDL^{+}23]. Echolocation [CLL^{+}24]. eckloni [ZZ24]. Eco [GZX^{+}22]. Eco-Substrate [GZX^{+}22]. Ecological [DCD^{+}24, Fra23, FCT19, KKB20, KYB^{+}23, LZL^{+}23, MWPS23, MWPS24, MBL20, PYP17, Poi24, RTBL^{+}18, WSL^{+}23, dlCRMB^{+}22]. Ecologically [AAB^{+}18]. Ecology [AFTÁPA^{+}23, BS23, GPD^{+}23, LCW23a, LKD22, LOBTL22, PCO^{+}23, VSH23]. Economic [ATEfA^{+}21, AAAF^{+}21, BÁD^{+}22, BASBW24, CBP^{+}24, HZL^{+}22, LV23, Tay19]. Economical [MBL20]. Economics [SNK^{+}23]. Economy [And23, JHLW24, LJP^{+}22]. Ecosystem [APD^{+}23a, APD^{+}23b, KAB^{+}20, SS23, ZTWK24]. Ecosystems [DD23]. Ecotourism [SLL22]. Ecotoxicology [AAB^{+}18]. Ectoparasite [CSSMV^{+}23]. Ectoparasitic [MNO^{+}22]. Edible [EBRS23, GMRJ22]. Editing [KEA^{+}23]. Editor [Hal23]. Editor-in-Chief [Hal23]. Editorial [Moy18, Piz22]. eDNA [PKSN23a, ATM^{+}24, HBL^{+}22, PKSN23b]. edodes [XYL^{+}23]. Edwardsiella [WSZ^{+}23, YLW^{+}23a]. Eel [BGMM^{+}24, JKP^{+}23, LPB^{+}24, MLK^{+}19, MDLA22, PRPW23, RCL^{+}23, ZWZ22]. Eels [DMA22, EFG^{+}23, HMVRFD19, SLYY23, WHX^{+}23]. EF [JML^{+}24]. EF-24 [JML^{+}24]. Effect [AAAF^{+}21, AHK^{+}23, ASM^{+}22, BSGMC^{+}22, BHP^{+}24, CWMX21, CFCE20, CZC^{+}22, CL21, CPJK23a, CPJK23b, DO24, EPKV17, GMBR^{+}21, GYH^{+}23, GSH^{+}24, HV24, HXS^{+}23, HYN^{+}24, HS18, JKP^{+}23, KKFC^{+}23, LHX^{+}23, LLS^{+}23, LCX^{+}23, MMMNAG23, MSB^{+}23, MIHH23, NWN^{+}22, NRÁGPM^{+}18, PYJ^{+}23, PSW^{+}23, PWCL23, dCPRG^{+}21, PKV^{+}22, QXM^{+}24, RRM^{+}20, RXL^{+}24, SZZ^{+}23, SAC23, SAB^{+}22, SIZ^{+}22, WGRM^{+}19, WZX^{+}24, WHM21, WCY^{+}24, ZWZ22]. Effective [AHLC19, HIO^{+}19, HCZ^{+}23, LXW^{+}24]. Effectively [NYS^{+}23]. Effectiveness [GZF^{+}22]. Effects [AEME^{+}23, BDŞ^{+}24, BHR^{+}23, BSH^{+}23, BEF^{+}23, CZW^{+}23, CSGE23, CSL24, CCP^{+}24, CFP^{+}23, DÁ23, DDN19, DK22, DPGFL^{+}22, DSC^{+}23, EAJ^{+}23, EAE^{+}23, EAJG24, FMXQ23, FMBPC^{+}23, FHF23, FLB^{+}21, GL23, GMDMT^{+}23, GGP^{+}23, dSGBdF23, GSHGE18, GZX^{+}22, HLK^{+}23, HHP^{+}24, HGSE23, HL24, HFEH^{+}23, IM24, JBdFS^{+}22, KJ22, LMC21, LMB^{+}23, LWZ^{+}22, LGZ^{+}23, LZH^{+}23, LV23, LSL^{+}24, LMMC^{+}22, LMNN21, MVPMAV^{+}22, MTHPJS^{+}23, MNO^{+}22, MTPK23, Näs18, NDC^{+}23, NSK^{+}23, PXH^{+}24, PHB^{+}23, PCH^{+}24, dAPAA^{+}24, PH23, RKHAMM22, RHU^{+}23, RCL^{+}23, SSSP21, SMO^{+}23, SBP23, SZF^{+}21, SPJ^{+}24, SBB^{+}19, SNK^{+}23, SAHS18, TJTV^{+}23, TCB^{+}24, ULR^{+}23, VAT^{+}23, WSZ^{+}23, WML^{+}21, WLL^{+}22, WZG^{+}23, WLW^{+}23, WSD^{+}23, WHX^{+}23, WZL^{+}24, WFZ^{+}24a, WLL^{+}24, WSA^{+}23, Web23, WMD^{+}24, WSE^{+}21, WLZ^{+}22b, XLX^{+}22, XYC^{+}22, XYC^{+}23, XYL^{+}23, XWP^{+}24, YLW^{+}23b, YHZ^{+}23, ZLZ^{+}24, ZYG^{+}23, ZWP^{+}23, ZYT^{+}24, dAdSCC^{+}23, vKRNL^{+}19]. Efficacy [CG20, IR22, LDD^{+}22, PSN18]. Efficiencies [APAHBMAG23]. Efficiency [BBF22, CGY^{+}23, KKFC^{+}23, MSB^{+}23, SNK^{+}23, TVL21, ZLF23]. Efficient [TPA^{+}24]. Effluent [LMC21]. Effort [LSJ24, SHT^{+}23, WLM^{+}20]. Efforts [HDW^{+}23]. EGb [CXW^{+}23]. Egg [RAR^{+}18]. Eggs [PAVCCJVV24, THS^{+}22, WMD^{+}24]. ehimensis [LCWH22]. EHP [HWX^{+}23]. Elasmobranch [LW18, PYP17]. Elasmobranchii [BMMD22, FGG^{+}22]. Elasmobranchs [ZSZSS^{+}22]. Eleginops [MPM^{+}18]. Element [XJC^{+}22]. Elements [XWR^{+}23]. Eleotridae [RMSPMC^{+}22]. Elephant [SAL19]. Elevated [DK22, DTS^{+}17, INCD23, PLC^{+}24, ZZW^{+}22]. Elizabethkingia [WWW^{+}24]. elongatus [PKY^{+}23]. Elopomorpha [KBK^{+}23]. Elucidate [BLA^{+}22]. Embryo [MMMNAG23]. Embryogenesis [DIL^{+}22, EAJG24, LLY^{+}24]. Embryonic [CZC^{+}22, CZC23, DK22, LXZ^{+}22, THS^{+}22, UDG^{+}19, YXS^{+}24, dlCRMB^{+}22]. Embryos [CGSBGN24, LL22, Ord19, XWR^{+}23].

Emerging [Näs18]. Emphasis [AJF23]. Empirical [CHW21, HWZ21]. Enclosed [SXQ^{+}22]. Encounter [AHLC19]. Encounters [FÅG^{+}23]. encrasicolus [DQC^{+}23]. Endangered [APD24, HCZ^{+}23, KSAB^{+}23, LDX^{+}23, MJL^{+}24, SKSL23, VV23]. Endangerment [YZL^{+}23]. Endemic [HVRCG18, VMBT24]. Endo [PKC^{+}19]. Endo- [PKC^{+}19]. Endo16 [XWR^{+}23]. Endocrine [RLB^{+}23, SAGG^{+}23, VCC^{+}18]. Endocrinology [KJ22]. Endogenous [WFZ^{+}24a]. Endotoxin [LS19]. Endurance [VF23]. Energetic [BB23]. Energy [HYN^{+}24, LWS^{+}23b, MdM^{+}23, ZLZ^{+}24]. Engagement [UGH^{+}24]. English [DMA22]. Engraulis [AAN22, DQC^{+}23, FZZ^{+}24]. Enhance [BFP^{+}23, CXW^{+}23]. Enhancement [GBT^{+}24, HHL^{+}20, JCL^{+}24, TOB^{+}23]. Enhancer [AHJ^{+}23]. Enhances [LCWH22, XCW^{+}23]. Enhancing [JBGCG^{+}24, OO24, WLW^{+}24]. Enriched [NSK^{+}23, VLCCA^{+}23]. Enrichment [CBCL23, WWW19]. Ensemble [WSL^{+}23]. Enteric [YLW^{+}23a]. Enteritis [SCSR22]. Enterococcus [XLX^{+}22, XYC^{+}22, XYC^{+}23]. Enterocytozoon [HWX^{+}23]. Enteromorpha [ZWD^{+}23]. Entire [WMD^{+}24]. Ento [SAB^{+}22]. Entosphenus [KSO^{+}23]. Entropy [KTT24]. Environment [HBL^{+}22, LXW^{+}24, LJP^{+}22, LWZ^{+}23, WLZ^{+}23, ZXY^{+}24]. Environmental [AJF23, AM24, AMK23, BÁD^{+}22, CL19, CMC^{+}24, DPGFL^{+}22, EM23, FHF23, GTC^{+}17, HJE^{+}23, HBG^{+}20, HMX^{+}21, HMVRFD19, HWZ21, JZX22, KKPL23, LKJ22, LWD^{+}23, LHL^{+}24, NKC^{+}21, RJMVC^{+}18, RJTVC^{+}19, ST17, VK23, WKB^{+}23, WM22, WGW^{+}23, WWY^{+}23, WFZ^{+}24b, YYW^{+}23, ZTWK24]. Environmentally [TPA^{+}24].

Environments [CPVMA^{+}24, LHZ^{+}24, VLCCA^{+}23]. Enzymatic [LYL^{+}24b, ULR^{+}23, WLW^{+}23]. Enzyme [DRFCL23, DZC^{+}22, ERE21, MBAM19, MLK^{+}19, MMVVCJ^{+}23, NRÁGPM^{+}18, PJPMMV^{+}22, WZL^{+}24, XYL^{+}23, YBL^{+}22, YHZ^{+}23, ZZL^{+}23, dAdSCC^{+}23]. Enzymes [FMBPC^{+}23, LGZ^{+}23, LYL^{+}24b, MVPMAV^{+}22, MMM^{+}24, SSSP21, TPC^{+}23]. Enzymology [FLB^{+}21]. Epidemiology [ŢPN^{+}23]. Epidermal [WML^{+}21]. Epigenetics [SMH^{+}22b]. Epinephelus [AGA^{+}21, AKM23, HAC^{+}23, SWS22, SZZ^{+}23, Tri23, WYG^{+}23, YGD^{+}23, ZHQ^{+}23]. Eradication [YWDP21]. Eriocheir [FLR^{+}22, GLX^{+}23, LWZ^{+}22, LXT^{+}22, LLS^{+}23, PLY^{+}24, PNW^{+}22, WLZ^{+}22a, WHL^{+}24, WWZ^{+}24, XXL^{+}24, XJC^{+}22, YMD^{+}21, ZGY^{+}23]. ERK [LWS^{+}23a]. Error [CBANCM^{+}21]. Erythrocyte [SSK^{+}22]. Erythrocytes [CSGE23, MTHPJS^{+}23, SSSS23]. erythropterus [LGZ^{+}23]. Esc [YLW^{+}23a]. Esox [FWJ21, Mee24]. Especially [APS^{+}23]. Essential [BSH^{+}23, GMBR^{+}21, JBdFS^{+}22, SBS^{+}23, dAdSCC^{+}23]. Establishment [CLJ^{+}23, LDZ^{+}24, MU21]. Ester [CWMX21]. Estimate [SHT^{+}23]. Estimated [TMPP23]. Estimates [BPO19, CSB^{+}20, CLL^{+}22, CBANCM^{+}21, FGBA^{+}23, HCZ^{+}23, LLWW23, MOW^{+}18, XWW^{+}24]. Estimating [JHNF24, LLW^{+}22]. Estimation [LJX^{+}21, LSJ24, Tri23]. Estradiol [CZC^{+}22, JSJ^{+}24, VMS^{+}23]. Estrogen [CZC^{+}22]. Estuarine [DMA22, FDM^{+}23a, Gre17, HBL^{+}22]. Estuarine-Dependent [FDM^{+}23a]. Estuary [Fra23, KBB^{+}21, LHL^{+}24, SFP17, GTC^{+}17, SLYY23, WSL^{+}23]. Ethmalosa [DTS^{+}17]. Ethyl [CWMX21]. Eupercaria [dAMPS^{+}23]. Europe [Fra23, EFQ23, NPT^{+}24, RRG22].

European [AHJ^{+}23, ANA^{+}23, AVT18, BFB^{+}23, BAS^{+}19, BBF22, BGMM^{+}24, CFCE20, CMC^{+}24, DMA22, EFG^{+}23, FGR19, FB22, FDM^{+}23b, GFDPSR22, HMVRFD19, LLKKV20, MDLA22, NZVB20, RGTPSCCC24, VPP^{+}22, XQLA18]. Eurozone [BBF22]. Euryhaline [WLT23]. eurystomus [KZC^{+}24]. Euthynnus [MLSC^{+}23]. Evaluate [FWJ21, MU21]. Evaluating [BDŞ^{+}24, FYH^{+}23, GLH^{+}23, KBB^{+}21, MD21, MGMG24, SCSR22, ZZZ^{+}23]. Evaluation [BPCC23, DRH18, EES^{+}23, EEdCSOPJ^{+}23, Fed23, FLFF23, GMCF^{+}22, GTSG18, GZF^{+}22, MBAM19, NWN^{+}21, PKSN23a, PKSN23b, RDI^{+}21, SPEGMC24, SHT^{+}23, VAT^{+}23, WHM21, XYF^{+}24, YLX^{+}22]. Evasion [QAC23]. Events [HBG^{+}20]. Evidence [CSJ^{+}23, GP17, KKPL23, LKJ22, PWCL23, VMBT24, WHM21, ZAM^{+}23]. Evolution [AMT^{+}24, Gae16, Gae17, MBZ^{+}21, SDA23, ZQL^{+}23]. Evolutionary [DNP^{+}23, Fed23]. Ex [CYL^{+}23]. Examination [HL24]. Examined [CHH^{+}23]. Examples [SBP23]. Excluder [CLL23a]. Exclusive [HZL^{+}22]. Excretion [JCR^{+}22, TU18]. Exercise [LWS^{+}21]. Exhibits [DTS^{+}17, LJR^{+}24]. Exogenous [PXH^{+}24]. Exoskeletal [PKC^{+}19]. Exotic [SGAC20]. Expansion [SKD^{+}23]. Expected [SBG^{+}24]. Experience [QAC23, YWDP21]. Experiences [DPGFL^{+}22, LOB^{+}23, SRBCGV^{+}21]. Experimental [CFLK21, HYXY23, NHR20, NNN23, NWN^{+}23, PLC^{+}24, PVY^{+}21, WXL^{+}23, WBK^{+}23]. Experimentally [ZZL^{+}23]. Expert [MSK^{+}22]. Expert-Based [MSK^{+}22]. Exploitation [FPRAT^{+}23]. Exploiting [SB20]. Exploration [LSJ^{+}22]. Exploring [ADM^{+}24, ZCX^{+}22]. Exporting [EN22]. Exposed [ADM^{+}24, AGA^{+}21, BAS^{+}19, dOCCH^{+}23, HLZ^{+}22]. Exposure [BMSGS^{+}18, BSH^{+}23, CGSBGN24, DLL^{+}23, DIL^{+}22, LZW^{+}24, RCL^{+}23, ŠT19, WKB^{+}23, WCY^{+}24]. Expressed [CAC^{+}17, ZLSB22]. Expression [ASM^{+}22, CZW^{+}23, CYL^{+}23, CZC^{+}22, ERE21, FCB^{+}21, FLR^{+}22, GWT^{+}24, GLY^{+}23, GRKC19, GGF^{+}22, JKP^{+}23, KJK^{+}23, LCC^{+}24, LLY^{+}22, LCWH22, LWL^{+}23, LXZ^{+}22, MKC^{+}22, MWZ^{+}23, MVPMAV^{+}22, NRÁGPM^{+}18, PLJ^{+}23, PJPMMV^{+}22, PSO^{+}19, QWY^{+}24, SML^{+}23a, TU18, WSZ^{+}23, WLW^{+}23, WYG^{+}23, WPLK23, YZH^{+}24, YLX^{+}22, YSG^{+}23, ZLW^{+}23, ZSZ^{+}23, ZSL^{+}24, ZWZ^{+}24, ZCT^{+}23, ZYT^{+}24, vKRNL^{+}19]. Expressions [ZHF^{+}22a]. Extended [CPJK23a, CPJK23b]. Extension [JROHVH^{+}23]. Extent [RKL24]. External [BHR^{+}23, RTBL^{+}18]. Extinct [EMFZ^{+}18]. Extract [AEME^{+}23, CXW^{+}23, CSV^{+}19, HFEH^{+}23, JROHVH^{+}23, VAT^{+}23, XCW^{+}23]. Extracting [ZXY^{+}24]. Extraction [NLTL23]. Extracts [CL21, PSS^{+}18, dCPRG^{+}21, WGRM^{+}19]. Extremely [BSB^{+}23].

F [Kim23a, Kim23b, SRL^{+}19]. F. [GWH^{+}22]. F1 [JFP^{+}18, WSI^{+}19]. Fabrication [EEE21]. Face [LSJ^{+}22]. Faceted [DCD^{+}24]. facetus [BSR^{+}20]. Facilities [CLZ^{+}24, WXL^{+}23]. Factor [EBH21, KJK^{+}23, SHW^{+}23, ŢPN^{+}23]. Factors [BÁD^{+}22, CH23, HWZ21, LWD^{+}23, LFH^{+}23, LHL^{+}24, NHC^{+}23, RAR^{+}18, SGAC20, VČ24, VK23, WM22, WHY^{+}24, WWY^{+}23, XZZ^{+}24, YZL^{+}23, ZWZ^{+}24]. faecalis [XLX^{+}22]. Faeces [MLH^{+}20]. faecium [XYC^{+}22, XYC^{+}23]. Fails [RDE^{+}23]. falciformis [KSWT22]. Fall [BMTR23]. Family [BAA^{+}19, GLY^{+}23, LLY^{+}22, OMC^{+}19, SHW^{+}23, WSZ^{+}23, WSON24, YLW^{+}23b, ZLW^{+}23, ZLZ^{+}22, ZVRH23]. Fang [TTT23]. Farm [JSRE^{+}24, TKF^{+}18]. Farm-Raised [JSRE^{+}24]. Farmed [AVT18, CVANRD^{+}21, DLL^{+}22, EES^{+}23, HLZ^{+}22, IR22, JSLE23, MGS^{+}23, MMY^{+}17, RDANPA^{+}24, RSA17, SALC^{+}19, XZD^{+}24]. Farming [GMRJ22, LXW^{+}24, MBL20, NNL^{+}23, RDG^{+}20, TVL21, TPA^{+}24]. Farms [JLT22, JKK^{+}20, KYB^{+}23, MDF^{+}23, NKC^{+}21, NHC^{+}23, SNZ^{+}23]. fasciatum [PCH^{+}24]. Fasting [FMBPC^{+}23, dAPAA^{+}24, ZSZ^{+}23]. Fat [WZX^{+}24]. Fathead [LMC21]. Fatty [ASM^{+}22, BKBR^{+}23, HZL^{+}22, LZM^{+}23, MNP^{+}16, PSW^{+}23, TMD^{+}19, VKP^{+}24, ZSZ^{+}23, ZYT^{+}24]. Favor [SCSS23]. Feasibility [LJX^{+}21]. Features [GWH^{+}22, KMSO18, ZXY^{+}24]. Fecundity [DTS^{+}17]. Fed [ATEfA^{+}21, BCG^{+}23, CWP^{+}21, DZC^{+}22, GMFG^{+}24, GBT^{+}24, MMVVCJ^{+}23, MMM^{+}24, PAVCCJVV24, SBS^{+}23, SEA^{+}23, SGG^{+}21, TMD^{+}19, WLW^{+}24, XYL^{+}23]. Feed [ATEfA^{+}21, CXW^{+}23, CSF^{+}23, CHZ^{+}24, DBP^{+}20, DRH18, EPKV17, HMN^{+}22, HIO^{+}19, HMP^{+}24, IM24, JKK^{+}20, KKFC^{+}23, MMVVCJ^{+}23, MLH^{+}20, OO24, Pao23, VH20, XCW^{+}23, YAEAB23, ZLX^{+}23, ZWD^{+}23, ZHHO23]. Feedback [LPB^{+}24]. Feeder [EEE21]. Feeders [CMC^{+}24]. Feeding [AHK^{+}23, Ang18, BTSK24, BWKS20, CSF^{+}23, HZL^{+}22, KSO^{+}23, KKPY22, KKFC^{+}23, LSS^{+}22, LOBTL22, MSRGCG^{+}23, RGTPSCCC24, TJTV^{+}23, WLL^{+}23, WLL^{+}24, WFZ^{+}23, XSZ^{+}23, dSSBdS23]. Feeds [GHS20, MBAM19, RRM^{+}20, RGVG19]. Feedstuffs [BSGMC^{+}22]. felis [OJC^{+}23]. Female [CBK^{+}21, DOB^{+}17, JSJ^{+}24, LYL^{+}24b, QAC23, RAR^{+}18, SBP23, SLY^{+}21, FHHC23, JCR^{+}22, PYJ^{+}23, SZZ^{+}23, WZL^{+}24, YGD^{+}23]. female-sign [FHHC23, WZL^{+}24]. Females [PAVCCJVV24]. FEMP [BBF22].

Fermentation [FDM^{+}23b, XYL^{+}23]. Fermented [WLW^{+}23, WZL^{+}24, XYC^{+}22, XYC^{+}23, ZYG^{+}23]. Fertility [RAR^{+}18, ZSL^{+}23]. Fertilized [LdRCV24]. Fidelity [SBT22]. Field [Ang18, GAR^{+}24]. Fifteen [APD24]. Fighting [LJR^{+}24]. Filial [CZC23]. Fillet [MKN^{+}16]. Fillets [JROHVH^{+}23, ŠT19]. Filter [CMC^{+}24, KSO^{+}23]. Filter-Feeding [KSO^{+}23]. fimbriata [DTS^{+}17]. Fin [HL24, VCL20]. Fin-Mounted [HL24]. Final [PYP17]. Finding [KMB18]. Findings [dCPFS^{+}24]. Fine [MSA24, SLY^{+}21]. Fine-Patterned [SLY^{+}21]. Fine-Scale [MSA24]. Finfish [JE18, NNL^{+}23, RAR^{+}18]. Fingerling [RJTVC^{+}19, SNK^{+}23]. Fingerlings [NDC^{+}23]. Finless [CYL^{+}23, CLL^{+}24, FLX^{+}22, LCZ^{+}23]. finmarchicus [BHP^{+}24]. Finned [PKY^{+}23]. Firm [AJF23]. Firms [AJF23]. First [DLL^{+}22, DCL^{+}23b, EBH21, GMCF^{+}22, JSRE^{+}24, JFP^{+}18, MKC^{+}22, MWPS23, MWPS24, dAMPS^{+}23, SQ23, TK24, VMDV^{+}22]. First-Generation [JFP^{+}18]. Fish [AM24, AMK23, AIŠRB22, APS^{+}23, AGE^{+}18, AJF^{+}22, BTSK24, BB23, BAA^{+}23, BCG^{+}23, BBCJ23, CL19, CSGE23, dRCCLRNRWK21, CXW^{+}23, CLL23a, CZC23, CP24, CFLK21, CSSMV^{+}23, DCD^{+}24, DCL^{+}23a, DMT^{+}19, DM24, DPGFL^{+}22, DTS^{+}17, DO24, EM23, EBM24, EPKV17, ENO21, EFQ23, EMFZ^{+}18, FDM^{+}23a, FLFF23, Fra23, GTC^{+}17, GBT^{+}24, Gre17, GZX^{+}22, HARB23b, HIO^{+}19, HHAG24, HYN^{+}24, HPJ^{+}23, Ims23, IBN^{+}23, ILA22, INCD23, ITG^{+}18, JBGCG^{+}24, JSLE23, JJK21, JZX22, JK24, JHNF24, KTT24, KJ22, KAB^{+}23, KFS23, KAB^{+}20, KYB^{+}23, LKU21, LMB^{+}23, LMLH22, LSJ^{+}22, LWD^{+}23, LFH^{+}23, LHZ^{+}24, LHL^{+}24, LWS^{+}21, LJX^{+}21, LL18a, LL18b, LWT^{+}24, LWZ^{+}23, LYY^{+}24, LBH^{+}24, LF24, MdM^{+}23, jMlL23, MD21, MPM^{+}18, MTM^{+}19, MDF^{+}23, Mil23, MHZ^{+}23, MZA^{+}23, MWPS23, MWPS24, MMD^{+}23, Moy18, NKC^{+}21, NdNFK^{+}24b, NdNFK^{+}24a, NWN^{+}23, NYS^{+}23]. Fish [OMC^{+}19, OOTC24, Ols19, Pao23, PKY^{+}23, PSP^{+}22, PWH^{+}23, PKSN23a, PKSN23b, QXM^{+}24, QPDGF^{+}23, RKHAMM22, RLB^{+}23, RLAE23, RJR^{+}22, RTBL^{+}18, RDG^{+}20, RKL24, RRG22, SGAC20, SRBCGV^{+}21, SAL19, SMO^{+}23, SMH^{+}24, SCSS23, SKD^{+}23, EEE21, SAGG^{+}23, SMAR24, SOW^{+}23, SXQ^{+}22, SNK^{+}23, SMH^{+}22b, SML^{+}23b, SLDC23, SAHS18, TMD^{+}19, ULR^{+}23, VM19, VCC^{+}18, WGRM^{+}19, WSL^{+}23, WMZ^{+}22, WLZ^{+}23, WLW^{+}23, WLW^{+}24, WMD^{+}24, WFZ^{+}23, WFZ^{+}24b, XYC^{+}22, XYC^{+}23, XCW^{+}23, XWP^{+}24, YLW^{+}23a, YJK23, Zac22, ZAM^{+}23, ZMA^{+}24, ZZW^{+}24, dSJC^{+}21, dSSBdS23]. Fish-Based [SXQ^{+}22]. Fished [MMAO22]. Fisher [JLT22, LV23, MFKS23, SKL^{+}23, TOB^{+}23, UGH^{+}24]. Fisherfolk [WM22]. Fisherie [FCT19]. Fisheries [BBF22, BFP^{+}23, BPCC23, CBP^{+}24, DÁ23, FM21b, FCT19, HWW^{+}24a, ISA^{+}22, KFS23, LKJ22, MSK^{+}22, MTPK23, MML22, OOGAS23, PMFBI22, QL22, SWUH23, SKT23, SRHCO23, SS23, SLL22, TOB^{+}23, XW24, XYT23, YYW^{+}23, YSY^{+}23]. Fishermen [SRHCO23]. Fishery [APAHBMAG23, CHW21, Fer23, HB24, Ho22, JHLW24, LLTM17, LJP^{+}22, Liu24, MBC^{+}24, MGMG24, MTPK23, MBP^{+}24, NWN^{+}21, NWN^{+}22, NIN^{+}19, PWCL23, dCPFS^{+}24, Sam24, ŢPN^{+}23, TDN^{+}22, WHM21, XW24, YLL22, ZXY^{+}24, ZCX^{+}22]. Fishery-Processing [NWN^{+}22]. Fishes [CGY^{+}23, CPJK23a, EHGS23, Fed23, FCF19, FLFF23, HMX^{+}21, Kim23a, KCCK23, KGP24, LBC^{+}24, LW18, LZL^{+}23, MS18, MWPS24, MBZ^{+}21, Piz22, PKSN23a, SALC^{+}19, Sor21, Sus20, TMM^{+}18, ZVRH23, Est16, Gae17, APD^{+}23a, EBH21, LL18a, Off18, Off19, Off23, XYC^{+}23].

FishEthoBase [SALC^{+}19]. Fishing [And23, AAN22, CHW21, CXL23, CZCW23, FYH^{+}23, FGBA^{+}23, dSCFQB^{+}23, LSJ24, MFKS23, OOG24, PMFBI22, PB24, TTT23, VK23, WYL23, XQLA18, XLC^{+}24, ZZW^{+}24]. Fishmeal [ATEfA^{+}21, BCG^{+}23, CSR22, FLB^{+}21, ZMLFS^{+}20]. Fishway [PAMG19]. Fistularia [TK24]. Fitness [MH23]. Fitting [MOW^{+}18]. Five [ZZZ^{+}23]. Fivespot [LPK^{+}23b]. Flagellar [SZZ^{+}23]. Flannelmouth [RHUJ24]. Flat [SRBCGV^{+}21]. Flat-V [SRBCGV^{+}21]. Flathead [EES^{+}23]. flavescens [CWP^{+}21, NHR20]. Flavobacterium [VČ24]. Flavor [HHP^{+}24, PBS^{+}22, WDL^{+}23]. flavus [EAJ^{+}23]. Flaws [LBC^{+}24]. Fleet [PMFBI22, YYW^{+}23]. Flesh [MGS^{+}23]. Flexible [MOW^{+}18]. fliL [SZZ^{+}23]. Floating [AHK^{+}23, HLK^{+}23, RXL^{+}24]. FLOCponics [HPJ^{+}23]. Flood [PCH^{+}24]. Flora [SJY^{+}22]. Florfenicol [SEA^{+}23]. Florida [LLTM17, BPCC23, CH23, EKL^{+}23, MPK^{+}23]. Flounder [CCP^{+}24, JKK^{+}20, JKK24, LPK^{+}23b, YLW^{+}23b, YSY^{+}23, ZLZ^{+}22]. Flow [PRPW23, TKF^{+}18, WMD^{+}24, ZLZ^{+}24]. Flow-Through [TKF^{+}18]. Fluctuating [MH23]. Fluctuation [SNSG^{+}19]. Fluctuations [LZZ^{+}22]. Fluid [RDE^{+}23, VAT^{+}23, WKB^{+}23]. Fluidized [SLSC^{+}24]. Flunixin [MSK^{+}21]. Fluorescence [LSY^{+}17]. Fluvial [dRCCLRNRWK21]. Fluvial-Lagoon-Deltaic [dRCCLRNRWK21]. Fly [CSR22, LWT^{+}24, NYS^{+}23, YAEAB23]. Flying [CLL^{+}22, ZXY^{+}24]. Focus [HDW^{+}23, Ols19, ZSZSS^{+}22]. Focused [WWD^{+}23]. Follicle [Web23]. Following [SBT22, ZSL^{+}24]. Food [AA23, LMLH22, LdRCV24, LLS^{+}23, LMMC^{+}22, QWR^{+}23, SMO^{+}23, SZF^{+}21, YJK23, ZLX^{+}23]. Foods [DDN19]. Foodsize [FHHC23]. Foraging [JBK^{+}23, LLS^{+}23, OSM23]. Force [CCCFE18]. Forecast [AAN22]. Form [HSZ^{+}22]. Formalin [LDD^{+}22]. Formalin-Killed [LDD^{+}22]. Formation [KSSI24, PSS^{+}18, RAR^{+}18]. Formulated [CWP^{+}21, DZC^{+}22]. Fortification [BF16]. Fostering [NNL^{+}23]. Fotemustine [DIL^{+}22]. Fouling [TPA^{+}24]. Foundation [LDW^{+}21]. Four [AIŠRB22, GAR^{+}24, SGAC20, ZHF^{+}22a, ZLL^{+}23]. FP17 [DLL^{+}22]. FPH [SNK^{+}23]. Fractal [DSC^{+}19, RDANPA^{+}24]. Fraction [HIO^{+}19]. Fractionation [BFM23]. Fragments [SLC^{+}22, ZZZ^{+}23]. Framework [XQLA18]. Frameworks [And23]. Frass [YAEAB23]. Free [BCG^{+}23]. Freeze [DSC^{+}23]. Freezing [HYXY23]. French [AKM23]. Frequency [CXT^{+}24, LGZ^{+}23, PWH^{+}23]. Frequency-Specific [CXT^{+}24]. Frequently [DMT^{+}19]. Fresh [DRH18, DMT^{+}19, RDG^{+}20]. Freshness [LSY^{+}17]. Freshwater [BPCC23, CLL^{+}23b, CGY^{+}23, DCR^{+}23, EFQ23, GAR^{+}24, INCD23, KSAB^{+}23, MJL^{+}24, MIHH23, MS18, MML22, NKC^{+}21, NSK^{+}23, Piz22, PWH^{+}23, RLB^{+}23, RRG22, SAGG^{+}23, VM19, XRX^{+}23]. Freshwaters [EFG^{+}23]. fridmani [CZC23]. Friendly [TPA^{+}24]. Frigate [ZCX^{+}22]. Fringe [BBN^{+}24]. Frog [WWW^{+}24]. Front [BS23]. Frontier [APAHBMAG23]. Frozen [CTY^{+}21, LJK^{+}22, RDI^{+}21]. Fructooligosaccharides [EAE^{+}23, PJPMMV^{+}22]. Fruit [HGSE23]. Fry [LLL^{+}23, LGZ^{+}23]. Full [TKF^{+}18]. fulvidraco [HWW^{+}24b, JLW^{+}24, PYJ^{+}23]. Function [JK24, LSY^{+}23, PJPMMV^{+}22]. Functional [CLJ^{+}23, Fra23, LMLH22, LHL^{+}24, LYL^{+}24a, Mil23, MPK^{+}23, OO24, YAEAB23, ZCT^{+}23, ZYH^{+}24]. Functionality [BSGMC^{+}22]. Functions [CL19, SRL^{+}19, TJW^{+}22]. Fund [BBF22]. Fungicide [NWN^{+}23]. Fungus [PSS^{+}18]. furcatus [FHHC23]. Furong [WZL^{+}24]. Fusarium [EAE^{+}23, PSS^{+}18]. fuscoguttatus [HAC^{+}23, SZZ^{+}23, WYG^{+}23, YGD^{+}23]. Future [APD^{+}23a, APD^{+}23b, EBM24, FCF19, MS20, SMO^{+}23].

G [Mil23]. G. [GP17, KLD^{+}23]. G720 [XLX^{+}22]. Gadiculus [GP17]. Gadidae [Gae17, Gae16, GP17]. Gadus [PH23, SCSS23, YHH^{+}20, vKRNL^{+}19]. Gafftopsail [OJC^{+}23]. Galápagos [JJB^{+}24]. Galaxaura [YZH^{+}24]. Galeus [BMMD22, DAF^{+}22]. galloprovincialis [LPF^{+}23]. Galvanizing [BASBW24]. Gaps [MMAO22]. Gar [FQÁGTR^{+}17, MPM^{+}21, MVPMAV^{+}22, NRÁGPM^{+}18, PJPMMV^{+}22, dlCRMB^{+}22]. gariepinus [AZY^{+}24, BHR^{+}23, Eny17, EPKV17, SBB^{+}19, WSE^{+}21]. Garlic [JROHVH^{+}23]. gasar [dOCCH^{+}23, SLSC^{+}24]. gasterostei [GMMNRS18]. Gasterosteus [BEMC23, DLL^{+}23, GMMNRS18, JBK^{+}23, KAJ^{+}24, SAL18]. Gastrointestinal [ATM^{+}24]. Gate [ZS21]. Gauging [SRBCGV^{+}21]. GC [PBS^{+}22]. GC-IMS [PBS^{+}22]. GCRV [YLX^{+}22]. Gear [MBC^{+}24]. Gehu [RJR^{+}22]. Gelatin [EEdCSOPJ^{+}23]. gen [CSJ^{+}23]. Gender [ZHF^{+}22b]. Gene [ASM^{+}22, CYL^{+}23, CMP^{+}23, ERE21, FCB^{+}21, GWT^{+}24, GLY^{+}23, GRKC19, GGF^{+}22, JKP^{+}23, KEA^{+}23, LCC^{+}24, LLY^{+}22, LYL^{+}24a, LWL^{+}23, MKC^{+}22, MWZ^{+}23, NRÁGPM^{+}18, PLJ^{+}23, PSO^{+}19, QWY^{+}24, SML^{+}23a, SZZ^{+}23, TU18, WLW^{+}23, WYG^{+}23, WPLK23, YZH^{+}24, YLX^{+}22, ZHF^{+}22a, ZQL^{+}23, ZWZ^{+}24, ZHQ^{+}23, ZCT^{+}23, ZYT^{+}24, ZZC^{+}22, ZZZ^{+}23, vKRNL^{+}19]. Generalist [BEG^{+}23]. Generation [BMTR23, HZG^{+}21, JFP^{+}18, PNW^{+}22]. Generational [DPGFL^{+}22]. Genes [CZW^{+}23, CBK^{+}21, CZC^{+}22, CWW^{+}23, CZL23, FLR^{+}22, JLW^{+}24, KJK^{+}23, LCWH22, MYY^{+}23, MVPMAV^{+}22, MBZ^{+}21, PJPMMV^{+}22, SLY^{+}21, SGG^{+}21, ZLW^{+}23, ZJJ^{+}22, ZYT^{+}24]. Genetic [ARH^{+}23, ANA^{+}23, APD^{+}23a, APD^{+}23b, CTTW23, CSB^{+}20, CMP^{+}23, FHF23, HCWH20, HHL^{+}20, HCZ^{+}23, HZZ^{+}23, JCL^{+}24, KBK^{+}23, LDBL19, LZC^{+}23b, MBPB24, NTP^{+}21, SWUH23, WTC^{+}22, WLZ^{+}22a, WFL^{+}23, ZLL^{+}23, ZGY^{+}23, ZZZ^{+}23]. Genetically [HLZ^{+}22, SML^{+}23a]. Genetics [BMOH23, FGHYCA23, Hal23, LYY^{+}24]. Genome [CZL23, DLL^{+}22, JJK21, LPK^{+}23a, LPK^{+}23b, LLY^{+}22, MBZ^{+}21, ZLW^{+}23]. Genome-Wide [LLY^{+}22, ZLW^{+}23]. Genomewide [GLY^{+}23]. Genomic [YZL^{+}23]. Genomics [KBK^{+}23]. Genotoxicity [AZY^{+}24, SAGG^{+}23]. Genotype [TBPJ23]. Genotype-1 [TBPJ23]. Genotypes [LBH^{+}24]. Gentian [FMXQ23, WLN^{+}23]. Genuine [XJC^{+}22]. Genus [APD24, MBD^{+}23, MLSC^{+}23, WLZ^{+}22a, WBK^{+}23]. Geographic [BGT^{+}20]. Geographic-Scale [BGT^{+}20]. Geography [RE21]. Geometric [KMSO18, XXL^{+}24]. Geosmin [LLKKV20]. Germ [YXS^{+}24, ZQLW23]. Gestation [UDG^{+}19]. GH [GRKC19, GLW^{+}22]. GH-Transgenic [GLW^{+}22]. GH/IGF [GRKC19]. Ghrelin [ZSZ^{+}23]. Giant [CWW^{+}23, HAC^{+}23, XRX^{+}23, ZHF^{+}22b]. Giants [TCD^{+}21]. Gibel [ULR^{+}23]. gibelio [ULR^{+}23]. GIFT [BRB^{+}23, HLZ^{+}22, ZZW^{+}22]. gigas [HZL^{+}22, SLSC^{+}24, ZXY^{+}24]. Gill [FMXQ23, FQÁGTR^{+}17, ZZT^{+}23, ZZW^{+}22]. Gillnet [FM21a, MGMG24, NIN^{+}19]. Gillnets [NNN23, YSY^{+}23]. Gills [CLL^{+}23b, CSJ^{+}23, HSZ^{+}22, MHZ^{+}23, SBS^{+}23, WLT23]. Gilthead [BSGMC^{+}22, CSB^{+}20, CCCFE18, CFCE20, ERE21, GMBR^{+}21, GMFG^{+}24, SNSG^{+}19, TJTV^{+}23, TMD^{+}19]. Giltheadsea [ZMLFS^{+}20]. Ginger [AEME^{+}23]. Ginkgo [CXW^{+}23]. Girella [KLD^{+}23]. Girellidae [KLD^{+}23]. Glacial [ANA^{+}23]. gladius [LLTM17, Poi24]. glanis [BBN^{+}24]. Glass [SLYY23]. glaucum [ADM^{+}24]. Glazed [CTY^{+}21]. Global [EN22, KKB20, PWCL23, SALC^{+}19, ZZW^{+}22]. Glochidia [MJL^{+}24]. glossodonta [KBK^{+}23]. GLP [ZSZ^{+}23]. GLP-1 [ZSZ^{+}23]. gls1 [XHC^{+}22]. Glucan [EAE^{+}23]. Glucans [NRÁGPM^{+}18]. Glucocorticoid [WLT23]. GLUT1 [ZLZL23]. Glutaminase [XHC^{+}22]. Glutamine [ZZC^{+}22]. Glycine [EPKV17]. Glycogen [WLC^{+}23, WLT23]. Glycolipid [HLZ^{+}22, YGD^{+}23]. Glycoprotein [JK24]. GnRHa [JFP^{+}18]. Goal [Liu24]. Gobiid [ZVRH23]. Gobiidae [KRAFO23, ZVRH23]. Gobiiform [KRAFO23]. Gobiiformes [KRAFO23]. Gobius [KRAFO23]. Goby [AAB^{+}18, CHH^{+}23, KAR^{+}23, LMMC^{+}22, PKSN23a, PKSN23b]. Golden [FŠS^{+}23, GGL^{+}23, WLW^{+}23]. Goldfish [HGSE23]. Goliath [AKM23, BC23, EKL^{+}23, MPK^{+}23, Tri23]. Gonad [BWKS20, LPB^{+}24, VCC^{+}18, ZQLW23]. Gonadal [CBK^{+}21, LXZ^{+}22, MKC^{+}22, YXS^{+}24, YYH^{+}24, ZLZ^{+}22, ZCT^{+}23]. Gonadotropin [PXH^{+}24, WSI^{+}19]. Gonadotropin-Releasing [WSI^{+}19]. Gonads [MYY^{+}23]. gonionotus [DRH18]. Gonochorism [MPK^{+}23]. Good [SMM^{+}18, SBG^{+}24]. Gourami [DAHM19, DM24]. Governance [CXL23]. Governing [And23]. Gracilaria [HSAF^{+}23]. Graded [SNK^{+}23]. Gradients [SML^{+}23a]. Gradual [TMM^{+}18]. grahami [ZXH^{+}23]. Grains [FDM^{+}23b]. Grande [CASMK23]. grandiflora [SBS^{+}23]. granosa [PLJ^{+}23]. Grape [BDŞ^{+}24, LMNN21]. Grapevine [HFEH^{+}23]. Grapsus [JJB^{+}24]. Grass [HHP^{+}24, HBG^{+}20, LWS^{+}21, LJR^{+}24, SFP17, WZX^{+}24, WWD^{+}23, YLX^{+}22, ZSQ^{+}21, ZSL^{+}24]. grayii [WTC^{+}22]. Grayling [ZLZ^{+}24]. Graysby [BPOS19]. Great [HL24, FLFF23]. Greater [GMA^{+}24, JFP^{+}18, NRKT19]. Greatest [EBM24]. Greece [TK24, CBP^{+}24, KMB18, KKB20]. Greek [GASS^{+}22, PB24, SKL^{+}23]. Green [CPJK23a, CPJK23b, GBT^{+}24, HAC^{+}23, LBH^{+}24, MCÁGHA^{+}17, YE20]. greenei [CTTW23]. gregaria [YMD^{+}21]. Grenadier [XSZ^{+}23]. Greville [HSAF^{+}23]. Grey [ATEfA^{+}21, EES^{+}23, GMCF^{+}22]. Ground [XLC^{+}24, ZXY^{+}24]. Grounds [AAN22, MMSK21]. Groundwater [BF16]. Group [ALNVDG^{+}22, CZL23, SKK^{+}23]. Grouper [AGA^{+}21, AKM23, BC23, EKL^{+}23, HWW^{+}24a, HAC^{+}23, MPK^{+}23, SWS22, SZZ^{+}23, Tri23, WYG^{+}23, WLN^{+}23, YGD^{+}23, LDZ^{+}24]. Grouper-Snapper [HWW^{+}24a]. Groupers [FMXQ23]. Grow [NdNFK^{+}24b]. Grow-Out [NdNFK^{+}24b]. Growth [ATEfA^{+}21, AEME^{+}23, ASM^{+}22, AGA^{+}21, BDŞ^{+}24, BFB^{+}23, BHR^{+}23, BRB^{+}23, BMTR23, BPOS19, BPO19, CSB^{+}20, CdOCH^{+}23, CVANRD^{+}21, CLL^{+}22, CXW^{+}23, CHJ^{+}23, CCFP19, CMC^{+}24, CBANCM^{+}21, CBCMRD^{+}23, DMA22, DLL^{+}23, DZC^{+}22, EAJ^{+}23, EAE^{+}23, EPKV17, EMFZ^{+}18, FZZ^{+}24, FB22, FM21b, FLB^{+}21, FCB^{+}21, GWH^{+}22, GL23, GFDPSR22, GMBR^{+}21, GMDMT^{+}23, GRKC19, GBT^{+}24, HGSE23, HFEH^{+}23, HS18, HPJ^{+}23, IM24, INCD23, JKP^{+}23, JKK24, JBdFS^{+}22, LLWW23, LdRCV24, LCWH22, LWT^{+}24, LSL^{+}24, LOT^{+}22, LCX^{+}23, LMNN21, MSB^{+}23, MWZ^{+}23, MPM^{+}21, MIHH23, MVPMAV^{+}22, MMVVCJ^{+}23, MOW^{+}18, MMM^{+}24, MPK^{+}23, Näs18, NRÁGPM^{+}18, PSS^{+}18, PNW^{+}22, PCH^{+}24, PJPMMV^{+}22, dAPAA^{+}24, QXM^{+}24, RRM^{+}20, RHU^{+}23, RDANPA^{+}24, RJTVC^{+}19, SSSP21, SNSVFL23, SBS^{+}23, SLY^{+}21, STZ^{+}23, SGG^{+}21, SBB^{+}19, SCT^{+}24, SJY^{+}22, SNK^{+}23, TJTV^{+}23, ULR^{+}23, VH20, WML^{+}21, WSD^{+}23, WYG^{+}23, WHL^{+}24, WZL^{+}24]. Growth [WFZ^{+}24a, WLW^{+}24, WLL^{+}24, WSA^{+}23, WSE^{+}21, WLZ^{+}22b, XLX^{+}22, XRX^{+}23, XCW^{+}23, XWP^{+}24, XWW^{+}24, YLW^{+}23b, YCR^{+}23, YHZ^{+}23, ZMLFS^{+}20, ZWZ22, ZSH^{+}23, ZYG^{+}23, ZWZ^{+}24, ZZ24, ZLX^{+}23, dAdSCC^{+}23]. Growth-Promoting [JBdFS^{+}22]. grubii [ZLZ^{+}24]. GSK3 [WLC^{+}23]. Guadiana [GTC^{+}17]. gualpensis [GLH^{+}23]. Guaratuba [CFM^{+}23]. Guaratuba-Babitonga-Itapocu [CFM^{+}23]. Gudgeon [ZHX^{+}24]. Guiana [AKM23]. guichenoti [ZHX^{+}24]. Guilds [Fra23, WSL^{+}23]. Gulf [AFTÁPA^{+}23, BMTR23, KMB18, LWD^{+}23, SWS22, USRDFO^{+}22, MSRGCG^{+}23, RGTPSCCC24, SXQ^{+}22]. Günther [AFTÁPA^{+}23]. Guppies [BEF^{+}23]. Guppy [GBT^{+}24]. Gut [ATEfA^{+}21, AML^{+}24, BSGMC^{+}22, CSL24, GMCF^{+}22, LHZ^{+}23, WLZ^{+}22b, XYC^{+}22, XYC^{+}23, YLW^{+}23a, YBL^{+}22, ZAM^{+}23]. guttatus [CVANRD^{+}21]. Gymnesigobius [KRAFO23]. Gymnocephalus [NHR20]. Gymnocypris [ZZ24]. Gyrodactylus [GMMNRS18]. GYS [WLC^{+}23].

H [WGRM^{+}19]. H. [Mil23]. Hāpuku [WSI^{+}19, WSS^{+}19]. Habit [LMLH22, WFZ^{+}23]. Habit-Specific [WFZ^{+}23]. Habitat [AKM23, Gre17, LHZ^{+}23, PEP24, SKF^{+}23, SMAR24, YWL^{+}24, ZCX^{+}22, ZVRH23]. Habitats [AM24, DMA22, MNP^{+}16, WSL^{+}23, YLH^{+}24]. Habits [BTSK24, MSRGCG^{+}23, OJC^{+}23, RGTPSCCC24, WLL^{+}23, XSZ^{+}23]. haematopterus [CHJ^{+}23, WSD^{+}23]. Haemocytes [ADM^{+}24]. Haemulon [BPO19]. Hagerman [CTTW23]. Hainan [LZL^{+}23]. Hairtail [FYL^{+}23, HWZ21, SKK^{+}23]. Hairy [LXT^{+}22]. Hake [AGC23, GFDPSR22]. Han [AMK23]. Handling [SBP23]. Happiness [RH19]. harak [MOW^{+}18]. Harbor [TTT23]. Hard [FCB^{+}21, JXW^{+}23]. Hardhead [OJC^{+}23]. harengus [BMTR23]. Harmful [JXW^{+}23, SAHS18]. Harnessing [LYY^{+}24]. Harvest [BGT^{+}20, MMSK21]. harveyi [LDZ^{+}24]. hasta [MMY^{+}17]. Hatcheries [BKJ^{+}24]. Hatchery [TRM^{+}23, VRKV24]. Hatching [CHJ^{+}23, Ord19, VKP^{+}24]. hCG [LPB^{+}24]. Head [JLW^{+}24, LLL^{+}23, MPM^{+}18, SMO^{+}22]. Head-Kidney [SMO^{+}22]. Head-to-Tail [LLL^{+}23]. Heads [NWN^{+}21, NWN^{+}23]. Health [ASM^{+}22, BSGMC^{+}22, DCD^{+}24, GMA^{+}24, LZH^{+}23, LW18, OOTC24, Sha19, SOW^{+}23, SNK^{+}23, ULR^{+}23, WHX^{+}23, WLW^{+}24, WFZ^{+}23, XWP^{+}24, ZWP^{+}23]. Heart [CLJ^{+}23, PLY^{+}24]. Hearts [ZMA^{+}24]. Heat [HZS^{+}21, ZHF^{+}22a]. Heatwaves [LPF^{+}23]. Heavy [MS18, WFZ^{+}23]. Heckel [APD24]. Heel [BAA^{+}23]. Helminths [EFG^{+}23, EFQ23]. Hematological [DCR^{+}23, JKK24, SCCM23]. Hematology [HS18, RHU^{+}23]. Hemorrhagic [CCP^{+}24, QXAY22]. Hepatic [HLZ^{+}22, KWK^{+}24, KSSI24, LYL^{+}24b, MMM^{+}24, YZH^{+}24]. Hepatopancreas [HWX^{+}23]. Hepatopancreatic [LJK^{+}22]. hepatopenaei [HWX^{+}23]. Hepatorenal [RIF^{+}23]. hepatus [MSRGCG^{+}23]. hepuensis [FLR^{+}22]. Herbicide [MTHPJS^{+}23]. Heritability [CSB^{+}20, XWW^{+}24]. Hermaphroditism [DAF^{+}22]. Hermetia [CSR22, FLB^{+}21, LWT^{+}24, NYS^{+}23, YAEAB23]. Herpesvirus [SLC^{+}22]. Herring [BMTR23, MNP^{+}16]. Herzenstein [LLWW23]. Heterodontus [PHB^{+}23].

Heterogeneous [CFLK21, WTC^{+}22]. Heterosigma [JXW^{+}23]. Heterospecific [LLY^{+}24]. Heterozygosity [MBPB24]. Hidden [TCD^{+}21]. Hierarchical [YJK23]. HIF [PLJ^{+}23]. HIF-1 [PLJ^{+}23]. High [BEMC23, dSGBdF23, HJE^{+}23, HYN^{+}24, HLZ^{+}22, LSL^{+}24, MBAM19, MDVM^{+}23, NDC^{+}23, NYS^{+}23, PLY^{+}24, QL22, RCR^{+}23, Tay19, TMD^{+}19, WGW^{+}23, WZX^{+}24, XYL^{+}23, SBG^{+}24]. High-Density [NYS^{+}23]. High-Fat [WZX^{+}24]. High-Resolution [BEMC23]. High-Seas [QL22]. High-Temperature [MDVM^{+}23, PLY^{+}24]. Highland [KYB^{+}23]. Hilsa [FM21a]. Hinders [RLB^{+}23]. Hippocampus [STZ^{+}23, ZQLW23]. hippurus [Poi24]. Histamine [WHX^{+}23]. Histoarchitecture [AEME^{+}23]. Histological [SMH^{+}24, SBP23]. Histology [HS18, PLY^{+}24, SGG^{+}21, TJTV^{+}23]. Histomorphology [FMXQ23, SBS^{+}23, WFZ^{+}24a]. Histone [LLY^{+}22]. Histopathological [AMT^{+}24, EAJ^{+}23, HYN^{+}24]. Histopathology [CLL^{+}23b, EES^{+}23]. Historical [APD^{+}23a, APD^{+}23b, LBC^{+}24, PB24]. History [SMO^{+}23, SMH^{+}22a, VH20, WTC^{+}22, YE20]. Hobbyists [VM19]. Holistic [FCT19]. Holocephali [SQ23]. Home [HMVRFD19]. Homeostasis [WLN^{+}23]. Hong [QWY^{+}24]. hongkongensis [LYL^{+}24a, QWY^{+}24]. Hook [YXS^{+}24]. Hookline [RJFCJC^{+}21]. Horizon [HDW^{+}23]. Horizontal [LCW23a]. Hormonal [PKV^{+}22]. Hormone [CMP^{+}23, CPJK23a, CPJK23b, KJ22, PXH^{+}24, WCY^{+}24, WSI^{+}19]. Hormones [CBK^{+}21, PXH^{+}24, ZWZ^{+}24]. Horse [MSA24]. Host [GMMNRS18, JKP^{+}23, LHZ^{+}23]. Host-Associated [JKP^{+}23]. Hosts [PYP17]. Hot [XYT23]. Hours [RDI^{+}21].

Housefly [HIO^{+}19]. HPG [CZW^{+}23]. HSP [FLR^{+}22]. Hsp47 [LWS^{+}21]. HSP60 [ZHF^{+}22a]. HSPs [ZHF^{+}22a]. Huanghe [LHL^{+}24, WSD^{+}23]. Hubbs [CSJ^{+}23]. Hucho [LDX^{+}23, WWY^{+}23]. Hue [LYY^{+}24]. Human [DCD^{+}24, LW18, RCL^{+}23]. Humoral [SSSP21]. Huron [PKSN23a, PKSN23b]. Husbandry [PLV^{+}19]. Huso [JCR^{+}22, Kim23a, Kim23b]. Hyalella [BSH^{+}23]. Hybrid [FHHC23, HAC^{+}23, JCR^{+}22, Kim23a, Kim23b, MMVVCJ^{+}23, PYJ^{+}23, SZZ^{+}23, WFX^{+}23, XWP^{+}24, YGD^{+}23, ZSL^{+}23]. hybridus [MBL20]. Hydrodynamic [WXL^{+}23, ZTWK24]. Hydrolagus [SQ23]. Hydrological [dSCFQB^{+}23]. Hydrologically [FWJ21, HBL^{+}22]. Hydrolysate [BHP^{+}24, SNK^{+}23]. Hydrolysates [NdNFK^{+}24b, NdNFK^{+}24a]. Hydrolysis [MABÁMSM22]. Hydropeaking [WSA^{+}23]. hydrophila [AEME^{+}23, GWT^{+}24, JLW^{+}24, LCWH22, MYW^{+}24, PYJ^{+}23, PSN18, dCPRG^{+}21, SSSP21, SNK^{+}23, ZLSB22]. Hydrophobic [HIO^{+}19]. Hydropower [BFLC19]. Hydroquinone [KSSI24]. Hyper [LYL^{+}24a]. Hyper-Salinity [LYL^{+}24a]. Hypersaline [DTS^{+}17]. Hypophthalmichthys [MGS^{+}23, PVY^{+}21, RMA^{+}18, ZLZL23, ZJC^{+}22b]. hypophthalmus [VH20]. Hypoxia [PLJ^{+}23, WLN^{+}23, ZLZL23, ZZW^{+}22].

I. [FHHC23]. Ib [AMT^{+}24]. Iberian [RGTPSCCC24, SGAC20, PAMG19, SRBCGV^{+}21]. Ice [FÅG^{+}23]. Iced [JROHVH^{+}23]. Icefish [TJW^{+}22]. Ichthyofauna [LKD22]. Ichthyoplankton [PWH^{+}23, SCBSSMA24]. Ichthyoplanktonic [AGC23]. ictaluri [YLW^{+}23a]. Ictalurus [FHHC23, KBCM19, QXAY22]. Idaho [CTTW23]. idella [LWS^{+}21]. idellus [HHP^{+}24, WLL^{+}22, WZX^{+}24, ZSQ^{+}21, ZSL^{+}24]. Identification [AAN22, CMP^{+}23, CWW^{+}23, EEdCSOPJ^{+}23, ENO21, FCB^{+}21, GLY^{+}23, KLD^{+}23, LSJ^{+}22, LSJ24, MMD^{+}23, OMC^{+}19, PWH^{+}23, QXAY22, QWY^{+}24, SBP23, SLY^{+}21, VAT^{+}23, WLC^{+}23, WFX^{+}23, XLP23, XXL^{+}24, XXL^{+}22, ZLW^{+}23, ZLSB22]. Identified [ARH^{+}23, UGH^{+}24, ZJJ^{+}22]. Identifies [RWF^{+}23]. Identify [EAJG24, YXS^{+}24]. Identifying [CLZ^{+}24, UGH^{+}24]. IGF [GRKC19]. IGF1 [Web23]. IGF2 [Web23]. Iglésias [KRAFO23]. IgM [ZLSB22]. II [FYL^{+}23, GLY^{+}23, KSSI24]. III [AMT^{+}24]. IL-17 [WSZ^{+}23]. Illegal [CXL23]. illucens [CSR22, FLB^{+}21, LWT^{+}24, NYS^{+}23, YAEAB23]. Illustrate [PMFBI22]. Image [CCCFE18, JTS^{+}24]. Images [JTS^{+}24, PSP^{+}22]. Imbalance [AZY^{+}24]. immaculata [SCT^{+}24]. Immortalized [CLJ^{+}23]. Immune [AEME^{+}23, CAC^{+}17, CFCE20, CSV^{+}19, CSSMV^{+}23, EAE^{+}23, GGL^{+}23, GSHGE18, HLZ^{+}22, JLW^{+}24, LWZ^{+}22, LCX^{+}23, MHZ^{+}23, MZA^{+}23, NRÁGPM^{+}18, PSO^{+}19, RIF^{+}23, RHU^{+}23, RXL^{+}24, SMO^{+}22, SMH^{+}24, SEA^{+}23, SPQ^{+}24, VCC^{+}18, VPP^{+}22, WML^{+}21, WLL^{+}22, WFZ^{+}24a, WBK^{+}23, XYF^{+}24, YMD^{+}21, ZAM^{+}23, ZMA^{+}24, ZHQ^{+}23]. Immune-Endocrine [VCC^{+}18]. Immune-Related [JLW^{+}24, PSO^{+}19]. Immunity [EES^{+}23, GBT^{+}24, HFEH^{+}23, HYN^{+}24, LCWH22, LDD^{+}22, LWT^{+}24, LHX^{+}23, MZA^{+}23, MMM^{+}24, QXM^{+}24, SSSP21, XLX^{+}22, YZH^{+}24, YHZ^{+}23, Zac22, ZYG^{+}23]. Immunization [PSN18]. Immunoassay [SLC^{+}22]. Immunohistochemical [SMH^{+}24, ZMA^{+}24]. Immunological [DCL^{+}23a, HGSE23, SAC23]. Immunoprotective [GYH^{+}23]. Immunoregulation [SCHT23]. Impact [AZY^{+}24, BAA^{+}23, CXT^{+}24, DPGFL^{+}22, EBH21, IBN^{+}23, INCD23, JHLW24, JKK24, MBC^{+}24, MFKS23, NKC^{+}21, RIF^{+}23, SRHCO23, SZSW21, VSH23]. Impacts [CBP^{+}24, CS23, DLL^{+}23, HJE^{+}23, KFS23, LFH^{+}23, MML22, NGMR23, ZSH^{+}23, ZTWK24]. Impaired [LMMC^{+}22]. Impairment [BB23]. Implementing [SS23]. Implication [BFB^{+}23]. Implications [AHLC19, APD^{+}23a, APD^{+}23b, BKBR^{+}23, ENO21, FCT19, TPC^{+}23, dlCRMB^{+}22]. Important [BAP22, Gre17, RKHAMM22]. Imported [SBS^{+}24]. Imprecise [PWH^{+}23]. Improve [BF16, NNN23, ZSQ^{+}21]. Improved [HARB23a, HLZ^{+}22, LWZ^{+}23, MOW^{+}18, WLZ^{+}23, WYL23, YZH^{+}24]. Improvement [GLX^{+}23, JROHVH^{+}23, MGS^{+}23]. Improves [FDM^{+}23b, MABÁMSM22, VH20]. IMS [PBS^{+}22]. IMTA [CMB^{+}24, CMC^{+}24]. In-Pond [FHHC23]. Inactivated [XYF^{+}24]. Inadequate [PWH^{+}23]. Including [GMFG^{+}24]. Inclusion [BHP^{+}24, MVPMAV^{+}22, MMVVCJ^{+}23, RRM^{+}20]. Income [MFKS23]. Incomes [WM22]. Incorporating [QL22]. Incorporation [PJPMMV^{+}22]. Increased [PHB^{+}23, WSA^{+}23]. Increases [BRB^{+}23, WWW19, ZSQ^{+}21]. Increasing [ATEfA^{+}21, SMM^{+}18]. Increments [MOW^{+}18]. Index [NIN^{+}19, SXQ^{+}22]. India [LBC^{+}24]. Indian [SQ23, SMM^{+}18, YWL^{+}24, SZSW21, WZG^{+}23]. Indicate [PYP17]. Indicator [BB23, LSY^{+}17, LCWH22, Liu24, MBP^{+}24]. Indices [AGC23, CMC^{+}24, CXT^{+}24, KŽM^{+}23, RIF^{+}23, TMPP23, XYC^{+}22, XYC^{+}23, YLH^{+}24]. indie [LCWH22]. Indigenous [GTC^{+}17]. Individual [CBANCM^{+}21, GFDPSR22, KCCK23, WSA^{+}23]. Individual-Based [WSA^{+}23]. Individuals [MGMG24]. Indo [WSON24]. Indonesia [HWW^{+}24a, MMSK21, MFKS23, NIN^{+}19, SZSW21, WLM^{+}20]. Induce [VMS^{+}23]. Induced [AZY^{+}24, CLL^{+}23b, ERE21, HAdM^{+}24, LPF^{+}23, SCSR22, SPQ^{+}24, WSI^{+}19, ZWZ^{+}24]. Induces [BWKS20, CGSBGN24, MDVM^{+}23]. Inducible [CAC^{+}17]. Induction [JFP^{+}18]. Industrial [ZWP^{+}23]. Industries [AJF23]. Industry [EvSCB23, HMN^{+}22, Ho22, XLP23]. Infected [CSSMV^{+}23, GGL^{+}23, SEA^{+}23, ZZL^{+}23]. Infection [AEME^{+}23, AMT^{+}24, EAE^{+}23, EES^{+}23, GYH^{+}23, HWX^{+}23, KJK^{+}23, LTZ^{+}22, PVY^{+}21, VCC^{+}18, VPP^{+}22, WWW^{+}24, WBK^{+}23, YMD^{+}21, YLX^{+}22, ZHF^{+}22a]. Infections [CCP^{+}24, LCWH22]. Infectivity [LJK^{+}22]. Infestation [PSO^{+}19]. Infestations [IR22, MNO^{+}22]. Inflammation [MYW^{+}24]. Inflammatory [WLW^{+}23]. Influence [BCG^{+}23, CdOCH^{+}23, CHW21, DCR^{+}23, DSC^{+}19, HAdM^{+}24, JBK^{+}23, MPM^{+}21, MCSB^{+}19, PLC^{+}24, PJPMMV^{+}22, PRPW23, SWUH23, SCSS23, SMAR24, SSK^{+}22, WSA^{+}23, YGD^{+}23]. Influenced [RMA^{+}18]. Influences [MWZ^{+}23]. Influencing [RAR^{+}18]. Inform [DOB^{+}17, EBM24]. Information [KTT24]. Informative [NTP^{+}21]. Informing [WWD^{+}23]. Infrastructures [SGANM^{+}24]. Ingredient [YAEAB23]. Ingredients [GHS20, MABÁMSM22, MBAM19, MLH^{+}20, ZHHO23]. Inhibited [PSS^{+}18]. Inhibiting [LWS^{+}23a]. Inhibitors [SSK^{+}22]. Inhibitory [DDN19, WZX^{+}24]. Inhibits [LWS^{+}23a, XCW^{+}23]. iniae [LBH^{+}24]. Initial [LHX^{+}23]. Injection [LLW^{+}22]. Inland [PB24]. Innate [LCWH22, MMM^{+}24, YMD^{+}21]. Innovation [JHLW24, LKJ22]. Innovative [MBC^{+}24]. Inorganic [MMMNAG23]. Insect [HGC^{+}23, RRM^{+}20, SAB^{+}22]. Insects [HMN^{+}22]. Insight [GTC^{+}17, LOBTL22, WTC^{+}22]. Insights [GMA^{+}24, HMN^{+}22, LCW^{+}23b, LCC^{+}24, MBPB24, PMFBI22, PB24, Poi24, PEP24, SDA23, Sus20, SAHS18, WLC^{+}23, YWDP21, ZQL^{+}23, ZSL^{+}23, FDM^{+}23a]. instead [WLL^{+}22]. Instream [FÅG^{+}23]. Insular [Tay19]. Insulin [ZWZ^{+}24]. Insulin-like [ZWZ^{+}24]. Insurance [WHM21]. Intake [LdRCV24]. Integrated [CdOCH^{+}23, CMB^{+}24, MBL20]. Integrating [GSK^{+}21, NPT^{+}24]. Integration [PB24]. Integrative [LBC^{+}24, dSSBdS23]. Integrity [ATEfA^{+}21, SXQ^{+}22]. Intelligence [CZCW23]. Intelligent [ILA22]. Intensification [VRKV24]. Intensity [FBSB24]. Intensive [WLZ^{+}23, ZTWK24]. Inter [ZJC^{+}22a]. Inter-Otolith [ZJC^{+}22a]. Interaction [MHZ^{+}23, YJK23]. Interactions [CFLK21, EFQ23, VCC^{+}18, Zac22]. Interactive [LSL^{+}24]. Interference [ZYT^{+}24]. Interferon [CWW^{+}23, KJK^{+}23]. Intergeneric [VMBT24]. Intermediary [FMBPC^{+}23]. Internal [ADM^{+}24]. International [CXL23, PWCL23]. Internationalization [NRKT19]. Intersected [APS^{+}21]. Intersexuality [DAF^{+}22]. Interspecific [LZL^{+}23]. Intervention [MTPK23]. Intestinal [CBCL23, DBP^{+}20, DCL^{+}23b, FLB^{+}21, GMDMT^{+}23, HWW^{+}24b, HLZ^{+}22, LCW^{+}23b, LZH^{+}23, MVPMAV^{+}22, MYW^{+}24, Näs18, NRÁGPM^{+}18, PJPMMV^{+}22, RCR^{+}23, SJY^{+}22, SNK^{+}23, TU18, WLW^{+}23, WHX^{+}23, WFZ^{+}24a, WLW^{+}24, XLX^{+}22, XRX^{+}23, XWP^{+}24, YGD^{+}23, YLH^{+}24, YHZ^{+}23, ZWP^{+}23]. Intestine [GMFG^{+}24, HSZ^{+}22, SBS^{+}23, WZL^{+}24, XYL^{+}23]. Intraovarian [UDG^{+}19]. Intraperitoneal [AMT^{+}24]. Introducing [XQLA18]. Introduction [EFQ23, GTC^{+}17, SBT22, Sor21]. Introgression [VMBT24]. Intubated [JCR^{+}22]. Invaded [FGR19]. Invader [KKB20]. Invasive [BEG^{+}23, dRCCLRNRWK21, CFLK21, DMB^{+}20, EFG^{+}23, Gre17, HVRCG18, KAB^{+}20, KAR^{+}23, LUM18, LV23, NHR20, Sor21, Sus20, WWD^{+}23, YWDP21, ZS21]. Invasiveness [BSR^{+}20]. Invasives [HDW^{+}23]. Invasivorism [BGT^{+}20]. Invertebrates [DPGFL^{+}22]. Investigating [AGC23, YSY^{+}23]. Investigation [JML^{+}24, SMH^{+}24, SNZ^{+}23, WXL^{+}23]. Ionian [LV23]. Ionic [Ord19]. Ionoregulatory [CLL^{+}23b]. Iowa [Mee24]. Ipomoea [RXL^{+}24]. IPS [LSJ^{+}23]. Iran [SKD^{+}23]. Iridovirus [AVT18, LDZ^{+}24]. Iridoviruses [JJK21]. Iron [LYW^{+}24, LSJ^{+}23]. irradians [HHM^{+}24]. Irradiated [HGSE23]. Irradiation [LSY^{+}23]. Island [EFG^{+}23, KSAB^{+}23, LZL^{+}23, TK24, SQ23, SLL22, XW24]. Islands [FPRAT^{+}23, JFP^{+}18, CLL^{+}22, CSSMV^{+}23, LSJ24, TCB^{+}24]. Isolated [BTSK24, JJK21]. Isolation [QXAY22, VCL20, WFX^{+}23, WCX^{+}24, XXL^{+}22]. Isothermal [LDZ^{+}24]. Isotope [DL23]. Isotopic [BFM23, FWJ21]. Israel [FGBA^{+}23]. Issue [Moy18, Sor21]. Issues [QL22, VMBT24]. Isurus [VV23]. itajara [AKM23, Tri23]. Italy [MDF^{+}23, SAL18]. Itapocu [CFM^{+}23]. IV [FYL^{+}23]. Ivory [DLL^{+}22].

Jack [BS23, FGHYCA23]. Jackson [PHB^{+}23]. Japan [LOT^{+}22]. Japanese [FZZ^{+}24, HXY^{+}23, IKT19, LPB^{+}24, MLK^{+}19, MU21, SLYY23, YLW^{+}23b, ZLZ^{+}22]. japonica [JKP^{+}23, LPB^{+}24, MLK^{+}19, SLYY23]. japonicus [CKMT23, CPJK23a, CPJK23b, FZZ^{+}24, FYL^{+}23, HXY^{+}23, KMSO18, SKK^{+}23, YBL^{+}22]. Java [MMSK21, ST17]. Jaw [TRM^{+}23]. Jeldes [GLH^{+}23]. Jellyfish [EEdCSOPJ^{+}23]. Jenyns [BSR^{+}20]. Jian [CXW^{+}23, XCW^{+}23]. Jigger [SZT^{+}23]. Jigging [FYH^{+}23].

Jinjiang [GLY^{+}23, WLC^{+}23]. Jinshaia [LZC^{+}23b]. Johnston [LL22]. Journal [Est16]. JTED [CLL23a]. Jumbo [HZL^{+}22, ZXY^{+}24]. June [LXT^{+}22]. Juvenile [ASM^{+}22, ALNVDG^{+}22, BF16, CSR22, DOB^{+}17, DZC^{+}22, EAJG24, FÅG^{+}23, GWH^{+}22, GZF^{+}22, HS18, IM24, KBCM19, KKFC^{+}23, LHX^{+}23, LWS^{+}23b, LZW^{+}24, MYW^{+}24, Näs18, RJMVC^{+}18, SBS^{+}23, SCCM23, SBB^{+}19, ULR^{+}23, WLW^{+}23, WLL^{+}23, WHX^{+}23, WFZ^{+}24a, WSA^{+}23, WSE^{+}21, WWY^{+}23, WWZ^{+}24, XWP^{+}24, ZWZ22, ZYG^{+}23, ZWP^{+}23, ZJC^{+}22a, dAdSCC^{+}23, CLL23a]. Juveniles [AASQPU^{+}23, BKBR^{+}23, FDM^{+}23b, GRKC19, INCD23, LXT^{+}22, MJL^{+}24, MdM^{+}23, MMVVCJ^{+}23, MLH^{+}20, NRÁGPM^{+}18, VPP^{+}22, WLZ^{+}22b].

kanagurta [JTS^{+}24]. Kariba [SNZ^{+}23]. Karun [SKD^{+}23]. Katsuwonus [NIN^{+}19]. Kelly [PKSN23a]. Kenai [FCT19]. kenojei [YCR^{+}23]. Kernel [WZX^{+}24]. Kessler [KZC^{+}24]. keta [WLL^{+}23]. Key [CZW^{+}23]. Keypoint [NLTL23]. Kidney [JLW^{+}24, MPM^{+}18, SMO^{+}22]. Kidneys [SPQ^{+}24]. Killed [LDD^{+}22]. Kim [Kim23a]. Kin [Tri23]. Kinase [BB23]. Kingfish [DBP^{+}20]. kisutch [ZYG^{+}23]. Knowledge [KSAB^{+}23, PLV^{+}19]. Known [HDW^{+}23]. Koi [LWT^{+}24, SLC^{+}22, LWT^{+}24]. Kokanee [SBT22]. Kong [QWY^{+}24]. Korea [AMK23, AML^{+}24, JJK21, JKK^{+}20, KKPY22, LPK^{+}23b, YCR^{+}23, YJK23]. Korean [SKK^{+}23]. Korill [NDC^{+}23]. Kovacić [KRAFO23]. Krefft [LCW23a]. Krill [BCG^{+}23]. Krøyer [CG20]. Kyoga [NGMR23].

L [HFEH^{+}23, RDE^{+}23, ZJC^{+}22b]. L. [AGC23, BAP22, CCCFE18, ERE21, GMBR^{+}21, STZ^{+}23, SSK^{+}22, SSSS23, WLL^{+}22, vKRNL^{+}19]. Lab [AJF^{+}22].

Labeo [RKHAMM22, SAGG^{+}23]. Labidochromis [SSSP21]. Laboratory [CBCMRD^{+}23, GGP^{+}23]. Laboratory-Conditioned [GGP^{+}23]. Labour [TTT23]. labrax [ALNVDG^{+}22, BGMM^{+}24, CFCE20, FDM^{+}23b, NZVB20, RRM^{+}20, RGTPSCCC24, RSA17, SAB^{+}22, VPP^{+}22]. labrosus [GMCF^{+}22]. Lacepède [TK24]. Lactobacillus [GCFA^{+}22, SPQ^{+}24]. lactuca [CdOCH^{+}23]. Lagocephalus [CBP^{+}24]. Lagoon [AHK^{+}23, AIŠRB22, dRCCLRNRWK21, KKB20, ISA^{+}22]. Lagoons [EFG^{+}23, XQLA18]. lagowskii [HYN^{+}24]. Lake [KAJ^{+}24, LXT^{+}22, PVY^{+}21, RJR^{+}22, SGAC20, SMH^{+}22a, VF23, XJC^{+}22, YWDP21, BWKS20, JSSD23, KAB^{+}20, MTPK23, NGMR23, OOGAS23, SAL18, SNZ^{+}23, SVMLP23, WFZ^{+}24b, YAS21, ZZW^{+}24]. Lakes [CH23, Mee24, SB20, PKSN23a, PKSN23b]. lalandi [DBP^{+}20, TRM^{+}23]. Lamniformes [VV23]. Lamprey [KSO^{+}23, RH19, YAS21]. lanceolatus [HAC^{+}23, SZZ^{+}23, YGD^{+}23]. Land [LUM18, NGMR23, PNW^{+}22]. Landings [RLAE23]. Landscape [LUM18]. Lanka [JSLE23]. lanzhouensis [YZL^{+}23]. Large [BFLC19, BSB^{+}23, CTY^{+}21, CXT^{+}24, LTZ^{+}22, Mee24, SBT22, VKP^{+}24, YWDP21]. Largehead [FYL^{+}23, SKK^{+}23]. Largemouth [KKPY22, SJY^{+}22, SHW^{+}23, XYL^{+}23, XXL^{+}22, ZZY^{+}23, ZHX^{+}24]. Larger [SCSS23]. Largest [WFZ^{+}24b]. Larvae [AJF^{+}22, BKBR^{+}23, CSR22, CFP^{+}23, CCFP19, DMT^{+}19, DDN19, FQÁGTR^{+}17, HIO^{+}19, JKP^{+}23, LWT^{+}24, MVPMAV^{+}22, Mil23, NYS^{+}23, PAVCCJVV24, PJPMMV^{+}22, RMSPMC^{+}22, SdSdOSSL23, ŠT19, TRM^{+}23, YAEAB23, Mil23]. Larval [BAA^{+}19, CZC23, CHH^{+}23, GP17, GWH^{+}22, GMA^{+}24, MMMNAG23, NYS^{+}23, PSN18, PLV^{+}19, PFM^{+}20, Poi24, RHUJ24, WSL^{+}23, dlCRMB^{+}22]. Larviculture [SIZ^{+}22]. Late [PH23]. Lateolabrax [HXY^{+}23, LZH^{+}23, WSZ^{+}23, YSG^{+}23]. Lateral [PRPW23]. Latest [HGC^{+}23]. latifrons [RMSPMC^{+}22]. latipes [MU21]. latipinnis [RHUJ24]. Latvian [KAR^{+}23]. lavaretus [LLKKV20]. Lawrence [BMTR23]. lazera [AAAF^{+}21]. LBB [SLS^{+}22]. LC [FYL^{+}23]. LC-MS [FYL^{+}23]. LCDV [CSB^{+}20]. Leads [FM21a]. Leaf [HFEH^{+}23]. Learned [Sus20]. Learning [BEF^{+}23, HARB23a, ILA22, JBGCG^{+}24, JTS^{+}24, SZT^{+}23, XLC^{+}24]. Learning-Based [XLC^{+}24]. Lease [XW24]. Leaves [CXW^{+}23]. Lecithin [BRB^{+}23, RGVG19]. Lectin [CSF^{+}23]. Leg [HSD^{+}24]. Legal [CXL23, XLP23]. Legislative [SKSL23]. lemuru [SZSW21]. Length [BLA^{+}22, CVANRD^{+}21, CP24, KZC^{+}24, KSWT22, LFH^{+}23, SGAC20, SLS^{+}22]. Length-Based [BLA^{+}22, KSWT22, SLS^{+}22]. lenok [PXH^{+}24]. lentillifera [LMNN21]. Lentinus [XYL^{+}23]. Leone [OOG24]. leonina [KLD^{+}23]. Lepeophtheirus [CG20, IR22, IBN^{+}23]. Leptin [CZW^{+}23]. Leptocephali [Mil23]. Lessons [OOGAS23, QWR^{+}23, Sus20, YWDP21]. Lethal [LMB^{+}23]. Lethrinus [MOW^{+}18]. Lettuce [CdOCH^{+}23, SAB^{+}22]. Leuciscus [ZLW^{+}23]. Level [ATEfA^{+}21, BAA^{+}23]. Levels [ASM^{+}22, CGSBGN24, IM24, JJB^{+}24, MLK^{+}19, MFKS23, NZVB20, PAVCCJVV24, SBS^{+}23, SNK^{+}23]. Liaohe [ZGY^{+}23]. Lice [IBN^{+}23, PSO^{+}19]. Life [CTY^{+}21, CMP^{+}20, CS23, DQC^{+}23, JROHVH^{+}23, MKH24, PMFBI22, QAC23, SMO^{+}23, SMH^{+}22a, VH20, YE20, ZMLFS^{+}20]. Life-Long [QAC23]. Light [CFCE20, FÅG^{+}23, FBSB24, GRO^{+}17, KBCM19]. Lightweight [WYL23]. like [LYL^{+}24a, YSG^{+}23, ZWZ^{+}24]. Limit [MDLA22, Tri23, ZZW^{+}22]. Limitations [Liu24]. Limited [PB24, SKT23]. Limiting [ŢPN^{+}23]. Limonene [dSJC^{+}21]. Lindseth [LL18a]. Line [CLJ^{+}23, DCL^{+}23b, MBPB24]. Lineages [KAJ^{+}24]. Linear [EM23]. Lines [CSGE23]. Linnaeus [CG20, DCR^{+}23, DMA22, FDM^{+}23a, FWJ21, KYB^{+}23, OGMG^{+}17, SAL18, WBK^{+}23, ZJC^{+}22a]. Linneaus [MSRGCG^{+}23]. Linseed [GMFG^{+}24]. Lionfish [BEG^{+}23]. Lipid [ASM^{+}22, BRB^{+}23, CXW^{+}23, DL23, DSC^{+}23, EPKV17, GRKC19, LHX^{+}23, PAVCCJVV24, RWF^{+}23, RGVG19, WGRM^{+}19, WSD^{+}23, WZX^{+}24, XCW^{+}23, YGD^{+}23, ZSZ^{+}23, ZHX^{+}24]. Lipids [MKN^{+}16, MNP^{+}16, SNSG^{+}19]. Lipopolysaccharide [SOW^{+}23, ZHF^{+}22a]. Lipped [GMCF^{+}22]. Liquid [JSLE23]. Literature [AHLC19, CG20, CS23, DDG^{+}22, GSK^{+}21, WWD^{+}23]. Lithobates [MTHPJS^{+}23]. Lithuanian [AIŠRB22, ISA^{+}22]. Litopenaeus [BHP^{+}24, CdOCH^{+}23, EAE^{+}23, HSAF^{+}23, HSD^{+}24, LXW^{+}24, LSL^{+}24, LMNN21, SCHT23, SEA^{+}23, WPLK23, YHZ^{+}23, ZLL^{+}23, ZWD^{+}23]. Live [DDN19, DZC^{+}22, LF24, VH20]. Livelihoods [SRHCO23]. Liver [CZW^{+}23, CBK^{+}21, FMBPC^{+}23, MMVVCJ^{+}23, MTHPJS^{+}23, MKN^{+}16, SBS^{+}23, SNK^{+}23, WLL^{+}22, XYL^{+}23, XWP^{+}24, ZJM^{+}23, ZZY^{+}23]. Livestock [GMRJ22, NLTL23]. lividus [SSSP21]. liza [MMM^{+}24]. Lizardfish [GMA^{+}24]. Loach [ZSL^{+}23]. Loaches [CSL24, FŠS^{+}23, GWH^{+}22]. Lobster [MMSK21]. Lobsters [ATM^{+}24]. Local [dSCFQB^{+}23, HLK^{+}23, KAB^{+}23, KYB^{+}23, MLH^{+}20]. Localization [ZAM^{+}23]. Locally [EMFZ^{+}18]. Loci [APD24]. Lock [ZS21]. Locks [ZS21]. Locks-and-Dams [ZS21]. Locomotor [OGMG^{+}17]. Logistic [RDANPA^{+}24]. Loins [EA18]. Long [BPCC23, CCP^{+}24, CTY^{+}21, DLL^{+}23, KJ22, KBB^{+}21, Näs18, PH23, QAC23, SXQ^{+}22, TRM^{+}23]. Long-Chain [TRM^{+}23]. Long-Short-Term [CTY^{+}21]. Long-Term [BPCC23, DLL^{+}23, KJ22, KBB^{+}21, Näs18, PH23, SXQ^{+}22, WLL^{+}24]. Longer [SAL19]. Longfin [AASQPU^{+}23, MLK^{+}19]. longicrura [KMB18]. Longline [WXL^{+}23]. Longliners [LSJ24]. Loop [LDZ^{+}24]. Loop-Mediated [LDZ^{+}24]. Lophiosilurus [SdSdOSSL23]. Loricariidae [HVRCG18]. Loss [LWZ^{+}23, LSY^{+}23]. Loss-of-Function [LSY^{+}23]. Losses [TVL21]. Louse [CG20]. Low [DRH18, HMP^{+}24, HWW^{+}24b, MD21, MBPB24, TMD^{+}19, WLW^{+}24]. Low-Cost [DRH18]. Low-Fish-Meal [WLW^{+}24]. Low-Heterozygosity [MBPB24]. Low-Temperature [HWW^{+}24b]. Lower [SLDC23, dSCFQB^{+}23, HHAG24, LHG^{+}23]. LPS [SOW^{+}23, WSZ^{+}23, ZHF^{+}22a]. lucerna [FDM^{+}23a]. lucidus [XZZ^{+}24]. Luciobarbus [APD24, SPQ^{+}24]. lucioperca [BKBR^{+}23]. lucius [FWJ21]. Lumpfish [BKJ^{+}24, CMP^{+}23, IR22]. lumpus [BKJ^{+}24, CMP^{+}23, IR22]. Lutjanus [CVANRD^{+}21, LGZ^{+}23, dCPFS^{+}24]. Lutraria [LXZ^{+}22]. luxR [ZHQ^{+}23]. Lymphocytes [RSA17]. Lymphoid [RSA17, SMO^{+}22].

M [Gae17]. M74 [VKP^{+}24]. macdonaldi [CBCMRD^{+}23]. Mackerel [BS23, CKMT23, CPJK23a, CPJK23b, FGHYCA23, MSA24, NNN23]. maclovinus [MPM^{+}18]. Maclura [dCPRG^{+}21]. Macroalgae [YZH^{+}24]. Macrobrachium [LSJ^{+}23, PAVCCJVV24, XRX^{+}23, ZCT^{+}23, ZYT^{+}24]. Macrocallista [PSW^{+}23]. Macronutrient [MMY^{+}17]. Macroparasite [EFG^{+}23]. Macrophage [LWS^{+}23a, ZAM^{+}23]. Macrophages [ZAM^{+}23]. macropomum [dAPAA^{+}24, SBS^{+}23]. Macroscopical [BWKS20]. Macrourus [XSZ^{+}23]. mactroides [GGP^{+}23]. maculata [ZWZ^{+}24]. maculatus [LZH^{+}23, WSZ^{+}23, YSG^{+}23]. Madagascar [And23]. Made [BEMC23]. maenas [YE20]. maeoticus [DDD^{+}23, ŢPN^{+}23]. magdalenae [MMMNAG23]. magister [LOT^{+}22]. Mahseer [INCD23]. Main [LZL^{+}23, MZA^{+}23]. Mainstream [JZX22]. Major [AIŠRB22, LWD^{+}23, TJW^{+}22, HIO^{+}19, KJK^{+}23]. Mako [VV23]. malabaricus [HAC^{+}23]. Male [CMP^{+}23, Fed23, FHF23, JSJ^{+}24, LYL^{+}24b, MBP^{+}24, PLY^{+}24, QAC23, SLY^{+}21, YMD^{+}21, FHHC23, JCR^{+}22, PYJ^{+}23, SZZ^{+}23, WZL^{+}24, YGD^{+}23]. Male-Only [MBP^{+}24]. male-sign [FHHC23]. Male-Specific [CMP^{+}23]. Malformations [BWKS20, TRM^{+}23]. Malmquist [ZLF23]. Manage [dSSBdS23]. Management [APD^{+}23a, APD^{+}23b, DOB^{+}17, DMB^{+}20, KAB^{+}20, LXW^{+}24, MTPK23, MBP^{+}24, PMFBI22, dCPFS^{+}24, QL22, SKT23, SKSL23, SS23, TOB^{+}23, XW24, YSY^{+}23]. Manager [TOB^{+}23]. Manchurian [PXH^{+}24]. Mandarin [LMLH22]. Mangrove [QWR^{+}23]. Manila [GWH21]. Manipulation [GRO^{+}17]. Mannan [MVPMAV^{+}22]. Mannan-Oligosaccharides [MVPMAV^{+}22]. Manta [LCW23a]. Manual [VPPF^{+}19]. Manure [RDG^{+}20]. MAPK [RWF^{+}23]. Mapping [CHJ^{+}23, ISA^{+}22]. Marbled [JCL^{+}24, WYG^{+}23]. Margate [BPO19]. marginatus [AGA^{+}21]. Marine [CSGE23, CWMX21, DPGFL^{+}22, FDM^{+}23a, HLK^{+}23, Ho22, JLT22, JHLW24, KGP24, LPF^{+}23, LKJ22, LJP^{+}22, LW18, MBC^{+}24, Mil23, MMAO22, MMD^{+}23, MFKS23, PKY^{+}23, RDANPA^{+}24, RGVG19, SKSL23]. marinus [OJC^{+}23]. marisalbi [MNP^{+}16]. Maritime [BBF22, CXL23]. Mark [Tri23, ZJC^{+}22b]. Marker [LMLH22]. Markers [ARH^{+}23, BAA^{+}19, SdSdOSSL23, WLZ^{+}22a, ZLL^{+}23]. Market [IFA^{+}23]. Marketplace [BFP^{+}23]. Marking [GZF^{+}22]. Markings [ZJC^{+}22a]. marmoratus [JCL^{+}24]. Marshall [LSJ24]. Marshes [GTC^{+}17]. Marteilia [LPF^{+}23]. Mask [PWH^{+}23]. masquinongy [Mee24]. Mass [JSLE23, Näs18, STZ^{+}23]. Mass-Scale [STZ^{+}23]. Materials [DBP^{+}20, DQC^{+}23, TJTV^{+}23]. Maternal [KAJ^{+}24]. Mating [QAC23]. Maturation [EMFZ^{+}18, FHF23, IM24, LPB^{+}24, PXH^{+}24, Web23, ZQLW23]. Maturing [MLK^{+}19, vKRNL^{+}19]. Maturity [CP24, EBRS23, SKK^{+}23]. max [EPKV17]. Maximum [CP24]. maximus [CWMX21, LWG^{+}23, LLW^{+}22, TMPP23, WBK^{+}23]. May [ZZW^{+}22]. Mayotte [SQ23]. Meagre [DDN19, GSHGE18, NTP^{+}21, RJMVC^{+}18, RJTVC^{+}19, TPC^{+}23, VMDV^{+}22, VPPF^{+}19]. Meal [ATEfA^{+}21, BCG^{+}23, CSR22, EPKV17, FLB^{+}21, HGC^{+}23, HIO^{+}19, HMP^{+}24, LWT^{+}24, NdNFK^{+}24b, NdNFK^{+}24a, NYS^{+}23, RRM^{+}20, SCSR22, SAB^{+}22, TMD^{+}19, ULR^{+}23, WGRM^{+}19, WLW^{+}23, WLW^{+}24, XYC^{+}22, XYC^{+}23, XWP^{+}24, ZYG^{+}23]. Meal-Induced [SCSR22]. Mean [CP24]. Means [Gae16, Gae17]. Measure [Liu24]. Measurement [NZVB20]. Measurements [BEMC23]. Measures [MDF^{+}23, MML22]. Measuring [MOW^{+}18]. Meat [BDŞ^{+}24, CXW^{+}23, LVB^{+}20]. Mechanism [JML^{+}24, LKU21, WWW^{+}24, ZSL^{+}23, ZJM^{+}23]. Mechanisms [ZZ24]. Mechanistic [LCC^{+}24]. Medaka [MU21]. Media [ŠT19]. Mediated [LDZ^{+}24, SFP17, XWR^{+}23]. Mediates [WLT23]. Medicago [CWP^{+}21]. Mediterranean [AA23, BFB^{+}23, DAF^{+}22, FPRAT^{+}23, GSK^{+}21, MMD^{+}23, PEP24, RKS^{+}24, SML^{+}23a, VV23, AGE^{+}18, BTSK24, DCL^{+}23a, EFG^{+}23, HMVRFD19, LUM18, MGMG24, MLSC^{+}23, RDI^{+}21, RDE^{+}23, TK24, TPA^{+}24, UGH^{+}24]. medits [KRAFO23]. Medium [YHZ^{+}23]. Medium-Sized [YHZ^{+}23]. Megalobrama [GWT^{+}24, MYW^{+}24]. megalops [USRDFO^{+}22]. Meglumine [MSK^{+}21]. Mekong [TVL21]. Melanocortin [KEA^{+}23]. Melanocortin-4 [KEA^{+}23]. melanopterus [PKV^{+}22]. melanostomus [Ang18, KAR^{+}23]. melas [SMH^{+}22a]. melastomus [BMMD22, DAF^{+}22]. meleagris [EEdCSOPJ^{+}23]. Meloxicam [MSK^{+}21]. Member [OMC^{+}19]. Members [WSZ^{+}23]. membranaceus [PYJ^{+}23, XCW^{+}23]. Memory [CTY^{+}21]. Mentha [dAdSCC^{+}23]. Mercenaria [GWH21]. Mercury [JJB^{+}24, MMMNAG23]. Meretrix [FCB^{+}21, JXW^{+}23]. Meristics [GP17]. Merits [MBP^{+}24]. Merluccius [AGC23]. Mesanophrys [ZZL^{+}23]. Mesh [CGY^{+}23, YSY^{+}23]. Mesocosms [NHR20]. Mesopelagic [BTSK24]. Mesophotic [HZS^{+}21]. Mesopredator [LOBTL22]. Metabarcoding [ATM^{+}24, AMK23, WFZ^{+}24b]. Metabolic [FYL^{+}23, HSZ^{+}22, Kim23a, Kim23b, LLS^{+}23, MMVVCJ^{+}23, MMM^{+}24, dAPAA^{+}24]. Metabolism [ASM^{+}22, FMBPC^{+}23, GMBR^{+}21, HYN^{+}24, HLZ^{+}22, LCWH22, RWF^{+}23, RJTVC^{+}19, SdSdOSSL23, SZF^{+}21, WSD^{+}23, WZX^{+}24, WLT23, YGD^{+}23, ZLZ^{+}24, ZYT^{+}24]. Metabolism-Related [ZYT^{+}24]. Metabolites [ITG^{+}18, MTM^{+}19, TPC^{+}23, YBL^{+}22]. Metabolome [TMM^{+}18]. Metabolomic [FYL^{+}23, HWX^{+}23]. Metabolomics [HWW^{+}24b, WGRM^{+}19]. Metabonomic [LSJ^{+}23]. Metal [WFZ^{+}23]. Metals [dRCCLRNRWK21, MS18, RKS^{+}24]. Metamorphosis [XWW^{+}24]. Metazoan [SVMLP23]. Metformin [SAGG^{+}23]. Methanesulfonate [CWMX21]. Method [AIŠRB22, ILA22, KLD^{+}23, LDZ^{+}24, LOBTL22, LLW^{+}22, SLS^{+}22, WMZ^{+}22]. Methods [BLA^{+}22, CG20, KSWT22, RMA^{+}18, TPA^{+}24]. methylisoborneol [LLKKV20]. Métier [LV23]. Metschnikowia [LYW^{+}24]. Mexico [APAHBMAG23, SWS22, USRDFO^{+}22]. mg [ZJC^{+}22b]. mg/L [ZJC^{+}22b]. Miamiensis [CCP^{+}24]. Michigan [JSSD23, PKSN23a, PKSN23b]. Micro [EBCM24]. Micro- [EBCM24]. Microalgae [GMDMT^{+}23]. Microalgae-Supplemented [GMDMT^{+}23]. Microalgal [DÁ23, PSW^{+}23]. Microbes [HWW^{+}24b, XRX^{+}23]. Microbial [HSZ^{+}22, HXS^{+}23, NWN^{+}22, RXL^{+}24, WGRM^{+}19, WZL^{+}24, YLH^{+}24]. Microbiological [HKS^{+}18, MAR^{+}18, RMA^{+}18, SUL^{+}23]. Microbiome [CBCL23]. Microbiomes [ATM^{+}24, HJE^{+}23]. Microbiota [BDŞ^{+}24, CSL24, DBP^{+}20, GMCF^{+}22, GMDMT^{+}23, LHZ^{+}23, LCW^{+}23b, RCR^{+}23, WLZ^{+}22b, XLX^{+}22, XYC^{+}22, XYC^{+}23, YLW^{+}23a, YBL^{+}22, YHZ^{+}23]. Microchemistry [FDM^{+}23a, FJL^{+}23, WTC^{+}22]. Microencapsulation [CBCL23]. Microflora [WLW^{+}23]. microphthalmum [SNSVFL23]. Microplastics [ADM^{+}24, TCB^{+}24]. Micropogonias [USRDFO^{+}22]. Micropterus [KKPY22, SOW^{+}23, SJY^{+}22, XYL^{+}23, XXL^{+}22, ZSH^{+}23, ZWP^{+}23, ZZY^{+}23]. Microsatellite [APD24, VMDV^{+}22, WLZ^{+}22a]. Microscopic [BMMD22]. Microscopy [CCCFE18]. Microsomal [KSSI24]. Microstructural [GWH^{+}22]. Microstructure [CLL^{+}22, DSC^{+}23, LOT^{+}22, YCR^{+}23]. Middle [CASMK23]. Midwestern [SB20]. Migrant [TTT23]. Migrating [WLL^{+}23]. Migration [LWS^{+}23a, PAMG19, WHY^{+}24].

Migrations [EKL^{+}23]. Migratory [EMFZ^{+}18, PWH^{+}23, WTC^{+}22, WLL^{+}23]. Military [LL22, MTPK23]. Mineral [XJC^{+}22]. Mini [GMRJ22]. Mini-Livestock [GMRJ22]. Mining [WHM21]. Minnows [LMC21]. miR [ZLZ^{+}22]. miR-200 [ZLZ^{+}22]. miricola [WWW^{+}24]. Misgurnus [ZSL^{+}23]. Mislabeling [SBS^{+}24]. Mississippi [ZS21]. Mitigate [NKC^{+}21]. Mitigating [TPA^{+}24]. Mitigation [MML22]. Mitochondrial [BBCJ23, CZL23, EAJG24, LPK^{+}23a, LPK^{+}23b, SSK^{+}22, WFL^{+}23]. Mitogenesis [SOW^{+}23]. Mitogenome [PKY^{+}23]. Mitogenomic [ZQL^{+}23]. Mitten [AML^{+}24, GLX^{+}23, LWZ^{+}22, LXT^{+}22, LLS^{+}23, PLY^{+}24, PNW^{+}22, WLZ^{+}22a, WHL^{+}24, WWZ^{+}24, XXL^{+}24, XJC^{+}22, YMD^{+}21, ZGY^{+}23]. Mixed [NNN23]. MMP [CL21]. MMP/TIMP [CL21]. MNPs [EBCM24]. Mobula [LCW23a]. Model [CKMT23, CTY^{+}21, FCT19, HARB23a, JTS^{+}24, VLCCA^{+}23, WSL^{+}23, XYT23, YHH^{+}20, ZLF23]. Modeling [AAN22, DOB^{+}17, FM21b, FBSB24, HWZ21, MMSK21, NRKT19, SWS22, WSL^{+}23, ZTWK24]. Modelling [CBCMRD^{+}23, WSA^{+}23]. Models [CMP^{+}20, MOW^{+}18, RDANPA^{+}24, SHT^{+}23]. Modifies [HLZ^{+}22]. Modify [BMSGS^{+}18]. Modulate [DBP^{+}20, WBK^{+}23]. Modulated [LS19]. Modulates [CSV^{+}19]. Modulation [PSO^{+}19, SNSG^{+}19, SMH^{+}22b, VPP^{+}22]. Molecular [CFM^{+}23, CASMK23, CSJ^{+}23, EES^{+}23, GP17, GWT^{+}24, LMLH22, LWS^{+}21, MLZ^{+}21, MMD^{+}23, NWN^{+}23, SHW^{+}23, VMBT24, WPLK23, XHC^{+}22, YSG^{+}23, ZHF^{+}22a, ZSZ^{+}23, ZLZL23, ZSL^{+}24, ZWZ^{+}24, ZZ24, ZZC^{+}22]. molitrix [MGS^{+}23, PVY^{+}21, RMA^{+}18, ZLZL23, ZJC^{+}22b]. Molly [MHZ^{+}23, ZMA^{+}24]. Molting [ZYH^{+}24]. Monahan [MWPS24]. Moniliformin [GTSG18]. Monitor [AMK23]. Monitored [JZX22]. Monitoring [BPCC23, DD23, HBL^{+}22, IKT19, KBB^{+}21, LSY^{+}17, LL18a, LL18b, NLTL23, TKF^{+}18]. Monogenea [GMMNRS18]. Monogeneans [CSJ^{+}23]. Montana [EHGS23]. Months [CHJ^{+}23]. moorei [DRFCL23]. morhua [PH23, SCSS23, YHH^{+}20, vKRNL^{+}19]. morio [SWS22]. Morone [HCWH20]. Morphine [CGSBGN24]. Morpho [CFM^{+}23]. Morpho-Molecular [CFM^{+}23]. Morphohistology [MMVVCJ^{+}23]. Morphological [CSJ^{+}23, KAB^{+}23, MJL^{+}24, MBPB24, XXL^{+}24, YYH^{+}24, Mil23]. Morphology [GP17, SDA23, XLX^{+}22, XYL^{+}23, YGD^{+}23, YHH^{+}20, ZZW^{+}22, ZXH^{+}23]. Morphometric [HHM^{+}24, KKB20, KYB^{+}23, MBD^{+}23, NZVB20, YJZ^{+}24]. Morphometrics [FB22, GP17]. Morphometry [FLB^{+}21]. Mortality [AVT18, BKJ^{+}24, BPOS19, BPO19, LLWW23, SCT^{+}24]. Moser [Mil23]. mossambicus [MMVVCJ^{+}23, SVMLP23, WLT23]. Motifs [CCP^{+}24]. Motivations [PB24]. Motives [SKL^{+}23]. Mounted [HL24]. Mouth [RMSPMC^{+}22]. Movement [BSB^{+}23, BBN^{+}24, DSC^{+}19, RGABD20, ZS21]. Movements [HMVRFD19, PEP24]. Mozambique [SVMLP23]. mrigala [PVY^{+}21]. mRNA [MLK^{+}19]. MS [BEMC23, CWMX21, FYL^{+}23, JSLE23]. MS-222 [BEMC23, CWMX21]. MS/MS [JSLE23]. MS4 [YBL^{+}22]. MSP [QWR^{+}23]. mTOR [LHX^{+}23, ZYG^{+}23]. Mucin [BMMD22]. Mucosal [CL19, CFCE20, Zac22]. Mucous [BMMD22]. Mucus [ERE21, GBT^{+}24, ITG^{+}18, LS19, RTBL^{+}18, TMD^{+}19, WML^{+}21]. Mud [RWF^{+}23, ZZL^{+}23]. Mugil [ATEfA^{+}21, EES^{+}23, GL23, GMDMT^{+}23, MABÁMSM22, MBAM19, MMM^{+}24].

Mugilids [SFP17]. Mukawa [AM24]. Mullerian [CMP^{+}23]. Mullet [ATEfA^{+}21, EES^{+}23]. Mullets [GMCF^{+}22]. Multi [CKMT23, CMB^{+}24, CGY^{+}23, DCD^{+}24, JSLE23, LUM18, LDZ^{+}24, MBP^{+}24, SMAR24, WMZ^{+}22, WWW^{+}24, XJC^{+}22]. Multi-Class [JSLE23]. Multi-Faceted [DCD^{+}24]. Multi-Indicator [MBP^{+}24]. Multi-Mesh [CGY^{+}23]. Multi-Mineral [XJC^{+}22]. Multi-Model [CKMT23]. Multi-Omics [WWW^{+}24]. Multi-Parameter [WMZ^{+}22]. Multi-Scale [SMAR24]. Multi-Trophic [CMB^{+}24]. Multi-Use [LUM18]. Multi-Visual [LDZ^{+}24]. Multidisciplinary [SCSR22]. Multilinear [KAB^{+}23]. Multiple [DSC^{+}23, KCCK23, LFH^{+}23, vSRB^{+}18]. Multiplex [KLD^{+}23, VMDV^{+}22]. Multispecies [RDG^{+}20]. muricatum [TCD^{+}21]. murphyi [BS23, FGHYCA23]. Murray [RKL24]. Musca [HIO^{+}19]. Muscle [BMSGS^{+}18, DZC^{+}22, FMBPC^{+}23, GRKC19, MTM^{+}19, MDVM^{+}23, WYG^{+}23, WHL^{+}24, ZSQ^{+}21]. Muskellunge [Mee24]. Mussel [CLL^{+}23b, MJL^{+}24, TPA^{+}24, WGRM^{+}19]. Mussels [LPF^{+}23]. mutica [SFP17]. Muxama [EA18]. Mycobacteriosis [HXY^{+}23]. Myeloid [SMO^{+}22]. mykiss [CBK^{+}21, CSV^{+}19, DSC^{+}23, HS18, LUM18, LLKKV20, PLC^{+}24, SMO^{+}22, SCSR22, VAT^{+}23, WFZ^{+}24a, WLL^{+}24, Web23, YZH^{+}24, vSRB^{+}18]. Myocarditis [CLJ^{+}23]. Mystery [KMS^{+}17]. Mystus [MAR^{+}18]. Mytella [ZZZ^{+}23]. Mytilus [LPF^{+}23, MYY^{+}23, WFL^{+}23].

n [CSJ^{+}23, BFM23]. NAChR [ZAM^{+}23]. Naming [RE21]. Nan [TTT23]. Nan-Fang-Ao [TTT23]. Nannochloropsis [SGG^{+}21]. Nano [EAJ^{+}23]. Nanocapsule [RIF^{+}23]. Nanoparticles [AEME^{+}23].

Nanoplastics [AZY^{+}24, CFP^{+}23, EBCM24]. Nardo [KMB18]. nardus [GMBR^{+}21]. Narrow [NNN23]. Narrow-Barred [NNN23]. Narrowing [HDW^{+}23, RJMVC^{+}18]. Naso [Tay19]. nasus [WHY^{+}24]. Natal [FWJ21, SBT22]. National [GASS^{+}22, MWPS24, MWPS23]. Native [APD^{+}23a, APD^{+}23b, AGE^{+}18, dRCCLRNRWK21, CFLK21, GTC^{+}17, HVRCG18, LMB^{+}23, NHR20, RDE^{+}23, SVMLP23, UH19]. Natural [BPOS19, BPO19, JROHVH^{+}23, LdRCV24, Mee24, MGS^{+}23, SCBSSMA24, ZLX^{+}23]. Naturally [MLK^{+}19, SEA^{+}23]. Nature [NPT^{+}24]. Nature-Based [NPT^{+}24]. Nauplii [SdSdOSSL23]. near [LSJ24]. Necrosis [LJK^{+}22]. Nectandra [SBS^{+}23]. Need [EBM24]. Needed [Sha19]. Negligible [TMM^{+}18]. Nematode [DCL^{+}23b, GMA^{+}24]. Nemipterus [AA23]. Neogobius [Ang18, KAR^{+}23]. Neolamprologus [JHNF24]. NEON [MWPS23, MWPS24]. Neophocaena [CYL^{+}23, CLL^{+}24, FLX^{+}22]. Neotropical [dSSBdS23]. Neotropics [PCO^{+}23]. Nephrotoxicity [AZY^{+}24]. Neptunea [ZZT^{+}23]. Net [CBP^{+}24, MFKS23]. Nets [CGY^{+}23, WYL23]. Netting [FBSB24]. Network [LHZ^{+}24, LYW^{+}24, MWPS23, MWPS24, WLZ^{+}23, WYL23]. Networks [AAN22, CTY^{+}21]. Neural [AAN22, CTY^{+}21, CMP^{+}20]. Neuro [MHZ^{+}23]. Neuro-Immune [MHZ^{+}23]. Neuroimmune [Zac22]. Neuronal [HAdM^{+}24]. Neutral [MLZ^{+}21]. Newly [FGR19]. News [SMM^{+}18]. Next [HZG^{+}21]. Next-Generation [HZG^{+}21]. Niche [LZL^{+}23, YYW^{+}23]. Nichols [FGHYCA23]. Nicobar [KSAB^{+}23]. Nile [AHK^{+}23, AMT^{+}24, BRB^{+}23, CLJ^{+}23, DCR^{+}23, dSGBdF23, dFBdSG^{+}19, GCFA^{+}22, JROHVH^{+}23, KYB^{+}23, LS19, LDD^{+}22, MTPK23, MBL20, NGMR23, NSK^{+}23, PSN18, RIF^{+}23, RHU^{+}23, SAC23, SNZ^{+}23, WCY^{+}24, dAdSCC^{+}23]. niloticus [PSN18, AEME^{+}23, AHK^{+}23, CLJ^{+}23, DCR^{+}23, dSGBdF23, GCFA^{+}22, HLZ^{+}22, KYB^{+}23, LS19, LDD^{+}22, LHX^{+}23, MMVVCJ^{+}23, MLH^{+}20, MBL20, SAC23, WCY^{+}24, ZZW^{+}22, dAdSCC^{+}23]. nimbosa [PSW^{+}23]. Ningaloo [TCD^{+}21]. Nitrate [PLC^{+}24]. Nitrite [LWS^{+}23b, SHW^{+}23]. Nitrogen [CSL24]. NLRC3 [YSG^{+}23]. NLRC3-like [YSG^{+}23]. NMR [WGRM^{+}19]. NMR-Based [WGRM^{+}19]. No [BEF^{+}23, Näs18, PKC^{+}19, SAL19]. Non [AGE^{+}18, EM23, GTC^{+}17, LMB^{+}23, LYL^{+}23, LOBTL22, LCX^{+}23, RXL^{+}24, ŠT19, UH19, ZHHO23]. Non-Conventional [ZHHO23]. Non-Destructive [LOBTL22]. Non-Dose-Dependent [LYL^{+}23]. Non-Indigenous [GTC^{+}17]. Non-Linear [EM23]. Non-Native [AGE^{+}18, LMB^{+}23, UH19]. Non-Specific [LCX^{+}23, RXL^{+}24]. Non-Thermal [ŠT19]. Nonspecific [LWZ^{+}22]. Nonstationary [HMX^{+}21]. Nonylphenol [LCC^{+}24]. Norepinephrine [LKU21]. North [HMX^{+}21, SMH^{+}22a, GP17, HS24, STZ^{+}23, WWD^{+}23]. Northeast [WLL^{+}23]. Northeastern [HKS^{+}18, AA23, SNSVFL23]. Northern [AGC23, BBN^{+}24, AAN22, LOBTL22, Mee24, YWL^{+}24, ZQLW23]. Northumberland [EBRS23]. Northwest [KMB18, CKMT23, FYH^{+}23, SHT^{+}23, SKF^{+}23]. Northwestern [HMX^{+}21]. Norwegian [AJF23, BKJ^{+}24, KAJ^{+}24]. Nose [PBS^{+}22]. Note [RE21]. Notes [KKB20, Näs18]. Notothenioid [MPM^{+}18]. novacula [CSSMV^{+}23]. Novel [CSF^{+}23, Dul23, NWN^{+}22, NWN^{+}23, OMC^{+}19, ZSL^{+}23]. Novo [ZSL^{+}23]. NP [EBCM24]. NPUST1 [LCWH22]. nrDNA [GWH21]. Nuclear [CZC^{+}22]. Nucleobase [OMC^{+}19]. Nucleobase-Ascorbate [OMC^{+}19]. Nucleotides [SCHT23]. nudiventris [HCZ^{+}23]. nudus [LCW^{+}23b]. Number [EAJG24]. Numeric [ZS21]. Nursery [FJL^{+}23, Gre17, NdNFK^{+}24a, NYS^{+}23, SLSC^{+}24]. Nusa [HWW^{+}24a]. Nutrient [GZX^{+}22, LCWH22, NSK^{+}23, YYH^{+}24]. Nutrients [DDG^{+}22, MdM^{+}23, YBL^{+}22]. Nutrition [BDŞ^{+}24, DZC^{+}22, HGC^{+}23, MLZ^{+}21, jMlL23, Pao23, PLV^{+}19, ZZC^{+}22]. Nutritional [HHP^{+}24, MABÁMSM22, MBAM19, Moy18, PNW^{+}22, RRM^{+}20, WHL^{+}24, YJZ^{+}24]. Nutritive [TJTV^{+}23]. NW [KKB20].

Object [LSJ^{+}22, LJX^{+}21, NLTL23].

Objective [SKL^{+}23]. Obligations [CXL23]. oblongata [YZH^{+}24]. obscurus [GZF^{+}22]. Observation [Ang18, MWPS23, MWPS24]. Observed [CBANCM^{+}21]. Occupies [YLW^{+}23a]. Occurrence [GAR^{+}24, SLYY23]. Occurring [THS^{+}22]. Ocean [DLL^{+}23, KWK^{+}24, KFS23, PFM^{+}20, ZSZSS^{+}22, ZXY^{+}24, FYH^{+}23, FGHYCA23, GPD^{+}23, KBK^{+}23, KSWT22, LL22, SML^{+}23a, SQ23, SZSW21, SHT^{+}23, SKF^{+}23, SMM^{+}18, WZG^{+}23, YYW^{+}23, YWL^{+}24]. Oceanic [KSAB^{+}23]. oceanica [SGG^{+}21]. Ocimum [VAT^{+}23]. Octopus [RGVG19, GRO^{+}17, RGVG19]. Odontesthes [GRKC19]. off [FGG^{+}22, HZL^{+}22, ST17, ZXY^{+}24]. officinalis [OGMG^{+}17]. Offs [vSRB^{+}18]. Offshore [JLT22, SGANM^{+}24, XZZ^{+}24]. Offspring [AJF^{+}22, LMMC^{+}22]. Oil [BCG^{+}23, GMBR^{+}21, GMFG^{+}24, MKN^{+}16, SBS^{+}23, WZX^{+}24, dAdSCC^{+}23]. Oil-Free [BCG^{+}23]. Oils [BSH^{+}23, JBdFS^{+}22]. Okamejei [YCR^{+}23]. oleivora [MJL^{+}24]. Oligomeric [WHX^{+}23]. Oligosaccharides [MVPMAV^{+}22]. Oligotrophic [SBT22]. olivaceus [CCP^{+}24, JKK^{+}20, YLW^{+}23b, YSY^{+}23]. Olive [CCP^{+}24, JKK^{+}20, YSY^{+}23]. Omics [EBCM24, WWW^{+}24]. Ommastrephes [XLC^{+}24]. Ompok [MSB^{+}23, SNK^{+}23]. Oncorhynchus [CBK^{+}21, CSV^{+}19, DSC^{+}23, HS18, LUM18, LLKKV20, PLC^{+}24, SMO^{+}22, SCSR22, VAT^{+}23, WLL^{+}23, WFZ^{+}24a, WLL^{+}24, Web23, YZH^{+}24, ZYG^{+}23, vSRB^{+}18]. One [CBP^{+}24, Sha19]. One-Health [Sha19]. Ongrowing [RGVG19]. Online [DD23]. Only [MBP^{+}24]. Onset [EBRS23]. Ontogeny [STZ^{+}23]. Oogenesis [RAR^{+}18, UDG^{+}19]. Oomycete [PSS^{+}18, XZD^{+}24]. Open [OOG24, PSP^{+}22, VCC^{+}18]. Open-Access [OOG24]. Open-Source [PSP^{+}22]. Operation [MTPK23]. Operations [PMFBI22, ZS21]. opercularis [KŽM^{+}23]. Ophidiidae [CSJ^{+}23]. Ophidiiformes [CSJ^{+}23]. opilio [BASBW24, FBSB24]. Opportunities [BKJ^{+}24, Ho22]. Opportunity [BFP^{+}23]. Opsariichthys [YXS^{+}24]. Optimization [EA18]. Optimized [LZC^{+}23a, VH20, ZS21]. Optimizing [HARB23b]. Options [TOB^{+}23]. Oral [GCFA^{+}22, HSD^{+}24]. oramin [YLH^{+}24]. orbignyanus [EMFZ^{+}18]. Orchestrates [SHW^{+}23]. Orchid [CZC23]. Order [MD21]. Ordines [KRAFO23]. Oregon [HHAG24, MD21]. Oreochromis [AEME^{+}23, AHK^{+}23, CLJ^{+}23, DCR^{+}23, EAJ^{+}23, dSGBdF23, GCFA^{+}22, HLZ^{+}22, KYB^{+}23, LS19, LDD^{+}22, LHX^{+}23, MSK^{+}21, MMVVCJ^{+}23, MLH^{+}20, MBL20, PSN18, SAC23, SVMLP23, WLT23, WCY^{+}24, ZZW^{+}22, ZHHO23, dAdSCC^{+}23]. Organic [CHZ^{+}24, DDD^{+}23, LZC^{+}23a, SML^{+}23b, TJTV^{+}23]. Organisms [CHZ^{+}24, EBCM24]. Organochlorine [HKS^{+}18]. Organs [YLX^{+}22]. Orientation [LLL^{+}23]. Origins [FWJ21]. Ornamental [LBH^{+}24]. Ortho [SBB^{+}19]. Ortho-Phosphate [SBB^{+}19]. orthoreovirus [TBPJ23]. Oryzias [MU21].

Osmoregulatory [RJMVC^{+}18]. Osteobrama [LBC^{+}24]. Osteogenic [TCV^{+}19]. Osteological [STZ^{+}23]. Other [AJF23, Gae16, Gae17, KZC^{+}24, Mil23, MMY^{+}17, RAR^{+}18, TOB^{+}23]. Otolith [FDM^{+}23a, FB22, FJL^{+}23, GP17, MBD^{+}23, MDLA22, MLSC^{+}23, PSP^{+}22, WTC^{+}22, ZXH^{+}23, ZJC^{+}22a]. Otoliths [Fer23, Gae16, Gae17, GWH^{+}22, GZF^{+}22]. oualaniensis [CLL^{+}22, HSZ^{+}22, YWL^{+}24]. Outbreak [AVT18, JSRE^{+}24]. Outbreaks [NHC^{+}23]. Output [TVL21]. Outputs [PWCL23]. Ovarian [FYL^{+}23, LWL^{+}23, MLK^{+}19, RWF^{+}23, RDE^{+}23, SMH^{+}24, Web23, WHY^{+}24, WCY^{+}24, WSS^{+}19, ZHX^{+}24]. Ovary [CBK^{+}21, FYL^{+}23, LWL^{+}23]. ovatus [GGL^{+}23, WLW^{+}23]. Overestimation [FM21a]. Overfishing [PWCL23]. Overlap [AGE^{+}18]. Overview [DMT^{+}19, HGC^{+}23, MDF^{+}23, MZA^{+}23, RCL^{+}23]. Overwinter [WSA^{+}23]. Overwintering [HBL^{+}22]. Oxidation [CXW^{+}23, DSC^{+}23, XCW^{+}23]. Oxidative [AZY^{+}24, BDŞ^{+}24, CSSMV^{+}23, FMXQ23, GMFG^{+}24, LGZ^{+}23, SSSP21, SPJ^{+}24, WWZ^{+}24, dAdSCC^{+}23]. Oxidative/Antioxidant [AZY^{+}24]. Oxygen [MPM^{+}21, RSA17]. oxygeneios [WSI^{+}19, WSS^{+}19]. oxyrinchus [VV23]. Oyster [dOCCH^{+}23, GLY^{+}23, IFA^{+}23, MSK^{+}22, SLSC^{+}24, WLC^{+}23]. Oysters [QWY^{+}24].

P [LL18a, JCR^{+}22, JK24, NZVB20].

P-Glycoprotein [JK24]. Pabda [SNK^{+}23]. Pacific [FYH^{+}23, FGHYCA23, KBK^{+}23, KSWT22, LL22, SS23, SHT^{+}23, SKF^{+}23, YYW^{+}23, BS23, CKMT23, CdOCH^{+}23, CSJ^{+}23, EAE^{+}23, HMX^{+}21, HSAF^{+}23, HMP^{+}24, KSO^{+}23, LZZ^{+}22, LSL^{+}24, SCHT23, SRHCO23, SHT^{+}23, SKF^{+}23, Tay19, TBPJ23, WGW^{+}23, WSON24, ZWD^{+}23]. Pacific-Wide [LZZ^{+}22]. Packaging [DQC^{+}23, SUL^{+}23]. Paenibacillus [LCWH22]. Pagellus [RJFCJC^{+}21]. Pagrus [BF16, HIO^{+}19, KJK^{+}23]. pagurus [EBRS23]. PAHs [SUL^{+}23]. Palabuhanratu [MMSK21]. Palaemon [HWX^{+}23]. Palizada [dRCCLRNRWK21]. pallasi [MNP^{+}16]. Palliative [RIF^{+}23]. Palm [GMFG^{+}24, RCR^{+}23]. Pampus [ZSZ^{+}23]. Panel [CHW21, VMDV^{+}22]. Pangasianodon [VH20]. Pangasius [VH20]. Papapouli [KKB20]. Para [SFP17]. Paradigm [JHNF24]. Paragalene [KMB18]. parahaemolyticus [FLR^{+}22, LJK^{+}22, LDZ^{+}24]. parahybae [HAdM^{+}24]. Paralarval [GRO^{+}17]. Paralichthyidae [LPK^{+}23b]. Paralichthys [CCP^{+}24, JKK^{+}20, YLW^{+}23b, YSY^{+}23]. Parallel [SML^{+}23a]. paramamosain [RWF^{+}23, ZZL^{+}23, ZYH^{+}24]. Parameter [HARB23b, WMZ^{+}22]. Parameters [AHK^{+}23, CZJ^{+}24, CSV^{+}19, CBANCM^{+}21, DRFCL23, DCR^{+}23, EHGS23, EAJ^{+}23, EPKV17, FMXQ23, GL23, GMDMT^{+}23, GBT^{+}24, HGSE23, JKK24, KAB^{+}23, MWZ^{+}23, MMM^{+}24, PKV^{+}22, TJTV^{+}23, TMPP23, VLCCA^{+}23, YJZ^{+}24, ZWZ22]. Paramisgurnus [CSL24]. Parasite [ENO21]. Parasites [SVMLP23, XZD^{+}24]. Parasitic [EFQ23, OSM23, Sha19]. pardalis [HVRCG18]. Parentage [LDBL19]. Parental [DPGFL^{+}22]. Parrotfish [LV23]. Part [HKS^{+}18, ISA^{+}22]. Partial [AAAF^{+}21, MCÁGHA^{+}17, NdNFK^{+}24b, NdNFK^{+}24a]. Participation [EN22]. Partitioning [BSB^{+}23]. Parts [AIŠRB22]. Parvalbumin [MBZ^{+}21]. parvum [BAS^{+}19, SAHS18]. Pass [PRPW23]. Passage [FLFF23, SRBCGV^{+}21]. Past [Ols19]. Patch [QWR^{+}23]. Path [SLL22]. Pathogen [LPF^{+}23]. Pathogenic [CSGE23, DO24, LYW^{+}24, PSS^{+}18]. Pathogenicity [DLL^{+}22, JJK21, QXAY22, WCX^{+}24, ZHQ^{+}23]. Pathogens [SCHT23, SNZ^{+}23, XZD^{+}24]. Pathology [JCY^{+}23]. Pathway [CGSBGN24, LHX^{+}23, RWF^{+}23, ZYG^{+}23, ZYH^{+}24]. Pathways [And23, RLB^{+}23]. Pattern [CLL^{+}24, Kim23a, Kim23b, SLYY23, ZLW^{+}23]. Patterned [SLY^{+}21]. Patterns [BBN^{+}24, CASMK23, DCD^{+}24, dSCFQB^{+}23, HMVRFD19, IFA^{+}23, MSA24, MWPS23, MWPS24, OSM23, dCPFS^{+}24, RGABD20, SFP17, SAGG^{+}23, WSZ^{+}23, ZWZ^{+}24, dSSBdS23]. PCR [CMP^{+}23, SPEGMC24, VMDV^{+}22]. Pearl [FMXQ23, GWH^{+}22, WLN^{+}23, SLYY23, WTC^{+}22]. Pejerrey [GRKC19]. Pelagic [KSWT22, PEP24, WGW^{+}23]. pelamis [NIN^{+}19]. Pelodiscus [CZC^{+}22]. Pelteobagrus [JLW^{+}24]. Pen [RJR^{+}22]. Penaeus [CBCL23, HXS^{+}23, HZG^{+}21, NdNFK^{+}24a, NYS^{+}23, RXL^{+}24, NdNFK^{+}24b]. Peninsula [FZZ^{+}24, LCZ^{+}23, APS^{+}23, RGTPSCCC24, SGAC20]. Pens [NWN^{+}22]. pentophthalmus [LPK^{+}23b]. Peppermint [dAdSCC^{+}23]. Peptide [ZAM^{+}23]. Peptides [CAC^{+}17]. Peracetic [HS18]. Perca [CWP^{+}21, NHR20]. Perception [JLT22]. Perceptions [OOG24]. Perch [CWP^{+}21, CCFP19, MTPK23, NGMR23, NHR20, Ols19, XZD^{+}24]. Percidae [BAA^{+}19]. Perciformes [LPK^{+}23a, PKY^{+}23, RMSPMC^{+}22, WSS^{+}19]. Percocypris [WLZ^{+}22b]. Performance [ATEfA^{+}21, ASM^{+}22, BCG^{+}23, CMC^{+}24, EAE^{+}23, EBM24, GZX^{+}22, HARB23b, HFEH^{+}23, JSLE23, LMMC^{+}22, LCX^{+}23, MWZ^{+}23, MMVVCJ^{+}23, PNW^{+}22, PAMG19, PJPMMV^{+}22, PRPW23, dAPAA^{+}24, RHU^{+}23, RGVG19, SPEGMC24, SBS^{+}23, SHT^{+}23, SBB^{+}19, SNK^{+}23, SMH^{+}22b, TPC^{+}23, ULR^{+}23, VH20, WML^{+}21, WLL^{+}22, WHL^{+}24, WFZ^{+}24a, WSE^{+}21, XCW^{+}23, XWP^{+}24, YLW^{+}23b, ZMLFS^{+}20, ZWZ22, dAdSCC^{+}23]. Performances [AAAF^{+}21, EAJ^{+}23, FLB^{+}21, PXH^{+}24, RDE^{+}23, SSSP21]. Perinereis [GLH^{+}23]. Period [MGMG24, SKK^{+}23, WMZ^{+}22]. perryi [YYH^{+}24, YYH^{+}24]. Persistent [DK22]. Perspective [Fed23, HWW^{+}24b, PWCL23, ZZ24]. Perspectives [LV23]. Perturbs [RE21]. Peru [ZXY^{+}24]. Peruvian [CPVMA^{+}24, HZL^{+}22]. Pesticide [HKS^{+}18, LMB^{+}23]. petimba [TK24]. phalaena [CHH^{+}23]. Pharmacokinetics [HSD^{+}24, MSK^{+}21]. Phase [NdNFK^{+}24b, NdNFK^{+}24a, KSSI24]. PHD [PLJ^{+}23]. Phenazines [NWN^{+}22]. Phenethylamine [LWG^{+}23]. Phenol [KSSI24]. Phenomena [SSSS23]. Phenotypes [FHF23]. Phenotypic [BAA^{+}19, EES^{+}23]. Phenylglucuronide [KSSI24]. Pheromone [LWG^{+}23, SRL^{+}19]. philippinarum [CSF^{+}23, GWH21, LSS^{+}22]. Philippines [MBD^{+}23]. Phoenix [HGSE23]. Phosphate [NSK^{+}23, SBB^{+}19]. Phosphate-Enriched [NSK^{+}23]. Phosphorescent [FBSB24]. Phosphorescent-Netting [FBSB24]. Phosphorus [JCR^{+}22]. Photobacterium [MMD^{+}23]. Photodynamic [DO24]. Photoperiod [CPJK23a, CPJK23b, LS19, MIHH23, WSD^{+}23]. Photoperiods [AGA^{+}21, SPJ^{+}24]. Photophore [DNP^{+}23]. Photoreceptor [KKPY22]. Photovoltaic [PNW^{+}22]. Phoxinus [HYN^{+}24]. Phragmites [WLL^{+}22]. Phylogenetic [CZL23, GLY^{+}23, LPK^{+}23a, LPK^{+}23b, PKY^{+}23, ZLW^{+}23]. Phylogenetical [PYP17]. Phylogeny [CASMK23]. Phylogeographic [VV23]. Phylogeography [WFL^{+}23]. Physical [WLW^{+}23]. Physical-Barrier-Related [WLW^{+}23]. Physicochemical [GLX^{+}23, MAR^{+}18, SUL^{+}23, WDL^{+}23]. Physiological [CPJK23a, CPJK23b, CXT^{+}24, DCL^{+}23a, JXW^{+}23, dAPAA^{+}24, RJFCJC^{+}21, ZSZSS^{+}22, vSRB^{+}18].

Physiology [BAS^{+}19, LSL^{+}24, jMlL23, PFM^{+}20, SDA23, WWD^{+}23, ZSH^{+}23]. Phytoplankton [LSS^{+}22]. Pieter [Gae17]. Pigmentation [GP17, IBN^{+}23, LYY^{+}24]. Pike [Ols19]. Pikeperch [BKBR^{+}23, CCFP19, Ols19]. pilchardus [BFB^{+}23, MKN^{+}16]. Pilot [BMSGS^{+}18, LXT^{+}22, RJMVC^{+}18, ZJC^{+}22b]. Pimelodidae [HAdM^{+}24]. Pimephales [LMC21]. pingi [WLZ^{+}22b]. pinnata [DRH18]. piperita [dAdSCC^{+}23]. Pisces [FGHYCA23, KMS^{+}17, LDBL19]. Piscidin [ZAM^{+}23]. Piscine [TBPJ23]. Piscivores [VF23]. Piscivorous [Gre17]. Pituitaries [vKRNL^{+}19]. Pituitary [CPJK23a, CPJK23b, LPB^{+}24, LWL^{+}23, SNSG^{+}19]. Pitx [AHJ^{+}23]. Place [RE21]. Plagioscion [dAMPS^{+}23]. Plaice [BAS^{+}19]. Plain [VLCCA^{+}23]. Plains [FLFF23]. Plant [BCG^{+}23, BSGMC^{+}22, LMC21, MABÁMSM22, MBAM19, XYL^{+}23]. Plant-Based [BCG^{+}23]. plantarum [GCFA^{+}22, SPQ^{+}24]. Plasma [CBK^{+}21, JKK24, PBS^{+}22, TPC^{+}23, TMM^{+}18, dAdSCC^{+}23]. Plasticity [vSRB^{+}18]. platessa [BAS^{+}19]. Platform [PSP^{+}22]. Platichthys [JKK24]. plecoglossicida [SZZ^{+}23, ZHQ^{+}23]. Pleopod [YYH^{+}24]. Pleuronectes [BAS^{+}19]. Pleuronectiformes [LPK^{+}23b]. Poecilia [BEF^{+}23, GBT^{+}24]. Poeciliidae [UDG^{+}19]. Poey [dCPFS^{+}24]. Points [BLA^{+}22]. Polar [KWK^{+}24, YHH^{+}20]. Polarization [ZAM^{+}23]. Policy [And23, FCT19, WHM21, XQLA18, XW24]. Pollutants [DDD^{+}23, Dul23, EM23]. Pollution [DCD^{+}24]. Poly [WSZ^{+}23]. Polyascus [YMD^{+}21]. Polychaete [GLH^{+}23]. Polyculture [GZX^{+}22, GSH^{+}24, HSAF^{+}23, LMNN21, NPT^{+}24, NSK^{+}23]. Polymerase [CMP^{+}23, KLD^{+}23]. Polymorphic [APD24]. Polyphenol [SFK^{+}23]. Polyphenol-Based [SFK^{+}23]. Polyphenols [CSL24]. Polyprion [PPAB^{+}18, PLV^{+}19, WSI^{+}19, WSS^{+}19]. Polysaccharide [MNO^{+}22]. Polysaccharides [LZH^{+}23]. Polystyrene [AZY^{+}24, CFP^{+}23, EBCM24]. Pomace [BDŞ^{+}24]. Pomacea [GMRJ22]. Pomacentridae [LDBL19]. Pomfret [BLA^{+}22, ZSZ^{+}23]. Pompano [GGL^{+}23, WLW^{+}23]. Pond [FHHC23, GZX^{+}22, GSH^{+}24]. Ponds [DCR^{+}23, GLX^{+}23, LdRCV24]. Ponics [SAB^{+}22]. Pontic [SCT^{+}24]. Pool [HZG^{+}21, PAMG19]. Poor [DOB^{+}17]. Poplar [RIF^{+}23]. Population [ARH^{+}23, AGC23, CTTW23, CH23, CLL^{+}22, FGHYCA23, GFDPSR22, HCWH20, HS24, HHL^{+}20, HCZ^{+}23, KBK^{+}23, KAR^{+}23, LCW23a, LDW^{+}21, LOT^{+}22, MDLA22, NTP^{+}21, WLZ^{+}22a, YE20, ZGY^{+}23]. Populations [AWCS23, APD^{+}23a, APD^{+}23b, HSZ^{+}22, SML^{+}23a, WWD^{+}23, ZAM^{+}23, ZZZ^{+}23]. porcus [SSSS23]. Porgy [MMY^{+}17, ZJM^{+}23]. Porphyra [LZH^{+}23]. Porpoise [CLL^{+}24, FLX^{+}22, LCZ^{+}23]. Porpoises [CYL^{+}23]. Port [PHB^{+}23]. Portugal [Fra23, BSR^{+}20, XQLA18]. Portuguese [RLAE23]. Portunidae [KKB20]. Portunus [LWS^{+}23b]. portusjacksoni [PHB^{+}23]. Pose [LJX^{+}21]. Poses [AA23]. Position [LPK^{+}23a, LLW^{+}22]. Positive [DDN19, FCF19]. Possession [MFKS23]. Possible [AGE^{+}18, Dul23, YAEAB23]. Post [BCG^{+}23, NYS^{+}23, Poi24, RDI^{+}21]. Post-Collection [RDI^{+}21]. Post-Larval [NYS^{+}23, Poi24]. Post-Smolt [BCG^{+}23]. Postbiotic [QPDGF^{+}23]. Poststocking [Mee24]. Pot [STZ^{+}23, ZQLW23]. Pot-Bellied [STZ^{+}23, ZQLW23]. Potamotrygoninae [GAR^{+}24]. Potassium [BF16, CHZ^{+}24, WSE^{+}21]. Potato [FQÁGTR^{+}17]. Potential [AA23, ATM^{+}24, AAN22, BSR^{+}20, CSF^{+}23, DCR^{+}23, DPGFL^{+}22, EAE^{+}23, EES^{+}23, EAJG24, EEdCSOPJ^{+}23, FJL^{+}23, GMDMT^{+}23, GTSG18, HDW^{+}23, JLT22, JLW^{+}24, LWG^{+}23, LCWH22, LW18, MHZ^{+}23, MMSK21, NKC^{+}21, NWN^{+}21, NWN^{+}22, RLAE23, SKF^{+}23, TCB^{+}24, UGH^{+}24, VSH23, YZL^{+}23]. Pots [FBSB24]. Powder [AEME^{+}23]. Power [PNW^{+}22]. Poyang [ZZW^{+}24]. Practical [CSR22, NYS^{+}23]. Practice [XW24]. Practices [LLTM17]. Prawn [LSJ^{+}23, NSK^{+}23, VRKV24]. Prawns [XRX^{+}23]. Praziquantel [IKT19]. Pre [HARB23b, SGG^{+}21]. Pre-Processed [SGG^{+}21]. Pre-Selection [HARB23b]. Precautionary [MBP^{+}24]. Precise [NLTL23]. Precision [LWZ^{+}23, Tri23]. Predator [AWCS23, AHLC19, DMB^{+}20, LLY^{+}24]. Predators [FÅG^{+}23]. Predatory [Ols19]. Predictability [Gre17]. Predicting [ZXY^{+}24, ZHF^{+}22b]. Prediction [CTY^{+}21, HARB23b, XLC^{+}24]. Predictors [FB22]. Prefabricated [FLFF23]. Preference [MU21, SCSS23, ST17]. Preferences [KKB20]. Preliminary [BPO19, FB22, GRO^{+}17, GTC^{+}17, GTSG18, IKT19, JSJ^{+}24, OGMG^{+}17, PEP24]. Premature [Ord19]. prenanti [ZHF^{+}22a]. Preparation [TDN^{+}22]. Preparations [HIO^{+}19]. Presence [LPF^{+}23, TCB^{+}24, WGW^{+}23]. Preservation [VM19, ZHF^{+}22b]. Preserved [WDL^{+}23]. Pressure [CZL23, HDW^{+}23, LMB^{+}23]. Prevalence [DMT^{+}19, FHF23]. Prevention [XLP23]. Previtellogenic [WSS^{+}19]. Prey [AHLC19, BTSK24, CCFP19, LLY^{+}24]. Prickly [WZX^{+}24]. Pristine [ADM^{+}24]. Proanthocyanidins [WHX^{+}23]. Probiotic [ATM^{+}24, CBCL23, HXS^{+}23, LCWH22, WBK^{+}23]. Probiotics [GCFA^{+}22, JE18, NKC^{+}21, SAC23, SMH^{+}22b, WFZ^{+}24a]. Problems [OOGAS23]. Procedures [ZWD^{+}23]. Process [FYL^{+}23, GGL^{+}23, WMD^{+}24]. Processed [SUL^{+}23, SGG^{+}21]. Processing [EA18, NWN^{+}21, NWN^{+}22, Sam24, ŠT19, ZWD^{+}23]. Prochilodus [MMMNAG23]. Prodigiosin [NWN^{+}21, NWN^{+}23]. Produced [VAT^{+}23]. Product [HB24, NDC^{+}23, NWN^{+}22, QPDGF^{+}23, SBB^{+}19]. Production [FHHC23, IFA^{+}23, LCX^{+}23, MdM^{+}23, NWN^{+}22, NWN^{+}23, PSW^{+}23, RDG^{+}20, WM22]. Productive [AAAF^{+}21]. Productivity [FM21a, HPJ^{+}23]. Products [AAAF^{+}21, Sam24, SBS^{+}24, ZWD^{+}23]. Proenkephalin [LMLH22]. Profile [ASM^{+}22, HZL^{+}22, PBS^{+}22, PSW^{+}23, RGVG19, ZSL^{+}24]. Profiles [DCR^{+}23, GLY^{+}23, LYL^{+}24b, PB24, SKL^{+}23, XJC^{+}22]. Profiling [ATM^{+}24, FYL^{+}23, GGF^{+}22, HZZ^{+}23, JCY^{+}23, MKC^{+}22, QPDGF^{+}23, SML^{+}23a, SMH^{+}24, ZMA^{+}24, ZLW^{+}23]. Profit [SKL^{+}23]. Profitability [AHK^{+}23]. Progeryonidae [KMB18]. Program [BPCC23, BGT^{+}20, KBB^{+}21, VM19]. Progress [HZZ^{+}23]. prolifera [YLW^{+}23b, ZWD^{+}23]. Proliferation [ZMA^{+}24]. Prolonged [dAPAA^{+}24]. promelas [LMC21]. Promise [PLV^{+}19]. Promising [CLJ^{+}23, CBCL23, LJR^{+}24]. Promote [BGT^{+}20, LWT^{+}24]. Promotes [BRB^{+}23]. Promoting [JBdFS^{+}22]. Prompts [Ord19]. Propagule [HDW^{+}23]. Properties [ATM^{+}24, BMSGS^{+}18, GLX^{+}23, INCD23, MAR^{+}18, MKN^{+}16, RMA^{+}18, YJZ^{+}24]. Propionate [SSSP21]. Proposal [CSJ^{+}23]. Proposed [BBCJ23, OOGAS23]. Prospective [BAA^{+}23]. Prospectives [HB24]. Prospects [LDW^{+}21, NPT^{+}24, SS23]. Prostaglandins [SRL^{+}19]. Protease [DDN19]. Proteases [MCÁGHA^{+}17, TDN^{+}22]. Protected [JLT22]. Protection [CL21, TTT23, XYF^{+}24, ZMA^{+}24]. Protective [AASQPU^{+}23, CCP^{+}24, NDC^{+}23]. Protects [HWW^{+}24b]. Protein [BSGMC^{+}22, CWP^{+}21, DSC^{+}23, Eny17, EPKV17, HLZ^{+}22, NdNFK^{+}24b, NdNFK^{+}24a, PAVCCJVV24, QWY^{+}24, SdSdOSSL23, SNK^{+}23, XYL^{+}23, XWP^{+}24, YHZ^{+}23, ZHF^{+}22a]. Protein/Carbohydrate [BSGMC^{+}22]. Proteins [CZW^{+}23, CGSBGN24, LS19]. Proteomic [YMD^{+}21]. Protogyny [MPK^{+}23]. Protosalanx [TJW^{+}22]. Prototype [EEE21]. Proven [LWT^{+}24]. Provide [DCL^{+}23b, WTC^{+}22]. Provides [CYL^{+}23, WLC^{+}23, ZSL^{+}23]. Province [MFKS23, XZZ^{+}24, XW24]. Provincial [CHW21]. Proximal [GMFG^{+}24]. Proximate [RGVG19]. PRV [TBPJ23]. PRV-1 [TBPJ23]. Prymnesium [BAS^{+}19, SAHS18]. przewalskii [ZZ24]. Przybyla [PKSN23a]. Przybyla-Kelly [PKSN23a]. PS [EBCM24]. Pseudecheneis [LHG^{+}23]. Pseudobagrus [PYJ^{+}23]. Pseudocaranx [MNO^{+}22]. Pseudochromis [CZC23]. Pseudogobionini [CZL23]. Pseudomonas [AEME^{+}23, SZZ^{+}23, ZHQ^{+}23]. Pseudoplatystoma [JSRE^{+}24, PCO^{+}23, PCH^{+}24]. Pseudopleuronectes [BFM23]. Pseudorhombus [LPK^{+}23b]. Pseudosciaena [CTY^{+}21]. Psychrobacter [WBK^{+}23]. psychrophilum [VČ24]. Pterioidea [ZQL^{+}23]. Pteriomorphia [ZQL^{+}23]. Pteroplatytrygon [PEP24, WGW^{+}23]. Pterygoplichthys [HVRCG18]. Puerto [QWR^{+}23]. PUFA [TRM^{+}23]. Puffer [SLY^{+}21]. Pufferfish [KMS^{+}17]. Puget [SOT^{+}23]. pulcher [JHNF24]. Pulse [PCH^{+}24]. punctata [KLD^{+}23]. punctatus [FHHC23, KBCM19, QXAY22, RGTPSCCC24]. Puntius [HHM^{+}18, HKS^{+}18]. Pupfish [AWCS23]. Purpleback [CLL^{+}22]. purpuratus [BÁD^{+}22, CPVMA^{+}24]. purpureus [dCPFS^{+}24]. Purse [XYT23, YYW^{+}23]. putida [AEME^{+}23]. Puzzle [RE21]. Pyloric [GMFG^{+}24]. PYY [ZSZ^{+}23].

qPCR [EAJG24]. QTL [CHJ^{+}23]. Quahog [GWH21]. Quality [BDŞ^{+}24, CWMX21, CMC^{+}24, DQC^{+}23, DCR^{+}23, DSC^{+}23, GL23, GZX^{+}22, HHP^{+}24, HXS^{+}23, HKS^{+}18, JHLW24, JROHVH^{+}23, KŽM^{+}23, LVB^{+}20, LCX^{+}23, LMNN21, MMMNAG23, MSB^{+}23, MIHH23, MGS^{+}23, PKV^{+}22, RAR^{+}18, RHU^{+}23, RXL^{+}24, SBB^{+}19, TKF^{+}18, ZSQ^{+}21]. Quantitative [DDG^{+}22, KTT24]. Queen [HS24, KŽM^{+}23]. quelen [dCPRG^{+}21]. quinqueradiata [IKT19, MNO^{+}22].

R [dSJC^{+}21]. R- [dSJC^{+}21]. R. [JTS^{+}24]. Rabbit [RDG^{+}20]. Rabbitfish [YLH^{+}24]. Rabbitfishes [LV23]. Races [SML^{+}23a]. Raceway [FHHC23]. Rachycentron [MKC^{+}22]. Rafinesque [DAF^{+}22, RDE^{+}23, VV23]. Raft [RXL^{+}24]. Rainbow [CBK^{+}21, CSV^{+}19, DSC^{+}23, FMBPC^{+}23, HS18, LUM18, LLKKV20, MDVM^{+}23, PLC^{+}24, SMO^{+}22, SCSR22, TKF^{+}18, VAT^{+}23, VF23, WFZ^{+}24a, WLW^{+}24, WLL^{+}24, Web23, YZH^{+}24, vSRB^{+}18]. Raised [BRB^{+}23, JSRE^{+}24]. Raja [FPRAT^{+}23, SSK^{+}22, SSSS23, TMPP23]. Ramirez [KRAFO23]. Ramirez-Amaro [KRAFO23]. Rana [XWW^{+}24]. Randall [AA23]. randalli [AA23]. Range [BFLC19, BBN^{+}24, HMVRFD19, RJMVC^{+}18]. Rapeseed [GMFG^{+}24]. Rapid [KLD^{+}23, LDZ^{+}24]. Rapidly [TMM^{+}18]. Rare [APD24, KMB18]. RAS [BHR^{+}23, BKBR^{+}23, LLKKV20]. Rastrelliger [JTS^{+}24]. Rate [CCFP19, JKP^{+}23, Kim23a, Kim23b, PRPW23, RMSPMC^{+}22, VH20, XWW^{+}24]. Rates [MTM^{+}19, SJY^{+}22, WYG^{+}23]. Rathbun [KKB20]. Rating [IM24]. Ratio [BSGMC^{+}22, EPKV17, SKK^{+}23]. Ration [IM24]. Rationale [FCF19]. Ratios [JCR^{+}22, YGD^{+}23]. Rats [GLW^{+}22]. Raw [DBP^{+}20, TJTV^{+}23]. Ray [AFTÁPA^{+}23, FPRAT^{+}23, PKY^{+}23, SNSVFL23, SSK^{+}22, SSSS23, TMPP23]. Ray-Finned [PKY^{+}23]. Rays [DOB^{+}17, LCW23a, SDA23]. Razor [LLY^{+}22]. Re [CFM^{+}23]. Re-Approaching [CFM^{+}23]. Reaches [LLWW23, SLDC23]. Reaction [CMP^{+}23, KLD^{+}23]. Reading [PSP^{+}22]. Ready [WDL^{+}23]. Ready-to-Eat [WDL^{+}23]. Real [WYL23]. Real-Time [WYL23]. Reared [AHK^{+}23, ALNVDG^{+}22, BF16, HMP^{+}24, NDC^{+}23, RCR^{+}23]. Rearing [BKBR^{+}23, GRO^{+}17, MSB^{+}23]. Rearrangements [ZQL^{+}23]. Reason [SMM^{+}18]. Recapture [Tri23]. Receivers [BFLC19]. Receives [EBH21]. Receiving [LMC21]. Receptor [KEA^{+}23, VAT^{+}23, WLT23, YLX^{+}22, ZSL^{+}24]. Receptors [CZC^{+}22, WSZ^{+}23]. Recirculating [SIZ^{+}22]. Recognition [GP17, HARB23a, LWZ^{+}23, SZT^{+}23, VPPF^{+}19]. Recombinant [CZW^{+}23, LTZ^{+}22, SLC^{+}22]. Recommendations [LXW^{+}24, dCPFS^{+}24]. Reconstruct [GSK^{+}21]. Reconstruction [LDW^{+}21]. Record [KRAFO23, dAMPS^{+}23, TK24]. Records [KRAFO23, SQ23, UGH^{+}24]. Recovery [RJFCJC^{+}21]. Recreational [FGBA^{+}23, LLTM17, PB24, TOB^{+}23]. Recruitment [SMO^{+}23]. Recycling [NWN^{+}23]. Red [EAJ^{+}23, HIO^{+}19, HSAF^{+}23, JJK21, KJK^{+}23, LCC^{+}24, LCX^{+}23, dCPFS^{+}24, SWS22, TK24, WZL^{+}24, YZH^{+}24, AWCS23, LSY^{+}17]. Redfin [MBD^{+}23]. Redox [EAE^{+}23, MMM^{+}24]. Reduce [ALNVDG^{+}22, KSO^{+}23]. Reduced [AHJ^{+}23, EvSCB23, SB20, ZZW^{+}22]. Reduces [dSGBdF23]. Reducing [MBC^{+}24]. Reduction [USRDFO^{+}22]. Reef [CZC23, CP24, LCW23a, QWR^{+}23, SMAR24, TCD^{+}21]. Reefs [HZS^{+}21, HLK^{+}23, LL22, SCBSSMA24, XW24]. Reeve [GGP^{+}23]. Refeeding [dAPAA^{+}24]. Reference [BLA^{+}22, LWL^{+}23]. Refinement [BKBR^{+}23]. Reflect [BBCJ23]. Reflection [RE21]. Reflections [CXL23]. Refrigerated [DQC^{+}23]. Refuge [ANA^{+}23]. Refuges [LOB^{+}23]. Regan [ZXH^{+}23]. Regarding [JLT22]. Regenerating [VCL20]. Regenerative [PKC^{+}19]. Regime [VH20, dlCRMB^{+}22]. Region [BAA^{+}23, FGR19, VV23, GAR^{+}24]. Regional [BBF22, GPD^{+}23, MDLA22]. Regions [IFA^{+}23, LKJ22]. regius [DDN19, GSHGE18, RJMVC^{+}18, RJTVC^{+}19, TPC^{+}23, VMDV^{+}22]. Regression [KAB^{+}23]. Regulated [OOG24]. Regulates [YBL^{+}22]. Regulating [RWF^{+}23]. Regulation [JK24, LPB^{+}24, LKJ22, LYW^{+}24, MLZ^{+}21, PLJ^{+}23, QWY^{+}24, WLC^{+}23, WLT23, XHC^{+}22, ZZC^{+}22]. Regulatory [EvSCB23, KJK^{+}23, XQLA18, XWR^{+}23]. Reintroduction [EMFZ^{+}18, ZHF^{+}22b]. Reinvasion [SFP17]. Related [CGSBGN24, CH23, CBK^{+}21, JLW^{+}24, MYY^{+}23, TU18, VF23, WLW^{+}23, WWD^{+}23, WLM^{+}20, ZJJ^{+}22, ZHF^{+}22b, ZYT^{+}24, vKRNL^{+}19, PSO^{+}19]. Relatedness [NTP^{+}21]. Relation [HB24]. Relationship [CP24, DBP^{+}20, HWZ21, KZC^{+}24, LCW^{+}23b, LYL^{+}23, LWD^{+}23, LFH^{+}23, LHL^{+}24, TRM^{+}23]. Relationships [CZL23, KKB20, SGAC20]. Relative [BEF^{+}23, EHGS23, Näs18]. Release [JCL^{+}24]. Released [LLY^{+}24]. Releasing [WSI^{+}19]. Relevant [AAB^{+}18]. Remediation [WCY^{+}24]. Remedies [XLP23]. Remote [BS23, CLZ^{+}24, KSAB^{+}23, ST17, ZCX^{+}22]. Removal [RJR^{+}22, ZS21]. Removing [HIO^{+}19]. Reovirus [ZSL^{+}24]. Repeat [GWH21]. Replace [NYS^{+}23]. Replaced [XWP^{+}24]. Replacement [AAAF^{+}21, CSR22, FLB^{+}21, NdNFK^{+}24b, NdNFK^{+}24a, WGRM^{+}19]. Replacing [ULR^{+}23, XYC^{+}22, XYC^{+}23]. Report [DLL^{+}22, JSRE^{+}24, UH19]. Reproduction [DLL^{+}23, GGP^{+}23, HZS^{+}21, LDX^{+}23, PLV^{+}19, vKRNL^{+}19]. Reproduction-Related [vKRNL^{+}19]. Reproductive [AFTÁPA^{+}23, CHH^{+}23, CPVMA^{+}24, DPGFL^{+}22, HHM^{+}18, HS24, KJ22, LMC21, LHG^{+}23, LMMC^{+}22, MGMG24, MH23, PPAB^{+}18, RLB^{+}23, SBP23, SMH^{+}22b, Tri23]. Republic [AHK^{+}23]. Requirements [Moy18, MMY^{+}17]. rerio [BGMM^{+}24, CAC^{+}17, HFEH^{+}23, LCWH22, MBPB24, Ord19, TU18]. Research [DDG^{+}22, LJX^{+}21, LW18, PB24, RE21, XLP23]. Reservoir [BFLC19, SBT22, WFZ^{+}23]. Resident [HCWH20]. Residual [CBANCM^{+}21, LXW^{+}24]. Residue [WZL^{+}24]. Residues [HKS^{+}18, JSLE23]. Resilience [BFP^{+}23, MSK^{+}22, OOGAS23]. Resistance [AEME^{+}23, EAE^{+}23, HLZ^{+}22, LDD^{+}22, SCHT23, SNK^{+}23, WBK^{+}23, XLX^{+}22]. Resistant [PVY^{+}21]. Resolution [BEMC23, MSA24]. Resolved [KBK^{+}23]. Resource [YSY^{+}23]. Resources [PNW^{+}22, dSSBdS23]. Respiratory [BAS^{+}19, SSK^{+}22]. Respirometry [BEMC23]. Response [BAP22, CPJK23a, CPJK23b, CSSMV^{+}23, EM23, EAE^{+}23, EES^{+}23, GSHGE18, GBT^{+}24, GGF^{+}22, HWX^{+}23, HLZ^{+}22, JXW^{+}23, KKPL23, KBCM19, KJK^{+}23, LWZ^{+}22, LWS^{+}21, LYL^{+}24a, LYW^{+}24, LCX^{+}23, PYJ^{+}23, QWY^{+}24, RIF^{+}23, SNSG^{+}19, SMO^{+}22, SEA^{+}23, SCCM23, SHW^{+}23, VPP^{+}22, WKB^{+}23, WSD^{+}23, WFZ^{+}24a, WBK^{+}23, YMD^{+}21, ZSZ^{+}23, ZLZL23, ZWZ^{+}24, ZHQ^{+}23, ZJM^{+}23].

Responses [AWCS23, ALNVDG^{+}22, CLL^{+}23b, dOCCH^{+}23, CXT^{+}24, GLX^{+}23, HMX^{+}21, JKK24, LHX^{+}23, LWS^{+}23b, MCSB^{+}19, MH23, RXL^{+}24, SAC23, SPQ^{+}24, ZAM^{+}23, dAdSCC^{+}23]. Responsive [CYL^{+}23]. Responsiveness [KWK^{+}24]. Restoration [KAB^{+}20]. Restored [SGAC20]. Restriction [LMMC^{+}22]. Results [GRO^{+}17, OGMG^{+}17, PWH^{+}23]. Rethinking [SYL^{+}24]. reticulata [BEF^{+}23, GBT^{+}24]. Retinal [KKPY22, SPJ^{+}24]. Reunion [SQ23]. Reveal [HBL^{+}22]. Revealed [JLW^{+}24, MDLA22, OJC^{+}23, SBS^{+}24, SPQ^{+}24, WGRM^{+}19, WWW^{+}24, YMD^{+}21]. Revealing [HZL^{+}22, LFH^{+}23]. Reveals [FYL^{+}23, LYW^{+}24, TJW^{+}22, WYG^{+}23, WLN^{+}23]. Revenue [ATEfA^{+}21]. Reversal [JSJ^{+}24, VMS^{+}23]. Review [AHLC19, BB23, BAA^{+}23, CG20, CLZ^{+}24, CS23, DMB^{+}20, DD23, EM23, EBCM24, EFQ23, GHS20, HMN^{+}22, HVRCG18, IR22, JE18, JK24, KFS23, LZC^{+}23a, LL18a, LL18b, LF24, MKH24, MS20, MML22, RLB^{+}23, RTBL^{+}18, SKSL23, SAC23, SML^{+}23b, TK24, VČ24, VRG^{+}24, WWD^{+}23, YCL^{+}23, ZSZSS^{+}22, ZHHO23]. Reviewers [Off18, Off19, Off20, Off21, Off23]. Reward [CGSBGN24]. Rewilding [HHP^{+}24]. Rhabdosargus [HHL^{+}20]. Rhabdovirus [JML^{+}24]. Rhabdoviruses [SML^{+}23b]. Rhamdia [dCPRG^{+}21]. Rhincodon [OSM23]. Rhinogobio [WCX^{+}24]. rhoifolium [PSS^{+}18]. Rhythm [WLN^{+}23]. Rhythms [OGMG^{+}17]. Ria [XQLA18]. Rich [SPQ^{+}24]. Rico [QWR^{+}23]. Riddle [SWUH23]. Rights [TTT23, XW24]. ringens [AAN22]. Rio [CASMK23]. Rise [BAA^{+}23]. Risk [ALNVDG^{+}22, BSR^{+}20, DMT^{+}19, DIL^{+}22, HDW^{+}23, NHC^{+}23, WFZ^{+}23, XLP23]. Risk-Taking [ALNVDG^{+}22]. Risks [Dul23, Ho22]. Risso [GMCF^{+}22]. River [AMK23, APS^{+}21, BAA^{+}23, BAA^{+}19, dRCCLRNRWK21, FWJ21, GWH^{+}22, GAR^{+}24, HCWH20, HMVRFD19, LLWW23, LHL^{+}24, LHG^{+}23, MS18, dAMPS^{+}23, SKD^{+}23, SLYY23, SCT^{+}24, WLL^{+}23, WSA^{+}23, WWY^{+}23, AM24, AWCS23, CHJ^{+}23, FJL^{+}23, FCT19, HCZ^{+}23, JZX22, LZC^{+}23b, RDE^{+}23, SLDC23, VRKV24, WTC^{+}22, WHY^{+}24, ZZW^{+}24, ZS21]. Riverine [RKL24]. Rivers [AML^{+}24, dSSBdS23, ZGY^{+}23]. rivoliana [AASQPU^{+}23]. rivulatus [LBH^{+}24]. RNA [AMK23, CCP^{+}24, ZYT^{+}24]. Roanoke [BAA^{+}19, HCWH20]. Robot [WMZ^{+}22]. Robotic [NLTL23]. Rockfish [JCL^{+}24, LZW^{+}24, SOT^{+}23, SZF^{+}21]. Rodlet [DCL^{+}23b]. rogercresseyi [PSO^{+}19]. rohita [RKHAMM22, SAGG^{+}23]. Rohu [RKHAMM22]. Role [ADM^{+}24, AASQPU^{+}23, CSF^{+}23, Fra23, LHZ^{+}23, LOB^{+}23, MHZ^{+}23, OO24, RIF^{+}23, VM19]. Roles [RTBL^{+}18]. Romania [SCT^{+}24]. Romanian [ŢPN^{+}23]. Room [SAL19]. rosenbergii [LSJ^{+}23, XRX^{+}23, ZCT^{+}23, ZYT^{+}24]. rostrata [WHX^{+}23, ZWZ22]. Rotated [LSJ^{+}22]. Rotating [LJX^{+}21]. Rotational [Ang18]. Round [AFTÁPA^{+}23, CZC23, KAR^{+}23, PKSN23a, PKSN23b, SNSVFL23]. Roundjaw [KBK^{+}23]. Routes [CFP^{+}23, LDD^{+}22]. RSIV [KJK^{+}23]. rubens [WKB^{+}23]. Ruditapes [CSF^{+}23, GWH21, LSS^{+}22, SML^{+}23a]. Ruff [RCR^{+}23]. Ruffe [NHR20]. Runoff [MH23]. Russell [AA23]. Russia [SKSL23]. Russian [AIŠRB22, DD23, RE21]. rustica [SPJ^{+}24]. ruthenus [Kim23a, Kim23b].

S [LL18a, dSJC^{+}21]. S- [dSJC^{+}21].

Saccular [FDM^{+}23a]. Safety

[HB24, Sam24]. sagax [SHT^{+}23, SKF^{+}23]. Sagmariasus [KKFC^{+}23]. Saharan [MML22]. saida [KWK^{+}24, YHH^{+}20]. Sailfin [HVRCG18]. Salaam [MBL20]. Salamander [CWW^{+}23, ZHF^{+}22b]. salar [CG20, IM24, SGG^{+}21, VKP^{+}24]. Saleh [HWW^{+}24a]. Saline [BF16]. Salinities [CHH^{+}23, RJMVC^{+}18]. Salinity [GLY^{+}23, LYL^{+}24a, LSL^{+}24, MPM^{+}21, MU21, QWY^{+}24, RJTVC^{+}19, ZSQ^{+}21]. Salmo [CG20, EAJG24, IM24, RDE^{+}23, SGG^{+}21, VKP^{+}24]. salmoides [KKPY22, SJY^{+}22, XYL^{+}23, XXL^{+}22, ZSH^{+}23, ZWP^{+}23, ZZY^{+}23]. Salmon [ANA^{+}23, AHLC19, BCG^{+}23, CG20, CS23, FJL^{+}23, FHF23, GLH^{+}23, IR22, IBN^{+}23, IM24, MIHH23, PSO^{+}19, TBPJ23, VKP^{+}24, WLL^{+}23, ZYG^{+}23, CG20, SGG^{+}21]. salmonicida [SMO^{+}22, XYF^{+}24]. Salmonid [MDF^{+}23, VSH23]. Salmonidae [KSSI24]. Salmonids [HJE^{+}23, LOB^{+}23, WSA^{+}23]. salmonis [CG20, IR22, IBN^{+}23]. Salt [GTC^{+}17, LF24]. Salts [CHZ^{+}24]. Salvelinus [WGRM^{+}19]. Sampling [AHLC19, CGY^{+}23, KBB^{+}21, PWH^{+}23, RRG22]. Sand [AAB^{+}18]. Sanda [KRAFO23]. Sander [BKBR^{+}23]. sapidus [KKB20]. saprophyticus [WFX^{+}23]. sarba [HHL^{+}20]. Sardina [BFB^{+}23, MKN^{+}16]. Sardine [BFB^{+}23, MKN^{+}16, SHT^{+}23, SKF^{+}23]. Sardinella [SZSW21]. Sardinops [SHT^{+}23, SKF^{+}23]. Satellite [CLZ^{+}24, LCW23a, ST17]. sativa [CWP^{+}21]. Saurida [GMA^{+}24]. saxatilis [HCWH20]. saylori [BMOH23]. SBM [ZLF23]. Scale [BFP^{+}23, BGT^{+}20, CBP^{+}24, HWW^{+}24a, KJ22, LSY^{+}17, MSA24, MFKS23, NWN^{+}22, PSP^{+}22, SKT23, SRHCO23, STZ^{+}23, SNZ^{+}23, SMAR24, ZJJ^{+}22]. Scale-Up [NWN^{+}22]. Scales [CCCFE18, XLC^{+}24]. Scallop [BÁD^{+}22, CPVMA^{+}24, HHM^{+}24, KŽM^{+}23]. scaphanocephalus [CSSMV^{+}23]. Scat [JSJ^{+}24, LWL^{+}23]. Scatophagus [JSJ^{+}24, LWL^{+}23]. Scatterplot [ST17]. Scavenger [ZSL^{+}24]. Scavenging [UH19]. sceleratus [CBP^{+}24]. Scenario [BAA^{+}23]. Scenarios [SMO^{+}23]. Schemes [NTP^{+}21]. Schizothorax [KZC^{+}24, ZHF^{+}22a, ZXH^{+}23]. Schlegel [FZZ^{+}24]. schlegelii [LZW^{+}24, SZF^{+}21, ZJM^{+}23]. Schliewen [KRAFO23]. schrenckii [JCR^{+}22]. Sciaenidae [CBCMRD^{+}23, dAMPS^{+}23, SLS^{+}22]. Science [Fer23, FM21b, GSK^{+}21]. Scientific [KKPL23, NNL^{+}23]. scleroptera [LLWW23]. Scomber [CKMT23, CPJK23a, CPJK23b, KMSO18]. Scomberomorus [NNN23]. Scophthalmus [CWMX21, DDD^{+}23, LWG^{+}23, LLW^{+}22, ŢPN^{+}23, TMPP23, WBK^{+}23]. Scoring [FM21a]. Scorpaena [SSSS23]. Scorpionfish [SSSS23]. Screening [DDD^{+}23, HDW^{+}23, KYB^{+}23]. scriba [LOBTL22]. Sculpin [CTTW23]. Scyliorhinus [BMSGS^{+}18]. Scylla [RWF^{+}23, ZZL^{+}23, ZYH^{+}24]. SE [AIŠRB22]. Sea [AIŠRB22, BAA^{+}23, BGMM^{+}24, CSB^{+}20, CdOCH^{+}23, CFCE20, CMC^{+}24, DAF^{+}22, GMFG^{+}24, GSK^{+}21, HHL^{+}20, JJK21, KMB18, KKB20, KJK^{+}23, LCZ^{+}23, LZH^{+}23, LMNN21, MBD^{+}23, PSO^{+}19, RH19, RSA17, SNSG^{+}19, SSK^{+}22, SAB^{+}22, TMD^{+}19, TK24, WKB^{+}23, WSZ^{+}23, WGW^{+}23, WLM^{+}20, XWR^{+}23, YCR^{+}23, YAS21, YSG^{+}23, ZLF23, AA23, BAA^{+}23, BFB^{+}23, BEG^{+}23, CLL^{+}22, DDD^{+}23, HB24, HWZ21, HSZ^{+}22, KAR^{+}23, LWD^{+}23, LSY^{+}17, LV23, LOBTL22, LOT^{+}22, MMD^{+}23, MNP^{+}16, NNL^{+}23, Ols19, Poi24, PEP24, RKS^{+}24, SML^{+}23a, SSSS23, ŢPN^{+}23, TK24, TMPP23, VV23, VKP^{+}24, XSZ^{+}23, YSY^{+}23, ZCX^{+}22]. Seabass [FDM^{+}23b, HXY^{+}23, NZVB20, RRM^{+}20, RGTPSCCC24, VPP^{+}22]. Seabream [BSGMC^{+}22, CCCFE18, CFCE20, ERE21, GMBR^{+}21, HIO^{+}19, MCSB^{+}19, PFM^{+}20, RJFCJC^{+}21, TJTV^{+}23]. Seafood [DDG^{+}22, EN22, SBS^{+}24, Sha19]. Seafood-Borne [Sha19]. Seafood-Exporting [EN22]. Seagrass [QWR^{+}23]. Seahorse [STZ^{+}23, ZQLW23]. Seas [LHL^{+}24, QL22]. Seascape [QWR^{+}23, SMAR24]. Season [SMH^{+}24, TRM^{+}23]. Seasonal [BBN^{+}24, CBK^{+}21, OGMG^{+}17, VPPF^{+}19, WSD^{+}23]. Seasons [ZZT^{+}23]. Seawater [MIHH23]. Seaweed [HSAF^{+}23]. Sebastes [LZW^{+}24, SZF^{+}21]. Sebastiscus [JCL^{+}24]. Second [KRAFO23]. Secretions [BMMD22]. Secretory [LS19]. Section [HCZ^{+}23, Hal23]. Sedative [dCPRG^{+}21]. Sedentary [HMVRFD19]. Sediment [GLX^{+}23]. Sediments [IKT19, LCW^{+}23b]. Seed [MSB^{+}23, WZX^{+}24]. Seeds [SLSC^{+}24]. Segmentation [LLW^{+}22]. Seine [XYT23, YYW^{+}23]. Selected [IFA^{+}23, KFS23, PKV^{+}22, ZLL^{+}23]. Selecting [TOB^{+}23]. Selection [BEF^{+}23, CSF^{+}23, CZL23, HARB23b, LWL^{+}23, MU21, ZHF^{+}22b]. Selectivity [LSS^{+}22, YSY^{+}23]. Selenium [AZY^{+}24, HWW^{+}24b, SPQ^{+}24]. Selenium-Rich [SPQ^{+}24]. Semantic [LLW^{+}22]. Semen [MMMNAG23, RDI^{+}21]. Semi [SXQ^{+}22]. Semi-Enclosed [SXQ^{+}22]. semilaevis [CL21, LZM^{+}23]. Seminal [Mil23]. Senegal [BC23]. Sensing [BS23, CLZ^{+}24, SGANM^{+}24, ST17, ZCX^{+}22]. Sensitivity [BBCJ23, DIL^{+}22, MKH24]. Sensitizes [LSY^{+}23]. Sensory [MAR^{+}18, PHB^{+}23, RMA^{+}18]. Sentinel [ADM^{+}24]. Sepia [OGMG^{+}17]. Septicemia [CCP^{+}24, QXAY22, YLW^{+}23a]. Sequences [JJK21, WFL^{+}23]. Sequencing [HZG^{+}21, MYY^{+}23, TJW^{+}22]. Sergeant [LDBL19]. Seriola [ASM^{+}22, AASQPU^{+}23, DBP^{+}20, IKT19, JFP^{+}18, MNO^{+}22, TRM^{+}23]. Seriolella [RCR^{+}23]. Serotypes [AMT^{+}24]. Serranus [LOBTL22, MSRGCG^{+}23]. Serum [LHX^{+}23, PXH^{+}24, XLX^{+}22, XYC^{+}22, XYC^{+}23, YLH^{+}24, ZWZ22]. Services [ISA^{+}22]. Sets [ZS21]. Setup [PLC^{+}24]. Seven [YJZ^{+}24]. Sex [BEF^{+}23, CMP^{+}23, CBK^{+}21, CPJK23a, CPJK23b, DM24, Fed23, JSJ^{+}24, LZM^{+}23, MYY^{+}23, MLK^{+}19, Näs18, PXH^{+}24, PKV^{+}22, SLY^{+}21, SKK^{+}23, SRL^{+}19, VMS^{+}23, ZWZ^{+}24, ZHF^{+}22b]. Sex-Differences [Näs18]. Sex-Related [CBK^{+}21, MYY^{+}23, ZHF^{+}22b]. Sex-Specific [BEF^{+}23]. Sexual [EBRS23, GBT^{+}24, PXH^{+}24, SKK^{+}23, VLCCA^{+}23]. Sexually [vKRNL^{+}19]. Shaanxi [WWY^{+}23]. Shad [SCT^{+}24]. Shading [DCR^{+}23]. Shallow [BSB^{+}23, HZS^{+}21, LOBTL22, SGAC20]. Shandong [FZZ^{+}24, LCZ^{+}23]. Shape [LFH^{+}23, MBD^{+}23, MDLA22, MLSC^{+}23]. Shaped [LLL^{+}23]. Shapes [QAC23]. Shaping [LHZ^{+}23]. Sharing [KKPL23]. Shark [HL24, KSWT22, PHB^{+}23, PKV^{+}22, RKS^{+}24, TCB^{+}24, UGH^{+}24]. Sharks [BSB^{+}23, DNP^{+}23, OSM23, SDA23, SMM^{+}18, WSON24, ZSZSS^{+}22]. Shelf [CTY^{+}21, DQC^{+}23, JROHVH^{+}23, dCPFS^{+}24, ZMLFS^{+}20]. Shelf-Life [CTY^{+}21, DQC^{+}23, JROHVH^{+}23, ZMLFS^{+}20]. Shell [CL21, CSV^{+}19, DLL^{+}22, HHM^{+}24]. Shelled [CZC^{+}22]. Shellfish [DD23, QXM^{+}24]. Shengsi [XW24]. shevtsovi [LOT^{+}22]. Shifts [AGE^{+}18, OSM23]. Ship [HCZ^{+}23]. Shock [ZHF^{+}22a]. Short [CTY^{+}21, MGS^{+}23, PH23, UH19, WKB^{+}23]. Short-Term [MGS^{+}23, UH19, WKB^{+}23]. Shortfin [VV23]. Shoshone [CTTW23]. Shrimp [AAAF^{+}21, APAHBMAG23, BHP^{+}24, CdOCH^{+}23, CBCL23, EAE^{+}23, HSAF^{+}23, HMP^{+}24, HSD^{+}24, JSLE23, LJK^{+}22, LXW^{+}24, LSL^{+}24, LMNN21, NWN^{+}21, PMFBI22, SCHT23, SEA^{+}23, TVL21, WPLK23, WDL^{+}23, ZLL^{+}23, ZWD^{+}23, ZTWK24]. Siavonga [SNZ^{+}23]. Siberian [AVT18]. Sichuan [LDX^{+}23]. Sicily [FGG^{+}22]. Side [RCL^{+}23]. sieboldii [LXZ^{+}22]. Sierra [OOG24]. Siganus [YLH^{+}24]. sign [FHHC23, PYJ^{+}23, SZZ^{+}23, WZL^{+}24, YGD^{+}23]. Signal [CLL^{+}24]. Signaling [RWF^{+}23, WPLK23, ZYG^{+}23, ZYH^{+}24]. Signalling [VPPF^{+}19]. Signature [MDLA22, VKP^{+}24]. Signatures [FYL^{+}23]. Significantly [NKC^{+}21, VH20]. Silage [MdM^{+}23]. Silico [SPEGMC24]. Silkrose [MNO^{+}22]. Silkworm [MNO^{+}22]. Silkworm-Derived [MNO^{+}22]. Silky [SMM^{+}18]. Siluriformes [CFM^{+}23, CASMK23, HAdM^{+}24]. Siluriforms [LHG^{+}23]. Silurus [BBN^{+}24, YZL^{+}23]. Silver [DRH18, HHL^{+}20, MGS^{+}23, dAMPS^{+}23, RMA^{+}18, SRL^{+}19, ZSZ^{+}23, ZLZL23]. Silvery [MMY^{+}17]. Silvery-Black [MMY^{+}17]. Simojoki [FJL^{+}23]. Simplified [MU21]. Simulated [CWMX21]. Simulating [SMO^{+}23, WSA^{+}23]. Simulation [HYXY23, ZS21]. sinensis [CZC^{+}22, CZJ^{+}24, GLX^{+}23, JCY^{+}23, LWZ^{+}22, LXT^{+}22, LLS^{+}23, LZC^{+}23b, PLY^{+}24, PNW^{+}22, WHL^{+}24, WWZ^{+}24, XXL^{+}24, XJC^{+}22, YMD^{+}21, ZLX^{+}23, ZGY^{+}23]. Singapore [LDZ^{+}24]. Single [GHS20, SSSP21, SLC^{+}22]. Single-Cell [GHS20]. Single-Chain [SLC^{+}22]. Siniperca [DZC^{+}22, JML^{+}24, LMLH22, XZD^{+}24]. Sink [AKM23, SB20]. Sinogastromyzon [GWH^{+}22]. Sinonovacula [LLY^{+}22]. Siphoning [MLH^{+}20]. Sipunculus [LCW^{+}23b]. Sisoridae [LHG^{+}23]. Site [BBCJ23, IKT19]. Sites [BEG^{+}23]. Sitosterol [TU18]. Situ [CYL^{+}23]. Six [BTSK24, HLK^{+}23, MWPS23, MWPS24, RDI^{+}21, SLS^{+}22]. Size [BHR^{+}23, BEF^{+}23, EBRS23, HMVRFD19, HCZ^{+}23, SCSS23, YSY^{+}23]. Sized [YHZ^{+}23]. Skeletal [MTM^{+}19, MDVM^{+}23]. Skin [CFCE20, CL21, ERE21, GBT^{+}24, HJE^{+}23, ITG^{+}18, LS19, LYY^{+}24, SBS^{+}23, TMD^{+}19, YHH^{+}20]. Skipjack [NIN^{+}19]. Slices [PBS^{+}22]. Slight [ZSQ^{+}21]. Slope [PRPW23]. Sludge [GLH^{+}23]. Smad3 [FCB^{+}21]. Small [BFP^{+}23, CBP^{+}24, DMA22, Gre17, HWW^{+}24a, ITG^{+}18, KAJ^{+}24, MFKS23, SKT23, SRHCO23, SNZ^{+}23, TCB^{+}24]. Small-Scale [BFP^{+}23, CBP^{+}24, HWW^{+}24a, MFKS23, SKT23, SRHCO23, SNZ^{+}23]. Smalleye [SNSVFL23]. Smallmouth [SOW^{+}23]. Smartphone [JTS^{+}24]. Smartphone-Captured [JTS^{+}24]. Smith [OSM23]. Smoking [SUL^{+}23]. Smolt [AHLC19, BCG^{+}23, MIHH23, VSH23]. Smoothers [ST17]. Snail [GMRJ22]. Snakehead [ZWZ^{+}24]. Snapper [BF16, HWW^{+}24a, LGZ^{+}23, dCPFS^{+}24]. Snappers [SBP23]. Snout [GWT^{+}24, MYW^{+}24, YXS^{+}24]. Snow [BASBW24, FBSB24, KZC^{+}24]. SNP [ARH^{+}23, HZG^{+}21]. SNPs [FCB^{+}21]. Social [AJF23, DCL^{+}23a, FCT19, dFBdSG^{+}19]. Social-Ecological [FCT19]. SOCS6 [WPLK23]. SOCS7 [WPLK23]. Sodium [SSSP21]. Soft [CZC^{+}22]. Soft-Shelled [CZC^{+}22]. solandri [GPD^{+}23]. solani [EAE^{+}23]. Soldier [CSR22, LWT^{+}24, NYS^{+}23, YAEAB23]. Sole [CZW^{+}23, LZM^{+}23]. Solea [SMO^{+}23]. Solenaia [MJL^{+}24]. Solenocera [WDL^{+}23]. Soles [SMO^{+}23]. Solid [FDM^{+}23b, MABÁMSM22]. Solid-State [FDM^{+}23b, MABÁMSM22]. Solids [CdOCH^{+}23]. Solubles [FDM^{+}23b]. Solution [NPT^{+}24]. Solutions [Dul23, OOGAS23]. Some [CHZ^{+}24, MKH24]. sophore [HHM^{+}18, HKS^{+}18]. sordidus [LDBL19]. Sorubim [JSRE^{+}24]. Sound [KCCK23, SOT^{+}23]. Soundscape [LL18a, LL18b]. Soundscapes [ZSH^{+}23]. Source [AKM23, Eny17, GSH^{+}24, PSP^{+}22, SB20, WFZ^{+}24b]. Source-Sink [SB20]. Sources [DCD^{+}24, RHU^{+}23, WLL^{+}24, YHZ^{+}23]. South [CTTW23, CASMK23, LWD^{+}23, MFKS23, YCR^{+}23, CLL^{+}22, FGHYCA23, HSZ^{+}22, KYB^{+}23, LLTM17, LV23, dAMPS^{+}23, WFZ^{+}24b, YJK23, ZCX^{+}22]. South-Central [CTTW23]. South-Eastern [CASMK23]. Southeast [BS23]. Southeastern [BPOS19, BPO19]. Southern [APD^{+}23a, APD^{+}23b, CFM^{+}23, Fed23, GFDPSR22, MSA24, MDLA22, NHC^{+}23, NIN^{+}19, PMFBI22, dCPFS^{+}24, ST17, ANA^{+}23, AFTÁPA^{+}23, BMTR23, DAF^{+}22, FGG^{+}22, HXY^{+}23, WLM^{+}20, WFZ^{+}23, ZVRH23]. Southwest [RGTPSCCC24]. Southwestern [SQ23]. Sox2 [LXZ^{+}22]. Sox9 [LXZ^{+}22]. Soy [BRB^{+}23]. Soybean [AAAF^{+}21, EPKV17, SCSR22, ULR^{+}23, WZL^{+}24, XYC^{+}22, XYC^{+}23, ZYG^{+}23]. sp [BMOH23, CSSMV^{+}23, EAJ^{+}23, HMP^{+}24, JSRE^{+}24, SGG^{+}21, ZZL^{+}23, ZHF^{+}22a]. Space [XYT23]. Space-Time [XYT23]. Spain [DMT^{+}19, JFP^{+}18]. Spanish [NNN23]. Sparid [MMY^{+}17]. Sparidentex [MMY^{+}17]. Sparus [BSGMC^{+}22, CSB^{+}20, CCCFE18, CFCE20, ERE21, GMBR^{+}21, GMDMT^{+}23, TJTV^{+}23, TMD^{+}19]. Sparusaurata [ZMLFS^{+}20]. Spatial [BEF^{+}23, DCL^{+}23a, DOB^{+}17, ENO21, JSSD23, LCW23a, LWD^{+}23, MWPS23, MMSK21, MLSC^{+}23, QAC23, RKL24, ZCX^{+}22, MWPS24]. Spatio [MSA24, SCBSSMA24, SKF^{+}23]. Spatio-Temporal [MSA24, SCBSSMA24, SKF^{+}23]. Spatiotemporal [XLC^{+}24, XYT23]. Spawning [BMTR23, CZC23, CHH^{+}23, EKL^{+}23, GFDPSR22, GGP^{+}23, HBG^{+}20, HAdM^{+}24, JFP^{+}18, PXH^{+}24, dCPFS^{+}24, SKK^{+}23, TRM^{+}23, WSI^{+}19]. Special [EFG^{+}23, Moy18, Sor21]. Species [APD24, AM24, AMK23, AIŠRB22, AGE^{+}18, AAB^{+}18, APS^{+}21, BTSK24, BBCJ23, dRCCLRNRWK21, CFLK21, DCL^{+}23a, DRFCL23, EBM24, FGBA^{+}23, GP17, HARB23a, HLK^{+}23, HVRCG18, JBGCG^{+}24, KZC^{+}24, KSSI24, KRAFO23, LWD^{+}23, LVB^{+}20, MD21, MMAO22, MBD^{+}23, MGMG24, Moy18, MMY^{+}17, PKC^{+}19, PLV^{+}19, Poi24, RKL24, RDANPA^{+}24, RKS^{+}24, SKSL23, SGAC20, SQ23, SKD^{+}23, SEA^{+}23, SOT^{+}23, SLS^{+}22, SLDC23, TCB^{+}24, TMPP23, VM19, VMBT24, VMDV^{+}22, VV23, VKP^{+}24, WZG^{+}23, WSON24, WWD^{+}23, WFZ^{+}24b, XZZ^{+}24, YJZ^{+}24, YJK23, ZZW^{+}24]. Specific [BEF^{+}23, CMP^{+}23, CXT^{+}24, LCX^{+}23, RXL^{+}24, SPJ^{+}24, WFZ^{+}23]. spectabilis [MTHPJS^{+}23]. Spectrometry [JSLE23]. Spectroscopy [KAB^{+}23]. Speed [ZLZ^{+}24]. speibonae [TDN^{+}22]. Sperm [PKV^{+}22, RDE^{+}23]. Spermatogonia [BGMM^{+}24]. Sphyraena [FB22]. Sphyraenidae [FB22]. Spillway [ZS21]. Spinach [LCX^{+}23]. Spinal [LKU21, dAMPS^{+}23]. Spindle [LLL^{+}23]. Spindle-Shaped [LLL^{+}23]. Spine [AHJ^{+}23]. Spine-Reduced [AHJ^{+}23]. Spined [AHJ^{+}23, BEMC23, DLL^{+}23, GMMNRS18, JBK^{+}23]. Spiny [ATM^{+}24]. Spleen [SHW^{+}23]. Splenic [JCY^{+}23]. Spore [PSS^{+}18]. Sporosarcina [YBL^{+}22]. Spot [XYT23, HL24]. Spotted [JSJ^{+}24, LZH^{+}23, LWL^{+}23, LMMC^{+}22, RGTPSCCC24, WSZ^{+}23, WWW^{+}24, YSG^{+}23]. spp [MSK^{+}21, SMO^{+}23, ŠT19, UGH^{+}24, ZHHO23]. Spread [Gre17]. Spring [BMTR23, ZCX^{+}22]. Squaliformes [ARH^{+}23]. Squalius [VMBT24]. Squalomorph [DNP^{+}23]. Squalus [ARH^{+}23]. squamosissimus [dAMPS^{+}23].

Squatina [UGH^{+}24]. Squid [CLL^{+}22, FYH^{+}23, HZL^{+}22, NWN^{+}22, SZT^{+}23, YWL^{+}24, ZXY^{+}24]. Squid-Jigging [FYH^{+}23]. Sri [JSLE23]. SSH [MABÁMSM22]. ssp [BWKS20]. SSR [ZLL^{+}23]. St [BMTR23]. Stable [DL23]. Stabling [GGF^{+}22]. Stage [CS23, JKP^{+}23, KMS^{+}17, Kim23a, Kim23b]. Stages [CZC^{+}22, CZJ^{+}24, CMP^{+}20, LWL^{+}23, MKH24, MPM^{+}21, VH20, WSS^{+}19, XRX^{+}23, ZCT^{+}23, ZHX^{+}24]. Stakeholders [OOG24]. Standard [CVANRD^{+}21]. Standardization [SHT^{+}23]. Standing [Tay19]. Staphylococcus [WFX^{+}23]. Star [WKB^{+}23]. Starch [FQÁGTR^{+}17]. Starry [JKK24]. State [FDM^{+}23b, MABÁMSM22]. States [CXL23, BPOS19, BPO19, HCWH20]. Stations [SRBCGV^{+}21]. Statistical [Liu24]. Statolith [LOT^{+}22]. Status [AEME^{+}23, ASM^{+}22, BDŞ^{+}24, EAE^{+}23, FPRAT^{+}23, GMFG^{+}24, HFEH^{+}23, LKD22, LWS^{+}23b, LSJ24, MMAO22, RIF^{+}23, RHU^{+}23, SBP23, SAL18]. Steindachner [DRFCL23]. Steindachneridion [HAdM^{+}24]. stellatus [JKK24]. Steroid [PXH^{+}24, ZWZ^{+}24]. Steroidogenic [MLK^{+}19]. Steroids [MLK^{+}19, PKV^{+}22]. Sterol [TU18]. Sthenoteuthis [CLL^{+}22, HSZ^{+}22, YWL^{+}24]. Stichaeidae [LPK^{+}23a]. Stickleback [BEMC23, DLL^{+}23, GMMNRS18, KAJ^{+}24, MH23, SAL18]. Sticklebacks [AHJ^{+}23, JBK^{+}23]. Stimulated [SMO^{+}22, SOW^{+}23]. Stimulation [dSGBdF23]. Stimulator [CWW^{+}23]. Stimuli [GGP^{+}23]. Stimulus [CXT^{+}24]. STING [CWW^{+}23]. Stingray [PEP24, RGABD20, WGW^{+}23]. Stingrays [GAR^{+}24]. Stochastic [APAHBMAG23]. Stock [AGC23, BLA^{+}22, BFB^{+}23, CKMT23, ENO21, FGG^{+}22, GFDPSR22, HHL^{+}20, SWS22, SLS^{+}22, TOB^{+}23, USRDFO^{+}22]. Stocked [MGS^{+}23]. Stocker [FHHC23]. Stocking [AHK^{+}23, BHR^{+}23, dSGBdF23, LSL^{+}24, LF24, MSB^{+}23, SIZ^{+}22, SLSC^{+}24, ZWP^{+}23, ZJM^{+}23]. Stocks [AIŠRB22, FM21a, NTP^{+}21, VKP^{+}24, ZXH^{+}23]. Stomach [OJC^{+}23]. Stomolophus [EEdCSOPJ^{+}23]. Storage [CTY^{+}21, MAR^{+}18, WDL^{+}23]. Story [TCD^{+}21]. Strain [AJF^{+}22, LJR^{+}24, WCX^{+}24]. Strains [HHM^{+}24, KYB^{+}23, ZLL^{+}23, ZZY^{+}23]. Strait [SZSW21]. Stranded [CCP^{+}24]. Strategies [AHLC19, GMMNRS18, HZL^{+}22, SKSL23]. Strategy [AHK^{+}23, NRKT19]. Stream [EBM24, LMC21, LOB^{+}23]. Streams [MD21, SBT22, UH19]. Strengths [CS23]. Streptococcus [AMT^{+}24, GGL^{+}23, GCFA^{+}22, GYH^{+}23, HLZ^{+}22, LCWH22, LDD^{+}22, LBH^{+}24, PSN18, RIF^{+}23]. Streptomyces [TDN^{+}22]. Stress [AASQPU^{+}23, BHR^{+}23, CSL24, CPJK23a, CPJK23b, CSSMV^{+}23, DCL^{+}23a, FMXQ23, FLR^{+}22, GLY^{+}23, GSHGE18, GGF^{+}22, HMP^{+}24, KJ22, LGZ^{+}23, LWG^{+}23, LCC^{+}24, LYL^{+}24a, LWS^{+}23b, LSL^{+}24, LZW^{+}24, LYW^{+}24, MWZ^{+}23, MDVM^{+}23, MCSB^{+}19, Ord19, PLY^{+}24, QWY^{+}24, SGANM^{+}24, SPJ^{+}24, SHW^{+}23, WMZ^{+}22, WWZ^{+}24, ZLZL23, ZYT^{+}24, dAdSCC^{+}23]. Stress-Protective [AASQPU^{+}23]. Stressed [LMMC^{+}22]. Stressors [HYN^{+}24, WKB^{+}23, vSRB^{+}18]. strigata [ZZZ^{+}23]. Striped [AHLC19, HCWH20, HBL^{+}22, SIZ^{+}22]. Stripping [MLH^{+}20]. Strobe [KBCM19]. Stroma [SMH^{+}24]. Strontium [GZF^{+}22, ZJC^{+}22a]. Structural [EEdCSOPJ^{+}23]. Structure [ARH^{+}23, ANA^{+}23, ATM^{+}24, CTTW23, CVANRD^{+}21, CLL^{+}22, CBANCM^{+}21, HSZ^{+}22, HCZ^{+}23, KMS^{+}17, LWZ^{+}22, LOT^{+}22, MDLA22, NTP^{+}21, RJR^{+}22, WLL^{+}22, WFL^{+}23, YHZ^{+}23, ZLL^{+}23, ZZW^{+}24, ZGY^{+}23]. Structures [ALNVDG^{+}22, UDG^{+}19, ZJC^{+}22b]. Studied [BAP22]. Studies [CLJ^{+}23, DD23, EBCM24, LWL^{+}23, PWH^{+}23, SOW^{+}23, SB20]. Study [AEME^{+}23, BMSGS^{+}18, BEG^{+}23, CZCW23, CMC^{+}24, FM21a, FCF19, GAR^{+}24, GLW^{+}22, Ho22, HAC^{+}23, Kim23a, Kim23b, KAB^{+}20, KSSI24, LLL^{+}23, LXT^{+}22, MSRGCG^{+}23, MDF^{+}23, MFKS23, NNL^{+}23, PAMG19, RJMVC^{+}18, RJFCJC^{+}21, SHT^{+}23, TMD^{+}19, TTT23, XLC^{+}24, XW24, XYT23, ZMLFS^{+}20, ZHF^{+}22b, ZYH^{+}24, ZJC^{+}22a, ZJC^{+}22b]. Studying [FCT19]. Sturgeon [AVT18, CZJ^{+}24, HCZ^{+}23, JCR^{+}22, Kim23a, Kim23b, LVB^{+}20, WFX^{+}23, ZLX^{+}23]. Sub [LMB^{+}23, MPM^{+}18, MML22]. Sub-Antarctic [MPM^{+}18]. Sub-Lethal [LMB^{+}23]. Sub-Saharan [MML22]. Subarctic [KAJ^{+}24, AHJ^{+}23]. Subchronic [GLW^{+}22]. Suboptimal [BRB^{+}23]. Subsidies [CHW21]. Substance [LYL^{+}23]. Substitution [EPKV17, WLW^{+}23, ZMLFS^{+}20]. Substitutions [ATEfA^{+}21]. Substrate [GZX^{+}22, GSH^{+}24, PRPW23]. subterranea [EPKV17]. subtilis [CBCL23, GCFA^{+}22, WFZ^{+}24a, XYC^{+}22, XYC^{+}23]. Subtropical [BS23, GL23]. Subunit [LTZ^{+}22]. Success [HMN^{+}22, Tri23]. Successful [EMFZ^{+}18, LDX^{+}23]. Sucker [RHUJ24]. Suffering [YLW^{+}23a]. Suggests [SML^{+}23a]. Suitability [FYH^{+}23, YWL^{+}24]. Suits [EEE21]. Sulawesi [MFKS23]. sulcatus [LHG^{+}23]. Sulu [MBD^{+}23]. Summer [VF23, ZCX^{+}22]. Sun [MAR^{+}18]. Sun-Dried [MAR^{+}18]. Sunray [PSW^{+}23]. Super [VAT^{+}23]. Supplemental [DRH18]. Supplementation [AASQPU^{+}23, BRB^{+}23, LHX^{+}23, QXM^{+}24, TPC^{+}23, WFZ^{+}24a, WLW^{+}24, XYL^{+}23, ZWZ22, ZYG^{+}23].

Supplemented [GMDMT^{+}23, GYH^{+}23, JKP^{+}23, MMM^{+}24, RGVG19]. Support [LSJ24]. Supposed [VČ24]. Suppressors [WPLK23]. Surf [RGTPSCCC24]. Surface [WLM^{+}20]. surinamensis [BPO19]. Surrounding [LCW^{+}23b]. Surveillance [RGABD20]. Survey [BPCC23, FGG^{+}22, TMPP23]. Survival [AGA^{+}21, BF16, CHH^{+}23, HXS^{+}23, JKP^{+}23, LWS^{+}23b, LMNN21, MPM^{+}21, Mee24, NRÁGPM^{+}18, PSS^{+}18, PYJ^{+}23, RJFCJC^{+}21, VF23, VH20, WSA^{+}23, ZZW^{+}22]. Susceptibility [CSB^{+}20, FM21a]. Suspended [ALNVDG^{+}22, CdOCH^{+}23, KSO^{+}23]. Sustainability [AJF23, BFP^{+}23, MGMG24, OO24]. Sustainable [LWT^{+}24, Liu24, PMFBI22, PWCL23, QL22]. Sustainably [SB20]. Sweden [AVT18]. Swept [RGABD20]. SWHAEZ [KYB^{+}23]. Swimbladder [DCL^{+}23b]. Swimming [BSH^{+}23, DK22, EBM24, LMB^{+}23, LWS^{+}23b, RDE^{+}23, VF23, ZLZ^{+}24]. Switzerland [BWKS20]. Swordfish [LLTM17, Poi24]. Swordtail [Fed23]. Swordtails [QAC23]. Symbiotic [EAE^{+}23]. Sympatric [HJE^{+}23]. Synbiotic [SSSP21]. Synbiotics [SFK^{+}23]. Synchrony [LZZ^{+}22]. Syndrome [LSJ^{+}23]. Synodontidae [GMA^{+}24]. Synthesis [PCO^{+}23, WLC^{+}23]. Synthetase [ZZC^{+}22]. Synthetic [WSI^{+}19]. System [ADM^{+}24, CdOCH^{+}23, dRCCLRNRWK21, CMC^{+}24, FHHC23, GRKC19, GZX^{+}22, GSH^{+}24, HCWH20, HAdM^{+}24, HPJ^{+}23, KEA^{+}23, LCX^{+}23, LMNN21, MSB^{+}23, MWZ^{+}23, MU21, MZA^{+}23, NLTL23, NRÁGPM^{+}18, NSK^{+}23, RXL^{+}24, SLC^{+}22, SIZ^{+}22, SLSC^{+}24, ZAM^{+}23]. Systematic [EFQ23, ZSZSS^{+}22, ZHHO23]. Systematics [CFM^{+}23, LBC^{+}24]. Systems [CMB^{+}24, JE18, JFP^{+}18, KTT24, LZC^{+}23a, MIHH23].

T [LPB^{+}24, RSA17]. Tachysurus [HWW^{+}24b, PYJ^{+}23]. Tactics [QAC23]. Tactile [dSGBdF23]. Tag [HBL^{+}22]. Taibai [WWY^{+}23]. Tail [LLL^{+}23]. taimen [LDX^{+}23]. Taiwan [Ho22, HHL^{+}20, HAC^{+}23, TTT23, YLL22]. Taiwanese [CLL23a]. Takalar [MFKS23]. Takifugu [GZF^{+}22]. Taking [ALNVDG^{+}22]. Tambaqui [LdRCV24, MdM^{+}23]. tambroides [INCD23]. Tandem [JSLE23]. Tanks [ALNVDG^{+}22, EEE21]. Tanzania [MBL20]. tarda [WSZ^{+}23]. Targeted [ITG^{+}18, MGMG24]. Targeting [ZLZ^{+}22]. Tasmanian [YWDP21]. Taste [LXT^{+}22]. Taxonomic [PWH^{+}23, SLDC23]. Taxonomy [CASMK23, Gae16, Gae17, LBC^{+}24]. Tea [CSL24]. Technical [BASBW24, TVL21]. Techniques [CLZ^{+}24, FBSB24]. Technological [BÁD^{+}22, LKJ22]. Technology [GL23, JHLW24, Pao23, RHU^{+}23, WFZ^{+}24b, YCL^{+}23]. Tegillarca [PLJ^{+}23]. Tegula [SPJ^{+}24]. Tejo [Fra23]. Telemetry [LCW23a]. Telencephalon [BEF^{+}23]. Teleost [MBZ^{+}21, PKC^{+}19, ZSZSS^{+}22]. Teleostei [CZL23, CBCMRD^{+}23, FŠS^{+}23, GMA^{+}24, KLD^{+}23, MBD^{+}23, SKK^{+}23]. Teleosts [GTSG18, Mil23, UDG^{+}19]. Temminck [FZZ^{+}24]. Temperate [HMX^{+}21, SMH^{+}22a]. Temperature [BRB^{+}23, DK22, FMXQ23, HZS^{+}21, HMP^{+}24, HYN^{+}24, HWW^{+}24b, HLZ^{+}22, MAR^{+}18, MPM^{+}21, MDVM^{+}23, PLY^{+}24, PLC^{+}24, PHB^{+}23, SML^{+}23a, SNSG^{+}19, SMO^{+}23, WSA^{+}23, WLM^{+}20, WLZ^{+}22b]. Temperatures [CHH^{+}23, RHUJ24, WML^{+}21, ZZW^{+}22]. Temporal [BMTR23, ENO21, MSA24, SCBSSMA24, SKF^{+}23, SLYY23, ZCX^{+}22]. Temporally [LS19]. Ten [LKJ22]. Tench [CSR22]. tenellum [PAVCCJVV24]. Tenggara [HWW^{+}24a]. Term [BPCC23, CTY^{+}21, DLL^{+}23, KJ22, KBB^{+}21, MGS^{+}23, Näs18, PH23, SXQ^{+}22, UH19, WKB^{+}23, WLL^{+}24]. Territorial [XW24]. Terror [LBH^{+}24]. Test [CHW21, JHNF24]. Testing [Dul23]. Tests [SPEGMC24]. Tetrabromobisphenol [vKRNL^{+}19]. Tetraodontidae [KMS^{+}17]. Tetraselmis [SGG^{+}21]. Texas [EBM24]. Text [WHM21]. Textural [YJZ^{+}24]. Texture [BMSGS^{+}18, KMSO18, KKFC^{+}23]. Thai [DRH18]. Thailand [SBS^{+}24]. Thaw [DSC^{+}23]. thazard [ZCX^{+}22]. Their [APD24, ATM^{+}24, BMMD22, Fer23, FBSB24, FLR^{+}22, KJK^{+}23, LWD^{+}23, LHL^{+}24, MMVVCJ^{+}23, MFKS23, PYP17, SDA23, TDN^{+}22, WSZ^{+}23]. Therapy [DO24]. Thermaikos [KMB18]. Thermal [EAJG24, GGF^{+}22, HMP^{+}24, LOB^{+}23, RHUJ24, SCSS23, SSSS23, ŠT19, dlCRMB^{+}22]. Thermally [CFLK21]. Thiamine [VKP^{+}24]. Thick [GMCF^{+}22]. Thick-Lipped [GMCF^{+}22]. thori [GP17]. Thornback [FPRAT^{+}23, SSK^{+}22, SSSS23, TMPP23]. Threadfin [AA23]. Threat [AA23, EES^{+}23, GTSG18, HVRCG18]. Threatened [SKSL23, SNSVFL23]. Three [AML^{+}24, BEMC23, BEG^{+}23, DLL^{+}23, GMMNRS18, JBK^{+}23, KSSI24, MBD^{+}23, RHUJ24, SHT^{+}23, SCCM23, WLL^{+}24, YLH^{+}24, YHZ^{+}23, ZWD^{+}23, ZTWK24]. Three-Dimensional [ZTWK24]. Three-Spined [BEMC23, DLL^{+}23, GMMNRS18, JBK^{+}23]. Threespine [KAJ^{+}24, SAL18]. Threshold [PWCL23]. Throughout [SNSG^{+}19]. Thun [BWKS20]. Thunnus [LYL^{+}24b]. Thymallus [ZLZ^{+}24]. Tibet [LHG^{+}23]. Tigris [MS18]. Tilapia [AHK^{+}23, AMT^{+}24, BRB^{+}23, CLJ^{+}23, EAJ^{+}23, EvSCB23, dSGBdF23, dFBdSG^{+}19, GCFA^{+}22, GYH^{+}23, HLZ^{+}22, JROHVH^{+}23, KYB^{+}23, LS19, LDD^{+}22, LHX^{+}23, LCX^{+}23, MSK^{+}21, MMVVCJ^{+}23, MLH^{+}20, MBL20, NGMR23, NHC^{+}23, NSK^{+}23, PSN18, PVY^{+}21, RIF^{+}23, RHU^{+}23, EEE21, SAC23, SNZ^{+}23, WLT23, WCY^{+}24, ZZW^{+}22, ZHHO23, dAdSCC^{+}23, DCR^{+}23, SVMLP23]. TiLV [PVY^{+}21]. Time [BMTR23, HHP^{+}24, HYXY23, MIHH23, SGAC20, WYL23, XYT23]. Times [HSD^{+}24]. TIMP [CL21]. Tinca [CSR22]. tinctoria [dCPRG^{+}21]. Tissue [AEME^{+}23, BB23, GRKC19, LWZ^{+}22, MSK^{+}21, Näs18, PNW^{+}22, dAPAA^{+}24, WLL^{+}22, WWZ^{+}24, ZZL^{+}23, ZZY^{+}23]. Tissues [LZM^{+}23, PAVCCJVV24, RSA17, RKS^{+}24, ZZT^{+}23]. Tocopherol [AASQPU^{+}23]. Todougba [AHK^{+}23]. Toho [AHK^{+}23]. Toho-Todougba [AHK^{+}23]. Tolerance [RHUJ24, ZZW^{+}22]. Tonga [VK23]. Tongue [CZW^{+}23, LZM^{+}23]. Tool [CLJ^{+}23, SMO^{+}23, TOB^{+}23]. Tool-Box [TOB^{+}23]. Tools [SKSL23]. Top [MMAO22]. Top-Fished [MMAO22]. Topic [Sor21]. Tor [INCD23]. Torquigener [KMS^{+}17]. Total [CdOCH^{+}23]. Totoaba [CBCMRD^{+}23]. Toxic [CFP^{+}23, GMBR^{+}21, JKK24, RKHAMM22]. Toxicity [BBCJ23, EBCM24, GLW^{+}22, HYN^{+}24, WWZ^{+}24]. Toxicology [CMP^{+}20, WWD^{+}23]. Trace [RKS^{+}24]. Tracers [OSM23]. Trachinotus [GGL^{+}23, WLW^{+}23]. Trachurus [BS23, FGHYCA23, MSA24]. Tracking [OOTC24]. Tract [TCB^{+}24]. Tracts [ATM^{+}24]. Trade [vSRB^{+}18]. Trade-Offs [vSRB^{+}18]. Traditional [EA18, MBC^{+}24, SUL^{+}23, SLL22, TKF^{+}18]. Trait [AGE^{+}18, SFP17, YYH^{+}24]. Trait-Mediated [SFP17]. Traits [CSB^{+}20, CHJ^{+}23, FCB^{+}21, SMO^{+}23, XWW^{+}24]. Trammel [CGY^{+}23]. Transcription [CBK^{+}21, SHW^{+}23]. Transcriptional [LYL^{+}24b, LYW^{+}24, PLJ^{+}23, SML^{+}23a]. Transcriptome [CYL^{+}23, CLL^{+}23b, DAHM19, FLX^{+}22, GGL^{+}23, JLW^{+}24, JCY^{+}23, LZW^{+}24, MYY^{+}23, MPM^{+}18, PYJ^{+}23, PLY^{+}24, RWF^{+}23, SDA23, SLY^{+}21, TJW^{+}22, WYG^{+}23, WHL^{+}24, WCY^{+}24, ZJJ^{+}22, ZZT^{+}23, ZZ24, ZSL^{+}23, ZJM^{+}23]. Transcriptome-Based [ZJM^{+}23]. Transcriptomes [DM24]. Transcriptomic [HWX^{+}23, JXW^{+}23, KWK^{+}24, SPQ^{+}24, WLN^{+}23, WSS^{+}19, ZCT^{+}23, ZZY^{+}23]. Transcriptomics [YHH^{+}20, ZHF^{+}22b]. Transcripts [ZLSB22, ZHF^{+}22b]. Transfer [JTS^{+}24, MIHH23]. Transgenic [GLW^{+}22]. Transient [HMVRFD19]. Transmissibility [LJK^{+}22]. Transmission [GMMNRS18, VCC^{+}18]. Transport [CWMX21, FMXQ23, LGZ^{+}23, LF24]. Transporter [MLZ^{+}21, OMC^{+}19]. Trap [RH19]. Traps [BASBW24, RH19]. Trash [CLL23a]. Trawl [CLL23a, FGG^{+}22, PMFBI22, TMPP23]. Treated [PBS^{+}22]. Treatment [LMC21, NDC^{+}23, PKV^{+}22]. Trematode [CSSMV^{+}23]. Trematomus [GGF^{+}22]. Trends [FM21b, Ols19, Piz22, RKL24, SOT^{+}23]. Trevally [MNO^{+}22]. Trial [JSJ^{+}24]. Tribe [KSAB^{+}23]. Tributary [SBT22]. Trichiuridae [SKK^{+}23]. Trichiurus [FYL^{+}23, SKK^{+}23]. Trichogaster [DM24]. Trichomycteridae [CFM^{+}23, CASMK23]. Trichomycterine [CASMK23]. trichopterus [DM24]. tridentatus [KSO^{+}23]. Triggered [RE21]. Triggerfish [HS24]. Trilostane [JSJ^{+}24]. Trionyx [JCY^{+}23]. Triple [RDANPA^{+}24]. Triploid [GLW^{+}22, MLZ^{+}21, WLW^{+}24, XHC^{+}22, ZZC^{+}22, ZSL^{+}23]. Triplophysa [LLWW23]. Trisopterine [Gae16, Gae17]. Triticum [WLL^{+}22]. trituberculatus [LWS^{+}23b]. Trophic [CMB^{+}24, GPD^{+}23]. Tropical [FQÁGTR^{+}17, MPM^{+}21, MVPMAV^{+}22, NRÁGPM^{+}18, PJPMMV^{+}22, PKV^{+}22, QWR^{+}23, SFP17, SBP23, WFZ^{+}23, dlCRMB^{+}22, KSWT22]. tropicus [FQÁGTR^{+}17, MPM^{+}21, MVPMAV^{+}22, NRÁGPM^{+}18, PJPMMV^{+}22, dlCRMB^{+}22]. Trout [APD^{+}23a, APD^{+}23b, CBK^{+}21, CSV^{+}19, DSC^{+}23, EAJG24, FMBPC^{+}23, FÅG^{+}23, HS18, KZC^{+}24, LUM18, LLKKV20, MD21, MDVM^{+}23, Näs18, PXH^{+}24, PLC^{+}24, RDI^{+}21, RDE^{+}23, SMO^{+}22, SUL^{+}23, SCSR22, TKF^{+}18, UH19, VAT^{+}23, VF23, VMS^{+}23, WFZ^{+}24a, WLW^{+}24, WLL^{+}24, Web23, YZH^{+}24, vSRB^{+}18]. TRPV1 [VAT^{+}23]. True [SMO^{+}23, FŠS^{+}23]. Trust [IR22]. trutta [EAJG24, RDE^{+}23]. Trypanorhynch [PYP17]. Tryptophan [GSHGE18, TPC^{+}23]. tubiashii [DLL^{+}22]. tulipa [MSK^{+}22]. tumbil [GMA^{+}24]. Tumor [CSGE23]. Tuna [DL23, EA18, HYXY23, LYL^{+}24b, LSJ24, NIN^{+}19, PBS^{+}22, ST17, THS^{+}22, VK23, WZG^{+}23, WLM^{+}20, XYT23, YYW^{+}23, ZCX^{+}22]. Turbidity [HV24, JBK^{+}23]. Turbot [CWMX21, DDD^{+}23, LWG^{+}23, LLW^{+}22, ŢPN^{+}23, TMPP23, WBK^{+}23, XYF^{+}24]. Turbulence [WMD^{+}24]. Türkiye [UGH^{+}24]. Turnover [BFM23]. Turtle [CZC^{+}22, KSAB^{+}23]. Twist [MPK^{+}23]. Two [APS^{+}21, BSB^{+}23, FLR^{+}22, GP17, HHM^{+}24, JJK21, KAJ^{+}24, LBC^{+}24, LMMC^{+}22, LBH^{+}24, Mee24, SLSC^{+}24, TCB^{+}24, TMPP23, VMBT24, VKP^{+}24, WZG^{+}23, WHL^{+}24, WBK^{+}23, ZVRH23]. Two-Spotted [LMMC^{+}22]. Type [CSF^{+}23, KKFC^{+}23, WPLK23, ZSL^{+}24, LWS^{+}21]. Types [SUL^{+}23, ZWD^{+}23]. typus [OSM23]. Tyrrhenian [DAF^{+}22, Poi24].

U.S. [EvSCB23, SWS22]. Uatumã [GAR^{+}24]. Uganda [KYB^{+}23, MTPK23, NGMR23]. Ultra [JSLE23]. Ultrafine [SIZ^{+}22]. Ultrasonography [LDX^{+}23]. Ultrasonography-Assisted [LDX^{+}23]. Ultrastructure [DNP^{+}23]. Ultraviolet [LSY^{+}23]. Ulva [CdOCH^{+}23, YLW^{+}23b].

Unconditioned [GGP^{+}23]. Undersized [MGMG24]. Understand [ST17]. Understanding [AGC23, BKJ^{+}24, HWW^{+}24a, SMO^{+}23, SRHCO23]. Underwater [HARB23a, LL18a, LL18b, LWZ^{+}23, SYL^{+}24, WLZ^{+}23, vSRB^{+}18]. Unfertilized [LdRCV24]. unguiculatus [MYY^{+}23, WFL^{+}23]. unicornis [Tay19]. Unit [GWH21, SHT^{+}23, WLM^{+}20]. United [BPOS19, BPO19, HCWH20]. Units [MLSC^{+}23]. Unraveling [PB24]. Unregulated [CXL23, OOG24]. Unreported [CXL23]. Untangling [HWW^{+}24a]. Untargeted [ITG^{+}18]. Unveils [YZL^{+}23]. Updated [GASS^{+}22]. UPLC [JSLE23]. UPLC-MS [JSLE23]. UPLC-MS/MS [JSLE23]. upon [TU18]. Upper [LLWW23, BAA^{+}19, GWH^{+}22, USRDFO^{+}22]. Upregulate [RDE^{+}23]. Upstream [LZC^{+}23b, SRBCGV^{+}21, ZS21]. Uptake [CFP^{+}23, TU18]. Upwelling [SZSW21]. Urchin [XWR^{+}23]. Urema [SVMLP23]. Urochloa [SFP17]. Urotrygon [AFTÁPA^{+}23, SNSVFL23]. USA [LLTM17, BAA^{+}19, EHGS23, HHAG24, MD21, SOT^{+}23, SMH^{+}22a]. Use [BDŞ^{+}24, BBF22, EBCM24, FQÁGTR^{+}17, GFDPSR22, HBG^{+}20, JBdFS^{+}22, KJ22, LUM18, LF24, MBAM19, PEP24, RCL^{+}23, RDI^{+}21, VRG^{+}24, WSL^{+}23, WYL23, XW24, XYT23, ZVRH23]. Used [GMMNRS18, NdNFK^{+}24b, PMFBI22, YHZ^{+}23]. Uses [LUM18]. Using [ATM^{+}24, AIŠRB22, AAN22, BEMC23, BAA^{+}19, CKMT23, CBANCM^{+}21, DQC^{+}23, EAJG24, FGG^{+}22, FWJ21, FBSB24, FCT19, HMN^{+}22, HDW^{+}23, HWZ21, ILA22, JFP^{+}18, JROHVH^{+}23, JTS^{+}24, JHNF24, KTT24, KEA^{+}23, LCW23a, LWZ^{+}23, Liu24, LLW^{+}22, MLSC^{+}23, NRKT19, PBS^{+}22, RDG^{+}20, RDANPA^{+}24, RGABD20, SLC^{+}22, SLS^{+}22, WLZ^{+}23, WSA^{+}23, WSI^{+}19, XLC^{+}24, XYC^{+}22, XYC^{+}23, XXL^{+}24, YCL^{+}23, ZLL^{+}23, ZCX^{+}22, ZS21].

Utilisation [MMVVCJ^{+}23]. Utility [FB22, ZZZ^{+}23]. Utilization [ATEfA^{+}21, LZC^{+}23a, NWN^{+}22, PNW^{+}22, Sam24, ST17].

V [SRBCGV^{+}21]. v1.0 [PSP^{+}22]. Vaccinated [GCFA^{+}22, ZLSB22]. Vaccine [LTZ^{+}22, LDD^{+}22, XYF^{+}24]. vachellii [PYJ^{+}23]. Valenciennes [CHH^{+}23]. Validation [JHNF24, MU21]. Valley [CTTW23]. Valproate [VLCCA^{+}23]. Value [EN22, MABÁMSM22, SCBSSMA24, Tay19, YLL22]. vannamei [BHP^{+}24, CdOCH^{+}23, CBCL23, EAE^{+}23, HSAF^{+}23, HXS^{+}23, HZG^{+}21, HSD^{+}24, LXW^{+}24, LSL^{+}24, LMNN21, NdNFK^{+}24b, NdNFK^{+}24a, NYS^{+}23, RXL^{+}24, SCHT23, SEA^{+}23, WPLK23, YHZ^{+}23, ZLL^{+}23, ZWD^{+}23]. var [CXW^{+}23, LCC^{+}24, XCW^{+}23]. var. [LWT^{+}24]. Variability [APD^{+}23a, APD^{+}23b, ALNVDG^{+}22, ENO21, GTC^{+}17, HZL^{+}22, NTP^{+}21, WZG^{+}23]. Variable [SLC^{+}22]. Variables [SMAR24]. Variants [AHJ^{+}23]. Variation [BBCJ23, BMTR23, dSCFQB^{+}23, GPD^{+}23, MKH24, MBPB24, SGAC20, VPPF^{+}19]. Variations [HMX^{+}21, KYB^{+}23, SKF^{+}23]. Varieties [WHL^{+}24]. Various [LYL^{+}24b, MTM^{+}19, RMA^{+}18, SPJ^{+}24, SIZ^{+}22, YLX^{+}22]. vasa [MKC^{+}22]. Vasotocin [HAdM^{+}24]. Vegetable [GMFG^{+}24]. velezensis [LJR^{+}24]. ventralis [WCX^{+}24]. Venus [PSW^{+}23]. Verification [HYXY23]. veronii [QXAY22, WCX^{+}24]. verreauxi [KKFC^{+}23]. versus [OOG24]. Vertebrae [NZVB20]. Vertebral [INCD23, YCR^{+}23]. Vertical [PEP24, VCC^{+}18, YCL^{+}23]. Vessel [CZCW23, VK23]. Vessels [FYH^{+}23, MFKS23]. vetula [HS24]. vgll3 [FHF23]. VHSV [CCP^{+}24]. VI [HYN^{+}24]. via [HWW^{+}24b, HZG^{+}21, NWN^{+}22, NWN^{+}23, SCBSSMA24, SYL^{+}24]. Viability [BASBW24, CHH^{+}23, EMFZ^{+}18, ŠT19, vKRNL^{+}19]. Vibration [FMXQ23, LGZ^{+}23]. Vibrio [DLL^{+}22, FLR^{+}22, JSRE^{+}24, LJK^{+}22, LTZ^{+}22, LDZ^{+}24, SEA^{+}23, XYF^{+}24]. Victoria [MTPK23]. Video [RGABD20]. Vietnam [TVL21]. View [Hal23, LV23]. Views [TOB^{+}23]. vinifera [HFEH^{+}23]. violacea [PEP24, RCR^{+}23, WGW^{+}23]. Viral [CCP^{+}24]. Virtual [BFP^{+}23]. Virulence [SZZ^{+}23, VČ24]. Virulent [ZLSB22]. Virus [CCP^{+}24, PVY^{+}21]. Viscera [MdM^{+}23]. Visibility [FBSB24]. Vision [JBGCG^{+}24]. Visual [FBSB24, LDZ^{+}24]. Vitamin [MYW^{+}24]. Vitellogenesis [DAHM19, MLK^{+}19, RWF^{+}23]. Vitellogenic [WSS^{+}19]. Vitis [HFEH^{+}23]. Vitro [CZW^{+}23, DDN19, EES^{+}23, MJL^{+}24, MKN^{+}16, PSS^{+}18, dCPRG^{+}21, SSK^{+}22, SSSS23, Web23, YHH^{+}20, dSJC^{+}21, vKRNL^{+}19]. vittatus [MAR^{+}18]. Viviparous [SMH^{+}24, UDG^{+}19]. Voandzeia [EPKV17]. Vocalization [CLL^{+}24]. Volatile [WDL^{+}23]. Volutharpa [YYH^{+}24]. Vukić [KRAFO23]. vulgaris [AZY^{+}24, Eny17, GRO^{+}17, RGVG19]. Vulnerability [FM21a, MSK^{+}22]. vulnificus [JSRE^{+}24, XYF^{+}24].

Wahoo [GPD^{+}23]. Walbaum [BFM23]. waleckii [ZLW^{+}23]. Wall [YBL^{+}22]. Wanlv [WFZ^{+}24b]. wanted [RLAE23]. Warm [VF23]. Warming [DLL^{+}23, KWK^{+}24, LMMC^{+}22, PFM^{+}20, ZZW^{+}22]. Washington [MD21, SOT^{+}23]. Waste [KKFC^{+}23, NWN^{+}21, TDN^{+}22]. Wastewater [LMC21]. Water [And23, CWMX21, CHH^{+}23, DCR^{+}23, GL23, GZX^{+}22, GSH^{+}24, HZS^{+}21, HXS^{+}23, IKT19, JE18, LHZ^{+}24, LOB^{+}23, LOBTL22, LCX^{+}23, LMNN21, MSB^{+}23, MGS^{+}23, PLC^{+}24, RDG^{+}20, RHU^{+}23, RSA17, RXL^{+}24, SLS^{+}22, TKF^{+}18, WLZ^{+}22b, WFZ^{+}24b, ZLZ^{+}24]. Waters [AKM23, BSB^{+}23, CLL^{+}22, DCD^{+}24, DTS^{+}17, FZZ^{+}24, FGHYCA23, HMX^{+}21, HS24, JCL^{+}24, LCZ^{+}23, LSJ24, LZL^{+}23, NIN^{+}19, SCSS23, ST17, SKK^{+}23, XZZ^{+}24, KAR^{+}23]. Wave [RGABD20, WXL^{+}23]. Wave-Swept [RGABD20]. Wavelength [CPJK23a, CPJK23b]. Wavelengths [SPJ^{+}24]. Weaknesses [CS23]. Weaning [BKBR^{+}23, PH23]. Web [AA23, QWR^{+}23]. Weight [CVANRD^{+}21, HARB23b, KZC^{+}24, LFH^{+}23, SGAC20, XWW^{+}24]. Weights [YWL^{+}24]. Weir [PAMG19]. Weirs [SRBCGV^{+}21]. Welcome [Est16]. Welfare [BHR^{+}23, DCL^{+}23a, DSC^{+}19, FCF19, dSGBdF23, dFBdSG^{+}19, IM24, OOTC24, SALC^{+}19, SAL19, SBB^{+}19, WSE^{+}21]. Wels [BBN^{+}24]. West [HWW^{+}24a, MMSK21, WSON24, BC23, MSK^{+}22, SML^{+}23a, YCR^{+}23]. Western [FPRAT^{+}23, KYB^{+}23, GPD^{+}23, MD21, NNL^{+}23, PEP24, SS23, TMPP23, YYW^{+}23]. Whale [OSM23]. Whelk [YYH^{+}24]. Where [BC23, RJMVC^{+}18]. White [CdOCH^{+}23, CBCL23, CHH^{+}23, EAE^{+}23, HSAF^{+}23, HL24, HMP^{+}24, HSD^{+}24, JLT22, LSL^{+}24, MNO^{+}22, MCSB^{+}19, RIF^{+}23, SCHT23, ZWD^{+}23, MNP^{+}16]. White-Barred [CHH^{+}23]. Whitefish [BWKS20, LLKKV20]. Whiteleg [BHP^{+}24, LMNN21, WPLK23, ZLL^{+}23]. Whitetail [ARH^{+}23]. Whole [MBZ^{+}21, SGG^{+}21, ULR^{+}23]. Whole-Body [ULR^{+}23]. Whole-Genome [MBZ^{+}21]. Wide [LZZ^{+}22, LLY^{+}22, ZLW^{+}23]. Wider [JLT22]. Wild [ATM^{+}24, AJF^{+}22, GMCF^{+}22, JHNF24, LDW^{+}21, LHZ^{+}23, MMD^{+}23, SOW^{+}23, TBPJ23]. Wild-Caught [AJF^{+}22]. Wind [JLT22]. Windows [MPM^{+}21]. Winter [FÅG^{+}23, GTC^{+}17]. Wistar [GLW^{+}22]. Withdrawal [HSD^{+}24]. within [APD24, And23, FJL^{+}23, JLT22]. woodiana [CLL^{+}23b]. Workers [TTT23]. Workflow [ITG^{+}18]. Worm [LCW^{+}23b]. Wound [ERE21]. Wound-Induced [ERE21]. Wreckfish [PPAB^{+}18, PLV^{+}19, WSI^{+}19, WSS^{+}19]. WSSV [HMP^{+}24, SPEGMC24]. wui [GWH^{+}22, LPK^{+}23a].

X [JCR^{+}22]. XC [YLX^{+}22]. XCR1 [YLX^{+}22]. Xenotransplantation [BGMM^{+}24]. XGBoost [HARB23b]. Xingu [dAMPS^{+}23]. Xiphias [LLTM17, Poi24]. Xisha [CLL^{+}22]. xoriguer [KRAFO23]. xpc [LSY^{+}23]. Xu [XYC^{+}23]. Xyrichthys [CSSMV^{+}23].

Yangcheng [LXT^{+}22, XJC^{+}22]. Yangtze [CYL^{+}23, CLL^{+}24, FLX^{+}22, JZX22, LZC^{+}23b, WSL^{+}23, WHY^{+}24, ZZW^{+}24, ZGY^{+}23]. Yaquina [HHAG24]. Yarlung [LHG^{+}23]. Year [APS^{+}21, CZC23, VKP^{+}24]. Year-Classes [VKP^{+}24]. Year-Round [CZC23]. Yearling [Mee24]. Years [Mil23, MWPS23, MWPS24]. Yeast [LYW^{+}24]. Yellow [CTY^{+}21, CWP^{+}21, CXT^{+}24, GGP^{+}23, HWW^{+}24b, JLW^{+}24, LTZ^{+}22, LCZ^{+}23, LHL^{+}24, NHR20, PYJ^{+}23, CHJ^{+}23, LLWW23, SLDC23]. Yellowfin [DL23, LYL^{+}24b, THS^{+}22, WLM^{+}20]. Yellowstone [KAB^{+}20]. Yellowtail [ASM^{+}22, AASQPU^{+}23, DBP^{+}20, MNO^{+}22]. yezoensis [LZH^{+}23]. YFI [XLX^{+}22]. YFI-G720 [XLX^{+}22]. Yields [CBANCM^{+}21]. Yolk [RAR^{+}18, RMSPMC^{+}22]. YOLOV5 [WLZ^{+}23, NLTL23]. YOLOv5s [WYL23].

Zambia  [NHC^{+}23,  SNZ^{+}23].

Zangbo [LHG^{+}23]. Zanthoxylum [PSS^{+}18, WZX^{+}24]. Zealand [MLK^{+}19]. Zebrafish [AJF^{+}22, BGMM^{+}24, CAC^{+}17, CGSBGN24, CFP^{+}23, CMP^{+}20, DSC^{+}19, DIL^{+}22, DK22, FLB^{+}21, HFEH^{+}23, LWS^{+}23a, LYL^{+}23, LCWH22, LLY^{+}24, LSY^{+}23, MBPB24, Ord19, TU18, TCV^{+}19, VLCCA^{+}23, VCL20, WWW19]. Zhejiang [XW24, XZZ^{+}24]. Zhoushan [JCL^{+}24]. Zone [KYB^{+}23, NNL^{+}23, RGTPSCCC24, RGABD20, HZL^{+}22]. Zoogeography [LKD22]. Zooplankton [ATEfA^{+}21, BWKS20, JSSD23]. Zootechnical [AHK^{+}23, DRFCL23, RDE^{+}23].

References

Akgun:2023:RTB

[AA23] Yagmur Akgun and Ekin Akoglu. Randall’s threadfin bream (Nemipterus randalli, Russell 1986) poses a potential threat to the Northeastern Mediterranean Sea food web. Fishes, 8(8):402, August 03, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/402.

Abu-Alya:2021:EPS

[AAAF^{+}21] Ibrahim S. Abu-Alya, Yousef M. Alharbi, Said I. Fathalla, Ibrahim S. Zahran, Saad M. Shousha, and Hassan A. Abdel-Rahman. Effect of partial soybean replacement by shrimp byproducts on the productive and economic performances in African catfish (Clarias lazera) diets. Fishes, 6(4):84, December 20, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/84.

Asnicar:2018:SGE

[AAB^{+}18] Davide Asnicar, Giedrė Ašmonaitė, Lina Birgersson, Charlotta Kvarnemo, Ola Svensson, and Joachim Sturve. Sand goby — an ecologically relevant species for behavioural ecotoxicology. Fishes, 3(1):13, February 20, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/1/13.

Armas:2022:IFP

[AAN22] Elier Armas, Hugo Arancibia, and Sergio Neira. Identification and forecast of potential fishing grounds for anchovy (Engraulis ringens) in Northern Chile using neural networks modeling. Fishes, 7(4): 204, August 15, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/204.

Asencio-Alcudia:2023:SPR

[AASQPU^{+}23] Gloria Gertrudys Asencio-Alcudia, Cesar Antonio Sepúlveda-Quiroz, Juan Carlos Pérez-Urbiola, María del Carmen Rodríguez-Jaramillo, Andressa Teles, Joan Sebastián Salas-Leiva, Rafael Martínez-García, Luis Daniel Jiménez-Martínez, Mario Galaviz, Dariel Tovar-Ramírez, and Carlos Alfonso Alvarez-González. Stressprotective role of dietary α-tocopherol supplementation in longfin yellowtail (Seriola rivoliana) juveniles. Fishes, 8(10):526, October 22, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/10/526.

Alesci:2024:EID

[ADM^{+}24] Alessio Alesci, Davide Di Paola, Sebastian Marino, Federica De Gaetano, Marco Albano, Silvana Morgante, Gabriele Rigano, Letterio Giuffrè, Magdalena Kotanska, Nunziacarla Spanò, and Eugenia Rita Lauriano. Exploring the internal defense system of Cerastoderma glaucum (Bruguière, 1789) exposed to pristine microplastics: The sentinel role of haemocytes as biomarkers. Fishes, 9 (6):241, June 20, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/241.

Ahmed:2023:CSG

[AEME^{+}23] Shaimaa A. A. Ahmed, Abdelhakeem El-Murr, Yasser Abd Elhakim, Mohamed M. Metwally, Amany Abd El Aziz Gharib, Shimaa A. Amer, Elsayed M. Younis, Abdel-Wahab A. Abdel-Warith, Simon J. Davies, and Enas N. M. Khalil. Comparative study on ginger powder and ginger extract nanoparticles: Effects on growth, immune– antioxidant status, tissue histoarchitecture, and resistance to Aeromonas hydrophila and Pseudomonas putida infection in Oreochromis niloticus. Fishes, 8 (5):259, May 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/5/259.

Alvarez-Fuentes:2023:REC

[AFTÁPA^{+}23] Carlos J. Alvarez-Fuentes, Javier Tovar-Ávila, Jorge Payan-Alejo, Darío A. Chávez-Arrenquín, Isaias H. Salgado-Ugarte, and Felipe Amezcua. Reproductive ecology of the Chilean round ray (Urotrygon chilensis, Günther, 1872) in the Southern Gulf of California. Fishes, 8 (4):193, April 04, 2023.

CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/193.

Aride:2021:DGE

[AGA^{+}21] Paulo Henrique R. Aride, Maria Fernanda S. Gomes, Darlan G. Azevedo, Gilson R. Sangali, Ana Cláudia F. Silva, Henrique D. Lavander, André B. Souza, Marcelo F. Polese, Douglas C. Mattos, Lucas A. Bassul, Leonardo D. Cardoso, Adriano T. Oliveira, and Caterina Faggio. Dusky grouper Epinephelus marginatus growth and survival when exposed to different photoperiods. Fishes, 6(3):31, August 12, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/31.

Alvarez:2023:IAI

[AGC23] Paula Alvarez, Dorleta Garcia, and Unai Cotano. Investigating the applicability of ichthyoplanktonic indices in better understanding the dynamics of the northern stock of the population of Atlantic hake Merluccius merluccius (L.). Fishes, 8(1): 50, January 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/1/50.

Arndt:2018:SEM

[AGE^{+}18] Erik Arndt, Or Givan, Dor Edelist, Oren Sonin, and Jonathan Belmaker. Shifts in Eastern Mediterranean fish communities: Abundance changes, trait overlap, and possible competition between native and non-native species. Fishes, 3(2): 19, April 20, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/2/19.

Adhikari:2023:PEV

[AHJ^{+}23] Dhurba Adhikari, Ida K. Hanssen, Steinar D. Johansen, Truls B. Moum, and Jarle T. Nordeide. Pitx 1 enhancer variants in spined and spine-reduced Subarctic European sticklebacks. Fishes, 8(3): 164, March 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/3/164.

Aizonou:2023:ESD

[AHK^{+}23] Romaric Aïzonou, Arsène Mathieu Houssou, Guillaume Koussovi, Comlan Eugène Dessouassi, Mardochée Ephraïm Achoh, Sessien Lionnel Hounhouedo, Clovis Idossou Hountchémè, and Hyppolite Agadjihouèdé. Effect of stocking density and feeding strategy on zootechnical parameters and profitability of Nile tilapia (Oreochromis niloticus) reared in floating cages in Toho-Todougba Complex Lagoon in Benin Republic. Fishes, 8 (4):192, April 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/192.

Andrews:2019:CAS

[AHLC19] Samuel N. Andrews, Sarah V. Hirtle, Tommi Linnansaari, and R. Allen Curry. Consumption of Atlantic salmon smolt by striped bass: a review of the predator–prey encounter literature and implications for the design of effective sampling strategies. Fishes, 4(4): 50, October 11, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/50.

Andrasunas:2022:AFM

[AIŠRB22] Vaidotas Andrašūnas, Edgaras Ivanauskas, Arvydas Švagždys, and Artūras Razinkovas-Baziukas. Assessment of four major fish species stocks in the Lithuanian and Russian parts of Curonian Lagoon (SE Baltic Sea) using CMSY method. Fishes, 7(1):9, January 03, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/9.

Axling:2022:BZL

[AJF^{+}22] Johanna Axling, Hampus Jakobsson, Natalia Frymus, Per-Ove Thörnqvist, Erik Petersson, and Svante Winberg. Boldness in zebrafish larvae— development and differences between a domesticated lab strain and offspring of wild-caught fish. Fishes, 7(4):197, August 08, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/197.

Aarstad:2023:NAF

[AJF23] Jarle Aarstad, Stig-Erik Jakobsen, and Arnt Fløysand. Norwegian aquaculture firms’ emphasis on environmental and social sustainability compared to firms in other industries. Fishes, 8(2):115, February 17, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/115.

Artero:2023:CWF

[AKM23] Céline Artero, Christopher C. Koenig, and Jessica E. Marsh. Are the coastal waters of French Guiana a source or sink habitat for Atlantic goliath grouper Epinephelus itajara? Fishes, 8(5):274, May 22, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/274.

Arechavala-Lopez:2022:SSR

[ALNVDG^{+}22] Pablo Arechavala-Lopez, Samira Nuñez-Velazquez, Carlos Diaz-Gil, Guillermo Follana-Berná, and João L. Saraiva. Suspended structures reduce variability of group risktaking responses of Dicentrarchus labrax juvenile reared in tanks. Fishes, 7(3):126, May 31, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/126.

Ahn:2024:EDC

[AM24] Hyojin Ahn and Michael J.

Miller. Environmental DNA characterization of the fish species composition in the Mukawa River and adjacent habitats. Fishes, 9(4):147, April 21, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/147.

An:2023:MCE

[AMK23] Hyung-Eun An, Min-Ho Mun, and Chang-Bae Kim. Metabarcoding by combining environmental DNA with environmental RNA to monitor fish species in the Han River, Korea. Fishes, 8(11):550, November 15, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/550.

An:2024:CBC

[AML^{+}24] Hyung-Eun An, Adeel Malik, Jeongho Lee, Min-Ho Mun, Kang Hyun Lee, Hah Young Yoo, and Chang-Bae Kim. Comparison of bacterial community composition in gut of Chinese mitten crabs from three distinct rivers in Korea. Fishes, 9(4):144, April 20, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/144.

AmorosoFerrari:2024:CHE

[AMT^{+}24] Natália Amoroso Ferrari, Leonardo Mantovani Favero, Cesar Toshio Facimoto, Alais Maria Dall Agnol, Marcos Letaif Gaeta, Thalita Evani Silva de Oliveira, Daniela Dib Gonçalves, Nelson Maurício Lopera-Barrero, Ulisses de Pádua Pereira, and Giovana Wingeter Di Santis. Clinical and histopathological evolution of acute intraperitoneal infection by Streptococcus agalactiae serotypes Ib and III in Nile tilapia. Fishes, 9(7):279, July 13, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/279.

Almodovar:2023:BAS

[ANA^{+}23] Ana Almodóvar, Graciela G. Nicola, Daniel Ayllón, Sheila Leal, Daniel F. Marchán, and Benigno Elvira. A benchmark for Atlantic salmon conservation: Genetic diversity and structure in a Southern European glacial refuge before the climate changed. Fishes, 8 (6):321, June 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/321.

Andriamahefazafy:2023:GDW

[And23] Mialy Andriamahefazafy.

Governing distant-water fishing within the blue economy in Madagascar: Policy frameworks, challenges and pathways. Fishes, 8(7):361, July 11, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/361.

Angradi:2018:FOR

[Ang18] Ted R. Angradi. A field observation of rotational feeding by Neogobius melanostomus. Fishes, 3 (1):5, January 22, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/5.

Aviles-Polanco:2023:CES

[APAHBMAG23] Gerzaín Avilés-Polanco, Marco Antonio Almendarez-Hernández, Luis Felipe Beltrán-Morales, and Fernando Aranceta-Garza. Cost efficiencies of the shrimp fishery in Mexico: a stochastic frontier analysis. Fishes, 8(9):472, September 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/472.

Antognazza:2023:CAA

[APD^{+}23a] Caterina M. Antognazza, Anja Palandaćić, Giovanni B. Delmastro, Giuseppe Crosa, and Serena Zaccara. Correction: Antognazza et al. Current and Historical Genetic Variability of Native Brown Trout Populations in a Southern Alpine Ecosystem: Implications for Future Management. Fishes 2023, 8, 411. Fishes, 8(11):561, November 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/561.  See

[APD^{+}23b].

Antognazza:2023:CHG

[APD^{+}23b] Caterina M. Antognazza, Anja Palandaćić, Giovanni B. Delmastro, Giuseppe Crosa, and Serena Zaccara. Current and historical genetic variability of native brown trout populations in a southern alpine ecosystem: Implications for future management. Fishes, 8(8): 411, August 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/411. See cor-

rection [APD^{+}23a].

Adyrbekova:2024:DCF

[APD24] Kamila Adyrbekova, Silvia Perea, and Ignacio Doadrio. Development and characterization of fifteen polymorphic microsatellite loci for rare and endangered species within Luciobarbus Heckel, 1843 genus in the Aral Basin and their conservation application. Fishes, 9 (5):169, May 7, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/169.

Azeiteiro:2021:DTA

[APS^{+}21] Ulisses M. Azeiteiro, Mário J. Pereira, Amadeu M. V. M. Soares, Heitor O. Braga, Fernando Morgado, Magda C. Sousa, João M. Dias, and Carlos Antunes. Dynamics of two anadromous species in a dam intersected river: Analysis of two 100-year datasets. Fishes, 6(2): 21, June 05, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/2/21.

Apostolou:2023:FBB

[APS^{+}23] Apostolos Apostolou, Luchezar Pehlivanov, Michael Schabuss, Horst Zornig, and Georg Wolfram. Fish biozonation in the Balkan Peninsula, especially in Bulgaria: a challenge. Fishes, 8(2): 91, February 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/91.

Adachi:2023:GPS

[ARH^{+}23] Aisni M. C. L. Adachi, Pollyana C. G. Roque, Fabio H. V. Hazin, Marcelo Vianna, Matheus M. Rotundo, Claudio Oliveira, Fausto Foresti, and Vanessa P. Cruz. Genetic population structure and diversity of the whitetail dogfish Squalus albicaudus (Chondrichthyes, Squaliformes) along the Brazilian Coast as identified by SNP markers. Fishes, 8(7):373, July 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/373.

Araujo:2022:DAA

[ASM^{+}22] Bruno Cavalheiro Araújo, Arleta Krystyna Skrzynska, Victor Hugo Marques, Aurora Tinajero, Oscar Basílio Del Rio-Zaragoza, Maria Teresa Viana, and José Antonio Mata-Sotres. Dietary arachidonic acid (20:4n-6) levels and its effect on growth performance, fatty acid profile, gene expression for lipid metabolism, and health status of juvenile California yellowtail (Seriola dorsalis). Fishes, 7 (4):185, July 25, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/185.

Abo-Taleb:2021:GPF

[ATEfA^{+}21] Hamdy A. Abo-Taleb, Mohamed M. M. Elfeky, Ahmad M. Azab, Mohamed M. Mabrouk, Mohamed A. Elokaby, Mohamed Ashour, Abdallah Tageldein Mansour, Othman F. Abdelzaher, Khamael M. Abualnaja, and Ahmed E. Sallam. Growth performance, feed utilization, gut integrity, and economic revenue of grey mullet, Mugil cephalus, fed an increasing level of dried zooplankton biomass meal as fishmeal substitutions. Fishes, 6(3): 38, September 10, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/3/38.

Amin:2024:SDM

[ATM^{+}24] Muhamad Amin, Hussein Taha, Laila Musdalifah, Muhamad Ali, Alimuddin Alimuddin, Sahrul Alim, and Takaomi Arai. Structure and diversity of microbiomes associated with the gastrointestinal tracts of wild spiny lobsters and profiling their potential probiotic properties using eDNA metabarcoding. Fishes, 9(7):264, July 4, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/264.

Axen:2018:OMA

[AVT18] Charlotte Axén, Niccolò Vendramin, and Anna Toffan. Outbreak of mortality associated with acipenser iridovirus European (AcIV-E) detection in Siberian sturgeon (Acipenser baerii) farmed in Sweden. Fishes, 3(4):42, October 16, 2018. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/3/4/42.

Anderson:2023:CAB

[AWCS23] Cody M. Anderson, Brian D. Wisenden, Cody A. Craig, and Craig A. Stockwell. Consistent antipredator behavioral responses among populations of Red River pupfish with disparate predator communities. Fishes, 8 (6):315, June 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/315.

Abdelbaky:2024:ICV

[AZY^{+}24] Shimaa A. Abdelbaky, Zakaria M. Zaky, Doha Yahia, Mohamed H. Kotob, Mohammed A. Ali, Mohammed Aufy, and Alaa El-Din H. Sayed. Impact of Chlorella vulgaris bioremediation and selenium on genotoxicity, nephrotoxicity and oxidative/ antioxidant imbalance induced by polystyrene nanoplastics in African catfish (Clarias gariepinus). Fishes, 9(2):76, February 15, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/2/76.

Buckwalter:2019:DLD

[BAA^{+}19] Joseph Buckwalter, Paul L. Angermeier, Jane Argentina, Skylar Wolf, Stephen Floyd, and Eric M. Hallerman. Drift of larval darters (family Percidae) in the Upper Roanoke River Basin, USA, characterized using phenotypic and DNA barcoding markers. Fishes, 4(4): 59, December 08, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/4/59.

Banaduc:2023:DDA

[BAA^{+}23] Doru Bănăduc, Sergey Afanasyev, John Robert Akeroyd, Aurel Năstase, Ion Năvodaru, Lucica Tofan, and Angela Curtean-Bănăduc. The Danube Delta: The Achilles heel of Danube River–Danube Delta– Black Sea region fish diversity under a Black Sea impact scenario due to sea level rise — a prospective review. Fishes, 8(7):355, July 07, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/355.

Bakit:2022:DEE

[BÁD^{+}22] José Bakit, Gonzalo Álvarez, Patricio A. Díaz, Eduardo Uribe, Rodrigo Sfeir, Sebastian Villasante, Tomas Gabriel Bas, Germán Lira, Hernán Pérez, Andrés Hurtado, Raúl González-Ávalos, and Jose Castillo-Venenciano. Disentangling environmental, economic, and technological factors driving scallop (Argopecten purpuratus) aquaculture in Chile. Fishes, 7(6): 380, December 08, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/380.

Baloch:2022:CSC

[BAP22] Ali Asghar Baloch, Ehdaa Eltayeb Eltigani Abdelsalam, and Veronika Piačková. Cytokines studied in carp (Cyprinus carpio L.) in response to important diseases. Fishes, 7(1): 3, December 24, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/3.

Bergsson:2019:RPE

[BAS^{+}19] Heidrikur Bergsson, Nikolaj Reducha Andersen, Morten Bo Søndergaard Svendsen, Per Juel Hansen, and John Fleng Steffensen. Respiratory physiology of European plaice (Pleuronectes platessa) exposed to Prymnesium parvum. Fishes, 4(2): 32, May 28, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/2/32.

Brown:2024:ATE

[BASBW24] Pete Brown, Tomas Araya-Schmidt, Terry Bungay, and Paul D. Winger. Assessing the technical and economic viability of galvanizing snow crab (Chionoecetes opilio) traps. Fishes, 9 (3):109, March 19, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/109.

Baldissera:2023:CKA

[BB23] Matheus D. Baldissera and Bernardo Baldisserotto. Creatine kinase activity as an indicator of energetic impairment and tissue damage in fish: a review. Fishes, 8 (2):59, January 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/59.

Baudoin:2023:AAV

[BBCJ23] Brooke A. Baudoin, Bonnie L. Brown, Robin D. Calfee, and Jill A. Jenkins. Amino acid variation at the mitochondrial binding site of Antimycin A is proposed to reflect sensitivity and toxicity differences among fish species. Fishes, 8(7):381, July 22, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/381.

Blanco:2022:ARE

[BBF22] Miguel Blanco, Lydia Bares, and Marcos Ferasso. Analysis of regional efficiency in the use of the European Maritime and Fisheries Fund (FEMP) in the Eurozone countries. Fishes, 7(2):53, April 22, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/2/53.

Bergstrom:2024:SDM

[BBN^{+}24] Kristofer Bergström, Hanna Berggren, Oscar Nordahl, Per Koch-Schmidt, Petter Tibblin, and Per Larsson. Seasonal and daily movement patterns by Wels catfish (Silurus glanis) at the northern fringe of its distribution range. Fishes, 9(7):280, July 14, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/7/280.

Barreiros:2023:WAG

[BC23] João Pedro Barreiros and Felicia C. Coleman. West African goliath grouper: Where are they between Senegal and Angola? Fishes, 8(6):318, June 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/6/318.

Barrows:2023:IKM

[BCG^{+}23] Frederick T. Barrows, Kelly B. Campbell, T. Gibson Gaylord, Rodrigo C. M. Sanchez, Sergio A. Castillo, and Ewen McLean. Influence of krill meal on the performance of postsmolt Atlantic salmon that are fed plant-based and animal-based fishmeal and fish oil-free diets. Fishes, 8(12): 590, November 30, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/590.

Barbacariu:2024:EUG

[BDŞ^{+}24] Cristian-Alin Barbacariu, Lenuţa Dîrvariu, Dana Andreea Şerban, Cristina Mihaela Rîmbu, Cristina Elena Horhogea, Gabriela Dumitru, Elena Todiraşcu-Ciornea, Constantin Lungoci, and Marian Burducea. Evaluating the use of grape pomace in Cyprinus carpio nutrition: Effects on growth, biochemistry, meat quality, microbiota, and oxidative status. Fishes, 9(6):219, June 7, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/219.

Boussard:2023:NSS

[BEF^{+}23] Annika Boussard, Stephanie Edlund, Stephanie Fong, David Wheatcroft, and Niclas Kolm. No sexspecific effects of artificial selection for relative telencephalon size during detour learning and spatial discrimination in guppies (Poecilia reticulata). Fishes, 8(11): 536, October 26, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/536.

Batjakas:2023:LDC

[BEG^{+}23] Ioannis E. Batjakas, Athanasios Evangelopoulos, Maria Giannou, Sofia Pappou, Eleftheria Papanikola, Maria Atsikvasi, Dimitris Poursanidis, and Chrysoula Gubili. Lionfish diet composition at three study sites in the Aegean Sea: an invasive generalist? Fishes, 8 (6):314, June 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/314.

Barnes:2023:CAM

[BEMC23] Megan Barnes, Brad Ebanks, Andrew Mac-Coll, and Lisa Chakrabarti. A common anaesthetic, MS-222, alters measurements made using highresolution respirometry in the three-spined stickleback (Gasterosteus aculeatus). Fishes, 8(1):

42, January 07, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/42.

Booth:2016:FAP

[BF16] Mark A. Booth and D. Stewart Fielder. Fortification of an aquafeed with potassium chloride does not improve survival of juvenile Australian snapper Pagrus auratus reared in potassium deficient saline groundwater. Fishes, 1(1):52–64, September 09, 2016. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/1/1/52.

Basilone:2023:AGE

[BFB^{+}23] Gualtiero Basilone, Rosalia Ferreri, Angelo Bonanno, Simona Genovese, Marco Barra, and Salvatore Aronica. Age and growth of European sardine (Sardina pilchardus) in the Central Mediterranean Sea: Implication for stock assessment. Fishes, 8 (4):202, April 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/202.

Babin:2019:DRA

[BFLC19] Amanda Babin, Lauren Fitzpatrick, Tommi Linnansaari, and R. Allen Curry. Detection range of acoustic receivers in a large hydropower reservoir. Fishes, 4(4):60, December 11, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/60.

Binstock:2023:ITF

[BFM23] Addie L. Binstock, Audrey S. Fox, and John A. Mohan. Isotopic turnover and fractionation of δ^{15}N and δ^{13}C in captive Pseudopleuronectes americanus (Walbaum). Fishes, 8(9):469, September 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/469.

Bolognini:2023:COD

[BFP^{+}23] Luca Bolognini, Cristina Frittelloni, Francesca Perretta, Martina Scanu, and Fabio Grati. From crisis to opportunity: Developing a virtual marketplace to enhance sustainability and resilience in small-scale fisheries. Fishes, 8(5): 272, May 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/272.

Blanes-Garcia:2024:XEE

[BGMM^{+}24] Marta Blanes-García, Zoran Marinović, Marina Morini, Alain Vergnet, Ákos Horváth, and Juan F. Asturiano. Xenotransplantation of European eel (Anguilla anguilla) spermatogonia in zebrafish (Danio rerio) and European sea bass (Dicentrarchus labrax). Fishes, 9(7): 290, July 21, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/290.

Bouska:2020:GSH

[BGT^{+}20] Wesley W. Bouska, David C. Glover, Jesse T. Trushenski, Silvia Secchi, James E. Garvey, Ruairi MacNamara, David P. Coulter, Alison A. Coulter, Kevin Irons, and Andrew Wieland. Geographicscale harvest program to promote invasivorism of bigheaded carps. Fishes, 5(3):29, September 01, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/3/29.

Bogwald:2024:ECF

[BHP^{+}24] Isak Bøgwald, Simon Herrig, Alice Marie Pedersen, Sileshi Gizachew Wubshet, and Karl-Erik Eilertsen. Effect of Calanus finmarchicus hydrolysate inclusion on diet attractiveness for whiteleg shrimp (Litopenaeus vannamei). Fishes, 9 (4):134, April 13, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/134.

Bassmann:2023:ESD

[BHR^{+}23] Björn Baßmann, Lisa Hahn, Alexander Rebl, Lisa Carolina Wenzel, Marc-Christopher Hildebrand, Marieke Verleih, and Harry Wilhelm Palm. Effects of stocking density, size, and external stress on growth and welfare of African catfish (Clarias gariepinus Burchell, 1822) in a commercial RAS. Fishes, 8 (2):74, January 26, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/74.

Bischoff:2023:DFA

[BKBR^{+}23] Adrian A. Bischoff, Melanie Kubitz, Laura Ballesteros-Redondo, Marcus Stüeken, Tobias Rapp, Patrick Fink, Wilhelm Hagen, and Harry Wilhelm Palm. Dynamics of fatty acids in pikeperch (Sander lucioperca) larvae and juveniles during early rearing and weaning in a commercial RAS — implications for dietary refinement. Fishes, 8(9): 444, August 31, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/444.

Boissonnot:2024:UCL

[BKJ^{+}24] Lauris Boissonnot, Camilla Karlsen, Thor Magne Jonassen, Silje Stensby-Skjærvik, Torolf Storsul, and Albert Kjartan Dagbjartarson Imsland. Understanding the causes of lumpfish (Cyclopterus lumpus) mortality in Norwegian hatcheries: Challenges and opportunities. Fishes, 9(7): 288, July 19, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/288.

Barua:2022:ALB

[BLA^{+}22] Suman Barua, Qun Liu, Mohammed Shahidul Alam, Petra Schneider, and Mohammad Mojibul Hoque Mozumder. Application of lengthbased assessment methods to elucidate biological reference points of black pomfret stock in the Bay of Bengal, Bangladesh. Fishes, 7 (6):384, December 11, 2022. CODEN ???? ISSN 2410-3888. URL

                                        https://www.mdpi.com/
                                        2410-3888/7/6/384.

Bosi:2022:MCM

[BMMD22] Giampaolo Bosi, Paolo Merella, Barbara J. Maynard, and Bahram Sayyaf Dezfuli. Microscopic characterization of the mucous cells and their mucin secretions in the alimentary canal of the blackmouth catshark Galeus melastomus (Chondrichthyes: Elasmobranchii). Fishes, 7(1):8, January 01, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/8.

Bourquin:2023:CGC

[BMOH23] Rebecca Bourquin, Michael J. Moore, Donald J. Orth, and Eric M. Hallerman. Conservation genetics of clinch dace Chrosomus sp. cf. saylori. Fishes, 8(7):365, July 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/365.

Barragan-Mendez:2018:AEC

[BMSGS^{+}18] Cristina Barragán-Méndez, Fini Sánchez-García, Ignacio Sobrino, Juan Miguel Mancera, and Ignacio Ruiz-Jarabo. Air exposure in catshark (Scyliorhinus canicula) modify muscle texture properties: a pilot study. Fishes, 3(3):34, September 04, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/34.

Burbank:2023:TVB

[BMTR23] Jacob Burbank, Jenni L.

McDermid, François Turcotte, and Nicolas Rolland. Temporal variation in Von Bertalanffy growth curves and generation time of Southern Gulf of St. Lawrence spring and fall spawning Atlantic herring (Clupea harengus). Fishes, 8 (4):205, April 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/205.

Bonvechio:2023:SEF

[BPCC23] Kimberly I. Bonvechio, Ramesh Paudyal, Chelsey Crandall, and Andrew K. Carlson. Survey evaluation of Florida’s freshwater fisheries long-term monitoring program. Fishes, 8 (4):216, April 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/216.

Burton:2019:PEA

[BPO19] Michael L. Burton, Jennifer C. Potts, and Andrew D. Ostrowski. Preliminary estimates of age, growth and natural mortality of margate, Haemulon album, and black margate, Anisotremus surinamensis, from the Southeastern United States. Fishes, 4(3):44, August 15, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/3/44.

Burton:2019:AGN

[BPOS19] Michael L. Burton, Jennifer C. Potts, Andrew D. Ostrowski, and Kyle W. Shertzer. Age, growth, and natural mortality of graysby, Cephalophilis cruentata, from the Southeastern United States. Fishes, 4(3):36, June 27, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/3/36.

Batista:2023:SLS

[BRB^{+}23] Rosana Oliveira Batista, Bianca Leticia Richter, Jorge Filipe Banze, Delano Dias Schleder, Maria Salhi, Renata Oselame Nobrega, Maria Fernanda Oliveira da Silva, Bruna Mattioni, James Eugene Pettigrew, and Débora Machado Fracalossi. Soy lecithin supplementation promotes growth and increases lipid digestibility in GIFT Nile tilapia raised at suboptimal temperature. Fishes, 8(8):404, August 03, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/404.

Belkin:2023:RSS

[BS23] Igor M. Belkin and Xin-Tang Shen. Remote sensing of the subtropical front in the Southeast Pacific and the ecology of Chilean jack mackerel Trachurus murphyi. Fishes, 8(1): 29, January 02, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/29.

Barash:2023:DPD

[BSB^{+}23] Adi Barash, Aviad Scheinin, Eyal Bigal, Ziv Zemah Shamir, Stephane Martinez, and Dan Tchernov. Depth partitioning and diel movement of two large carcharhinid sharks in extremely shallow waters. Fishes, 8(2):85, January 31, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/85.

Basto-Silva:2022:EDP

[BSGMC^{+}22] Catarina Basto-Silva, Irene García-Meilán, Ana Couto, Cláudia R. Serra, Paula Enes, Aires Oliva-Teles, Encarnación Capilla, and Inês Guerreiro. Effect of dietary plant feedstuffs and protein/ carbohydrate ratio on gilthead seabream (Sparus aurata) gut health and functionality. Fishes, 7 (2):59, April 07, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/2/59.

Becker:2023:EHB

[BSH^{+}23] Alessandra Janaína Becker, Sandro Santos, Berta Maria Heinzmann, Braulio Otomar Caron, and Bernardo Baldisserotto. Exposure of Hyalella bonariensis (Crustacea, Amphipoda) to essential oils: Effects on anesthesia and swimming activity. Fishes, 8(3): 149, March 02, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/3/149.

Baduy:2020:DRA

[BSR^{+}20] Flavia Baduy, João L.

Saraiva, Filipe Ribeiro, Adelino V. M. Canario, and Pedro M. Guerreiro. Distribution and risk assessment of potential invasiveness of Australoheros facetus (Jenyns, 1842) in Portugal. Fishes, 5(1): 3, December 27, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/1/3.

Badouvas:2024:FHP

[BTSK24] Nicholas Badouvas, Konstantinos Tsagarakis, Stylianos Somarakis, and Paraskevi K. Karachle. Feeding habits and prey composition of six mesopelagic fish species from an isolated central Mediterranean basin. Fishes, 9(7):277, July 12, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/277.

Bernet:2020:ZFI

[BWKS20] Daniel Bernet, Thomas Wahli, Christoph Küng, and Helmut Segner. Zooplankton feeding induces macroscopical gonad malformations in whitefish (Coregonus ssp.) from Lake Thun, Switzerland. Fishes, 5 (3):26, August 20, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/3/26.

Caccia:2017:APE

[CAC^{+}17] Elisabetta Caccia, Maria Agnello, Marcello Ceci, Patricia Strickler Dinglasan, Gerardo R. Vasta, and Nicla Romano. Antimicrobial peptides are expressed during early development of zebrafish (Danio rerio) and are inducible by immune challenge. Fishes, 2(4):20, November 08, 2017. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/4/20.

Costa:2023:MPT

[CASMK23] Wilson J. E. M. Costa, Valter M. Azevedo-Santos, José Leonardo O. Mattos, and Axel M. Katz. Molecular phylogeny, taxonomy and distribution patterns of trichomycterine catfishes in the Middle Rio Grande Drainage, South-Eastern Brazil (Siluriformes: Trichomycteridae). Fishes, 8(4): 206, April 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/206.

Curiel-Bernal:2021:UOR

[CBANCM^{+}21] Marcelo V. Curiel-Bernal, E. Alberto Aragón-Noriega, Miguel Á. Cisneros-Mata, Laura Sánchez-Velasco, S. Patricia A. Jiménez-Rosenberg, and Alejandro Parés-Sierra. Using observed residual error structure yields the best estimates of individual growth parameters. Fishes, 6(3): 35, September 02, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/35.

Cheng:2023:MBS

[CBCL23] Ann-Chang Cheng, Rolissa Ballantyne, Shieh-Tsung Chiu, and Chun-Hung Liu. Microencapsulation of Bacillus subtilis E20 probiotic, a promising approach for the enrichment of intestinal microbiome in white shrimp, Penaeus vannamei. Fishes, 8 (5):264, May 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/264.

Curiel-Bernal:2023:MEG

[CBCMRD^{+}23] Marcelo V. Curiel-Bernal, Miguel Á. Cisneros-Mata, Guillermo Rodríguez-Domínguez, Laura Sánchez-Velasco, S. Patricia A. Jiménez-Rosenberg, Alejandro Parés-Sierra, and E. Alberto Aragón-Noriega. Modelling early growth of Totoaba macdonaldi (Teleostei: Sciaenidae) under laboratory conditions. Fishes, 8(3):155, March 04, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/3/155.

Chen:2021:SCP

[CBK^{+}21] Huiqin Chen, Baoliang Bi, Lingfu Kong, Hua Rong, Yanhua Su, and Qing Hu. Seasonal changes in plasma hormones, sex-related genes transcription in brain, liver and ovary during gonadal development in female rainbow trout (Oncorhynchus mykiss). Fishes, 6(4):62, November 12, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/62.

Christidis:2024:AOB

[CBP^{+}24] Georgios Christidis, Stratos Batziakas, Panagiota Peristeraki, Evangelos Tzanatos, Stylianos Somarakis, and George Tserpes. Another one bites the net: Assessing the economic impacts of Lagocephalus sceleratus on small-scale fisheries in Greece. Fishes, 9 (3):104, March 07, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/104.

Casado:2018:ASG

[CCCFE18] Francisco Casado, Santiago Casado, Diana Ceballos-Francisco, and María Ángeles Esteban. Assessment of the scales of gilthead seabream (Sparus aurata L.) by image analysis and atomic force microscopy. Fishes, 3 (1):9, January 30, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/1/9.

Cortay:2019:DAP

[CCFP19] Aurore Cortay, Tatiana Colchen, Pascal Fontaine, and Alain Pasquet. Does addition of perch larvae as prey affect the growth, development and cannibalism rate of pikeperch larvae? Fishes, 4(1): 21, March 18, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/21.

Choi:2024:PEL

[CCP^{+}24] Hee-Jae Choi, Da-Yeon Choi, Jung-Jin Park, Hye Jin Jung, Bo-Seong Kim, Sung-Chul Hong, Jun-Hwan Kim, and Yue Jai Kang. Protective effects of long double-stranded RNA with different CpG motifs against Miamiensis avidus and Viral Hemorrhagic Septicemia Virus (VHSV) infections in olive flounder (Paralichthys olivaceus). Fishes, 9(6): 227, June 13, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/6/227.

Carvalho:2023:ITS

[CdOCH^{+}23] Andrezza Carvalho, Léa Carolina de Oliveira Costa, Mariana Holanda, Luís H. Poersch, and Gamze Turan. Influence of total suspended solids on the growth of the sea lettuce Ulva lactuca integrated with the Pacific white shrimp Litopenaeus vannamei in a biofloc system. Fishes, 8(3):163, March 15, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/3/163.

Ceballos-Francisco:2020:ELD

[CFCE20] Diana Ceballos-Francisco, Alberto Cuesta, and María Ángeles Esteban. Effect of light–dark cycle on skin mucosal immune activities of gilthead seabream (Sparus aurata) and European sea bass (Dicentrarchus labrax). Fishes, 5(1): 10, February 24, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/1/10.

Ciepiela:2021:BIB

[CFLK21] Lindsy R. Ciepiela, Ryan M. Fitzpatrick, Samuel T. Lewis, and Yoichiro Kanno. Behavioral interactions between a native and an invasive fish species in a thermally heterogeneous experimental chamber. Fishes, 6(4):75, December 07, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/75.

Costa:2023:MMD

[CFM^{+}23] Wilson J. E. M. Costa, Caio R. M. Feltrin, José Leonardo O. Mattos, Roger H. Dalcin, Vinicius Abilhoa, and Axel M. Katz. Morphomolecular discordance? Re-approaching systematics of Cambeva (Siluriformes: Trichomycteridae) from the Guaratuba-Babitonga-Itapocu Area, Southern Brazil. Fishes, 8(2):63, January 20, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/2/63.

Contino:2023:URB

[CFP^{+}23] Martina Contino, Greta Ferruggia, Roberta Pecoraro, Elena Maria Scalisi, Gianfranco Cavallaro, Carmela Bonaccorso, Cosimo Gianluca Fortuna, Antonio Salvaggio, Fabiano Capparrucci, Teresa Bottari, and Maria Violetta Brundo. Uptake routes and biodistribution of polystyrene nanoplastics on zebrafish larvae and toxic effects on development. Fishes, 8(3): 168, March 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/168.

Cerbule:2020:SLB

[CG20] Kristine Cerbule and Jacques Godfroid. Salmon louse (Lepeophtheirus salmonis (Krøyer)) control methods and efficacy in Atlantic Salmon (Salmo salar (Linnaeus)) aquaculture: a literature review. Fishes, 5 (2):11, March 31, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/2/11.

Calderon-Garcia:2024:EZE

[CGSBGN24] Andres Angel Calderon-Garcia, Mercedes Sánchez-Barba, and Veronica Gonzalez-Nunez. Exposure of zebrafish embryos to morphine and cocaine induces changes in the levels of dopamine and of proteins related to the reward pathway. Fishes, 9(7):268, July 5, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/268.

Cheng:2023:CEM

[CGY^{+}23] Zhaohai Cheng, Lei Gao, Lixiong Yu, Xinbin Duan, Fengyue Zhu, Huiwu Tian, Daqing Chen, and Mingdian Liu. Catch efficiency of multi-mesh trammel nets for sampling freshwater fishes. Fishes, 8(9):464, September 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/464.

Carlson:2023:BPD

[CH23] Andrew K. Carlson and Mark V. Hoyer. Bluegill population demographics as related to abiotic and biotic factors in Florida lakes. Fishes, 8(2):100, February 07, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/100.

Chiu:2023:CRB

[CHH^{+}23] Pei-Sheng Chiu, Shine-Wei Ho, Cheng-Hsuan Huang, Yen-Chun Lee, and Yu-Hung Lin. Captive reproductive behavior, spawning, and early development of whitebarred goby Amblygobius phalaena (Valenciennes, 1837) and examined larval survival and viability at different water temperatures and salinities. Fishes, 8 (7):364, July 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/364.

Chen:2023:QMG

[CHJ^{+}23] Yuhan Chen, Jintai Huang, Zhan Jin, Junping Chen, Meng Zhang, Miao Yu, Hongxia Jiang, Lei Wang, and Zhigang Qiao. QTL mapping of growth traits in Yellow River carp (Cyprinus carpio haematopterus) at 5–17 months after hatching. Fishes, 8(2):

79, January 29, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/79.

Chai:2021:IFS

[CHW21] Panfeng Chai, Qiuguang Hu, and Xinyi Wei. Influence of fishery subsidies on fishing: Empirical test based on China’s provincial panel data. Fishes, 6(3):40, September 13, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/40.

Chen:2024:AOA

[CHZ^{+}24] Junxiang Chen, Shilong He, Zelong Zhang, Jiajun Li, Xiuxia Zhang, Juntao Li, Jiarui Xu, Peihua Zheng, Jianan Xian, and Yaopeng Lu. Application of organic acid salts as feed additives in some aquatic organisms: Potassium diformate. Fishes, 9(3): 85, February 24, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/85.

Cai:2023:SAC

[CKMT23] Kai Cai, Richard Kindong, Qiuyun Ma, and Siquan Tian. Stock assessment of chub mackerel (Scomber japonicus) in the Northwest Pacific using a multi-model approach. Fishes, 8(2): 80, January 30, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/80.

Cabillon:2019:MBF

[CL19] Nikko Alvin R. Cabillon and Carlo C. Lazado. Mucosal barrier functions of fish under changing environmental conditions. Fishes, 4(1): 2, January 10, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/1/2.

Choi:2021:EMT

[CL21] Soo-Cheol Choi and In-Ah Lee. Effect of MMP/TIMP balancing of Cynoglossus semilaevis shell extracts on skin protection. Fishes, 6 (3):34, August 24, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/3/34.

Chen:2023:NCI

[CLJ^{+}23] Yanghui Chen, Yuan Li, Dongneng Jiang, Defeng Zhang, Yu Huang, Jia Cai, Jichang Jian, and Bei Wang. A new conditionally immortalized Nile tilapia (Oreochromis niloticus) heart cell line: Establishment and functional characterization as a promising tool for tilapia myocarditis studies. Fishes, 8(3):167, March 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/167.

Chen:2022:BME

[CLL^{+}22] Ziyue Chen, Huajie Lu, Wei Liu, Kai Liu, and Xinjun Chen. Beak microstructure estimates of the age, growth, and population structure of purpleback flying squid (Sthenoteuthis oualaniensis) in the Xisha Islands waters of the South China Sea. Fishes, 7(4):187, July 26, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/187.

Chen:2023:JTF

[CLL23a] Chao-Ching Chen, Hsin-Zong Lin, and Kuo-Wei Lan. Juvenile and Trash Fish Excluder Device (JTED) for Taiwanese bottom trawl. Fishes, 8 (4):189, March 30, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/189.

Chen:2023:CII

[CLL^{+}23b] Xiubao Chen, Hongbo Liu, Karsten Liber, Tao Jiang, and Jian Yang. Copper-induced ionoregulatory disturbance, histopathology, and transcriptome responses in freshwater mussel (Anodonta woodiana) gills. Fishes, 8 (7):368, July 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/368.

Chen:2024:VPE

[CLL^{+}24] Jia Chen, Haiying Liang, Danqing Lin, Jialu Zhang, Dong Li, Kun Ye, Wenfei Lu, and Kai Liu. Vocalization pattern and echolocation signal characteristics of Yangtze finless porpoise (Neophocaena asiaeorientalis asiaeorientalis) in captivity. Fishes, 9 (4):119, March 28, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/119.

Chen:2024:RAS

[CLZ^{+}24] Ao Chen, Zehua Lv, Junbo Zhang, Gangyi Yu, and Rong Wan. Review of the accuracy of satellite remote sensing techniques in identifying coastal aquaculture facilities. Fishes, 9(2): 52, January 27, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/52.

Checa:2024:CAA

[CMB^{+}24] Daniel Checa, Brett M.

Macey, John J. Bolton, Marissa Brink-Hull, Pauline O’Donohoe, Alessandro Cardozo, Luis Henrique Poersch, and Inmaculada Sánchez. Circularity assessment in aquaculture: The case of integrated multi-trophic aquaculture (IMTA) systems. Fishes, 9(5):165, May 4, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/165.

Cotou:2024:GPE

[CMC^{+}24] Efthimia Cotou, Helen Miliou, Evanthia Chatzoglou, Eirini Schoina, Nektarios Politakis, Dimitra Kogiannou, Eleni Fountoulaki, Afrodite Androni, Aggeliki Konstantinopoulou, Georgia Assimakopoulou, and Cosmas Nathanailides. Growth performance and environmental quality indices and biomarkers in a co-culture of the European sea bass with filter and deposit feeders: a case study of an IMTA system. Fishes, 9(2):69, February 08, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/2/69.

Costa:2020:ZMN

[CMP^{+}20] Bruna Patricia Dutra Costa, Layana Aquino Moura, Sabrina Alana Gomes Pinto, Monica Lima-Maximino, and Caio Maximino. Zebrafish models in neural and behavioral toxicology across the life stages. Fishes, 5(3):23, July 31, 2020. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/3/23.

Chaiyasut:2023:DAP

[CMP^{+}23] Kasidis Chaiyasut, Paul Merviel, Peter Palma, Johanna Perschthaler, Eduardo Jimenez-Fernandez, Andrew Davie, and Alejandro P. Gutierrez. Development of an accurate polymerase chain reaction (PCR) assay for genetic sex identification in lumpfish (Cyclopterus lumpus) based on male-specific anti-mullerian hormone (amh) gene. Fishes, 8 (6):327, June 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/327.

Chu:2024:RBM

[CP24] Elaine Chu and Daniel Pauly. The relationship between mean length at maturity and maximum length in coral reef fish. Fishes, 9 (4):130, April 9, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/130.

Choi:2023:CCA

[CPJK23a] Young Jae Choi, Seul Gi Na Ra Park, A-Hyun Jo, and Jun-Hwan Kim. Correction: Choi et al. Physiological Effect of Extended Photoperiod and Green Wavelength on the Pituitary Hormone, Sex Hormone and Stress Response in Chub Mackerel, Scomber japonicus. Fishes 2023, 8, 77. Fishes, 8(5):263, May 16, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/263.  See

[CPJK23b].

Choi:2023:PEE

[CPJK23b] Young Jae Choi, Seul Gi Na Ra Park, A-Hyun Jo, and Jun-Hwan Kim. Physiological effect of extended photoperiod and green wavelength on the pituitary hormone, sex hormone and stress response in chub mackerel, Scomber japonicus. Fishes, 8(2): 77, January 29, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/77. See cor-

rection [CPJK23a].

Crisostomo:2024:RCP

[CPVMA^{+}24] Rafael Octavio Crisóstomo, Renzo Pepe-Victoriano, Sheda Méndez-Ancca, Abel Walter Zambrano-Cabanillas, Olegario Marín-Machuca, Hernan Mauricio Perez, Víctor Yana-Mamani, and Mario Ruiz-Choque. Reproductive conditioning of the Peruvian scallop Argopecten purpuratus in different environments. Fishes, 9(1): 9, December 24, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/1/9.

Crozier:2023:CRI

[CS23] Lisa G. Crozier and Jared E. Siegel. A comprehensive review of the impacts of climate change on salmon: Strengths and weaknesses of the literature by life stage. Fishes, 8 (6):319, June 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/319.

Carballo:2020:HEG

[CSB^{+}20] Carlos Carballo, Hyun Suk Shin, Concepción Berbel, Maria Jesús Zamorano, Juan Jose Borrego, Eva Armero, Juan Manuel Afonso, and Manuel Manchado. Heritability estimates and genetic correlation for growth traits and LCDV susceptibility in gilthead sea bream (Sparus aurata). Fishes, 5(1): 2, December 25, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/1/2.

Chen:2023:NCT

[CSF^{+}23] Sentao Chen, Peng Shi, Qingkai Feng, Xiaoting Qiu, Jilin Xu, Xiaojun Yan, and Chengxu Zhou. A novel C-type lectin and its potential role in feeding and feed selection in Ruditapes philippinarum. Fishes, 8(2): 62, January 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/62.

Campos-Sanchez:2023:ECF

[CSGE23] Jose Carlos Campos-Sánchez, Francisco A. Guardiola, and María Ángeles Esteban. Effects of cantharidin on fish erythrocytes, tumor cell lines, and marine pathogenic bacteria. Fishes, 8(5): 270, May 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/5/270.

Cruces:2023:PBG

[CSJ^{+}23] Celso Luis Cruces, Raquel Simões, Arnaldo Maldonado Júnior, Ruperto Severino, Jhon Darly Chero, and José Luis Luque. Proposal of Brotulella n. gen. for monogeneans from the gills of the Pacific bearded brotula Brotula clarkae Hubbs, 1944 (Ophidiiformes: Ophidiidae) based on morphological and molecular evidence. Fishes, 8(12): 588, November 30, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/588.

Chai:2024:AED

[CSL24] Yuqiao Chai, Shuhao Sun, and Yingdong Li. Assessing the effects of dietary tea polyphenols on the gut microbiota of loaches (Paramisgurnus dabryanus) under chronic ammonia nitrogen stress. Fishes, 9 (5):180, May 15, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/180.

Carral:2022:RDF

[CSR22] Jose M. Carral and María Sáez-Royuela. Replacement of dietary fishmeal by black soldier fly larvae (Hermetia illucens) meal in practical diets for juvenile tench (Tinca tinca). Fishes, 7(6):390, December 15, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/390.

Cohen-Sanchez:2023:IOS

[CSSMV^{+}23] Amanda Cohen-Sánchez, Antoni Gabriel Sánchez-Mairata, José María Valencia, Antonio Box, Samuel Pinya, Silvia Tejada, and Antoni Sureda. Immune and oxidative stress response of the fish Xyrichthys novacula infected with the trematode Ectoparasite scaphanocephalus sp. in the Balearic Islands. Fishes, 8(12): 600, December 06, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/600.

Coccia:2019:CSE

[CSV^{+}19] Elena Coccia, Francesco Siano, Maria Grazia Volpe, Ettore Varricchio, Orhan Tufan Eroldogan, and Marina Paolucci. Chestnut shell extract modulates immune parameters in the rainbow trout Oncorhynchus mykiss. Fishes, 4(1):18, March 12, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/1/18.

Campbell:2023:GDP

[CTTW23] Matthew R. Campbell, Eric D. Tretter, James C. Trainer, and Richard A. Wilkison. Genetic diversity and population structure of Shoshone sculpin Cottus greenei in the Hagerman Valley of South-Central Idaho. Fishes, 8 (1):55, January 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/1/55.

Chu:2021:SLP

[CTY^{+}21] Yuanming Chu, Mingtang Tan, Zhengkai Yi, Zhaoyang Ding, Dazhang Yang, and Jing Xie. Shelf-life prediction of glazed large yellow croaker (Pseudosciaena crocea) during frozen storage based on Arrhenius model and long-short-term memory neural networks model. Fishes, 6(3):39, September 10, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/39.

Castillo-Vargasmachuca:2021:SDS

[CVANRD^{+}21] Sergio G. Castillo-Vargasmachuca, Eugenio Alberto Aragón-Noriega, Guillermo Rodríguez-Domínguez, Leonardo Martínez-Cárdenas, Eulalio Arámbul-Muñoz, and Álvaro J. Burgos Arcos. The standard deviation structure as a new approach to growth analysis in weight and length data of farmed Lutjanus guttatus. Fishes, 6(4): 60, November 12, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/60.

Cao:2021:EAA

[CWMX21] Jie Cao, Qi Wang, Jun Mei, and Jing Xie. Effect of 3-aminobenzoic acid ethyl ester methanesulfonate (MS-222) on quality of marine cultured turbot (Scophthalmus maximus) during simulated transport in water. Fishes, 6(2):20, May 19, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/2/20.

Coburn:2021:APC

[CWP^{+}21] Jessica Coburn, M. Scott Wells, Nicholas B. D. Phelps, T. Gibson Gaylord, and Deborah A. Samac. Acceptance of a protein concentrate from alfalfa (Medicago sativa) by yellow perch (Perca flavescens) fed a formulated diet. Fishes, 6 (2):9, March 25, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/2/9.

Chen:2023:ICS

[CWW^{+}23] Zhongyuan Chen, Ronghua Wang, Ruisong Wang, Naitong Yu, and Pinhong Yang. Identification and characterization of the stimulator of interferon genes (STING) in Chinese giant salamander Andrias davidianus. Fishes, 8(12):597, December 04, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/597.

Chen:2023:IUU

[CXL23] Xidi Chen, Qi Xu, and Lun Li. Illegal, unreported, and unregulated fishing governance in disputed maritime areas: Reflections on the international legal obligations of states. Fishes, 8 (1):36, January 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/36.

Cui:2024:FSR

[CXT^{+}24] Xiaojie Cui, Pengxiang Xu, Tao Tian, Mingyuan Song, Xuyang Qin, Dehua Gong, Yan Wang, Xuguang Zhang, Binbin Xing, Mingzhi Li, and Leiming Yin. Frequencyspecific responses: The impact of an acoustic stimulus on behavioral and physiological indices in large yellow croaker. Fishes, 9(6): 217, June 7, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/217.

Chen:2023:EGB

[CXW^{+}23] Gangfu Chen, Jing Xu, Min Wu, Huatao Li, Qihui Yang, and Lin Feng. Extract of Ginkgo biloba leaves (EGb) decrease lipid oxidation in fish feed and meat and enhance growth and antioxidant capacity in Jian carp (Cyprinus carpio var. Jian). Fishes, 8(11):564, November 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/564.

Cao:2023:BTA

[CYL^{+}23] Zhichen Cao, Denghua Yin, Zhanwei Li, Yan Yan, Peng Zhang, Sigang Zhang, Danqing Lin, Zhong Hua, Jialu Zhang, Congping Ying, Han Zhang, Pao Xu, Guixin Dong, and Kai Liu. Blood transcriptome analysis provides responsive changes in gene expression between Ex Situ and captive Yangtze finless porpoises (Neophocaena asiaeorientalis asiaeorientalis). Fishes, 8(12): 593, November 30, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/593.

Chen:2022:EEE

[CZC^{+}22] Guobin Chen, Tong Zhou, Meng Chen, Guiwei Zou, and Hongwei Liang. Effect of estradiol on estrogen nuclear receptors genes expression on embryonic development stages in Chinese soft-shelled turtle (Pelodiscus sinensis). Fishes, 7(5):223, October 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/5/223.

Chen:2023:YRS

[CZC23] Jun Yu Chen, Chaoshu Zeng, and Jennifer M. Cobcroft. Year-round spawning, filial cannibalism, and embryonic and larval development of the coral reef fish orchid dottyback, Pseudochromis fridmani. Fishes, 8(9): 451, September 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/451.

Cheng:2023:AAI

[CZCW23] Xin Cheng, Fan Zhang, Xinjun Chen, and Jintao Wang. Application of artificial intelligence in the study of fishing vessel behavior. Fishes, 8(10): 516, October 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/516.

Chen:2024:BCP

[CZJ^{+}24] Zhaowen Chen, Yueping Zheng, Kaile Ji, Yueyong Shang, Youji Wang, and Menghong Hu.

Blood-chemistry parameters comparison among different age stages of Chinese sturgeon Acipenser sinensis. Fishes, 9(6):218, June 7, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/218.

Chen:2023:PRP

[CZL23] Lin Chen, Xiaoyu Zhang, and Huanzhang Liu. Phylogenetic relationships of the pseudogobionini group (Teleostei: Cyprinidae) with selection pressure analyses to genes of mitochondrial genome. Fishes, 8 (4):201, April 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/201.

Cai:2023:ERL

[CZW^{+}23] Xin Cai, Yaxing Zhang, Bin Wang, Aijun Cui, Yan Jiang, Zhaojun Meng, Yuting Li, and Yongjiang Xu. Effects of recombinant leptin proteins on the expression of key genes in the HPG axis and liver of tongue sole in vitro. Fishes, 8(12):608, December 16, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/608.

Diaz:2023:EMB

[DÁ23] Patricio A. Díaz and Gonzalo Álvarez. Effects of microalgal blooms on aquaculture and fisheries. Fishes, 8(9): 461, September 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/461.

deAguiar:2023:EDP

[dAdSCC^{+}23] Gustavo Augusto Carvalho Costacurta de Aguiar, Cristiana Leonor da Silva Carneiro, Daniel Abreu Vasconcelos Campelo, Rafael Costa Teixeira Rusth, João Felipe Ribeiro Maciel, Bernardo Baldisserotto, Jener Alexandre Sampaio Zuanon, Alexmiliano Vogel de Oliveira, Maria Goreti de Almeida Oliveira, Mariella Bontempo Duca de Freitas, Wilson Massamitu Furuya, and Ana Lúcia Salaro. Effects of dietary peppermint (Mentha piperita) essential oil on growth performance, plasma biochemistry, digestive enzyme activity, and oxidative stress responses in juvenile Nile tilapia (Oreochromis niloticus). Fishes, 8(7):374, July 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/374.

DIglio:2022:BIA

[DAF^{+}22] Claudio D’Iglio, Marco Albano, Sergio Famulari, Nunziacarla Spanò, Paola Rinelli, Serena Savoca, and Gioele Capillo. Basic intersexuality (abnormal hermaphroditism) in the blackmouth catshark, Galeus melastomus, (Rafinesque, 1810), from the Southern Tyrrhenian Sea (Central Mediterranean Sea). Fishes, 7 (3):120, May 28, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/120.

Degani:2019:VBG

[DAHM19] Gad Degani, Amir Alon, Akram Hajouj, and Ari Meerson. Vitellogenesis in blue gourami is accompanied by brain transcriptome changes. Fishes, 4(4):54, October 29, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/4/54.

Montag:2023:FRS

[dAMPS^{+}23] Luciano Fogaça de Assis Montag, Luiz Antônio Wanderley Peixoto, Lidia Brasil Seabra, Liziane Amaral Barbosa Gonçalves, Cleonice Maria Cardoso Lobato, Marina Barreira Mendonça, Tiago Octavio Begot, Erival Gonçalves Prata, and Tiago Magalhães da Silva Freitas. First record of spinal deformity in the South American silver croaker Plagioscion squamosissimus (Eupercaria: Sciaenidae) in the Xingu River, Brazil. Fishes, 8 (7):363, July 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/363.

Porto:2024:EPF

[dAPAA^{+}24] Lívia de A. Porto, Yhago P. A. S. Assis, Matheus P. S. Amorim, Paulo E. C. M. de Oliveira, Alessandro L. Paschoalini, Nilo Bazzoli, Ronald K. Luz, and Gisele C. Favero. Effects of prolonged fasting and refeeding on metabolic, physiological, tissue, and growth performance adjustments in Colossoma macropomum. Fishes, 9(2): 71, February 10, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/71.

Dam:2020:AFR

[DBP^{+}20] Chinh Thi My Dam, Mark Booth, Igor Pirozzi, Michael Salini, Richard Smullen, Tomer Ventura, and Abigail Elizur. Alternative feed raw materials modulate intestinal microbiota and its relationship with digestibility in yellowtail kingfish Seriola lalandi. Fishes, 5(2):14, May 11, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/2/14.

Damir:2024:WFM

[DCD^{+}24] Nicoleta Damir, Valentina Coatu, Diana Danilov, Luminita Lazăr, and Andra Oros. From waters to fish: a multi-faceted analysis of contaminants’ pollution sources, distribution patterns, and ecological and human health consequences. Fishes, 9 (7):274, July 11, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/7/274.

Dara:2023:FWA

[DCL^{+}23a] Mariano Dara, Pierluigi Carbonara, Claudia La Corte, Daniela Parrinello, Matteo Cammarata, and Maria Giovanna Parisi. Fish welfare in aquaculture: Physiological and immunological activities for diets, social and spatial stress on Mediterranean aqua cultured species. Fishes, 8(8): 414, August 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/414.

Dezfuli:2023:RCP

[DCL^{+}23b] Bahram Sayyaf Dezfuli, Giuseppe Castaldelli, Massimo Lorenzoni, Antonella Carosi, Mykola Ovcharenko, and Giampaolo Bosi. Rodlet cells provide first line of defense against swimbladder nematode and intestinal coccidian in Anguilla anguilla. Fishes, 8(2):66, January 21, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/2/66.

Pinheiro:2024:RFS

[dCPFS^{+}24] Larissa da Costa Pinheiro, Beatrice Padovani Ferreira, Paulo Emílio Santos, Ualerson Peixoto, and Bianca Bentes. Recent findings on spawning patterns and recommendations for the fishery management of the southern red snapper — Lutjanus purpureus (Poey, 1866) — on the Amazon Continental Shelf of Brazil. Fishes, 9 (4):136, April 15, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/136.

Pires:2021:MTE

[dCPRG^{+}21] Luana da Costa Pires, Patricia Rodrigues, Quelen Iane Garlet, Luisa Barichello Barbosa, Bibiana Petri da Silveira, Guerino Bandeira Junior, Lenise de Lima Silva, Amanda Gindri, Rodrigo Coldebella, Cristiane Pedrazzi, Agueda Palmira Castagna de Vargas, Bernardo Baldisserotto, and Berta Maria Heinzmann. Maclura tinctoria extracts: In vitro antibacterial activity against Aeromonas hydrophila and sedative effect in Rhamdia quelen. Fishes, 6(3):25, July 27, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/3/25.

Dayrit:2023:PIS

[DCR^{+}23] Geraldine B. Dayrit, Emmanuel M. Vera Cruz, Channarong Rodkhum, Mahmoud Mabrok, Pattareeya Ponza, and Mudjekeewis D. Santos. Potential influence of shading in freshwater ponds on the water quality parameters and the hematological and biochemical profiles of Nile tilapia (Oreochromis niloticus Linnaeus, 1758). Fishes, 8 (6):322, June 17, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/322.

Dvoretsky:2023:SBO

[DD23] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Shellfish as biosensors in online monitoring of aquatic ecosystems: a review of Russian studies. Fishes, 8(2):102, February 08, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/102.

Danilov:2023:SOP

[DDD^{+}23] Diana Danilov, Lorena Dediu, Nicoleta Alexandra Damir, Valentina Coatu, and Luminita Lazar. Screening for organic pollutants in the Black Sea turbot (Scophthalmus maeoticus). Fishes, 8(5):265, May 17, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/265.

Durazzo:2022:NBC

[DDG^{+}22]  Alessandra  Durazzo,

Gabriella Di Lena, Paolo Gabrielli, Antonello Santini, Ginevra Lombardi-Boccia, and Massimo Lucarini. Nutrients and bioactive compounds in seafood: Quantitative literature research analysis. Fishes, 7(3):132, June 05, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/132.

Diken:2019:IEP

[DDN19] Gürkan Diken, Orhan Demir, and Mehmet Naz. The inhibitory effects and positive contributions of live foods on protease activities of meagre, Argyrosomus regius (Asso 1801), larvae in vitro assay. Fishes, 4(1):5, February 04, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/1/5.

Goncalves-de-Freitas:2019:SBW

[dFBdSG^{+}19] Eliane Gonçalves de Freitas, Marcela Cesar Bolognesi, Ana Carolina dos Santos Gauy, Manuela Lombardi Brandão, Percilia Cardoso Giaquinto, and Marisa Fernandes-Castilho. Social behavior and welfare in Nile tilapia. Fishes, 4 (2):23, March 27, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/2/23.

DiPaola:2022:SZE

[DIL^{+}22] Davide Di Paola, Carmelo Iaria, Giovanni Lanteri, Marika Cordaro, Rosalia Crupi, Rosalba Siracusa, Ramona D’Amico, Roberta Fusco, Daniela Impellizzeri, Salvatore Cuzzocrea, Nunziacarla Spanò, Enrico Gugliandolo, and Alessio Filippo Peritore. Sensitivity of zebrafish embryogenesis to risk of fotemustine exposure. Fishes, 7(2): 67, April 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/2/67.

Dimitriadi:2022:EET

[DK22] Anastasia Dimitriadi and George Koumoundouros. Elevated embryonic temperature has persistent adverse effects on zebrafish swimming capacity. Fishes, 7(6): 373, December 06, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/373.

Dance:2023:LCC

[DL23] Michael A. Dance and Mitchell S. Lovell. Lipid correction for carbon stable isotope analysis of yellowfin tuna. Fishes, 8(9):446, September 01, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/446.

laCruz:2022:LDT

[dlCRMB^{+}22] Simrith E. Cordova de la Cruz, Marta F. Riesco, Gil Martínez-Bautista, Daniel Calzada-Ruiz, Talhia Martínez-Burguete, Emyr S. Peña-Marín, Carlos Alfonso Álvarez-Gonzalez, and Ignacio Fernández. Larval development in tropical gar (Atractosteus tropicus) is dependent on the embryonic thermal regime: Ecological implications under a climate change context. Fishes, 7(1): 16, January 11, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/16.

Dai:2022:FRG

[DLL^{+}22] Chen Dai, Xiaoxin Li, Dapeng Luo, Qingming Liu, Yun Sun, Zhigang Tu, and Minghui Shen. First report on genome analysis and pathogenicity of Vibrio tubiashii FP17 from farmed ivory shell (Babylonia areolata). Fishes, 7(6):396, December 17, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/396.

Devergne:2023:ILT

[DLL^{+}23] Jimmy Devergne, Véronique Loizeau, Christophe Lebigre, Anne Bado-Nilles, Sophie Collet, Olivier Mouchel, Ugo Iaria, Marie-Madeleine Le Gall, Lauriane Madec, Cyril Turiès, and Arianna Servili. Impacts of long-term exposure to ocean acidification and warming on three-spined stickleback (Gasterosteus aculeatus) growth and reproduction. Fishes, 8(10): 523, October 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/523.

Degani:2024:SDB

[DM24] Gad Degani and Ari Meerson. Sex differences in the brain transcriptomes of adult blue gourami fish (Trichogaster trichopterus). Fishes, 9(7):287, July 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/7/287.

Denis:2022:AGE

[DMA22] Jérémy Denis, Kélig Mahé, and Rachid Amara. Abundance and growth of the European eels (Anguilla anguilla Linnaeus, 1758) in small estuarine habitats from the Eastern English Channel. Fishes, 7(5): 213, October 23, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/5/213.

Dunker:2020:DRA

[DMB^{+}20] Kristine Dunker, Robert Massengill, Parker Bradley, Cody Jacobson, Nicole Swenson, Andy Wizik, and Robert DeCino. A decade in review: Alaska’s adaptive management of an invasive apex predator. Fishes, 5(2):12, April 21, 2020. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/2/12.

Debenedetti:2019:PRA

[DMT^{+}19] Ángela L. Debenedetti, Elena Madrid, María Trelis, Francisco J. Codes, Florimar Gil-Gómez, Sandra Sáez-Durán, and Màrius V. Fuentes. Prevalence and risk of anisakid larvae in fresh fish frequently consumed in Spain: an overview. Fishes, 4(1): 13, February 21, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/13.

Duchatelet:2023:ECP

[DNP^{+}23] Laurent Duchatelet, Charlotte Nuyt, Nathan Puozzo, Jérôme Mallefet, and Jérôme Delroisse. Evolutionary conservation of photophore ultrastructure in sharks: The case of a dalatiid squalomorph. Fishes, 8(2): 87, February 01, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/87.

Dube:2024:AAP

[DO24] Edith Dube and Grace Emily Okuthe. Applications of antimicrobial photodynamic therapy in aquaculture: Effect on fish pathogenic bacteria. Fishes, 9(3):99, March 05, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/99.

Dedman:2017:ASM

[DOB^{+}17] Simon Dedman, Rick Officer, Deirdre Brophy, Maurice Clarke, and David G. Reid. Advanced spatial modeling to inform management of data-poor juvenile and adult female rays. Fishes, 2(3):12, August 04, 2017. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/3/12.

Costa:2023:BRO

[dOCCH^{+}23] Léa Carolina de Oliveira Costa, Andrezza Carvalho, Mariana Holanda, Jorge Santos, Lucélia Borges, Bruna Guterres, Je Nam Junior, Virginia Fonseca, Larissa Muller, Luis Romano, Silvia Botelho, Marcelo Pias, Juliane Ventura, and Luís H. Poersch. Biological responses of oyster Crassostrea gasar exposed to different concentrations of biofloc. Fishes, 8(12):586, November 29, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/586.

Dominguez-Petit:2022:PER

[DPGFL^{+}22] Rosario Domínguez-Petit, Cristina García-Fernández, Ezequiel Leonarduzzi, Karina Rodrigues, and Gustavo Javier Macchi. Parental effects and reproductive potential of fish and marine invertebrates: Crossgenerational impact of environmental experiences. Fishes, 7(4):188, July 27, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/188.

Dambrosio:2023:SLQ

[DQC^{+}23] Angela Dambrosio, Nicoletta Cristiana Quaglia, Maria Antonietta Colonna, Flavia Capuozzo, Francesco Giannico, Simona Tarricone, Anna Caputi Jambrenghi, and Marco Ragni. Shelf-life and quality of anchovies (Engraulis encrasicolus) refrigerated using different packaging materials. Fishes, 8(5):268, May 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/268.

Castaneda-Chavez:2021:CMN

[dRCCLRNRWK21] María del Refugio Castañeda-Chávez, Fabiola Lango-Reynoso, Gabycarmen Navarrete-Rodríguez, and Armando Toyokazu Wakida-Kusunoki. Concentration of metals in native and invasive species of fish in the fluviallagoon-deltaic system of the Palizada River, Campeche. Fishes, 6(4): 72, December 03, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/4/72.

David-Ruales:2023:ZPE

[DRFCL23] Carlos David-Ruales, Débora Machado Fracalossi, and Felipe Collazos-Lasso. Zootechnical parameters and enzyme activity in the species Brycon moorei (Steindachner 1878). Fishes, 8(12):592, November 30, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/592.

Das:2018:EFA

[DRH18] Mousumi Das, Ferdous Ibn Rahim, and Md. Amzad Hossain. Evaluation of fresh Azolla pinnata as a lowcost supplemental feed for Thai silver barb Barbonymus gonionotus. Fishes, 3(1):15, March 05, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/15.

Deakin:2019:WCI

[DSC^{+}19] Anthony G. Deakin, Joseph W. Spencer, Andrew R. Cossins, Iain S. Young, and Lynne U.

Sneddon. Welfare challenges influence the complexity of movement: Fractal analysis of behaviour in zebrafish. Fishes, 4(1):8, February 07, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/8.

Du:2023:EMF

[DSC^{+}23] Ningning Du, Yanchun Sun, Zhongxiang Chen, Xiaoli Huang, Chenhui Li, Lei Gao, Shuyan Bai, Peng Wang, and Qirui Hao. Effects of multiple freeze–thaw cycles on protein and lipid oxidation, microstructure and quality characteristics of rainbow trout (Oncorhynchus mykiss). Fishes, 8(2):108, February 13, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/108.

Furtado:2023:HCL

[dSCFQB^{+}23] Maura da Silva Costa Furtado, Joaquim Carlos Barbosa Queiroz, Bianca Bentes, Edson Koiti Kudo Yasojima, Débora de Oliveira Thomaz, Larissa da Costa Pinheiro, Mauro Luis Ruffino, and Victoria Isaac. The hydrological cycle of the Lower Amazon in Brazil determines the variation in local fishing patterns. Fishes, 8(7):371, July 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/371.

Gauy:2023:BTS

[dSGBdF23] Ana Carolina dos Santos Gauy, Marcela Cesar Bolognesi, and Eliane Gonçalves de Freitas. Body tactile stimulation reduces the effects of high stocking density on the welfare of Nile tilapia (Oreochromis niloticus). Fishes, 8(6):320, June 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/6/320.

daSilva:2021:VAA

[dSJC^{+}21] Elisia Gomes da Silva, Guerino Bandeira Junior, Juliana Felipetto Cargnelutti, Roberto Christ Vianna Santos, André Gündel, and Bernardo Baldisserotto. In vitro antimicrobial and antibiofilm activity of S-(-)-limonene and R-(+)-limonene against fish bacteria. Fishes, 6 (3):32, August 19, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/3/32.

daSilveira:2023:IAM

[dSSBdS23] Estevan Luiz da Silveira, Nabil Semmar, Eduardo Luis Cupertino Ballester, and André Martins Vaz dos Santos. Integrative analysis to manage aquatic resources based on fish feeding patterns in neotropical rivers. Fishes, 8(3): 157, March 06, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/157.

Doring:2017:EFB

[DTS^{+}17] Julian Döring, Maik Tiedemann, Moritz Stäbler, Hans Sloterdijk, and Werner Ekau. Ethmalosa fimbriata (Bowdich 1825), a clupeid fish that exhibits elevated batch fecundity in hypersaline waters. Fishes, 2(3): 13, August 18, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/3/13.

DulceEstevao:2023:APR

[Dul23] Maria Dulce Estêvão.

Aquatic pollutants: Risks, consequences, possible solutions and novel testing approaches. Fishes, 8(2):97, February 06, 2023. CODEN ???? ISSN 2410-3888. URL

                                        https://www.mdpi.com/
                                        2410-3888/8/2/97.

Ding:2022:GMN

[DZC^{+}22] Liyun Ding, Yanping Zhang, Jiacheng Chen, Wenjing Chen, Shouqi Xie, and Qingtang Chen. Growth, muscle nutrition composition, and digestive enzyme activities of the juvenile and adult Siniperca chuatsi fed on live baits and a formulated diet. Fishes, 7(6):379, December 08, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/379.

Esteves:2018:OPC

[EA18] Eduardo Esteves and Jaime Aníbal. Optimization of processing conditions of traditional cured tuna loins — muxama. Fishes, 3 (1):3, January 09, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/3.

Eissa:2023:PSE

[EAE^{+}23] El-Sayed Hemdan Eissa, Ragaa A. Ahmed, Nadia A. Abd Elghany, Amal Elfeky, Saadea Saadony, Norhan H. Ahmed, Salah El-Sayed Sakr, Geraldine B. Dayrit, Charlene Princess S. Tolenada, Adlene Anne C. Atienza, Mahmoud Mabrok, and Hala F. Ayoub. Potential symbiotic effects of β-1,3 glucan, and fructooligosaccharides on the growth performance, immune response, redox status, and resistance of Pacific white shrimp, Litopenaeus vannamei to Fusarium solani infection. Fishes, 8(2): 105, February 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/105.

Eissa:2023:DEN

[EAJ^{+}23] El-Sayed Hemdan Eissa, Bothaina A. Alaidaroos, Samyah D. Jastaniah, Mohammad Bodrul Munir, Manal E. Shafi, Yasmin M. Abd El-Aziz, Walaa K. Bazina, Saadiah binti Ibrahim, Moaheda E. H. Eissa, Marina Paolucci, Fatima S. Alaryani, Nadia N. B. Abd El-Hamed, Mohamed E. Abd El-Hack, and Saadea Saadony. Dietary effects of nano curcumin on growth performances, body composition, blood parameters and histopathological alternation in red tilapia (Oreochromis sp.) challenged with Aspergillus flavus. Fishes, 8(4):208, April 17, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/208.

Erlandsson:2024:UQI

[EAJG24] Ann Erlandsson, Giedrė Ašmonaitė, Bror Jonsson, and Larry Greenberg. Using qPCR to identify potential effects of thermal conditions during embryogenesis on mitochondrial DNA copy number in juvenile brown trout Salmo trutta. Fishes, 9 (4):142, April 19, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/142.

Eliso:2024:PMN

[EBCM24] Maria Concetta Eliso, Barbara Billè, Tiziana Cappello, and Maria Maisano. Polystyrene micro- and nanoplastics (PS MNPs): a review of recent advances in the use of -omics in PS NP toxicity studies on aquatic organisms. Fishes, 9(3):98, March 05, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/98.

Esteban:2021:FRF

[EBH21] Maria Angeles Esteban, Bernardo Baldisserotto, and Eric Hallerman. Fishes receives its first impact factor. Fishes, 6 (3):29, August 11, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/29.

Emadi:2024:SPA

[EBM24] Cameron M. Emadi, Preston T. Bean, and Edward M. Mager. Swimming performance assessments of fish species of greatest conservation need to inform future stream crossing designs in Texas. Fishes, 9 (6):234, June 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/234.

Easton:2023:CSO

[EBRS23] Blair Alexander Andrew Easton, Andrew Boon, Joe Richards, and Kevin Scott. Comparing the size at onset of sexual maturity of edible crab (Cancer pagurus, Cancridae) in Berwickshire and Northumberland. Fishes, 8(5):260, May 13, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/260.

Esparza-Espinoza:2023:CSI

[EEdCSOPJ^{+}23] Dania Marisol Esparza-Espinoza, Hisila del Carmen Santacruz-Ortega, Maribel Plascencia-Jatomea, Santiago P. Aubourg, Jesús Aarón Salazar-Leyva, Francisco Rodríguez-Felix, and Josafat Marina Ezquerra-Brauer. Chemical-structural identification of crude gelatin from jellyfish (Stomolophus meleagris) and evaluation of its potential biological activity. Fishes, 8(5):246, May 08, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/246.

Shal:2021:DFA

[EEE21] Ahmed Mohamed El Shal, Faisal Mohamed El Sheikh, and Atef Mohamed Elsbaay. Design and fabrication of an automatic fish feeder prototype suits tilapia tanks. Fishes, 6(4): 74, December 06, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/74.

Elbahnaswy:2023:BCI

[EES^{+}23] Samia Elbahnaswy, Gehad E. Elshopakey, Medhat S.

Shakweer, Elsayed A. A. Eldessouki, Abdelwahab A. Abdelwarith, Elsayed M. Younis, Simon J. Davies, and Mai A. M. El-Son. Bacterial co-infection as a potential threat to farmed flathead grey mullet (Mugil cephalus): Phenotypic and molecular diagnosis, histopathology, immunity response, and in vitro antibacterial evaluation. Fishes, 8(7):357, July 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/357.

Esposito:2023:MCS

[EFG^{+}23] Anaïs Esposito, Jean-José Filippi, Charlotte Gerbaud, Quentin Godeaux, Rémi Millot, Paul-Jean Agostini, Camille Albertini, Eric Durieux, Joséphine Foata, and Yann Quilichini. Macroparasite communities with special attention to invasive helminths in European eels Anguilla anguilla from freshwaters and brackish lagoons of a Mediterranean island. Fishes, 8(7):375, July 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/375.

Esposito:2023:PHF

[EFQ23] Anaïs Esposito, Joséphine Foata, and Yann Quilichini. Parasitic helminths and freshwater fish introduction in Europe: a systematic review of dynamic interactions. Fishes, 8(9): 450, September 07, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/450.

Eckelbecker:2023:RCP

[EHGS23] Robert W. Eckelbecker, Nathaniel M. Heili, Christopher S. Guy, and David A. Schmetterling. Relative condition parameters for fishes of Montana, USA. Fishes, 8(1):28, December 31, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/1/28.

Ellis:2023:SMA

[EKL^{+}23] Robert D. Ellis, Christopher C. Koenig, James V. Locascio, Christopher R. Malinowski, and Felicia C. Coleman. Spawning migrations of the Atlantic goliath grouper along the Florida Atlantic Coast. Fishes, 8(8):398, August 01, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/8/398.

Eguiraun:2023:NLA

[EM23] Harkaitz Eguiraun and Iciar Martinez. Nonlinear analyses of fish behaviours in response to aquatic environmental pollutants — a review. Fishes, 8(6):311, June 12, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/311.

Esquivel-Muelbert:2018:AVR

[EMFZ^{+}18] Juan R. Esquivel-Muelbert, Luisa Fontoura, Éverton Zardo, Danilo P. Streit, Jr., Adriane Esquivel-Muelbert, and Juan R. E. Garcia. Assessing the viability of reintroduction of locally extinct migratory fish Brycon orbignyanus: Successful growth, dispersal and maturation. Fishes, 3 (4):39, October 01, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/4/39.

Eegunjobi:2022:DGV

[EN22] Ruth Eegunjobi and Nicholas Ngepah. The determinants of global value chain participation in developing seafood-exporting countries. Fishes, 7(4):186, July 26, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/186.

Espinola-Novelo:2021:STV

[ENO21] Juan F. Espínola-Novelo and Marcelo E. Oliva. Spatial and temporal variability of parasite communities: Implications for fish stock identification. Fishes, 6(4): 71, December 03, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/4/71.

Enyidi:2017:CVP

[Eny17] Uchechukwu D. Enyidi. Chlorella vulgaris as protein source in the diets of African catfish Clarias gariepinus. Fishes, 2(4):17, October 16, 2017. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/2/4/17.

Enyidi:2017:EFP

[EPKV17] Uchechukwu D. Enyidi, Juhani Pirhonen, Juhani Kettunen, and Jouni Vielma. Effect of feed protein:lipid ratio on growth parameters of African catfish Clarias gariepinus after fish meal substitution in the diet with bambaranut (Voandzeia subterranea) meal and soybean (Glycine max) meal. Fishes, 2(1):1, January 30, 2017. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/1/1.

Espinosa-Ruiz:2021:WIC

[ERE21] Cristóbal Espinosa-Ruíz and María Ángeles Esteban. Wound-induced changes in antioxidant enzyme activities in skin mucus and in gene expression in the skin of gilthead seabream (Sparus aurata L.). Fishes, 6(2):15, April 18, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/2/15.

Esteban:2016:WNJ

[Est16] Maria Angeles Esteban.

Welcome to the new journal Fishes. Fishes, 1 (1):16–17, May 31, 2016. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/1/1/16.

Engle:2023:RCB

[EvSCB23] Carole R. Engle, Jonathan van Senten, Charles Clark, and Noah Boldt. Has the regulatory compliance burden reduced competitiveness of the U.S. tilapia industry? Fishes, 8(3):151, March 02, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/3/151.

Filipsson:2023:WBJ

[FÅG^{+}23] Karl Filipsson, Veronika Åsman, Larry Greenberg, Martin Österling, Johan Watz, and Eva Bergman. Winter behavior of juvenile brown trout in a changing climate: How do light and ice cover affect encounters with instream predators? Fishes, 8 (10):521, October 20, 2023. CODEN ???? ISSN 2410-3888. URL

                                        https://www.mdpi.com/
                                        2410-3888/8/10/521.

Ferri:2022:AGU

[FB22] Josipa Ferri and An– dela Brzica. Age, growth, and utility of otolith morphometrics as predictors of age in the European barracuda, Sphyraena sphyraena (Sphyraenidae): Preliminary data. Fishes, 7 (2):68, April 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/2/68.

Frank:2024:LIP

[FBSB24] Colin Frank, Shannon Bayse, Rioghnach Steiner, and Pierre-Paul Bitton. Light intensity of phosphorescentnetting pots and determining their visibility to snow crab (Chionoecetes opilio) using visual modeling techniques. Fishes, 9(5):185, May 17, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/185.

Fu:2021:IEC

[FCB^{+}21] Lulu Fu, Qiudie Chi, Yongbo Bao, Hanhan Yao, Zhihua Lin, and Yinghui Dong. Identification and expression characterization of the Smad3 gene and SNPs associated with growth traits in the hard clam (Meretrix meretrix). Fishes, 6(4):83, December 16, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/83.

Fife-Cook:2019:PWF

[FCF19] Isabel Fife-Cook and Becca Franks. Positive welfare for fishes: Rationale and areas for future study. Fishes, 4 (2):31, May 23, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/2/31.

Franklin:2019:SKR

[FCT19] Maxwell Franklin, Martin Cenek, and E. Jamie Trammell. Studying Kenai River fisheries’ social-ecological drivers using a holistic fisheries agent-based model: Implications for policy and adaptive capacity. Fishes, 4(2):33, May 31, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/2/33.

Ferreira:2023:CLL

[FDM^{+}23a] Inês Ferreira, Felippe A. Daros, Cláudia Moreira, Diana Feijó, Alberto Rocha, Ana Mendez-Vicente, Jorge Pisonero, and Alberto Teodorico Correia. Is Chelidonichthys lucerna (Linnaeus, 1758) a marine estuarine-dependent fish? Insights from saccular otolith microchemistry. Fishes, 8(7):383, July 24, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/383.

Filipe:2023:SSF

[FDM^{+}23b] Diogo Filipe, Mário Dias, Rui Magalhães, Helena Fernandes, José Salgado, Isabel Belo, Aires Oliva-Teles, and Helena Peres. Solidstate fermentation of distiller’s dried grains with solubles improves digestibility for European seabass (Dicentrarchus labrax) juveniles. Fishes, 8(2):90, February 03, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/90.

Fedder:2023:SDM

[Fed23] Jens Fedder. Sex determination and male differentiation in southern swordtail fishes: Evaluation from an evolutionary perspective. Fishes, 8(8):407, August 06, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/407.

Ferri:2023:OTA

[Fer23] Josipa Ferri. Otoliths and their applications in fishery science. Fishes, 8 (1):35, January 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/35.

Frid:2023:CES

[FGBA^{+}23] Ori Frid, Tal Gavriel, Yigael Ben-Ari, Adi Weinberger, Hagar Yancovich-Shalom, and Jonathan Belmaker. Catch estimates and species composition of recreational fishing in Israel. Fishes, 8(2):69, January 23, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/69.

Falsone:2022:ASD

[FGG^{+}22] Fabio Falsone, Vita Gancitano, Michele Luca Geraci, Giacomo Sardo, Danilo Scannella, Fabrizio Serena, Sergio Vitale, and Fabio Fiorentino. Assessing the stock dynamics of Elasmobranchii off the Southern Coast of Sicily by using trawl survey data. Fishes, 7(3):136, June 07, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/136.

Fuentes:2023:PGC

[FGHYCA23] Sandra Ferrada Fuentes, Ricardo Galleguillos, Victoria Herrera-Yáñez, and Cristian B. Canales-Aguirre. Population genetics of Chilean jack mackerel, Trachurus murphyi Nichols, 1920, (Pisces, Carangidae), in waters of the South Pacific Ocean. Fishes, 8 (3):162, March 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/162.

Ferreira:2019:DEC

[FGR19] Marco Ferreira, João Gago, and Filipe Ribeiro. Diet of European catfish in a newly invaded region. Fishes, 4(4): 58, December 08, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/58.

Fraser:2023:EGG

[FHF23] Thomas W. K. Fraser, Tom J. Hansen, and Per Gunnar Fjelldal. Environmental and genetic (vgll3) effects on the prevalence of male maturation phenotypes in domesticated Atlantic salmon. Fishes, 8(5): 275, May 22, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/275.

Fantini-Hoag:2023:BAP

[FHHC23] Leticia Fantini-Hoag, Terry Hanson, and Jesse Chappell. Bioeconomic analysis of in-pond raceway system production of foodsize and stocker hybrid catfish (channel catfish Ictalurus punctatus [female-sign]× blue catfish, I. furcatus [malesign]). Fishes, 8(2):

96, February 06, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/96.

Finnas:2023:POM

[FJL^{+}23] Viktor Finnäs, Erkki Jokikokko, Jan-Olof Lill, Yann Lahaye, Henry Hägerstrand, and Kai Lindström. Potential of otolith microchemistry to distinguish nursery areas of salmon within River Simojoki. Fishes, 8(6):332, June 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/332.

Fronte:2021:FRH

[FLB^{+}21] Baldassare Fronte, Rosario Licitra, Carlo Bibbiani, Lucia Casini, Mahanama De Zoysa, Vincenzo Miragliotta, Simona Sagona, Francesca Coppola, Letizia Brogi, and Francesca Abramo. Fishmeal replacement with Hermetia illucens meal in aquafeeds: Effects on zebrafish growth performances, intestinal morphometry, and enzymology. Fishes, 6(3): 28, August 05, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/28.

Fitzpatrick:2023:EPF

[FLFF23] Ryan M. Fitzpatrick, David W. Longrie, Ryan J. Friebertshauser, and H. Paul Foutz. Evaluation of a prefabricated fish passage design for Great Plains fishes. Fishes, 8(8): 403, August 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/403.

Fu:2022:CTH

[FLR^{+}22] Qianni Fu, Jinxia Liu, Tianjiao Ren, Zining Zhang, Zihang Ma, Zhenyu Lan, Yitao Duan, Ziwei Liang, Boyu Chen, Yan Zhang, Peng Zhu, and Yongyan Liao. Cloning of two HSP genes of Eriocheir hepuensis and their expression under Vibrio parahaemolyticus stress. Fishes, 7(6):372, December 05, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/372.

Fang:2022:CAB

[FLX^{+}22] Di-An Fang, Kai Liu, Dong-Po Xu, Yin-Ping Wang, and Pao Xu. Comparative analysis of blood transcriptome in the Yangtze finless porpoise (Neophocaena asiaeorientalis). Fishes, 7 (2):61, April 10, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/2/61.

Faruque:2021:CSA

[FM21a] Hasan Faruque and Hiroyuki Matsuda. Conservative scoring approach in productivity susceptibility analysis leads to an overestimation of vulnerability: a study from the hilsa gillnet bycatch stocks of Bangladesh. Fishes, 6 (3):33, August 24, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/3/33.

Flinn:2021:TGM

[FM21b] Shane A. Flinn and Stephen R. Midway. Trends in growth modeling in fisheries science. Fishes, 6(1):1, January 19, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/1/1.

Fernandez-Muela:2023:EFI

[FMBPC^{+}23] Montserrat Fernández-Muela, Rubén Bermejo-Poza, Almudena Cabezas, Concepción Pérez, Elisabet González de Chavarri, María Teresa Díaz, Fernando Torrent, Morris Villarroel, and Jesús De la Fuente. Effects of fasting on intermediary metabolism enzymes in the liver and muscle of rainbow trout. Fishes, 8 (1):53, January 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/53.

Fang:2023:ETS

[FMXQ23] Dan Fang, Jun Mei, Jing Xie, and Weiqiang Qiu. The effects of transport stress (temperature and vibration) on blood biochemical parameters, oxidative stress, and gill histomorphology of pearl gentian groupers. Fishes, 8 (4):218, April 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/218.

Ferragut-Perello:2023:ECS

[FPRAT^{+}23] Francesca Ferragut-Perello, Sergio Ramírez-Amaro, Athanassios C. Tsikliras, Natalia Petit-Marty, Donna Dimarchopoulou, Enric Massutí, Alba Serrat, and Francesc Ordines. Exploitation and conservation status of the thornback ray (Raja clavata) in the Balearic Islands (Western Mediterranean). Fishes, 8(2):

117, February 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/117.

Frias-Quintana:2017:UPS

[FQÁGTR^{+}17] Carlos A. Frías-Quintana, Carlos A. Álvarez-González, Dariel Tovar-Ramírez, Rafael Martínez-García, Susana Camarillo-Coop, Emyr Peña, and Mario A. Galaviz. Use of potato starch in diets of tropical gar (Atractosteus tropicus, gill 1863) larvae. Fishes, 2(1):3, March 07, 2017. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/2/1/3.

Franca:2023:CFR

[Fra23] Susana França. Changes in the functional role of the Tejo Estuary (Portugal, Europe) according to fish ecological guilds. Fishes, 8(11): 545, November 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/11/545.

Fedorcak:2023:TCG

[FŠS^{+}23] Jakub Fedorčák, Radek Šanda, Tihomir Stefanov, Jan Mendel, Michal Nowak, Peter Križek, Anabel Perdices, Jasna Vukić, and Ján Koščo. The “True colours” of golden loaches (Teleostei: Cobitidae). Fishes, 8(2): 119, February 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/119.

Fitzpatrick:2021:UID

[FWJ21] Ryan M. Fitzpatrick, Dana L. Winkelman, and Brett M. Johnson. Using isotopic data to evaluate Esox lucius (Linnaeus, 1758) natal origins in a hydrologically complex river basin. Fishes, 6(4): 67, November 22, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/67.

Fei:2023:ESF

[FYH^{+}23] Yingjie Fei, Shenglong Yang, Mengya Huang, Xiaomei Wu, Zhenzhen Yang, Jiangyue Zhao, Fenghua Tang, Wei Fan, and Sanling Yuan. Evaluating suitability of fishing areas for squidjigging vessels in the Northwest Pacific Ocean derived from AIS data. Fishes, 8(10):530, October 23, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/10/530.

Feng:2023:LMB

[FYL^{+}23] Liu-Ying Feng, Li-Ping Yan, Run-Wei Li, Sheng-Fa Li, Jia-Hua Cheng, and Yan Jin. LC-MS based metabolomic profiling of largehead hairtail (Trichiurus japonicus) ovary reveals metabolic signatures of ovarian developmental process (II– IV). Fishes, 8(5):262, May 14, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/5/262.

Fellatami:2024:AGJ

[FZZ^{+}24] Kalysta Fellatami, Wenchao Zhang, Chi Zhang, Shude Liu, and Yongjun Tian. Age and growth of Japanese anchovy (Engraulis japonicus, Temminck & Schlegel, 1846) in coastal waters around Shandong Peninsula, China. Fishes, 9 (4):124, March 30, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/124.

Gaemers:2016:TDE

[Gae16] Pieter A. M. Gaemers. Taxonomy, distribution and evolution of trisopterine Gadidae by means of otoliths and other characteristics. Fishes, 1(1):18– 51, July 17, 2016. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/1/1/18. See cor-

rection [Gae17].

Gaemers:2017:CPM

[Gae17] Pieter A. M. Gaemers.

Correction: Pieter A. M. Gaemers. Taxonomy, Distribution and Evolution of Trisopterine Gadidae by Means of Otoliths and Other Characteristics. Fishes 2016, 1, 18–51. Fishes, 2(3):11, July 17, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/3/11.  See

[Gae16].

Gomes:2024:OFF

[GAR^{+}24] Maria Fernanda S. Gomes, Paulo Henrique R. Aride, Maiko Willas S. Ribeiro, Cristiane C. Guimarães, Tiago C. Nóbrega, Rayana M. Paixão, Adriene Núzia A. Santos, and Adriano Teixeira de Oliveira. Occurrence of four freshwater stingrays (Chondrichthyes: Potamotrygoninae) in the Uatumã River Basin, Amazon Region: a field study. Fishes, 9(5):186, May 17, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/186.

Giovos:2022:UGN

[GASS^{+}22] Ioannis Giovos, Roxani Naasan Aga-Spyridopoulou, Fabrizio Serena, Alen Soldo, Adi Barash, Nikolaos Doumpas, Georgios A. Gkafas, Dimitra Katsada, George Katselis, Periklis Kleitou, Vasileios Minasidis, Yannis P. Papastamatiou, Eleana Touloupaki, and Dimitrios K. Moutopoulos. An updated Greek national checklist of chondrichthyans. Fishes, 7(4):199, August 09, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/199.

Govindharajan:2024:ESM

[GBT^{+}24] Sattanathan Govindharajan, Balamuralikrishnan Balasubramanian, Vivi Thapo, Sournamanikam Venkatalakshmi, and Wen-Chao Liu. Enhancement of skin mucus immunity, carotenoid content, sexual parameters, and growth response in guppy fish (Poecilia reticulata) fed with green algae (Chaetomorpha aerea) diets. Fishes, 9(3):101, March 06, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/101.

Guimaraes:2022:OAP

[GCFA^{+}22] Mateus Cardoso Guimarães, Isabel M. Cerezo, Miguel Frederico Fernandez-Alarcon, Mariene Miyoko Natori, Luciana Yuri Sato, Camila A. T. Kato, Miguel Angel Moriñigo, Silvana Tapia-Paniagua, Danielle de Carla Dias, Carlos Massatoshi Ishikawa, Maria José T. Ranzani-Paiva, Luara Lucena Cassiano, Erna Elisabeth Bach, Patrícia B. Clissa, Daniele P. Orefice, and Leonardo Tachibana. Oral administration of probiotics (Bacillus subtilis and Lactobacillus plantarum) in Nile tilapia (Oreochromis niloticus) vaccinated and challenged with Streptococcus agalactiae. Fishes, 7(4):211, August 22, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/211.

Garcia-Fernandez:2022:UDG

[GFDPSR22] Cristina García-Fernández, Rosario Domínguez-Petit, and Fran Saborido-Rey. The use of daily growth to analyze individual spawning dynamics in an asynchronous population: The case of the European hake from the southern stock. Fishes, 7(4):208, August 18, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/208.

Greco:2022:GEP

[GGF^{+}22] Samuele Greco, Anastasia Serena Gaetano, Gael Furlanis, Francesca Capanni, Chiara Manfrin, Piero Giulio Giulianini, Gianfranco Santovito, Paolo Edomi, Alberto Pallavicini, and Marco Gerdol. Gene expression profiling of Trematomus bernacchii in response to thermal and stabling stress. Fishes, 7(6):387, December 13, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/387.

Gao:2023:TAI

[GGL^{+}23] Jie Gao, Hua-Yang Guo, Ming-Jian Liu, Ke-Cheng Zhu, Bo Liu, Bao-Suo Liu, Nan Zhang, Shi-Gui Jiang, and Dian-Chang Zhang. Transcriptome analysis of the immune process of golden pompano (Trachinotus ovatus) infected with Streptococcus agalactiae. Fishes, 8 (1):52, January 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/52.

Gauthier:2023:TCR

[GGP^{+}23] Nathália Byrro Gauthier, Filipe Scavone Góes, Virgínia Fonseca Pedrosa, Fábio Roselet, Luis Alberto Romano, and Ronaldo Olivera Cavalli. Towards the control of the reproduction of the yellow clam Amarilladesma mactroides (Reeve, 1854) in captivity: Effects of different stimuli on the spawning of laboratoryconditioned and unconditioned breeders. Fishes, 8(1):37, January 04, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/37.

Glencross:2020:ASC

[GHS20] Brett D. Glencross, David Huyben, and Johan W. Schrama. The application of singlecell ingredients in aquaculture feeds — a review. Fishes, 5(3):22, July 16, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/3/22.

Garces:2023:ABT

[GL23] Sara Garcés and Gabriele Lara. Applying biofloc technology in the culture of Mugil cephalus in subtropical conditions: Effects on water quality and growth parameters. Fishes, 8(8):420, August 17, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/8/420.

Gomez:2023:EBC

[GLH^{+}23] Silvia Gómez, Gabriele Lara, Carlos Felipe Hurtado, René Espinoza Alvarado, Jayro Gutiérrez, José Carlos Huechucoy, Guillermo Valenzuela-Olea, and Alice Turner. Evaluating the bioremediation capacity of the polychaete Perinereis gualpensis (Jeldes, 1963) for Atlantic salmon aquaculture sludge. Fishes, 8(8): 417, August 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/8/417.

Guo:2022:DST

[GLW^{+}22] Jingya Guo, Yongming Li, Yaping Wang, Boyong Chen, Yingxin Hu, Yasheng Musha, Xiaoyun He, Tao Tong, and Kunlun Huang. A 90-day subchronic toxicity study of consumption of GH-transgenic triploid carp in Wistar rats. Fishes, 7(1): 10, January 05, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/10.

Gao:2023:RSB

[GLX^{+}23] Tianheng Gao, Nannan Li, Wenlei Xue, Yuning Hu, and Hai Lin. The responses of sediment bacterial communities in Chinese mitten crab (Eriocheir sinensis) culture ponds to changes in physicochemical properties caused by sediment improvement. Fishes, 8(2): 98, February 07, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/98.

Ge:2023:ACI

[GLY^{+}23] Guangyu Ge, Zhihong Liu, Tao Yu, Liqing Zhou, Xiujun Sun, Zhuanzhuan Li, and Yanxin Zheng. Aminotransferase class I and II gene family in the Jinjiang oyster (Crassostrea ariakensis): Genomewide identification, phylogenetic analysis and expression profiles after salinity stress. Fishes, 8(9):459, September 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/459.

Ghahvei:2024:HIN

[GMA^{+}24] Younes Ghahvei, Mohammad Mirzaei, Shahrzad Azizi, Shadi Hashemnia, and Shokoofeh Shamsi. Health insights from nematode larval characterization in greater lizardfish, Saurida tumbil (Bloch, 1795) (Teleostei, Synodontidae). Fishes, 9 (4):143, April 20, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/143.

Garcia-Marquez:2021:ATA

[GMBR^{+}21] Jorge García-Márquez, Andre Barany, Álvaro Broz Ruiz, Benjamín Costas, Salvador Arijo, and Juan Miguel Mancera. Antimicrobial and toxic activity of citronella essential oil (Cymbopogon nardus), and its effect on the growth and metabolism of gilthead seabream (Sparus aurata L.). Fishes, 6(4): 61, November 12, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/61.

Garcia-Marquez:2022:FEA

[GMCF^{+}22] Jorge García-Márquez, Isabel M. Cerezo, Félix L. Figueroa, Roberto Teófilo Abdala-Díaz, and Salvador Arijo. First evaluation of associated gut microbiota in wild thick-lipped grey mullets (Chelon labrosus, Risso 1827). Fishes, 7(4): 209, August 19, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/209.

Garcia-Marquez:2023:PEM

[GMDMT^{+}23] Jorge García-Márquez, Marta Domínguez-Maqueda, Miguel Torres, Isabel M. Cerezo, Eva Ramos, Francisco Javier Alarcón, Juan Miguel Mancera, Juan Antonio Martos-Sitcha, Miguel Ángel Moriñigo, and María Carmen Balebona. Potential effects of microalgaesupplemented diets on the growth, blood parameters, and the activity of the intestinal microbiota in Sparus aurata and Mugil cephalus. Fishes, 8(8):409, August 09, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/8/409.

Garcia-Meilan:2024:OSP

[GMFG^{+}24]  Irene  García-Meilán,

Ramon Fontanillas, Joaquim Gutiérrez, Encarnación Capilla, Isabel Navarro, and Ángeles Gallardo. Oxidative status of the pyloric caeca and proximal intestine in gilthead sea bream fed diets including different vegetable oil blends from palm, rapeseed and linseed. Fishes, 9(6):228, June 14, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/6/228.

Grano-Maldonado:2018:TSU

[GMMNRS18] Mayra I. Grano-Maldonado, Juan Moreno-Navas, and Maria Amparo Rodriguez-Santiago. Transmission strategies used by Gyrodactylus gasterostei (Monogenea) on its host, the three-spined stickleback Gasterosteus aculeatus. Fishes, 3 (2):20, May 29, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/2/20.

Ghosh:2022:FEA

[GMRJ22] Sampat Ghosh, Victor Benno Meyer-Rochow, and Chuleui Jung. Farming the edible aquatic snail Pomacea canaliculata as a mini-livestock. Fishes, 7(1):6, December 28, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/6.

Gaemers:2017:RDT

[GP17] Pieter A. M. Gaemers and Jan Y. Poulsen. Recognition and distribution of two North Atlantic Gadiculus species, G. argenteus and G. thori (Gadidae), based on otolith morphology, larval pigmentation, molecular evidence, morphometrics and meristics. Fishes, 2 (3):15, August 29, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/3/15.

Gough:2023:RVT

[GPD^{+}23] Brendan Gough, Alexandra Prouse, Michael A. Dance, R. J. David Wells, and Jay R. Rooker. Regional variation in the trophic ecology of wahoo (Acanthocybium solandri) in the Western Atlantic Ocean. Fishes, 8(10):519, October 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/10/519.

Greenwood:2017:DSH

[Gre17] M. F. D. Greenwood. Distribution, spread, and habitat predictability of a small, invasive, piscivorous fish in an important estuarine fish nursery. Fishes, 2(2):6, May 13, 2017. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/2/2/6.

Gomez-Requeni:2019:DLC

[GRKC19] Pedro Gómez-Requeni, Mauricio Nestor Kraemer, and Luis Fabián Canosa. The dietary lipid content affects the tissue gene expression of muscle growth biomarkers and the GH/IGF system of pejerrey (Odontesthes bonariensis) juveniles. Fishes, 4(3): 37, June 27, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/3/37.

Garrido:2017:PRL

[GRO^{+}17] Diego Garrido, Diana B. Reis, Diego Orol, Rui A. Gonçalves, M. Virginia Martín, António V. Sykes, Covadonga Rodríguez, Beatriz C. Felipe, Xiaodong Zheng, Luis Lagos, and Eduardo Almansa. Preliminary results on light conditions manipulation in Octopus vulgaris (Cuvier, 1797) paralarval rearing. Fishes, 2(4): 21, November 24, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/4/21.

Guo:2024:EAS

[GSH^{+}24] Kun Guo, Mumin Shi, Xiaoli Huang, Liang Luo, Shihui Wang, Rui Zhang, Wei Xu, Guoliang Ruan, and Zhigang Zhao. The effect of artificial substrate and carbon source addition on bacterial diversity and community composition in water in a pond polyculture system. Fishes, 9(3): 80, February 20, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/80.

Gonzalez-Silvera:2018:EDT

[GSHGE18] Daniel Gonzalez-Silvera, Marcelino Herrera, Inmaculada Giráldez, and María Ángeles Esteban. Effects of the dietary tryptophan and aspartate on the immune response of meagre (Argyrosomus regius) after stress. Fishes, 3 (1):6, January 24, 2018.

CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/6.

Giovos:2021:ILB

[GSK^{+}21] Ioannis Giovos, Fabrizio Serena, Dimitra Katsada, Athanasios Anastasiadis, Adi Barash, Charis Charilaou, Jason M. Hall-Spencer, Fabio Crocetta, Alexander Kaminas, Demetris Kletou, Mary Maximiadi, Vasileios Minasidis, Dimitrios K. Moutopoulos, Roxani Naasan Aga-Spyridopoulou, Ioannis Thasitis, and Periklis Kleitou. Integrating literature, biodiversity databases, and citizenscience to reconstruct the checklist of chondrichthyans in Cyprus (Eastern Mediterranean Sea). Fishes, 6(3):24, July 26, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/3/24.

Goncalves:2017:PIW

[GTC^{+}17] Renata Gonçalves, Maria Alexandra Teodósio, Joana Cruz, Radhouan Ben-Hamadou, Ana Dulce Correia, and Luís Chícharo. Preliminary insight into winter native fish assemblages in Guadiana Estuary salt marshes coping with environmental variability and nonindigenous fish introduction. Fishes, 2(4): 19, October 26, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/4/19.

Goncalves:2018:PEM

[GTSG18] Rui A. Gonçalves, Marco Tarasco, Dian Schatzmayr, and Paulo Gavaia. Preliminary evaluation of moniliformin as a potential threat for teleosts. Fishes, 3(1): 4, January 16, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/1/4.

Guo:2021:CAN

[GWH21] Zhansheng Guo, Zhen Wang, and Xuguang Hou. Comparative analysis of the nrDNA repeat unit of Manila clam Ruditapes philippinarum and quahog Mercenaria mercenaria. Fishes, 6(3):42, September 17, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/42.

Gao:2022:GMF

[GWH^{+}22] Minghui Gao, Zhiqiang Wu, Liangliang Huang, Xichang Tan, Mingsi Li, and Haibo Huang. Growth and microstructural features in otoliths of larval and juvenile Sinogastromyzon wui (F. Balitoridae, River Loaches) of the Upper Pearl River, China. Fishes, 7(2):57, April 01, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/2/57.

Gao:2024:CGB

[GWT^{+}24] Jinwei Gao, Hao Wu, Xing Tian, Jiayu Wu, Min Xie, Zhenzhen Xiong, Dongsheng Ou, Zhonggui Xie, and Rui Song. Calmodulin gene of blunt snout bream (Megalobrama amblycephala): Molecular characterization and differential expression after Aeromonas hydrophila and cadmium challenges. Fishes, 9 (5):182, May 16, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/182.

Guo:2023:IEC

[GYH^{+}23] Ruijie Guo, Kai Yu, Kai Huang, Qiang Lin, and Ting Liu. Immunoprotective effect of Coptis chinensis-supplemented diet on Streptococcus agalactiae infection in tilapia. Fishes, 8(7): 370, July 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/370.

Gu:2022:EEM

[GZF^{+}22] Lingling Gu, Hui Zhang, Guangpeng Feng, Yong Liu, Zhiqiang Han, Feng Zhao, Qing Ye, Wangjiao Hu, and Chao Song. Evaluation of the effectiveness of marking juvenile Takifugu obscurus otoliths with strontium. Fishes, 7(6):371, December 05, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/371.

Guo:2022:CEE

[GZX^{+}22] Kun Guo, Zhigang Zhao, Jun Xie, Liang Luo, Shihui Wang, Rui Zhang, Wei Xu, and Xiaoli Huang. Combined effects of eco-substrate and carbon addition on water quality, fish performance and nutrient budgets in the pond polyculture system. Fishes, 7(5):212, October 23, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/5/212.

Huang:2023:BAH

[HAC^{+}23] Po-Lin Huang, Farok Afero, Yao Chang, Bo-Ying Chen, Hsun-Yu Lan, Yen-Lung Hou, and Cheng-Ting Huang. The bioeconomic analysis of hybrid giant grouper (Epinephelus fuscoguttatus × Epinephelus lanceolatus) and green grouper (Epinephelus malabaricus): a case study in Taiwan. Fishes, 8(12):610, December 17, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/610.

Honji:2024:AVN

[HAdM^{+}24] Renato M. Honji, Bruno C. Araújo, Paulo H. de Mello, Martín R. Ramallo, Leonel Morandini, Danilo Caneppele, and Renata G. Moreira. Argininevasotocin neuronal system in Steindachneridion parahybae (Siluriformes: Pimelodidae) and its influence on artificially induced spawning in captivity. Fishes, 9(6):235, June 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/6/235.

Hallerman:2023:GBS

[Hal23] Eric M. Hallerman. Genetics and biotechnology — the Section Editor-in-Chief’s view. Fishes, 8(2):112, February 15, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/112.

Hamzaoui:2023:IDL

[HARB23a] Mahdi Hamzaoui, Mohamed Ould-Elhassen Aoueileyine, Lamia Romdhani, and Ridha Bouallegue. An improved deep learning model for underwater species recognition in aquaculture. Fishes, 8(10):514, October 16, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/514.

Hamzaoui:2023:OXP

[HARB23b] Mahdi Hamzaoui, Mohamed Ould-Elhassen Aoueileyine, Lamia Romdhani, and Ridha Bouallegue. Optimizing XG-Boost performance for fish weight prediction through parameter preselection. Fishes, 8(10): 505, October 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/505.

Hala:2024:ASF

[HB24] Edmond Hala and Rigers Bakiu. Adriatic Sea fishery product safety and prospectives in relation to climate change. Fishes, 9(5):160, April 28, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/160.

Hayer:2020:UED

[HBG^{+}20] Cari-Ann Hayer, Michael F. Bayless, Amy George, Nathan Thompson, Catherine A. Richter, and Duane C. Chapman. Use of environmental DNA to detect grass carp spawning events. Fishes, 5 (3):27, August 27, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/3/27.

Harris:2022:EAT

[HBL^{+}22] Michaela Harris, Nathalie Brodeur, Francis LeBlanc, Scott Douglas, Paul Chamberland, Thomas Guyondet, Royce Steeves, and Nellie Gagné. eDNA and acoustic tag monitoring reveal congruent overwintering distributions of striped bass in a hydrologically complex estuarine environment. Fishes, 7(4):183, July 23, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/183.

Harris:2020:PGA

[HCWH20] Sheila C. Harris, W. Robert Cope, Isaac Wirgin, and Eric M. Hallerman. Population genetic assessment of anadromous and resident striped bass (Morone saxatilis) in the Roanoke River System, Eastern United States. Fishes, 5(3): 24, August 07, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/3/24.

Hu:2023:EEP

[HCZ^{+}23] Guo Hu, Feng Chen, Ying Zhang, Peixian Luan, Zhiyuan Luo, Jiangong Niu, Peng Zheng, Sai Wang, Tao Zhang, Yongjun Shu, and Feng Ji. Estimates of the effective population size and genetic structure of the critically endangered ship sturgeon (Acipenser nudiventris) in the Chinese section of the Ili River. Fishes, 8(7):354, July 07, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/354.

Hill:2023:NHU

[HDW^{+}23] Jeffrey E. Hill, Allison Durland Donahou, Emily S. Wooley, Lauren N. Lapham, and Quenton M. Tuckett. Narrowing the horizon: Using known invasives and propagule pressure to focus risk screening efforts on potential invasives. Fishes, 8(5): 266, May 17, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/266.

Hoseinifar:2023:EGV

[HFEH^{+}23] Seyed Hossein Hoseinifar, Zohreh Fazelan, Ehab El-Haroun, Morteza Yousefi, Metin Yazici, Hien Van Doan, and Marina Paolucci. The effects of grapevine (Vitis vinifera L.) leaf extract on growth performance, antioxidant status, and immunity of zebrafish (Danio rerio). Fishes, 8 (6):326, June 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/326.

Hasan:2023:CIM

[HGC^{+}23] Imam Hasan, Francesco Gai, Simona Cirrincione, Simona Rimoldi, Giulio Saroglia, and Genciana Terova. Chitinase and insect meal in aquaculture nutrition: a comprehensive overview of the latest achievements. Fishes, 8(12):607, December 14, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/607.

Heidarieh:2023:EID

[HGSE23] Marzieh Heidarieh, Amin Gholamhosseini, Najmeh Sheikhzadeh, and Maria Angeles Esteban. Effects of γ-irradiated date (Phoenix dactylifera) fruit on growth, immunological and antioxidant parameters of goldfish (Carassius auratus). Fishes, 8(5): 251, May 09, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/251.

Heppell:2024:CDC

[HHAG24] Scott A. Heppell, Selina S.

Heppell, N. Scarlett Arbuckle, and M. Brett Gallagher. A crossdecadal change in the fish and crustacean community of Lower Yaquina Bay, Oregon, USA. Fishes, 9(4):125, March 30, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/125.

Hsu:2020:PGA

[HHL^{+}20] Te-Hua Hsu, Chang-Wen Huang, Hung-Tai Lee, Yi-Hsuan Kuo, Kwang-Ming Liu, Cheng-Hui Lin, and Hong-Yi Gong. Population genetic analysis for stock enhancement of silver sea bream (Rhabdosargus sarba) in Taiwan. Fishes, 5(2):19, June 16, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/2/19.

Hasan:2018:RBP

[HHM^{+}18] Tajmira Hasan, Md. Foysul Hossain, Md. Mamun, Md. Jahangir Alam, Mohammad Abdus Salam, and S. M. Rafiquzzaman. Reproductive biology of Puntius sophore in Bangladesh. Fishes, 3(2):22, June 11, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/2/22.

Huang:2024:MAT

[HHM^{+}24] Xianpeng Huang, Zhenlin Hao, Junxia Mao, Luo Wang, Xubo Wang, and Ying Tian. Morphometric analysis of two shell color strains of the bay scallop Argopecten irradians. Fishes, 9 (7):267, July 5, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/267.

Han:2024:ERA

[HHP^{+}24] Qianyun Han, Jiajie Hu, Weicong Pan, Jin Yu, Xiaoguo Ying, Jinpeng Weng, Weiye Li, and Xudong Weng. Effects of rewilding aquaculture time on nutritional quality and flavor characteristics of grass carp (Ctenopharyngodon idellus). Fishes, 9(7):275, July 11, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/275.

Hashizume:2019:HMD

[HIO^{+}19] Atsushi Hashizume, Atsushi Ido, Takashi Ohta, Serigne Thierno Thiaw, Ryusaku Morita, Munenori Nishikawa, Takayuki Takahashi, Chiemi Miura, and Takeshi Miura. Housefly (Musca domestica) larvae preparations after removing the hydrophobic fraction are effective alternatives to fish meal in aquaculture feed for red seabream (Pagrus major). Fishes, 4(3):38, June 27, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/38.

Hamilton:2023:EIS

[HJE^{+}23] Erin F. Hamilton, Collin L. Juurakko, Katja Engel, Josh D. Neufeld, John M. Casselman, Charles W. Greer, and Virginia K. Walker. Environmental impacts on skin microbiomes of sympatric high Arctic salmonids. Fishes, 8 (4):214, April 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/214.

Hussain:2018:OPR

[HKS^{+}18] Md. Ashraf Hussain, Md. Lutful Kabir, Md. Abu Sayeed, A. T. M. Mahbub-E-Elahi, Md. Sultan Ahmed, and Md Jakiul Islam. Organochlorine pesticide residues and microbiological quality assessment of dried barb, Puntius sophore, from the northeastern part of Bangladesh. Fishes, 3(4):44, November 09, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/4/44.

Hicks:2024:EED

[HL24] Grace L. Hicks and Phillip S. Lobel. An examination of the effects of dorsal fin-mounted SPOT on Carcharodon carcharias, the great white shark. Fishes, 9 (6):231, June 17, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/231.

Han:2023:AAE

[HLK^{+}23] Chenglong Han, Kefeng Liu, Toshihisa Kinoshita, Biao Guo, Yifan Zhao, Yuhang Ye, Yufei Liu, Osamu Yamashita, Debin Zheng, Wenhui Wang, and Xueqiang Lu. Assessing the attractive effects of floating artificial reefs and combination reefs on six local marine species. Fishes, 8(5):248, May 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/248.

Huang:2022:DPM

[HLZ^{+}22] Dongyu Huang, Hualiang Liang, Jian Zhu, Mingchun Ren, and Xianping Ge. Dietary protein modifies hepatic glycolipid metabolism, intestinal immune response, and resistance to Streptococcus agalactiae of genetically improved farmed tilapia (GIFT: Oreochromis niloticus) exposed to high temperature. Fishes, 7(4): 202, August 12, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/202.

Hameed:2022:SAI

[HMN^{+}22] Amna Hameed, Waqar Majeed, Muhammad Naveed, Uzma Ramzan, Matteo Bordiga, Maryam Hameed, Saud Ur Rehman, and Naureen Rana. Success of aquaculture industry with new insights of using insects as feed: a review. Fishes, 7(6): 395, December 17, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/395.

Hoffling:2024:ASM

[HMP^{+}24] Flávia Banderó Hoffling, Alex Silva Marquezi, Isabela Pinheiro, Cedric Simon, Artur Nishioka Rombenso, Walter Quadros Seiffert, Felipe do Nascimento Vieira, and Delano Dias Schleder. Aurantiochytrium sp. meal as feed additive for Pacific white shrimp reared under low temperature and challenged by WSSV in association with thermal stress. Fishes, 9(3):108, March 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/108.

Herrera:2019:TSC

[HMVRFD19] Mercedes Herrera, Raquel Moreno-Valcárcel, Ramón De Miguel Rubio, and Carlos Fernández-Delgado. From transient to sedentary? Changes in the home range size and environmental patterns of movements of European eels (Anguilla anguilla) in a Mediterranean river. Fishes, 4 (3):43, August 13, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/43.

He:2021:NRD

[HMX^{+}21] Qingsong He, Shuyang Ma, Qinwang Xing, Wenchao Zhang, Haiqing Yu, and Zhenjiang Ye. Nonstationary responses of demersal fishes to environmental variations in temperate waters of the Northwestern North Pacific under a changing climate. Fishes, 6 (3):22, July 22, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/22.

Ho:2022:CRO

[Ho22] Ching-Hsien Ho. Climate risks and opportunities of the marine fishery industry: a case study in Taiwan. Fishes, 7(3):116, May 21, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/116.

Hwang:2023:PFC

[HPJ^{+}23] Ju-Ae Hwang, Jun Seong Park, Hae Seung Jeong, Hyeongsu Kim, and Seung-Yoon Oh. Productivity of fish and crop growth and characteristics of bacterial communities in the FLOCponics system. Fishes, 8(8): 422, August 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/8/422.

Hushangi:2018:EPA

[HS18] Ramtin Hushangi and Seyed Pezhman Hosseini Shekarabi. Effect of a peracetic acid-based disinfectant on growth, hematology and histology of juvenile rainbow trout (Oncorhynchus mykiss). Fishes, 3(1): 10, February 06, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/10.

Hernandez:2024:QTB

[HS24] Jesús M. Rivera Hernández and Virginia R. Shervette. Queen triggerfish Balistes vetula age-based population demographics and reproductive biology for waters of the North Caribbean. Fishes, 9(5):162, April 29, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/162.

Henry-Silva:2023:PPW

[HSAF^{+}23] Gustavo Gonzaga Henry-Silva, Joseanna Alves, Dallas Flickinger, Renata Gomes-Rebouças, and Ambrosio Bessa-Junior. Polyculture of Pacific white shrimp Litopenaeus vannamei (Boone) and red seaweed Gracilaria birdiae (Greville) under different densities. Fishes, 8 (1):54, January 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/54.

Huynh:2024:PWT

[HSD^{+}24] Thi Kim Duyen Huynh, Marie-Louise Scippo, Mathias Devreese, Siska Croubels, Quoc Thinh Nguyen, Caroline Douny, Thi Hoang Oanh Dang, Quoc Viet Le, and Minh Phu Tran. Pharmacokinetics and withdrawal times of cefotaxime in white leg shrimp (Litopenaeus vannamei) after oral administration. Fishes, 9 (6):232, June 17, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/6/232.

Hu:2022:MCS

[HSZ^{+}22] Xiaojuan Hu, Haochang Su, Peng Zhang, Zuozhi Chen, Yu Xu, Wujie Xu, Jie Li, Guoliang Wen, and Yucheng Cao. Microbial community structure and metabolic characteristics of intestine and gills of dwarf-form populations of Sthenoteuthis oualaniensis in South China Sea. Fishes, 7(4): 191, August 04, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/191.

Horka:2024:ETB

[HV24] Petra Horka and Monika Vlachova. The effect of turbidity on the behavior of bleak (Alburnus alburnus). Fishes, 9(1): 3, December 20, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/1/3.

Hossain:2018:ASC

[HVRCG18] Mohammad Y. Hossain, Robert L. Vadas, Jr., Ramon Ruiz-Carus, and Shams M. Galib. Amazon sailfin catfish Pterygoplichthys pardalis (Loricariidae) in Bangladesh: a critical review of its invasive threat to native and endemic aquatic species. Fishes, 3(1):14, February 22, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/1/14.

Herdiana:2024:UCS

[HWW^{+}24a] Yudi Herdiana, Budy Wiryawan, Sugeng Hari Wisudo, James Richard Tweedley, Irfan Yulianto, Heidi Retnoningtyas, and Neil Randell Loneragan. Untangling the complexity of small-scale fisheries: Building an understanding of groupersnapper fisheries dynamics in Saleh Bay, West Nusa Tenggara, Indonesia. Fishes, 9(1):2, December 19, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/1/2.

Hu:2024:SPY

[HWW^{+}24b] Junru Hu, Lei Wang, Guoxia Wang, Hongxia Zhao, Huijie Lu, Kai Peng, Wen Huang, Zhenxing Liu, Ding Liu, and Yuping Sun. Selenium protects yellow catfish (Tachysurus fulvidraco) from lowtemperature damage via the perspective analysis of metabolomics and intestinal microbes. Fishes, 9(2):56, January 30, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/56.

Hu:2023:TMA

[HWX^{+}23] Guangwei Hu, Weili Wang, Kai Xu, Chao Wang, Dexue Liu, Jing Xu, Binlun Yan, Nanjing Ji, and Huan Gao. Transcriptomic and metabolomic analyses of Palaemon carinicauda hepatopancreas in response to Enterocytozoon hepatopenaei (EHP) infection. Fishes, 8(2):92, February 05, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/92.

Hu:2021:RBE

[HWZ21] Jinfei Hu, Ping Wang, and Hailong Zhang. The relationship between environmental factors and catch abundance of hairtail in the East China Sea using empirical dynamic modeling. Fishes, 6(4):80, December 13, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/80.

Hu:2023:EBP

[HXS^{+}23] Xiaojuan Hu, Yu Xu, Haochang Su, Wujie Xu, Guoliang Wen, Chuangwen Xu, Keng Yang, Song Zhang, and Yucheng Cao. Effect of a bacillus probiotic compound on Penaeus vannamei survival, water quality, and microbial communities. Fishes, 8 (7):362, July 11, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/362.

Huang:2023:CMC

[HXY^{+}23] Zengchao Huang, Liwen Xu, Shiping Yang, Shuanghu Cai, Jichang Jian, and Yucong Huang. A case of mycobacteriosis in cultured Japanese seabass (Lateolabrax japonicus) in Southern China. Fishes, 8(1):33, January 03, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/33.

Hu:2024:ETC

[HYN^{+}24] Tingting Hu, Cunrun Ye, Zhaoyang Ning, Tianmei Liu, and Weijie Mu. Effect of toxicity of chromium (VI) stressors alone and combined to high temperature on the histopathological, antioxidation, immunity, and energy metabolism in fish Phoxinus lagowskii. Fishes, 9(5):168, May 5, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/168.

Huo:2023:SAE

[HYXY23] Yilin Huo, Dazhang Yang, Jing Xie, and Zhikang Yang. Simulation analysis and experimental verification of freezing time of tuna under freezing conditions. Fishes, 8(9): 470, September 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/470.

Huang:2021:SDP

[HZG^{+}21] Yongyu Huang, Lili Zhang, Hui Ge, Guodong Wang, Shiyu Huang, and Zhangwu Yang. SNP development in Penaeus vannamei via next-generation sequencing and DNA pool sequencing. Fishes, 6(3): 36, September 07, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/36.

Hu:2022:FAP

[HZL^{+}22] Guanyu Hu, Zhenfang Zhao, Bilin Liu, Dongming Lin, Jiawei Liang, Zhou Fang, and Xinjun Chen. Fatty acid profile of jumbo squid (Dosidicus gigas) off the Peruvian Exclusive Economic Zone: Revealing the variability of feeding strategies. Fishes, 7(5):221, October 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/5/221.

Haguenauer:2021:DHC

[HZS^{+}21] Anne Haguenauer, Frédéric Zuberer, Gilles Siu, Daphne Cortese, Ricardo Beldade, and Suzanne C. Mills. Deep heat: a comparison of water temperature, anemone bleaching, anemonefish density and reproduction between shallow and mesophotic reefs. Fishes, 6(3):37, September 09, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/37.

Huang:2023:PGB

[HZZ^{+}23] Lishi Huang, Guosong Zhang, Yupeng Zhang, Xinlan Li, Zhong Luo, Wenyu Liu, Fu Luo, Haifeng Liu, Shaowu Yin, Jun Jiang, Xia Liang, and Quanquan Cao. Profiling genetic breeding progress in Bagrid catfishes. Fishes, 8(8):426, August 21, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/426.

Imsland:2023:CFD

[IBN^{+}23] Albert Kjartan Dagbjartarson Imsland, Jon Petter Berg, Velimir Nola, Lena Geitung, and Tina Oldham. Cleaner fish do not impact the pigmentation of salmon lice (Lepeophtheirus salmonis) in commercial aquaculture cages. Fishes, 8(9):455, September 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/455.

Iitembu:2023:CDA

[IFA^{+}23] Johannes A. Iitembu, Daniel Fitzgerald, Themistoklis Altintzoglou, Pierre Boudry, Peter Britz, Carrie J. Byron, Daniel Delago, Sophie Girard, Colin Hannon, Marcia Kafensztok, Francisco Lagreze, Jefferson Francisco Alves Legat, Angela Puchnick Legat, Adriane K. Michaelis, Ingelinn Eskildsen Pleym, Simone Sühnel, William Walton, and Åsa Strand. Comparative description and analysis of oyster aquaculture in selected Atlantic regions: Production, market dynamics, and consumption patterns. Fishes, 8(12): 584, November 29, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/584.

Ido:2019:PMP

[IKT19] Atsushi Ido, Motohisa Kanemaru, and Yoshiharu Tanioka. Preliminary monitoring of praziquantel in water and sediments at a Japanese amberjack (Seriola quinqueradiata) aquaculture site. Fishes, 4(2):24, March 28, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/2/24.

Iqbal:2022:IDF

[ILA22] Usama Iqbal, Daoliang Li, and Muhammad Akhter. Intelligent diagnosis of fish behavior using deep learning method. Fishes, 7(4): 201, August 11, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/201.

Imsland:2024:EDF

[IM24] Albert Kjartan Dagbjartarson Imsland and Hjörtur Methúsalemsson. The effects of different feed ration levels on growth, welfare rating, and early maturation in juvenile Atlantic salmon (Salmo salar). Fishes, 9(2): 70, February 10, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/2/70.

Imsland:2023:CFA

[Ims23] Albert Kjartan Dagbjartarson Imsland. Cleaner fish in aquaculture. Fishes, 8(2):83, January 31, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/83.

Iskandar:2023:ECD

[INCD23] Nur Syuhada Iskandar, Noorashikin Md Noor, Zaidi Che Cob, and Simon Kumar Das. Elevated carbon dioxide and its impact on growth, blood properties, and vertebral column of freshwater fish mahseer, Tor tambroides juveniles. Fishes, 8 (6):307, June 07, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/307.

Imsland:2022:LWT

[IR22] Albert Kjartan Dagbjartarson Imsland and Patrick Reynolds. In lumpfish we trust? The efficacy of lumpfish Cyclopterus lumpus to control Lepeophtheirus salmonis infestations on farmed Atlantic salmon: a review. Fishes, 7(5): 220, October 25, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/5/220.

Ivanauskas:2022:MAC

[ISA^{+}22] Edgaras Ivanauskas, Andrius Skersonas, Vaidotas Andrašūnas, Soukaina Elyaagoubi, and Artūras Razinkovas-Baziukas. Mapping and assessing commercial fisheries services in the Lithuanian part of the Curonian Lagoon. Fishes, 7(1):19, January 14, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/19.

Ivanova:2018:WTU

[ITG^{+}18] Lada Ivanova, Haitham Tartor, Søren Grove, Anja B. Kristoffersen, and Silvio Uhlig. Workflow for the targeted and untargeted detection of small metabolites in fish skin mucus. Fishes, 3 (2):21, June 07, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/2/21.

Junior:2022:UCE

[JBdFS^{+}22] Guerino Bandeira Junior, Adriane Erbice Bianchini, Carine de Freitas Souza, Sharine Nunes Descovi, Liana da Silva Fernandes, Lenise de Lima Silva, Juliana Felipetto Cargnelutti, and Bernardo Baldisserotto. The use of cinnamon essential oils in aquaculture: Antibacterial, anesthetic, growthpromoting, and antioxidant effects. Fishes, 7 (3):133, June 06, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/133.

Jareno:2024:EFA

[JBGCG^{+}24] Javier Jareño, Guillermo Bárcena-González, Jairo Castro-Gutiérrez, Remedios Cabrera-Castro, and Pedro L. Galindo. Enhancing fish auction with deep learning and computer vision: Automated caliber and species classification. Fishes, 9 (4):133, April 13, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/133.

Jensen:2023:ITF

[JBK^{+}23] Lasse Lange Jensen, Thomas Bjørn, Andreas Hein Korsgaard, Cino Pertoldi, and Niels Madsen. Influence of turbidity on foraging behaviour in three-spined sticklebacks (Gasterosteus aculeatus). Fishes, 8(12):609, December 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/609.

Jiang:2024:DAE

[JCL^{+}24] Senping Jiang, Xinyi Chen, Pengfei Li, Haoxue Wang, Jiji Li, Kaida Xu, and Yingying Ye. Does the artificial enhancement and release activity affect the genetic diversity of marbled rockfish Sebastiscus marmoratus in Zhoushan waters? Fishes, 9 (5):161, April 28, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/161.

Jin:2022:PAE

[JCR^{+}22] Jiali Jin, Zhipeng Chu, Rui Ruan, Wei Liu, Xihua Chen, and Chuangju Li. Phosphorus absorption and excretion in hybrid sturgeon (Huso dauricus [female] X Acipenser schrenckii [male]) intubated with different Ca/P ratios. Fishes, 7(3):138, June 10, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/138.

Jiao:2023:PST

[JCY^{+}23] Jinbiao Jiao, Jing Chen, Jiayun Yao, Yanli Li, Xuemei Yuan, Lei Huang, Shengqi Su, and Haiqi Zhang. The pathology and splenic transcriptome profiling of Trionyx sinensis challenged with Bacillus cereus. Fishes, 8(2):84, January 31, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/84.

Jahangiri:2018:APW

[JE18] Ladan Jahangiri and María Ángeles Esteban. Administration of probiotics in the water in finfish aquaculture systems: a review. Fishes, 3 (3):33, August 22, 2018.

CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/33.

Jerez:2018:SIF

[JFP^{+}18] Salvador Jerez, Ioannis Fakriadis, Maria Papadaki, M. Virginia Martín, Juana Rosa Cejas, and Constantinos C. Mylonas. Spawning induction of firstgeneration (F1) greater amberjack Seriola dumerili in the Canary Islands, Spain using Gn-RHa delivery systems. Fishes, 3(3):35, September 05, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/35.

Jiang:2024:IDD

[JHLW24] Yiying Jiang, Lei Huang, Yang Liu, and Shuang Wang. Impact of digital development and technology innovation on the marine fishery economy quality. Fishes, 9(7): 266, July 5, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/266.

Jungwirth:2024:ECA

[JHNF24] Arne Jungwirth, Anna Horsfield, Paul Nührenberg, and Stefan Fischer. Estimating cognitive ability in the wild: Validation of a detour test paradigm using a cichlid fish (neolamprologus pulcher). Fishes, 9(2): 50, January 27, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/50.

Jackson:2024:MLC

[JJB^{+}24] George D. Jackson, Christine H. Jackson, Maura Brandao, Adam K. Jackson, Eduardo Espinoza, and Monica Soria-Carvajal. Mercury levels in the crab Grapsus grapsus across the Galápagos Archipelago. Fishes, 9(6):233, June 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/233.

Jeong:2021:CGS

[JJK21] Min-A Jeong, Ye-Jin Jeong, and Kwang-Il Kim. Complete genome sequences and pathogenicity analysis of two red sea bream iridoviruses isolated from cultured fish in Korea. Fishes, 6(4):82, December 15, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/82.

Johnston:2024:RPG

[JK24] Christina U. Johnston and Christopher J. Kennedy. A review of P-glycoprotein function and regulation in fish. Fishes, 9(2):51, January 27, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/51.

Jung:2020:FDO

[JKK^{+}20] Joo-Young Jung, Soohwan Kim, Kyochan Kim, Bong-Joo Lee, Kang-Woong Kim, and Hyon-Sob Han. Feed and disease at olive flounder (Paralichthys olivaceus) farms in Korea. Fishes, 5(3):21, July 14, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/3/21.

Jung:2024:TID

[JKK24] Hyeok-Chan Jung, Jun-Hwan Kim, and Ju-Chan Kang. Toxic impact of dietary cadmium on bioaccumulation, growth, hematological parameters, plasma components, and antioxidant responses in starry flounder (Platichthys stellatus). Fishes, 9(2): 59, January 30, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/59.

Jang:2023:EHA

[JKP^{+}23] Won Je Jang, Shin-Kwon Kim, So Young Park, Dong Pil Kim, Yun-Jy Heo, Haham Kim, Su-Jeong Lee, Min Gyu Shin, Eun-Woo Lee, Seunghyung Lee, and Jong Min Lee. Effect of host-associated bacillus-supplemented artificial diets on growth, survival rate, and gene expression in early-stage eel larvae (Anguilla japonica). Fishes, 8 (5):247, May 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/247.

Jhan:2022:PCD

[JLT22] Hao-Tang Jhan, Hsin-Ta Lee, and Kuo-Huan Ting. The potential compatibility of designating offshore wind farms within wider marine protected areas— conservation of the Chinese white dolphin regarding fishers’ perception. Fishes, 7(4):195, August 05, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/195.

Jiang:2024:TAR

[JLW^{+}24] Senhao Jiang, Yuting Lei, Ti Wang, Ruiting Ma, Chunqiang Hou, and Qiuning Liu. Transcriptome analysis revealed potential immune-related genes of head kidney in the yellow catfish (Pelteobagrus fulvidraco) challenged with Aeromonas hydrophila. Fishes, 9 (3):100, March 06, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/100.

Ju:2024:IAM

[JML^{+}24] Pei-Mo Ju, Shi-Wei Ma, Yi-Yang Li, Shi-Fan Zhang, Jun Li, and Guang-Zhou Zhou. Investigation of the antiviral mechanism of curcumin analog EF-24 against Siniperca cachuatsi rhabdovirus. Fishes, 9(5):179, May 15, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/179.

Mao:2023:FNP

[jMlL23] Xiao jiang Mao and Kang le Lu. Fish nutrition and physiology. Fishes, 8(8):401, August 02, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/401.

Jimenez-Ruiz:2023:QIS

[JROHVH^{+}23] Edgar Iván Jiménez-Ruíz, Víctor Manuel Ocaño-Higuera, Santiago Valdez-Hurtado, José Alberto Cruz-Guzmán, Cesar Benjamín Otero-León, Saúl Ruíz-Cruz, Alba Mery Garzón-García, Hebert Jair Barrales-Cureño, Dalila Fernanda Canizales-Rodríguez, Cinthia Jhovanna Pérez-Martínez, and María Teresa Sumaya-Martínez. Quality improvement and shelf-life extension of iced Nile tilapia fillets using natural garlic extract. Fishes, 8(6):325, June 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/325.

Jiang:2024:PTM

[JSJ^{+}24] Zheng-Ting Jiang, Gang Shi, Dong-Neng Jiang, Yu Li, Yuan-Qing Huang, Hong-Juan Shi, and Guang-Li Li. Preliminary trial of male to female sex reversal by 17β-estradiol in combination with trilostane in spotted scat (Scatophagus argus). Fishes, 9(1): 1, December 19, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/1/1.

Jayasinghe:2023:DMC

[JSLE23] G. D. T. M. Jayasinghe, Joanna Szpunar, Ryszard Lobinski, and E. M. R. K. B. Edirisinghe. Determination of multi-class antibiotics residues in farmed fish and shrimp from Sri Lanka by ultra performance liquid chromatographytandem mass spectrometry (UPLC-MS/MS). Fishes, 8(3):154, March 04, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/154.

Janampa-Sarmiento:2024:FRV

[JSRE^{+}24] Peter C. Janampa-Sarmiento, Francisco Y. T. Reis, Renata C. Egger, Santiago B. de Pádua, Sóstenes A. C. Marcelino, João L. R. Cunha, Felipe Pierezan, Henrique C. P. Figueiredo, and Guilherme C. Tavares. First report of Vibrio vulnificus outbreak in farmraised sorubim (Pseudoplatystoma sp.) from Brazil. Fishes, 9(2): 54, January 29, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/2/54.

Johnston:2023:SDB

[JSSD23] Samuel J. Johnston, Jason B. Smith, Brady D. Slater, and Jonathan P. Doubek. Spatial density, biomass, and composition of crustacean zooplankton on Lake Michigan beaches. Fishes, 8(12):599, December 06, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/599.

Jongjaraunsuk:2024:TLM

[JTS^{+}24] Roongparit Jongjaraunsuk, Wara Taparhudee, Soranuth Sirisuay, Methee Kaewnern, Varunthat Dulyapurk, and Sommai Janekitkarn. Transfer learning model application for Rastrelliger brachysoma and R. kanagurta image classification using smartphonecaptured images. Fishes, 9(3):103, March 07, 2024. CODEN ???? ISSN 2410-3888. URL

                                        https://www.mdpi.com/
                                        2410-3888/9/3/103.

Ji:2023:PTR

[JXW^{+}23] Nanjing Ji, Mingyang Xu, Junyue Wang, Junjia Li, Shishi Liu, Xueyao Yin, Xin Shen, and Yuefeng Cai. Physiological and transcriptomic response of Asiatic hard clam Meretrix meretrix to the harmful alga Heterosigma akashiwo. Fishes, 8(2): 67, January 22, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/67.

Jia:2022:FDM

[JZX22] Hui Jia, Hui Zhang, and Weiwei Xian. Fish diversity monitored by environmental DNA in the Yangtze River mainstream. Fishes, 7(1): 1, December 22, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/1.

Koel:2020:YLE

[KAB^{+}20] Todd M. Koel, Jeffery L. Arnold, Patricia E. Bigelow, Travis O. Brenden, Jeffery D. Davis, Colleen R. Detjens, Philip D. Doepke, Brian D. Ertel, Hayley C. Glassic, Robert E. Gresswell, Christopher S. Guy, Drew J. MacDonald, Michael E. Ruhl, Todd J. Stuth, David P. Sweet, John M. Syslo, Nathan A. Thomas, Lusha M. Tronstad, Patrick J. White, and Alexander V. Zale. Yellowstone Lake ecosystem restoration: a case study for invasive fish management. Fishes, 5 (2):18, June 12, 2020. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/2/18.

Kerzerho:2023:MRA

[KAB^{+}23] Vincent Kerzérho, Florence Azaïs, Serge Bernard, Sylvain Bonhommeau, Blandine Brisset, Laurent De Knyff, Mohan Julien, Michel Renovell, Tristan Rouyer, Claire Saraux, and Fabien Soulier. Multilinear regression analysis between local bioimpedance spectroscopy and fish morphological parameters. Fishes, 8(2):88, February 01, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/88.

Karlsen:2024:TDM

[KAJ^{+}24] Bård Ove Karlsen, Dhurba Adhikari, Tor Erik Jørgensen, Ida Klykken Hanssen, Truls Borg Moum, Jarle Tryti Nordeide, and Steinar Daae Johansen. Two distinct maternal lineages of threespine stickleback (Gasterosteus aculeatus) in a small Norwegian subarctic lake. Fishes, 9(7):285, July 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/285.

Kruze:2023:PDI

[KAR^{+}23] Eriks Kruze, Andris Avotins, Loreta Rozenfelde, Ivars Putnis, Ivo Sics, Laura Briekmane, and Jens Olsson. The population development of the invasive round goby Neogobius melanostomus in Latvian Waters of the Baltic Sea. Fishes, 8 (6):305, June 07, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/305.

Kidd:2021:ESD

[KBB^{+}21] Jess A. Kidd, Monica Boudreau, Robert C. Bailey, Michael R. van den Heuvel, Mark R. Servos, and Simon C. Courtenay. Evaluating the sampling design of a longterm community-based estuary monitoring program. Fishes, 6(3): 27, August 02, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/27.

Kim:2019:BRJ

[KBCM19] Jaewoo Kim, Caitlyn Bondy, Catherine M. Chandler, and Nicholas E. Mandrak. Behavioural response of juvenile common carp (Cyprinus carpio) and juvenile channel catfish (Ictalurus punctatus) to strobe light. Fishes, 4 (2):29, May 04, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/2/29.

Kamikawa:2023:GCR

[KBK^{+}23] Keith Kamikawa, Brian W. Bowen, Donald Kobayashi, Kimberly Peyton, and Elizabeth Wallace. Genetic connectivity of roundjaw bonefish Albula glossodonta (Elopomorpha, Albulidae) in the Central Pacific Ocean resolved through ddRADbased population genomics. Fishes, 8(12): 585, November 29, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/585.

Kim:2023:CSA

[KCCK23] Hansoo Kim, Sungho Cho, Jee Woong Choi, and Donhyug Kang. Characteristics of sound attenuation by individual and multiple fishes. Fishes, 8(3):161, March 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/161.

Khalil:2023:EMR

[KEA^{+}23] Karim Khalil, Ahmed Elaswad, Hisham Abdelrahman, Maximillian Michel, Wenbiao Chen, Shikai Liu, Ramjie Odin, Zhi Ye, David Drescher, Khoi Vo, William S. Bugg, Guyu Qin, Yujia Yang, Nathan J. C. Backenstose, Zhanjiang Liu, Roger D. Cone, and Rex Dunham. Editing the melanocortin-4 receptor gene in channel catfish using the CRISPR-Cas9 system. Fishes, 8(2):116, February 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/116.

Kim:2023:SRI

[KFS23] Hongsik Kim, Ana C.

Franco, and U. Rashid Sumaila. A selected review of impacts of ocean deoxygenation on fish and fisheries. Fishes, 8 (6):316, June 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/316.

Kovacic:2024:FMC

[KGP24] Marcelo Kovačić, Vasilis Gerovasileiou, and Robert A.

Patzner. Fishes in marine caves. Fishes, 9 (6):243, June 20, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/243.

Kim:2023:CKD

[Kim23a] Dong In Kim. Correction: Kim, D. I. A Study on the Metabolic Rate Change Pattern in F_{2} Hybrid Sturgeon, the Bester (Huso huso × Acipenser ruthenus), during the Early Developmental Stage. Fishes 2023, 8, 113. Fishes, 8 (5):257, May 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/257.  See

[Kim23a].

Kim:2023:SMR

[Kim23b]  Dong In Kim.  A

study on the metabolic rate change pattern in F_{2} hybrid sturgeon, the bester (Huso huso × Acipenser ruthenus), during the early developmental stage. Fishes, 8(2):113, February 15, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/2/113. See

correction [Kim23a].

Kennedy:2022:UFS

[KJ22] Emily K. C. Kennedy and David M. Janz. The use of fish scale hormone concentrations in the assessment of longterm stress and associated adverse effects on reproductive endocrinology. Fishes, 7(6):393, December 16, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/393.

Kim:2023:CRS

[KJK^{+}23] Kyung-Ho Kim, Min-Soo Joo, Gyoungsik Kang, Won-Sik Woo, Min-Young Sohn, Ha-Jeong Son, and Chan-Il Park. Characterization of red sea bream (Pagrus major) interferon regulatory factor 5 and 6 genes and their expression in response to RSIV infection. Fishes, 8(2): 114, February 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/114.

Kampouris:2020:MRG

[KKB20] Thodoros E. Kampouris, Emmanouil Kouroupakis, and Ioannis E. Batjakas. Morphometric relationships of the global invader Callinectes sapidus Rathbun, 1896 (Decapoda, Brachyura, Portunidae) from Papapouli Lagoon, NW Aegean Sea, Greece. With notes on its ecological preferences. Fishes, 5(1):5, January 14, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/1/5.

Kropielnicka-Kruk:2023:EFT

[KKFC^{+}23] Katarzyna Kropielnicka-Kruk, Quinn P. Fitzgibbon, Mohamed B. Codabaccus, Andrew J. Trotter, Chris G. Carter, and Gregory G. Smith. Effect of feed texture and dimensions, on feed waste type and feeding efficiency in juvenile Sagmariasus verreauxi. Fishes, 8(11): 553, November 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/553.

Kang:2023:SSE

[KKPL23] Chang-Keun Kang, Hyun-Woo Kim, Hyun Je Park, and Chung Il Lee. Sharing scientific evidence of the response of aquatic animals to environmental change. Fishes, 8(1): 40, January 05, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/40.

Kim:2022:CBF

[KKPY22] Jae Goo Kim, Su Hwan Kim, Jong Young Park, and Su-Hyang Yoo. Correlation between feeding behaviors and retinal photoreceptor cells of largemouth bass, Micropterus salmoides, in Korea. Fishes, 7(1): 25, January 19, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/25.

Kim:2023:DMP

[KLD^{+}23] Eun-Mi Kim, Mi-Nan Lee, Chun-Mae Dong, Jae-Koo Noh, Eun-Soo Noh, Woo-Jin Kim, Bo-Hye Nam, and Young-Ok Kim. Development of a multiplex polymerase chain reaction method for rapid and accurate identification of Girella punctata and G. leonina (Teleostei: Girellidae). Fishes, 8(8): 415, August 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/415.

Kampouris:2018:FRC

[KMB18] Thodoros E. Kampouris, Debora Milenkova, and Ioannis E. Batjakas. On the finding of the rare crab Paragalene longicrura (Nardo, 1868) (Crustacea, Decapoda, Brachyura, Progeryonidae) from Thermaikos Gulf, Northwest Aegean Sea, Greece. Fishes, 3(3):30, August 08, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/30.

Kawase:2017:DES

[KMS^{+}17] Hiroshi Kawase, Ryo Mizuuchi, Hirofumi Shin, Yuki Kitajima, Koh Hosoda, Masahiro Shimizu, Daisuke Iwai, and Shigeru Kondo. Discovery of an earliest-stage “Mystery Circle” and development of the structure constructed by pufferfish, Torquigener albomaculosus (Pisces: Tetraodontidae). Fishes, 2(3): 14, August 23, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/3/14.

Kitasato:2018:ADB

[KMSO18] Airi Kitasato, Tomo Miyazaki, Yoshihiro Sugaya, and Shinichiro Omachi. Automatic discrimination between Scomber japonicus and Scomber australasicus by geometric and texture features. Fishes, 3 (3):26, June 27, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/26.

Kovacic:2023:SRG

[KRAFO23] Marcelo Kovačić, Sergio Ramírez-Amaro, Maria Teresa Farriols, and Francesc Ordines. The second record of Gymnesigobius medits Kovačić, Ordines, Ramirez-Amaro & Schliewen, 2019, the deepest benthic gobiiform species, and the additional records of Gobius xoriguer Iglésias, Vukić & Šanda, 2021 (Actinopterygii: Gobiiformes: Gobiidae). Fishes, 8(6):331, June 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/331.

Kiruba-Sankar:2023:ATK

[KSAB^{+}23] Rajendran Kiruba-Sankar, Sirisha Adamala, Jessica Barman, Kandasamy Saravanan, Jayasimhan Praveenraj, Eswaran Yuvaraj, Gokhlesh Kumar, and Sheikh Zamir Ahmed. Aboriginal tribe’s knowledge of the endangered freshwater turtle Cuora amboinensis in Car Nicobar, a remote oceanic island in the Bay of Bengal. Fishes, 8(10): 517, October 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/10/517.

Kalan:2023:FFP

[KSO^{+}23] Parker Kalan, John Steinbeck, Freddy Otte, Sean C. Lema, and Crow White. Filterfeeding Pacific lamprey (Entosphenus tridentatus) ammocetes can reduce suspended concentrations of E. coli bacteria. Fishes, 8(2):101, February 08, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/101.

Kolanczyk:2024:CSP

[KSSI24] Richard C. Kolanczyk, Laura E. Solem, Patricia K. Schmieder, and James M. McKim III. A comparative study of Phase I and II hepatic microsomal biotransformation of phenol in three species of Salmonidae: Hydroquinone, catechol, and phenylglucuronide formation. Fishes, 9(7): 284, July 17, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/7/284.

Kindong:2022:LBA

[KSWT22] Richard Kindong, Ousmane Sarr, Feng Wu, and Siquan Tian. Lengthbased assessment methods for the conservation of a pelagic shark, Carcharhinus falciformis from the Tropical Pacific Ocean. Fishes, 7 (4):184, July 25, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/184.

Kadota:2024:QAA

[KTT24] Minoru Kadota, Shinsuke Torisawa, and Tsutomu Takagi. Quantitative assessment and analysis of fish behavior in closed systems using information entropy. Fishes, 9(6):224, June 12, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/224.

Kempf:2024:HTR

[KWK^{+}24] Sarah Kempf, Heidrun Sigrid Windisch, Kristina Lore Kunz, Hans-Otto Pörtner, Felix Christopher Mark, and Magnus Lucassen. Hepatic transcriptomic responsiveness of polar cod, Boreogadus saida, to ocean acidification and warming. Fishes, 9 (7):271, July 10, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/271.

Kwikiriza:2023:MVN

[KYB^{+}23] Gerald Kwikiriza, Mourine J. Yegon, Nelson Byamugisha, Apulnal Beingana, Faith Atukwatse, Alex Barekye, Juliet K. Nattabi, and Harald Meimberg. Morphometric variations of Nile tilapia (Oreochromis niloticus) (Linnaeus, 1758) local strains collected from different fish farms in South Western Highland Agro-Ecological Zone (SWHAEZ), Uganda: Screening strains for aquaculture. Fishes, 8 (4):217, April 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/217.

Karimov:2024:LWA

[KZC^{+}24] Erkin Karimov, Bernhard Zeiringer, Johan Coeck, Pieterjan Verhelst, Bakhtiyor Karimov, Otabek Omonov, Martin Schletterer, and Daniel S. Hayes. Length– weight–age relationship of Schizothorax eurystomus Kessler, 1872 and comparison to other snow trout species in Central Asia. Fishes, 9(3):94, February 29, 2024. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/9/3/94.

Kovacic:2023:CQB

[KŽM^{+}23] Ines Kovačić, Ante Žunec, Mauro Matešković, Petra Burić, Neven Iveša, Mauro Štifanić, and Jadranka Frece. Commercial quality, biological indices and biochemical composition of queen scallop Aequipecten opercularis in culture. Fishes, 8(1): 48, January 11, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/48.

Laskar:2024:ITC

[LBC^{+}24] Boni Amin Laskar, Dhriti Banerjee, Sangdeok Chung, Hyun-Woo Kim, Ah Ran Kim, and Shantanu Kundu. Integrative taxonomy clarifies the historical flaws in the systematics and distributions of two osteobrama fishes (Cypriniformes: Cyprinidae) in India. Fishes, 9(3): 87, February 27, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/87.

Luo:2024:TGS

[LBH^{+}24] Zhang Luo, Xiaohui Bai, Shuang Hao, Mengyu Wang, Yongjiang Wu, and Hanchang Sun. Two genotypes of Streptococcus iniae are the causative agents of diseased ornamental fish, green terror cichlid (Aequidens rivulatus). Fishes, 9(4):140, April 17, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/140.

Li:2024:MIN

[LCC^{+}24] Die Li, Xiaojuan Cui, Shuailin Chen, Jia Xu, Yujing Li, Qiongyu Zhang, and Yuandong Sun. Mechanistic insights into nonylphenol stress on BMP2 and BMP4 gene expression in red crucian carp (Carassius auratus red var.). Fishes, 9(5):159, April 28, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/5/159.

Lassauce:2023:SEP

[LCW23a] Hugo Lassauce, Olivier Chateau, and Laurent Wantiez. Spatial ecology of the population of reef manta rays, Mobula alfredi (Krefft, 1868), in New Caledonia using satellite telemetry 1 — horizontal behaviour. Fishes, 8(6):328, June 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/328.

Li:2023:IRB

[LCW^{+}23b] Junwei Li, Suwen Chen, Peng Wu, Changbo Zhu, Ruiping Hu, Ting Li, and Yongjian Guo. Insights into the relationship between intestinal microbiota of the aquaculture worm Sipunculus nudus and surrounding sediments. Fishes, 8 (1):32, January 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/32.

Lin:2022:APP

[LCWH22] Pei-Hui Lin, Sai-Wei Chen, Zhi-Hong Wen, and Shao-Yang Hu. Administration of the potential probiotic Paenibacillus ehimensis NPUST1 enhances expression of indicator genes associated with nutrient metabolism, growth and innate immunity against Aeromonas hydrophila and Streptococcus indie infections in zebrafish (Danio rerio). Fishes, 7(6):386, December 12, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/386.

Luo:2023:EWS

[LCX^{+}23] Yuan-Yuan Luo, Xian-Can Chen, Rui-Lin Xie, Zhuo-Hao Ruan, Zhi-Qiang Lu, Liang-Sen Jiang, Yi-Fu Li, and Wen-Sheng Liu. The effect of water spinach on the water quality, antioxidant system, non-specific immune response, growth performance, and carbon balance in red tilapia production. Fishes, 8(10): 515, October 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/10/515.

Li:2023:DAE

[LCZ^{+}23] Yongtao Li, Zhaolong Cheng, Tao Zuo, Mingxiang Niu, Ruisheng Chen, and Jun Wang. Distribution and abundance of the East Asian finless porpoise in the coastal waters of Shandong Peninsula, Yellow Sea, China. Fishes, 8(8): 410, August 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/8/410.

Lobel:2019:GAP

[LDBL19] Lisa Kerr Lobel, Devin M.

Drown, Paul H. Barber, and Phillip S. Lobel. A genetic assessment of parentage in the blackspot sergeant damselfish, Abudefduf sordidus (Pisces: Pomacentridae). Fishes, 4(4): 53, October 24, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/53.

Linh:2022:EDR

[LDD^{+}22] Nguyen Vu Linh, Le Thanh Dien, Ha Thanh Dong, Nuttapon Khongdee, Seyed Hossein Hoseinifar, Mohamed Saiyad Musthafa, Mahmoud A. O. Dawood, and Hien Van Doan. Efficacy of different routes of formalin-killed vaccine administration on immunity and disease resistance of Nile tilapia (Oreochromis niloticus) challenged with Streptococcus agalactiae. Fishes, 7(6):398, December 19, 2022. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/7/6/398.

Lima:2024:NFI

[LdRCV24] Adriana Ferreira Lima, Anderson Guilherme Pereira dos Reis, Vladimir Eliodoro Costa, and Wagner Cotroni Valenti. Natural food intake and its contribution to tambaqui growth in fertilized and unfertilized ponds. Fishes, 9 (4):139, April 16, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/139.

Li:2021:FPW

[LDW^{+}21] Junyi Li, Hao Du, Jinming Wu, Hui Zhang, Li Shen, and Qiwei Wei. Foundation and prospects of wild population reconstruction of acipenser dabryanus. Fishes, 6(4):55, October 30, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/55.

Leng:2023:SUA

[LDX^{+}23] Xiaoqian Leng, Hao Du, Wei Xiong, Peilin Cheng, Jiang Luo, and Jinming Wu. Successful ultrasonographyassisted artificial reproduction of critically endangered Sichuan taimen (Hucho bleekeri). Fishes, 8(3):152, March 03, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/3/152.

Li:2024:EMV

[LDZ^{+}24] Tao Li, Ronggang Ding, Jing Zhang, Yongcan Zhou, Chunsheng Liu, Zhenjie Cao, and Yun Sun. The establishment of the multi-visual loopmediated isothermal amplification method for the rapid detection of Vibrio harveyi, Vibrio parahaemolyticus, and Singapore grouper iridovirus. Fishes, 9(6): 225, June 13, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/225.

Luz:2024:USA

[LF24] Ronald Kennedy Luz and Gisele Cristina Favero. Use of salt, anesthetics, and stocking density in transport of live fish: a review. Fishes, 9(7):286, July 18, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/286.

Li:2023:WLR

[LFH^{+}23] Yushan Li, Mingjun Feng, Liangliang Huang, Peiyu Zhang, Hongxia Wang, Jingwei Zhang, Yuehan Tian, and Jun Xu. Weight–length relationship analysis revealing the impacts of multiple factors on body shape of fish in China. Fishes, 8(5):269, May 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/269.

Li:2023:EVF

[LGZ^{+}23] Jiayang Li, Yu Guo, Xinye Zhao, Shengjie Zhou, Zhenhua Ma, Gang Yu, Chuanxin Qin, and Xingqiang Wang. The effects of vibration frequency on oxidative stress, digestive enzymes and ATPases of crimson snapper (Lutjanus erythropterus) fry during transport. Fishes, 8(12): 603, December 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/603.

Lin:2023:RCP

[LHG^{+}23] Pengcheng Lin, Huaming Hu, Zheng Gong, Jian Wang, and Xin Gao. Reproductive characteristics of Pseudecheneis sulcatus (Siluriforms: Sisoridae) in the Lower Yarlung Zangbo River, Tibet. Fishes, 8(2):106, February 10, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/2/106.

Li:2024:FAB

[LHL^{+}24] Shaowen Li, Yufang Huang, Fan Li, Yanfen Liu, Hongliang Ma, Xiaomin Zhang, Xiuxia Wang, Wei Chen, Guangxin Cui, and Tiantian Wang. Functional alpha and beta diversity of fish communities and their relationship with environmental factors in the Huanghe River (Yellow River) Estuary and Adjacent Seas, China. Fishes, 9 (6):222, June 12, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/222.

Liu:2023:EDL

[LHX^{+}23] Yongqiang Liu, Enhao Huang, Yi Xie, Liuqing Meng, Dongsheng Liu, Ziqi Zhang, Jiang Zhou, Qin Zhang, and Tong Tong. The effect of dietary lipid supplementation on the serum biochemistry, antioxidant responses, initial immunity, and mTOR pathway of juvenile tilapia (Oreochromis niloticus). Fishes, 8(11):535, October 26, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/535.

Li:2023:HHD

[LHZ^{+}23] Xinghao Li, Xueli Huang, Liya Zhao, Wei Cai, Yuhe Yu, and Jin Zhang. Host habitat as a dominant role in shaping the gut microbiota of wild crucian carp (Carassius auratus). Fishes, 8 (7):369, July 17, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/369.

Li:2024:DND

[LHZ^{+}24] Yinjia Li, Zeyuan Hu, Yixi Zhang, Jihang Liu, Wan Tu, and Hong Yu. DDEYOLOv9: Network for detecting and counting abnormal fish behaviors in complex water environments. Fishes, 9 (6):242, June 20, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/6/242.

Liu:2024:USD

[Liu24] Suyu Liu. Using sustainable development goal indicator 14.7.1 to measure sustainable fishery: The statistical limitations. Fishes, 9(2): 57, January 30, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/2/57.

Lee:2022:ITA

[LJK^{+}22] Chorong Lee, Hye Jin Jeon, Bum Keun Kim, Seong-Kyoon Choi, Sumi Kim, Gwang Il Jang, Ji Hyung Kim, and Jee Eun Han. Infectivity and transmissibility of acute hepatopancreatic necrosis disease associated Vibrio parahaemolyticus in frozen shrimp archived at −80^{∘}C. Fishes, 7 (3):125, May 30, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/125.

Liu:2022:CDM

[LJP^{+}22] Yang Liu, Yiying Jiang, Zhaobin Pei, Limin Han, Hongrun Shao, Yang Jiang, Xiaomeng Jin, and Saihong Tan. Coordinated development of the marine environment and the marine fishery economy in China, 2011–2020. Fishes, 7(6): 391, December 15, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/391.

Liu:2024:BVC

[LJR^{+}24] Lijuan Liu, Dongdong Jiang, Yan Ren, Cunbin Shi, Yajun Wang, Jiyuan Yin, Qing Wang, and Defeng Zhang.

The Bacillus velezensis CYS06 strain exhibits promising applications in fighting grass carp bacterial diseases. Fishes, 9(1):7, December 22, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/1/7.

Lin:2021:FRF

[LJX^{+}21] Bin Lin, Kailin Jiang, Zhiqi Xu, Feiyi Li, Jiao Li, Chaoli Mou, Xinyao Gong, and Xuliang Duan. Feasibility research on fish pose estimation based on rotating box object detection. Fishes, 6(4): 65, November 19, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/4/65.

Lipej:2022:AAI

[LKD22] Lovrenc Lipej, Marcelo Kovačić, and Jakov Dulčić. An analysis of Adriatic ichthyofauna— ecology, zoogeography, and conservation status. Fishes, 7(2):58, April 04, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/2/58.

Li:2022:ERT

[LKJ22] Yutong Li, Jiaojiao Kong, and Jianyue Ji. Environmental regulation, technological innovation and development of marine fisheries—evidence from ten coastal regions in China. Fishes, 7(1): 20, January 14, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/20.

Lee:2021:MDC

[LKU21] Cheng-Linn Lee, Yuri Kominami, and Hideki Ushio. Mechanism of delayed convulsion in fish: the actions of norepinephrine in spinal cord. Fishes, 6(2):12, March 31, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/2/12.

Lindseth:2018:ALL

[LL18a] Adelaide V. Lindseth and Phillip S. Lobel. Addendum: Lindseth, A. and Lobel, P. S. Underwater Soundscape Monitoring and Fish Bioacoustics: a Review. Fishes 2018, 3,

36. Fishes, 3(4):40, October 08, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/4/40.

Lindseth:2018:USM

[LL18b] Adelaide V. Lindseth and Phillip S. Lobel. Underwater soundscape monitoring and fish bioacoustics: a review. Fishes, 3(3):36, September 12, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/3/36.

Lobel:2022:DEB

[LL22] Lisa Kerr Lobel and Phillip S. Lobel. Damselfish embryos as a bioindicator for military contamination on coral reefs at Johnston Atoll, Pacific Ocean. Fishes, 7(2):55, April 26, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/2/55.

Lindholm-Lehto:2020:AGM

[LLKKV20] Petra Lindholm-Lehto, Juha Koskela, Janne Kaseva, and Jouni Vielma. Accumulation of geosmin and 2-methylisoborneol in European whitefish Coregonus lavaretus and rainbow trout Oncorhynchus mykiss in RAS. Fishes, 5(2):13, May 11, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/2/13.

Li:2023:DSS

[LLL^{+}23] Jianping Li, Chen Li, Congcong Li, Wei Luo, Kang Wu, Songming Zhu, and Zhangying Ye. Design and study of a spindle-shaped fry headto-tail orientation device. Fishes, 8(3):143, February 28, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/143.

Long:2023:EFD

[LLS^{+}23] Qiaolin Long, Jian Liu, Yunfei Sun, Zhigang Yang, Boping Tang, and Yongxu Cheng. The effect of food deprivation on foraging behavior and digestive and metabolic capacities of the Chinese mitten crab, Eriocheir sinensis. Fishes, 8(1): 47, January 11, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/47.

Lerner:2017:RSX

[LLTM17] Justin Lerner, Juan C.

Levesque, and Liana Talaue-McManus. Recreational swordfish (Xiphias gladius) fishery: Angler practices in South Florida (USA). Fishes, 2 (4):18, October 27, 2017. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/2/4/18.

Luo:2022:MEI

[LLW^{+}22] Wei Luo, Chen Li, Kang Wu, Songming Zhu, Zhangying Ye, and Jianping Li. A method for estimating the injection position of turbot (Scophthalmus maximus) using semantic segmentation. Fishes, 7(6):385, December 11, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/385.

Li:2023:EAG

[LLWW23] Peilun Li, Jiacheng Liu, Tai Wang, and Jilong Wang. Estimates of the age, growth, and mortality of Triplophysa scleroptera (Herzenstein, 1888) in the upper reaches of the Yellow River, China. Fishes, 8(9):457, September 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/457.

Lian:2022:GWC

[LLY^{+}22] Jiaying Lian, Liyuan Lv, Hanhan Yao, Zhihua Lin, and Yinghui Dong. Genome-wide characterization and analysis of expression of the histone gene family in razor clam, Sinonovacula constricta. Fishes, 7(1): 5, December 25, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/5.

Lin:2024:CCR

[LLY^{+}24] Ainuo Lin, Yaxi Li, Zhi Yan, Huilin Zhang, Baozhen Jiang, Jingjing Chen, Derui Wang, Huan Wang, Xiaodong Li, Zhen Lu, and Ke Li. Chemical cues released by predators’ consumption of heterospecific prey alter the embryogenesis of zebrafish. Fishes, 9(3):95, March 01, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/95.

Leite:2023:SUP

[LMB^{+}23] Tamara Leite, Daniel Mameri, Paulo Branco, Inês Vieira, Margarida Oliveira, and José Maria Santos. Swimming under pressure: The sub-lethal effects of a pesticide on the behaviour of native and non-native cypriniformes fish. Fishes, 8(9): 462, September 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/462.

Leese:2021:EWT

[LMC21] Joseph M. Leese, Julia McMahon, and Joseph C. Colosi. Effects of wastewater treatment plant effluent in a receiving stream on reproductive behavior of fathead minnows (Pimephales promelas). Fishes, 6 (2):14, April 12, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/2/14.

Li:2022:FCM

[LMLH22] Yingcen Li, Yunliang Miao, Xufang Liang, and Shan He. Functional characterization and molecular marker development of the proenkephalin as biomarker of food addiction in food habit domestication of Mandarin fish (Siniperca chuatsi). Fishes, 7(3): 118, May 27, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/118.

Lopes:2022:DEF

[LMMC^{+}22] Ana F. Lopes, Robyn Murdoch, Sara Martins-Cardoso, Carolina Madeira, Pedro M. Costa, Ana S. Félix, Rui F. Oliveira, Narcisa M. Bandarra, Catarina Vinagre, Ana R. Lopes, Emanuel J. Gonçalves, and Ana Margarida Faria. Differential effects of food restriction and warming in the two-spotted goby: Impaired reproductive performance and stressed offspring. Fishes, 7(4): 194, August 05, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/194.

Ly:2021:EDD

[LMNN21] Khanh Van Ly, David Kamau Murungu, Dung Phuong Nguyen, and Ngoc Anh Thi Nguyen. Effects of different densities of sea grape Caulerpa lentillifera on water quality, growth and survival of the whiteleg shrimp Litopenaeus vannamei in polyculture system. Fishes, 6(2): 19, May 05, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/2/19.

Linnansaari:2023:RCW

[LOB^{+}23] Tommi Linnansaari, Antóin M.

O’Sullivan, Cindy Breau, Emily M. Corey, Elise N. Collet, R. Allen Curry, and Richard A. Cunjak. The role of coldwater thermal refuges for stream salmonids in a changing climate — experiences from Atlantic Canada. Fishes, 8(9): 471, September 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/471.

Lokovsek:2022:IFE

[LOBTL22] Ana Lokovšek, Martina Orlando-Bonaca, Domen Trkov, and Lovrenc Lipej. An insight into the feeding ecology of Serranus scriba, a shallow water mesopredator in the Northern Adriatic Sea, with a non-destructive method. Fishes, 7(4): 210, August 20, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/210.

Lu:2022:AGP

[LOT^{+}22] Huajie Lu, Yuzhe Ou, Yurong Teng, Ziyue Chen, and Xinjun Chen. Age, growth and population structure analyses of the Berryteuthis magister shevtsovi in the Japan Sea by statolith microstructure. Fishes, 7(5):215, October 24, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/5/215.

Lai:2024:DFR

[LPB^{+}24] Xiaojian Lai, Shuai Peng, Zhaoren Bai, Le Cao, Huixuan Huang, Yonghua Jiang, and Yilei Wang. Direct feedback regulation of E2, t, and hCG in the brain– pituitary–gonad axis of Japanese eel (Anguilla japonica) during artificial maturation. Fishes, 9(7):265, July 5, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/265.

Lattos:2023:ADM

[LPF^{+}23] Athanasios Lattos, Dimitrios K. Papadopoulos, Konstantinos Feidantsis, Ioannis A. Giantsis, Ioannis Georgoulis, Dimitrios Karagiannis, and Basile Michaelidis. Antioxidant defense of Mytilus galloprovincialis mussels induced by marine heatwaves in correlation with Marteilia pathogen presence. Fishes, 8(8):408, August 08, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/408.

Lee:2023:CMGa

[LPK^{+}23a] Yong-Suk Lee, Maheshkumar Prakash Patil, Jong-Oh Kim, Yu-Jin Lee, Yong Bae Seo, Jin-Koo Kim, Kiran R. Mahale, and Gun-Do Kim. Complete mitochondrial genome and phylogenetic position of Chirolophis wui (Perciformes: Stichaeidae). Fishes, 8(3):165, March 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/165.

Lee:2023:CMGb

[LPK^{+}23b] Yong-Suk Lee, Maheshkumar Prakash Patil, Jong-Oh Kim, Yu-Jin Lee, Yong Bae Seo, Jin-Koo Kim, Rahul K. Suryawanshi, and Gun-Do Kim. The complete mitochondrial genome of the fivespot flounder, Pseudorhombus pentophthalmus (Pleuronectiformes: Paralichthyidae), from Korea and its phylogenetic analysis. Fishes, 8(3):150, March 02, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/150.

Lazado:2019:SPS

[LS19] Carlo C. Lazado and Peter Vilhelm Skov. Secretory proteins in the skin mucus of Nile tilapia (Oreochromis niloticus) are modulated temporally by photoperiod and bacterial endotoxin cues. Fishes, 4(4):57, December 05, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/57.

Li:2022:FFI

[LSJ^{+}22] Danyang Li, Houcheng Su, Kailin Jiang, Dan Liu, and Xuliang Duan. Fish face identification based on rotated object detection: Dataset and exploration. Fishes, 7(5):219, October 25, 2022. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/7/5/219.

Li:2023:MAM

[LSJ^{+}23] Xi-Lian Li, Pei-Jing Shen, Wen-Ping Jiang, Ji-Lun Meng, Hai-Hua Cheng, and Qiang Gao. Metabonomic analysis of Macrobrachium rosenbergii with Iron Prawn Syndrome (IPS). Fishes, 8(4):196, April 09, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/196.

Lu:2024:SIS

[LSJ24] Zhengwei Lu, Liming Song, and Keji Jiang. Status identification in support of fishing effort estimation for tuna longliners in waters near the Marshall Islands based on AIS data. Fishes, 9(2):66, February 08, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/66.

Liu:2024:AIE

[LSL^{+}24] Fei Liu, Jinfeng Sun, Jinnan Long, Lichao Sun, Chang Liu, Xiaofan Wang, Long Zhang, Pengyuan Hao, Zhongkai Wang, Yanting Cui, Renjie Wang, and Yuquan Li. Assessing the interactive effects of high salinity and stocking density on the growth and stress physiology of the Pacific white shrimp Litopenaeus vannamei. Fishes, 9(2):62, January 31, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/62.

Liu:2022:FSR

[LSS^{+}22] Yin Liu, Lun Song, Guangjun Song, Jinhao Wu, Kun Wang, Zhaohui Wang, and Suxuan Liu. Feeding selectivity of Ruditapes philippinarum on phytoplankton. Fishes, 7 (5):222, October 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/5/222.

Liao:2017:MRS

[LSY^{+}17] Qiuhong Liao, Tetsuhito Suzuki, Kohno Yasushi, Dimas Firmanda Al Riza, Makoto Kuramoto, and Naoshi Kondo. Monitoring Red Sea bream scale fluorescence as a freshness indicator. Fishes, 2 (3):10, July 10, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/3/10.

Liu:2023:LFG

[LSY^{+}23] Kai Liu, Zhaoxiang Sun, Chun Yang, Li Jan Lo, and Jun Chen. Lossof-function of xpc sensitizes zebrafish to ultraviolet irradiation. Fishes, 8(4):191, April 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/191.

Li:2022:DRD

[LTZ^{+}22] Xiaomeng Li, Yuanzhen Tan, Zheng Zhang, Yupeng Huang, Pengfei Mu, Zhengwei Cui, and Xinhua Chen. Development of recombinant dihydrolipoamide dehydrogenase subunit vaccine against Vibrio infection in large yellow croaker. Fishes, 7(1): 17, January 11, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/17.

Lacy:2018:IRT

[LUM18] Shaw Lacy, Fernando Ugalde, and Luca Mao. Invasive rainbow trout (Oncorhynchus mykiss) are not affected by different land uses in a multi-use, Mediterranean climate landscape. Fishes, 3(4): 37, September 26, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/4/37.

Liontakis:2023:EVE

[LV23] Angelos Liontakis and Vassiliki Vassilopoulou. An economic view on the effects of invasive rabbitfishes based on fishers’ perspectives: The case of the parrotfish métier in the South Ionian Sea. Fishes, 8(9): 447, September 04, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/447.

Lopez:2020:SMC

[LVB^{+}20] Annalaura Lopez, Mauro Vasconi, Federica Bellagamba, Tiziana Mentasti, and Vittorio Maria Moretti. Sturgeon meat and caviar quality from different cultured species. Fishes, 5 (1):9, February 07, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/1/9.

Luer:2018:PHH

[LW18] Carl A. Luer and Catherine J. Walsh. Potential human health applications from marine biomedical research with elasmobranch fishes. Fishes, 3(4):47, December 06, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/4/47.

Li:2023:SDM

[LWD^{+}23] Menghui Li, Xuehui Wang, Feiyan Du, Shuai Peng, Lianggen Wang, Dianrong Sun, Yuezhong Wang, Pimao Chen, and Yongsong Qiu. Spatial distribution of major fish species catches and their relationship with environmental factors in the Beibu Gulf, South China Sea. Fishes, 8(11):559, November 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/559.

Li:2023:PPD

[LWG^{+}23] Jiyuan Li, Yanfeng Wang, Teng Guo, Shihong Xu, Guang Gao, Feng Liu, Xiaoyang Guo, Yanduo Wu, Haixia Zhao, and Jun Li. Phenethylamine is a potential density stress pheromone in turbot (Scophthalmus maximus). Fishes, 8(10):506, October 11, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/10/506.

Liu:2023:SRG

[LWL^{+}23] Zhilong Liu, Tuo Wang, Peng Liu, Dongneng Jiang, Xi Liu, Siping Deng, Tianli Wu, Yang Huang, Chunhua Zhu, Guangli Li, and Mouyan Jiang. Selection of reference gene for expression studies in the ovary and pituitary of spotted scat (Scatophagus argus) at different ovarian stages. Fishes, 8(2): 120, February 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/120.

Liang:2021:MCG

[LWS^{+}21] Xiao Liang, Ying Wan, Zhiyuan Shen, Yanmei Liu, Dapeng Li, Li Li, Rong Tang, and Xi Zhang. Molecular characterization of Hsp47 in grass carp (Ctenopharyngodon idella) and its correlation with Type I collagen in response to fish aerobic exercise. Fishes, 6(2): 17, April 23, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/2/17.

Li:2023:DDI

[LWS^{+}23a] Yu-Jiao Li, Xin-Liang Wang, Ling-Yu Shi, Zong-Yi Wang, Zi-Ang Zhao, Shu-Chao Ge, and Bing Hu. DUSP2 deletion inhibits macrophage migration by inhibiting ERK activation in zebrafish. Fishes, 8(6): 310, June 12, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/310.

Liu:2023:SES

[LWS^{+}23b] Xiaochen Liu, Daixia Wang, Yan Shang, Xuee Yu, Baoquan Gao, Jianjian Lv, Jitao Li, Ping Liu, Jian Li, and Xianliang Meng. Survival, energy status, and cellular stress responses of the juvenile swimming crab Portunus trituberculatus under acute nitrite stress. Fishes, 8 (4):215, April 19, 2023.

CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/215.

Linh:2024:BSF

[LWT^{+}24] Nguyen Vu Linh, Supreya Wannavijit, Khambou Tayyamath, Nguyen Dinh-Hung, Thitikorn Nititanarapee, Md Afsar Ahmed Sumon, Orranee Srinual, Patima Permpoonpattana, Hien Van Doan, and Christopher L. Brown. Black soldier fly (Hermetia illucens) larvae meal: a sustainable alternative to fish meal proven to promote growth and immunity in koi carp (Cyprinus carpio var. koi). Fishes, 9(2):53, January 28, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/2/53.

Li:2022:EAA

[LWZ^{+}22] Mingshuai Li, Shihui Wang, Zhigang Zhao, Liang Luo, Rui Zhang, Kun Guo, Lanlan Zhang, and Yuhong Yang. Effects of alkalinity on the antioxidant capacity, nonspecific immune response and tissue structure of Chinese mitten crab Eriocheir sinensis. Fishes, 7(4): 206, August 18, 2022.

CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/206.

Liu:2023:FRU

[LWZ^{+}23] Liang Liu, Junfeng Wu, Tao Zheng, Haiyan Zhao, Han Kong, Boyu Qu, and Hong Yu. Fish recognition in the underwater environment using an improved ArcFace loss for precision aquaculture. Fishes, 8(12): 591, November 30, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/591.

Liu:2022:TAJ

[LXT^{+}22] Hongbo Liu, Junren Xue, Jing Tang, Tao Jiang, Xiubao Chen, and Jian Yang. Taste attributes of the “June Hairy Crab” juveniles of Chinese mitten crab (Eriocheir sinensis) in Yangcheng Lake, China— a pilot study. Fishes, 7 (3):128, May 31, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/128.

Li:2024:DMR

[LXW^{+}24] Feipeng Li, Siyu Xie, Mingzhu Wang, Ling Chen, and Haixiang Yu. Distribution and management of residual antibiotics in the Litopenaeus vannamei shrimp farming environment: Recommendations for effective control. Fishes, 9(3):84, February 23, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/84.

Lu:2022:CES

[LXZ^{+}22] Min Lu, Zenghou Xing, Yurui Zhou, Youhou Xu, Huijing Peng, Jie Zou, Solomon Felix Dan, Zhicai She, Pengliang Wang, Jinfeng Liu, Shaomin Qin, Jialin Yang, and Peng Zhu. Cloning and expression of Sox2 and Sox9 in embryonic and gonadal development of Lutraria sieboldii. Fishes, 7(6): 392, December 16, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/392.

Li:2023:NDD

[LYL^{+}23] Yaxi Li, Zhi Yan, Ainuo Lin, Xiaodong Li, and Ke Li. Non-dose-dependent relationship between antipredator behavior and conspecific alarm substance in zebrafish. Fishes, 8(2): 76, January 28, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/76.

Lin:2024:CCF

[LYL^{+}24a] Jinji Lin, Ziqi Yu, Yang Leng, Jiexiong Zhu, Feifei Yu, Yishan Lu, Jiayu Chen, Wenhao He, Yixin Zhang, and Yaoshen Wen. Cloning, characterization and functional analysis of caspase 8-like gene in apoptosis of crassostrea hongkongensis response to hyper-salinity stress. Fishes, 9(5):172, May 9, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/172.

Liu:2024:CAE

[LYL^{+}24b] Dongge Liu, Hao Yang, Shuisheng Li, Hai Huang, Guangli Li, and Huapu Chen. Comparative analysis of enzymatic activities and transcriptional profiles of various hepatic enzymes between male and female yellowfin tuna (Thunnus albacares). Fishes, 9 (5):184, May 17, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/184.

Liu:2024:TAR

[LYW^{+}24] Jun Liu, Songyue You, Yuting Wang, Jie Bao, and Hongbo Jiang. Transcriptional analysis reveals the iron regulation network of the pathogenic yeast Metschnikowia bicuspidata in response to iron stress. Fishes, 9(6):236, June 18, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/236.

Liu:2024:HHA

[LYY^{+}24] Jialong Liu, Miaomiao Yin, Zhi Ye, Jingjie Hu, and Zhenmin Bao. Harnessing hue: Advances and applications of fish skin pigmentation genetics in aquaculture. Fishes, 9(6): 220, June 10, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/220.

Li:2023:OUO

[LZC^{+}23a] Changwei Li, Xiaoyu Zhang, Yu Chen, Shiyu Zhang, Limin Dai, Wenjing Zhu, and Yuan Chen. Optimized utilization of organic carbon in aquaculture biofloc systems: a review. Fishes, 8(9): 465, September 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/465.

Luo:2023:GDJ

[LZC^{+}23b] Yang Luo, Yufeng Zhang, Ruli Cheng, Qinghua Li, Yu Zhang, Yingwen Li, and Yanjun Shen. Genetic diversity of Jinshaia sinensis (Cypriniformes, Balitoridae) distributed upstream of the Yangtze River. Fishes, 8 (2):75, January 28, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/75.

Lin:2023:EPY

[LZH^{+}23] Hao Lin, Sishun Zhou, Zhangfan Huang, Jianrong Ma, Lumin Kong, Yi Lin, Zhongying Long, Huihui Qin, Longhui Liu, Yanbo Zhao, and Zhongbao Li. The effects of Porphyra yezoensis polysaccharides on intestinal health of spotted sea bass, Lateolabrax maculatus. Fishes, 8(8): 419, August 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/419.

Luo:2023:ENI

[LZL^{+}23] Zhengli Luo, Yanbo Zhou, Ning Liu, Liangming Wang, Yan Liu, Binbin Shan, Manting Liu, Cheng Chen, Changping Yang, and Dianrong Sun. Ecological niche and interspecific association of the main fishes in the coastal waters of Hainan Island, China. Fishes, 8(10): 511, October 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/511.

Liu:2023:SDF

[LZM^{+}23] Jiahao Liu, Feiran Zhang, Qiang Ma, Yuliang Wei, Mengqing Liang, and Houguo Xu. Sex differences in fatty acid composition of Chinese tongue sole (Cynoglossus semilaevis) tissues. Fishes, 8(8):421, August 17, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/421.

Liu:2024:TAJ

[LZW^{+}24] Changlin Liu, Zheng Zhang, Shouyong Wei, Wenjie Xiao, Chao Zhao, Yue Wang, and Liguo Yang. Transcriptome analysis of juvenile black rockfish Sebastes schlegelii under air exposure stress. Fishes, 9 (6):239, June 19, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/239.

Li:2022:CDS

[LZZ^{+}22] Haoyu Li, Xiaonan Zhang, Yang Zhang, Qi Liu, Fengwen Liu, Donglin Li, and Hucai Zhang. Climatedriven synchrony in anchovy fluctuations: A Pacific-wide comparison. Fishes, 7(4):193, August 05, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/193.

Martinez-Antequera:2022:SSH

[MABÁMSM22] Francisca P. Martínez-Antequera, Isabel Barranco-Ávila, Juan A. Martos-Sitcha, and Francisco J. Moyano. Solid-state hydrolysis (SSH) improves the nutritional value of plant ingredients in the diet of Mugil cephalus. Fishes, 7(1):4, December 25, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/4.

Majumdar:2018:CPM

[MAR^{+}18] Bhaskar Chandra Majumdar, Faria Afrin, Md. Golam Rasul, Dinesh Chandra Shaha, and A. K. M. Azad Shah. Changes in physicochemical, microbiological, and sensory properties of sun-dried Mystus vittatus during storage at ambient temperature. Fishes, 3(3): 32, August 20, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/32.

Martinez:2019:EEA

[MBAM19] Francisca P. Martínez, Laura Bermúdez, María J. Aznar, and Francisco J. Moyano. Evaluation of enzyme additives on the nutritional use of feeds with a high content of plant ingredients for Mugil cephalus. Fishes, 4(4):56, December 04, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/56.

Madia:2024:RAF

[MBC^{+}24] Manfredi Madia, Massimiliano Bottaro, Tiziana Cillari, Andrea Li Vorsi, Luca Castriota, Maria Rita Amico, Sergio Bizzarri, Teresa Maggio, Manuela Falautano, Michele Gristina, Ilaria Di Lauro, Floriana Trova, Patrizia Perzia, Roberto D’Ambra, Enrico Casola, and Mauro Sinopoli. Reducing artisanal fishery impact on marine community: New data from comparison of innovative and traditional gear. Fishes, 9 (5):171, May 8, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/171.

Morales:2023:OMS

[MBD^{+}23] Christian James C.

Morales, Kyle Dominic E. Barnuevo, Emmanuel S. Delloro, Jr., Roxanne A. Cabebe-Barnuevo, Jenylle Kate S. Calizo, Sanny David P. Lumayno, and Ricardo P. Babaran. Otolith morphometric and shape distinction of three redfin species under the genus Decapterus (Teleostei: Carangidae) from Sulu Sea, Philippines. Fishes, 8(2): 95, February 05, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/95.

Mulokozi:2020:EEA

[MBL20] Deogratias Pius Mulokozi, Håkan Berg, and Torbjörn Lundh. An ecological and economical assessment of integrated amaranth (Amaranthus hybridus) and Nile tilapia (Oreochromis niloticus) farming in Dar es Salaam, Tanzania. Fishes, 5(3): 30, September 18, 2020. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/3/30.

Mullowney:2024:MMI

[MBP^{+}24] Darrell Mullowney, Krista Baker, Julia Pantin, William Coffey, Kyle Lefort, and Katie Morrissey. Merits of multiindicator precautionary approach management in a male-only crab fishery. Fishes, 9(5):187, May 18, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/5/187.

Martinez-Bautista:2024:GBM

[MBPB24] Gil Martinez-Bautista, Pamela Padilla, and Warren W. Burggren. Genetic basis for morphological variation in the zebrafish Danio rerio: Insights from a low-heterozygosity line. Fishes, 9(5):164, May 2, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/5/164.

Mukherjee:2021:EPG

[MBZ^{+}21] Subham Mukherjee, Oldřich Bartoš, Kamila Zdeňková, Petr Hanák, Petra Horká, and Zuzana Musilova. Evolution of the parvalbumin genes in teleost fishes after the wholegenome duplication. Fishes, 6(4):70, December 01, 2021. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/6/4/70.

Martinez-Cardenas:2017:PCD

[MCÁGHA^{+}17] Leonardo Martínez-Cárdenas, Carlos A. Álvarez-González, Oscar U. Hernández-Almeida, Carlos A. Frías-Quintana, Jesús T. Ponce-Palafox, and Sergio Castillo-Vargasmachuca. Partial characterization of digestive proteases in the green cichlid, Cichlasoma beani. Fishes, 2 (1):4, March 08, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/1/4.

Moreira:2019:IAS

[MCSB^{+}19] Márcio Moreira, Anaísa Cordeiro-Silva, Marisa Barata, Pedro Pousão-Ferreira, and Florbela Soares. Influence of age on stress responses of white seabream to amyloodiniosis. Fishes, 4 (2):26, April 08, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/2/26.

Martens:2021:ECF

[MD21] Kyle D. Martens and Jason Dunham. Evaluating coexistence of fish species with coastal cutthroat trout in low order streams of Western Oregon and Washington, USA. Fishes, 6 (1):4, January 30, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/1/4.

Mazzucato:2023:OBM

[MDF^{+}23] Matteo Mazzucato, Tiziano Dorotea, Eleonora Franzago, Paolo Mulatti, Giulio Marchetti, Claudia Casarotto, Andrea Fabris, Nicola Ferrè, Anna Toffan, Andrea Marsella, Azzurra Callegaro, Grazia Manca, and Manuela Dalla Pozza. Overview on the biosecurity measures of salmonid fish farms: a case study in Italy. Fishes, 8(11): 554, November 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/554.

Moura:2022:RPS

[MDLA22] Ana Moura, Ester Dias, Rodrigo López, and Carlos Antunes. Regional population structure of the European eel at the southern limit of its distribution revealed by otolith shape signature. Fishes, 7(3):135, June 07, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/135.

MacedoSantana:2023:FVS

[MdM^{+}23] Thiago Macedo Santana, Francisco de Matos Dantas, Driely Kathriny Monteiro Dos Santos, Juliana Tomomi Kojima, Yugo Moraes Pastrana, Rogério Souza De Jesus, and Ligia Uribe Gonçalves. Fish viscera silage: Production, characterization, and digestibility of nutrients and energy for tambaqui juveniles. Fishes, 8(2): 111, February 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/111.

Molina:2023:HTS

[MDVM^{+}23] Alfredo Molina, Phillip Dettleff, Valentina Valenzuela-Muñoz, Cristian Gallardo-Escarate, and Juan Antonio Valdés. Hightemperature stress induces autophagy in rainbow trout skeletal muscle. Fishes, 8(6):303, June 06, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/303.

Meerbeek:2024:PBS

[Mee24] Jonathan R. Meerbeek.

Poststocking behavior and survival of large yearling muskellunge (Esox masquinongy) in two Northern Iowa natural lakes. Fishes, 9 (6):216, June 6, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/216.

Muslim:2023:SSM

[MFKS23] Ahmad Imam Muslim, Miho Fujimura, Tsuji Kazunari, and Muslim Salam. Small-scale marine fishers’ possession of fishing vessels and their impact on net income levels: a case study in Takalar District, South Sulawesi Province, Indonesia. Fishes, 8(9): 463, September 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/463.

Mouchlianitis:2024:ESE

[MGMG24] Foivos A. Mouchlianitis, Maria Garagouni, George Minos, and Kostas Ganias. Evaluating the sustainability of an Eastern Mediterranean gillnet fishery based on the catches of undersized individuals and the reproductive period of targeted species. Fishes, 9(4): 122, March 29, 2024.

CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/122.

Miao:2023:IFQ

[MGS^{+}23] Xiaomin Miao, Hui Guo, Yong Song, Chunying Du, Jingyun Feng, Yixi Tao, Hao Xu, and Yun Li. Improvement of flesh quality of farmed silver carp (Hypophthalmichthys molitrix) by short-term stocked in natural water. Fishes, 8(3):142, February 28, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/3/142.

Mundahl:2023:CRF

[MH23] Neal D. Mundahl and Kelsey A. Hoffmann. Condition, reproductive fitness, and fluctuating asymmetry in brook stickleback: Responses to anthropogenic runoff. Fishes, 8(11):557, November 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/557.

Mokhtar:2023:GMF

[MHZ^{+}23] Doaa M. Mokhtar, Marwa M.

Hussein, Giacomo Zaccone, Alessio Alesci, Eugenia Rita Lauriano, and Ramy K. A. Sayed. Gills of molly fish: a potential role in neuro-immune interaction. Fishes, 8 (4):195, April 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/195.

Martinez:2023:EPT

[MIHH23] Enrique Pino Martinez, Albert Kjartan Dagbjartarson Imsland, Anne-Camilla Diesen Hosfeld, and Sigurd Olav Handeland. Effect of photoperiod and transfer time on Atlantic salmon smolt quality and growth in freshwater and seawater aquaculture systems. Fishes, 8(4):212, April 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/212.

Miller:2023:YAH

[Mil23] Michael J. Miller. 43 years after H. G. Moser’s seminal “Morphological and Functional Aspects of Marine Fish Larvae”: The commonalities of leptocephali and larvae of other marine teleosts. Fishes, 8(11): 548, November 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/548.

Ma:2024:VCG

[MJL^{+}24] Xueyan Ma, Wu Jin, Guohua Lv, Wanwen Chen, Dongpo Xu, Pao Xu, Dan Hua, and Haibo Wen. In vitro culture of glochidia and morphological changes in juveniles of the endangered freshwater mussel Solenaia oleivora. Fishes, 9 (2):49, January 27, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/49.

Ma:2022:CEP

[MKC^{+}22] Qian Ma, Jiehua Kuang, Gang Chen, Jiandong Zhang, Jiansheng Huang, Feifan Mao, and Qiling Zhou. Cloning and expression profiling of the gene vasa during first annual gonadal development of cobia (Rachycentron canadum). Fishes, 7(2):60, April 10, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/2/60.

Marteinson:2024:DRV

[MKH24] Sarah Marteinson, Melanie Kingsbury, and Dounia Hamoutene. Data review on the variation in sensitivity to aquaculture chemotherapeutants in some crustacean life stages. Fishes, 9 (7):273, July 11, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/273.

Morphis:2016:AVA

[MKN^{+}16] Gregory Morphis, Aggeliki Kyriazopoulou, Constantina Nasopoulou, Eleni Sioriki, Constantinos A. Demopoulos, and Ioannis Zabetakis. Assessment of the in vitro antithrombotic properties of sardine (Sardina pilchardus) fillet lipids and cod liver oil. Fishes, 1(1):1– 15, September 28, 2016. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/1/1/1.

Mmanda:2020:DLF

[MLH^{+}20] Francis Pius Mmanda, Jan Erik Lindberg, Anna Norman Haldén, Matern S. P. Mtolera, Rukia Kitula, and Torbjörn Lundh. Digestibility of local feed ingredients in tilapia Oreochromis niloticus juveniles, determined on faeces collected by siphoning or stripping. Fishes, 5(4):32, October 15, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/4/32.

Matsubara:2019:CSS

[MLK^{+}19] Hajime Matsubara, P. Mark Lokman, Yukinori Kazeto, Hiromi Okumura, Shigeho Ijiri, Toshiaki Hirai, Graham Young, Shinji Adachi, and Kohei Yamauchi. Changes in sex steroids and ovarian steroidogenic enzyme mRNA levels in artificially maturing Japanese eel (Anguilla japonica) and naturally maturing new Zealand longfin eel (Anguilla dieffenbachii) during vitellogenesis. Fishes, 4(4): 52, October 17, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/52.

Munoz-Lechuga:2023:DSU

[MLSC^{+}23] Rubén Muñoz-Lechuga, Fambaye Ngom Sow, Diaha N’Guessan Constance, Davy Angueko, David Macías, Alexia Massa-Gallucci, Guelson Batista da Silva, Jorge M. S. Gonçalves, and Pedro G. Lino. Differentiation of spatial units of genus Euthynnus from the Eastern Atlantic and the Mediterranean using otolith shape analysis. Fishes, 8(6):317, June 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/6/317.

Mao:2021:MCN

[MLZ^{+}21] Zhuangwen Mao, Shengwei Luo, Dafang Zhao, Xiang Zhou, Zilong Zhang, Yangbo Xiao, Shenping Cao, Yonghua Zhou, Shaojun Liu, Jianzhou Tang, and Zhen Liu. Molecular characterization and nutrition regulation of the neutral amino acid transporter ASCT2 in triploid crucian carp. Fishes, 6(4):77, December 09, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/4/77.

Miqueleiz:2022:CSG

[MMAO22] Imanol Miqueleiz, Rafael Miranda, Arturo Hugo Ariño, and Elena Ojea. Conservation-status gaps for marine top-fished commercial species. Fishes, 7(1):2, December 23, 2022. CODEN ???? ISSN 2410-3888. URL

                                        https://www.mdpi.com/
                                        2410-3888/7/1/2.

Morick:2023:MIP

[MMD^{+}23] Danny Morick, Yuval Maron, Nadav Davidovich, Ziv Zemah-Shamir, Yaarit Nachum-Biala, Peleg Itay, Natascha Wosnick, Dan Tchernov, and Shimon Harrus. Molecular identification of Photobacterium damselae in wild marine fish from the Eastern Mediterranean Sea. Fishes, 8(2):60, January 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/60.

Muringai:2022:SSA

[MML22] Rodney Tatenda Muringai, Paramu Mafongoya, and Romano Trent Lottering. Sub-Saharan Africa freshwater fisheries under climate change: a review of impacts, adaptation, and mitigation measures. Fishes, 7 (3):131, June 02, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/131.

Mori:2024:GMH

[MMM^{+}24] Natacha C. Mori, Bruna T.

Michelotti, Caio C. F. Magnotti, Caroline A. Bressan, Letícia B. Bianchin, Fernando J. Sutili, Ana Paula G. Almeida, Luiz C. Kreutz, Maria A. Pavanato, Vinicius R. Cerqueira, Silvio T. da Costa, Bernardo Baldisserotto, and Berta M. Heinzmann. Growth, metabolic, hepatic and redox parameters, digestive enzymes and innate immunity in Mugil liza fed a citral-supplemented diet. Fishes, 9(6):240, June 20, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/6/240.

Madariaga-Mendoza:2023:EIM

[MMMNAG23] Diana Madariaga-Mendoza, José Marrugo-Negrete, and Víctor Atencio-García. Effect of inorganic mercury on semen quality, embryo and larval development of bocachico Prochilodus magdalenae. Fishes, 8(9):445, August 31, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/445.

Mukhtar:2021:SMP

[MMSK21] Mutia Kamalia Mukhtar, Masita Dwi Mandini Manessa, Supriatna Supriatna, and Liya Tri Khikmawati. Spatial modeling of potential lobster harvest grounds in Palabuhanratu Bay, West Java, Indonesia. Fishes, 6(2):16, April 21, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/2/16.

Mendez-Martinez:2023:GPF

[MMVVCJ^{+}23] Yuniel Méndez-Martínez, Alan Rodrigo Vera-Veliz, Edilmar Cortés-Jacinto, Yanis Cruz-Quintana, Aroldo Botello-Leon, Pedro Daniel Mendoza-Carranza, and Natalia S. Calvo. Growth performance, feed utilisation, digestive and metabolic enzyme activity, and liver morphohistology in hybrid tilapia (Oreochromis mossambicus × Oreochromis niloticus) juveniles fed with the inclusion of chitosan in their diet. Fishes, 8(11): 546, November 09, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/546.

Mozanzadeh:2017:MRS

[MMY^{+}17] Mansour Torfi Mozanzadeh, Jasem G. Marammazi, Morteza Yaghoubi, Naser Agh, Esmaeil Pagheh, and Enric Gisbert. Macronutrient requirements of silveryblack porgy (Sparidentex hasta): a comparison with other farmed sparid species. Fishes, 2(2):5, May 13, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/2/5.

Miura:2022:ESD

[MNO^{+}22] Takeshi Miura, Munenori Nishikawa, Yuki Otsu, Muhammad Fariz Zahir Ali, Atsushi Hashizume, and Chiemi Miura. The effects of silkwormderived polysaccharide (silkrose) on ectoparasitic infestations in yellowtail (Seriola quinqueradiata) and white trevally (Pseudocaranx dentex). Fishes, 7(1): 14, January 09, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/14.

Murzina:2016:LFA

[MNP^{+}16] Svetlana A. Murzina, Zinaida A. Nefedova, Svetlana N. Pekkoeva, Tatiana R. Ruokolainen, Pauli O. Ripatti, Andrey V. Semushin, and Nina N. Nemova. Lipids and fatty acids of the White Sea herring Clupea pallasi marisalbi Berg (Clupeiformes, Clupeidae) from different habitats of the White Sea. Fishes, 1(1):65–76, October 13, 2016. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/1/1/65.

Midway:2018:IGE

[MOW^{+}18] Stephen R. Midway, Andrew Ostrowski, Lindsey West, Mario Hernandez, and Matthew D. Robertson. Improved growth estimates for Lethrinus harak: Measuring increments, adjusting ages, and fitting flexible growth models. Fishes, 3 (3):31, August 14, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/31.

Moyano:2018:ESI

[Moy18] Francisco Javier Moyano. Editorial for the special issue on nutritional requirements in new fish species under culture. Fishes, 3(2): 18, March 23, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/2/18.

Murie:2023:AGF

[MPK^{+}23] Debra J. Murie, Daryl C.

Parkyn, Christopher C. Koenig, Felicia C. Coleman, Christopher R. Malinowski, Jessica A. Cusick, and Robert D. Ellis. Age, growth, and functional gonochorism with a twist of diandric protogyny in goliath grouper from the Atlantic Coast of Florida.

Fishes, 8(8):412, August 11, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/412.

Martinez:2018:HKT

[MPM^{+}18] Danixa Martínez, Juan Pablo Pontigo, Francisco J. Morera, Alejandro. Yañéz, and Luis Vargas-Chacoff. Head kidney transcriptome analysis and characterization for the sub-Antarctic notothenioid fish Eleginops maclovinus. Fishes, 3(1):8, January 25, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/1/8.

Martinez:2021:SGD

[MPM^{+}21] Gil Martínez, Emyr Peña, Rafael Martínez, Susana Camarillo, Warren Burggren, and Alfonso Álvarez. Survival, growth, and development in the early stages of the tropical gar Atractosteus tropicus: Developmental critical Windows and the influence of temperature, salinity, and oxygen availability. Fishes, 6(1): 5, February 12, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/1/5.

Mensoor:2018:DHM

[MS18] Montazer Mensoor and Ali Said. Determination of heavy metals in freshwater fishes of the Tigris River in Baghdad. Fishes, 3(2):23, June 13, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/2/23.

McColl:2020:BCC

[MS20] Kenneth A. McColl and Agus Sunarto. Biocontrol of the common carp (Cyprinus carpio) in Australia: a review and future directions. Fishes, 5(2):17, June 02, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/2/17.

Mendes:2024:SHM

[MSA24] Hugo Mendes, Cristina Silva, and Manuela Azevedo. Southern horse mackerel (Trachurus trachurus) spatio-temporal distribution patterns based on fine-scale resolution data. Fishes, 9(3): 93, February 29, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/93.

Majhi:2023:ESD

[MSB^{+}23] Snigdha S. Majhi, Soibam Khogen Singh, Pradyut Biswas, Reshmi Debbarma, Janmejay Parhi, Soibam Ngasotter, Gusheinzed Waikhom, Dharmendra Kumar Meena, Ayam Gangarani Devi, Sudhanshu S. Mahanand, K. A. Martin Xavier, and Arun Bhai Patel. Effect of stocking density on growth, water quality changes and cost efficiency of butter catfish (Ompok bimaculatus) during seed rearing in a biofloc system. Fishes, 8(2):61, January 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/61.

Martin:2021:CPT

[MSK^{+}21] Miriam Martin, Stephen Smith, Michael Kleinhenz, Geraldine Magnin, Zhoumeng Lin, David Kuhn, Shawnee Montgomery, and Johann Coetzee. Comparative pharmacokinetics and tissue concentrations of flunixin meglumine and meloxicam in tilapia (Oreochromis spp.). Fishes, 6(4): 68, November 25, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/68.

Mahu:2022:CRA

[MSK^{+}22] Edem Mahu, Salieu Sanko, Allieubakarr Kamara, Ernest Obeng Chuku, Elizabeth Effah, Zacharie Sohou, Yaovi Zounon, Victoria Akinjogunla, Ruth Oluwatoyin Akinnigbagbe, Hamet Diaw Diadhiou, and Robert Marchant. Climate resilience and adaptation in West African oyster fisheries: An expert-based assessment of the vulnerability of the oyster Crassostrea tulipa to climate change. Fishes, 7(4):205, August 16, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/205.

Madera-Santana:2023:DDC

[MSRGCG^{+}23] Sara Madera-Santana, Carlos Rodríguez-García, Jairo Castro-Gutiérrez, Ángel Rafael Domínguez-Bustos, and Remedios Cabrera-Castro. Discarded but not dismissed: a comprehensive study of the feeding habits of the brown comber (Serranus hepatus, (Linneaus 1758)) in the Gulf of Cádiz (NE Atlantic). Fishes, 8(11): 541, November 02, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/11/541.

Mendez-Tepepa:2023:CEA

[MTHPJS^{+}23] Maribel Méndez-Tepepa, Karla Hernández-Pérez, Libertad Juárez-Santacruz, Senobia Rosalia Cruz-Lumbreras, Edelmira García-Nieto, Arely Anaya-Hernández, and Cuauhtémoc Morales-Cruz. Cytotoxic effects of the atrazine herbicide on erythrocytes and liver damage in Lithobates spectabilis. Fishes, 8(4):207, April 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/207.

Maslov:2019:CAS

[MTM^{+}19] Dmitry L. Maslov, Oxana P. Trifonova, Anton N. Mikhailov, Konstantin V. Zolotarev, Kirill V. Nakhod, Valeriya I. Nakhod, Nataliya F. Belyaeva, Marina V. Mikhailova, Petr G. Lokhov, and Alexander I. Archakov. Comparative analysis of skeletal muscle metabolites of fish with various rates of aging. Fishes, 4 (2):25, March 31, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/2/25.

Mpomwenda:2023:COC

[MTPK23] Veronica Mpomwenda, Tumi Tómasson, Jón Geir Pétursson, and Dadi Mar Kristófersson. From co-operation to coercion in fisheries management: The effects of military intervention on the Nile perch fishery on Lake Victoria in Uganda. Fishes, 8(11): 563, November 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/563.

Miyanishi:2021:ESS

[MU21] Hiroshi Miyanishi and Katsuhisa Uchida. Establishment of a simplified system to evaluate salinity preference and validation of behavioral salinity selection in the Japanese medaka, Oryzias latipes. Fishes, 6(2):18, April 28, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/2/18.

Maytorena-Verdugo:2022:IMO

[MVPMAV^{+}22] Claudia I. Maytorena-Verdugo, Emyr S. Peña-Marín, Carina S. Alvarez-Villagómez, Graciela M. Pérez-Jiménez, César A. Sepúlveda-Quiroz, and Carlos A. Alvarez-González. Inclusion of mannanoligosaccharides in diets for tropical gar Atractosteus tropicus larvae: Effects on growth, digestive enzymes, and expression of intestinal barrier genes. Fishes, 7 (3):127, May 31, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/127.

Monahan:2023:SPF

[MWPS23] Dylan Monahan, Jeff S. Wesner, Stephanie M. Parker, and Hannah Schartel. Spatial patterns in fish assemblages across the National Ecological Observation Network (NEON): the first six years. Fishes, 8(11): 552, November 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/552.  See

correction [MWPS24].

Monahan:2024:CMA

[MWPS24] Dylan Monahan, Jeff S. Wesner, Stephanie M. Parker, and Hannah Schartel. Correction: Monahan et al., Spatial Patterns in Fish Assemblages across the National Ecological Observation Network (NEON): The First Six Years. Fishes 2023, 8, 552. Fishes, 9(2):68, February 08, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/68.  See

[MWPS23].

Mao:2023:IAS

[MWZ^{+}23] Hanping Mao, Bin Wang, Jian Zhao, Yafei Wang, Xiaoxue Du, and Qiang Shi. Influences of aquaponics system on growth performance, antioxidant parameters, stress parameters and gene expression of Carassius auratus. Fishes, 8(7):360, July 11, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/360.

Muhammad:2024:VCA

[MYW^{+}24] Abdullateef Mukhtar Muhammad, Chang Yang, Jingyuan Wang, Xianping Ge, Bo Liu, Linghong Miao, Guodong Gao, and Qunlan Zhou. Vitamin C alleviates intestinal inflammation caused by Aeromonas hydrophila in juvenile blunt snout bream (Megalobrama amblycephala). Fishes, 9(4):129, April 8, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/129.

Ma:2023:TSA

[MYY^{+}23] Yanwen Ma, Yingying Ye, Ronghui Yao, Pengzhi Qi, and Jiji Li. Transcriptome sequencing analysis of sexrelated genes in the gonads of Mytilus unguiculatus. Fishes, 8(9): 456, September 11, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/456.

Mokhtar:2023:MCF

[MZA^{+}23] Doaa M. Mokhtar, Giacomo Zaccone, Alessio Alesci, Michal Kuciel, Manal T. Hussein, and Ramy K. A. Sayed. Main components of fish immunity: an overview of the fish immune system. Fishes, 8(2):93, February 05, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/93.

Naslund:2018:RMB

[Näs18] Joacim Näslund. Relative mass of brainand intestinal tissue in juvenile brown trout: No long-term effects of compensatory growth; with additional notes on emerging sex-differences. Fishes, 3(4):38, September 28, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/4/38.

Nastac:2023:PEK

[NDC^{+}23] Lacrămioara (Grădinariu)

Năstac, Lorena Dediu, Mirela Creţu, Cristian Rîmniceanu, Angelica Docan, Iulia Grecu, Floricel Maricel Dima, Maria Desimira Stroe, and Camelia Vizireanu. The protective effects of korill product on carp fingerlings reared in high densities and challenged with albendazole treatment. Fishes, 8(3): 153, March 04, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/153.

Negrini:2024:PRF

[NdNFK^{+}24a] Celma Negrini, Caio Henrique do Nascimento Ferreira, Rafael Ortiz Kracizy, Rosane Lopes Ferreira, Luana Costa, Marlise Teresinha Mauerwerk, Milena Cia Retcheski, Luisa Helena Cazarolli, Wilson Rogério Boscolo, and Eduardo Luis Cupertino Ballester. Partial replacement of fish meal with protein hydrolysates in the diet of Penaeus vannamei (Boone, 1934) during the nursery phase.

Fishes, 9(2):75, February 15, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/75.

Negrini:2024:DPH

[NdNFK^{+}24b] Celma Negrini, Caio Henrique do Nascimento Ferreira, Rafael Ortiz Kracizy, Rosane Lopes Ferreira, Luana Cardoso dos Santos, Milena Cia Retcheski, Marlise Teresinha Mauerwerk, Luisa Helena Cazarolli, Wilson Rogério Boscolo, and Eduardo Luis Cupertino Ballester. Different protein hydrolysates can be used in the Penaeus vannamei (Boone, 1934) diet as a partial replacement for fish meal during the growout phase. Fishes, 9(2): 73, February 12, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/73.

Nambi:2023:ILC

[NGMR23]  Rebecca  Walugembe

Nambi, Abebe Getahun, Fredrick Jones Muyodi, and Edward Rukuunya. Impacts of land cover changes on catches of Nile perch and Nile tilapia on Lake Kyoga, Uganda. Fishes, 8(3): 158, March 10, 2023.

CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/158.

Ndashe:2023:ARF

[NHC^{+}23] Kunda Ndashe, Bernard Mudenda Hang’ombe, Katendi Changula, John Yabe, Mulemba Tillika Samutela, Mwansa M. Songe, Alexander Shula Kefi, Loziwe Njobvu Chilufya, and Martin Sukkel. An assessment of the risk factors associated with disease outbreaks across tilapia farms in Central and Southern Zambia. Fishes, 8(1):49, January 12, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/49.

Newman:2020:CBI

[NHR20] Raymond M. Newman, Fred G. Henson, and Carl Richards. Competition between invasive ruffe (Gymnocephalus cernua) and native yellow perch (Perca flavescens) in experimental mesocosms. Fishes, 5(4):33, October 17, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/4/33.

Novianto:2019:DAI

[NIN^{+}19] Dian Novianto, Ilham, Chandara Nainggolan, Syarif Syamsuddin, Arief Efendi, Sugianto Halim, Yaser Krisnafi, Muhamad Handri, Abdul Basith, Yusrizal, Erick Nugraha, Suciadi Catur Nugroho, and Bram Setyadji. Developing an abundance index of skipjack tuna (Katsuwonus pelamis) from a coastal drifting gillnet fishery in the southern waters of Indonesia. Fishes, 4(1): 10, February 11, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/1/10.

Nathanailides:2021:PPS

[NKC^{+}21] Cosmas Nathanailides, Markos Kolygas, Konstantina Choremi, Theodoros Mavraganis, Evangelia Gouva, Kosmas Vidalis, and Fotini Athanassopoulou. Probiotics have the potential to significantly mitigate the environmental impact of freshwater fish farms. Fishes, 6(4):76, December 08, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/4/76.

Natesan:2023:ARS

[NLTL23] Balaji Natesan, Chuan-Ming Liu, Van-Dai Ta, and Raymond Liao. Advanced robotic system with keypoint extraction and YOLOv5 object detection algorithm for precise livestock monitoring. Fishes, 8(10): 524, October 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/524.

Nenciu:2023:FDW

[NNL^{+}23] Magda Nenciu, Victor Niţă, Luminiţa Lazăr, Alina Spînu, and Elena Vlăsceanu-Mateescu. Fostering the development of Western Black Sea aquaculture: a scientific case study for finfish cage farming allocated zone designation. Fishes, 8(2): 104, February 09, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/104.

Nguyen:2023:EMG

[NNN23] Luong Trong Nguyen, Khanh Quoc Nguyen, and Toan Phi Nguyen. Experimental mixed gillnets improve catches of narrow-barred Spanish mackerel (Scomberomorus commerson). Fishes, 8(4):210, April 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/210.

Nekrasova:2024:API

[NPT^{+}24] Oksana Nekrasova, Mihails Pupins, Volodymyr Tytar, Leonid Fedorenko, Oleksandr Potrokhov, Arturs Škute, Andris Čeirāns, Kathrin Theissinger, and Jean-Yves Georges. Assessing prospects of integrating Asian carp polyculture in Europe: a nature-based solution under climate change? Fishes, 9(4):148, April 22, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/148.

Nieves-Rodriguez:2018:EGD

[NRÁGPM^{+}18] Karen N. Nieves-Rodríguez, Carlos Alfonso Álvarez-González, Emyr S. Peña-Marín, Fernando Vega-Villasante, Rafael Martínez-García, Susana Camarillo-Coop, Dariel Tovar-Ramírez, Laura T. Guzmán-Villanueva, Karl B. Andree, and Enric Gisbert. Effect of β-glucans in diets on growth, survival, digestive enzyme activity, and immune system and intestinal barrier gene expression for tropical gar (Atractosteus tropicus) juveniles. Fishes, 3(3):27, July 10, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/27.

Nijssen:2019:DIS

[NRKT19] Edwin J. Nijssen, Machiel J.

Reinders, Athanasios Krystallis, and Gemma Tacken. Developing an internationalization strategy using diffusion modeling: the case of greater amberjack. Fishes, 4(1):12, February 16, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/12.

Nuswantoro:2023:EPE

[NSK^{+}23] Soko Nuswantoro, Tzu-Yuan Sung, Meki Kurniawan, Tsung-Meng Wu, Bonien Chen, and Ming-Chang Hong. Effects of phosphate-enriched nutrient in the polyculture of Nile tilapia and freshwater prawn in an aquaponic system. Fishes, 8(2):81, January 30, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/81.

Nousias:2021:GVP

[NTP^{+}21] Orestis Nousias, Konstantinos Tzokas, Leonidas Papaharisis, Katerina Ekonomaki, Dimitrios Chatziplis, Costas Batargias, and Costas S. Tsigenopoulos. Genetic variability, population structure, and relatedness analysis of meagre stocks as an informative basis for new breeding schemes. Fishes, 6(4): 78, December 10, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/4/78.

Nguyen:2021:BPF

[NWN^{+}21] Van Bon Nguyen, San-Lang Wang, Anh Dzung Nguyen, Tu Quy Phan, Kuaanan Techato, and Siriporn Pradit. Bioproduction of prodigiosin from fishery processing waste shrimp heads and evaluation of its potential bioactivities. Fishes, 6(3):30, August 11, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/3/30.

Nguyen:2022:UFP

[NWN^{+}22] Thi Hanh Nguyen, San-Lang Wang, Thi Huyen Nguyen, Manh Dung Doan, Thi Ha Trang Tran, Van Anh Ngo, Nhat Duoc Ho, Thi Ngoc Tran, Chien Thang Doan, Van Chung Do, Anh Dzung Nguyen, and Van Bon Nguyen. Utilization of fisheryprocessing by-product squid pens for scale-up production of phenazines via microbial conversion and its novel potential antinematode effect. Fishes, 7(3):113, May 15, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/113.

Nguyen:2023:RFH

[NWN^{+}23] Van Bon Nguyen, San-Lang Wang, Thi Hanh Nguyen, Tu Quy Phan, Thi Huyen Nguyen, Thi Ha Trang Tran, Manh Dung Doan, Van Anh Ngo, and Anh Dzung Nguyen. Recycling fish heads for the production of prodigiosin, a novel fungicide via experimental and molecular docking characterization. Fishes, 8(9): 468, September 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/468.

Nunes:2023:BSF

[NYS^{+}23] Alberto J. P. Nunes, Hiroshi Yamamoto, João Paulo Simões, João Luiz Pisa, Nelson Miyamoto, and Jordana Sampaio Leite. The black soldier fly (Hermetia illucens) larvae meal can costeffectively replace fish meal in practical nursery diets for post-larval Penaeus vannamei under high-density culture. Fishes, 8(12):605, December 10, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/605.

Nikiforidou:2020:VMM

[NZVB20] Vasiliki Nikiforidou, Stefanos Zaoutsos, Nikolaos Vlahos, and Panagiotis Berillis. Vertebrae morphometric measurement and Ca/P levels of different age European seabass (Dicentrarchus labrax). Fishes, 5(4): 37, December 08, 2020. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/4/37.

Office:2018:ARF

[Off18] Fishes Editorial Office. Acknowledgement to reviewers of Fishes in 2017. Fishes, 3(1):7, January 25, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/7.

Office:2019:ARF

[Off19] Fishes Editorial Office.

Acknowledgement to reviewers of Fishes in 2018. Fishes, 4(1):4, January 29, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/4.

Office:2020:ARF

[Off20] Fishes Editorial Office.

Acknowledgement to reviewers of fishes in 2019. Fishes, 5(1):8, February 04, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/1/8.

Office:2021:ARF

[Off21] Fishes Editorial Office.

Acknowledgment to reviewers of fishes in 2020. Fishes, 6(1):3, January 27, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/1/3.

Office:2023:ARF

[Off23] Fishes Editorial Office.

Acknowledgment to the reviewers of Fishes in 2022. Fishes, 8(2):58, January 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/58.

Oliveira:2017:PRD

[OGMG^{+}17] Catarina C. V. Oliveira, Mayra I. Grano-Maldonado, Rui A. Gonçalves, Paulo A. Frias, and António V. Sykes. Preliminary results on the daily and seasonal rhythms of cuttlefish Sepia officinalis (Linnaeus, 1758) locomotor activity in captivity. Fishes, 2(3):9, June 28, 2017. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/3/9.

Osowski:2023:DHH

[OJC^{+}23] Ana R. Osowski, Matthew B. Jargowsky, Pearce T. Cooper, Sean P. Powers, and J. Marcus Drymon. Dietary habits of hardhead (Ariopsis felis) and gafftopsail (Bagre marinus) catfish revealed through DNA barcoding of stomach contents. Fishes, 8(11):539, October 31, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/539.

Olsson:2019:PCT

[Ols19] Jens Olsson. Past and current trends of coastal predatory fish in the Baltic Sea with a focus on perch, pike, and pikeperch. Fishes, 4(1):

7, February 06, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/7.

Oliveira:2019:INN

[OMC^{+}19] Diogo Oliveira, André M.

Machado, Tiago Cardoso, Mónica Lopes-Marques, L. Filipe C. Castro, and Raquel Ruivo. Identification of a novel nucleobaseascorbate transporter family member in fish and amphibians. Fishes, 4(1):1, January 01, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/1.

Onomu:2024:RFF

[OO24] Abigail John Onomu and Grace Emily Okuthe. The role of functional feed additives in enhancing aquaculture sustainability. Fishes, 9 (5):167, May 5, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/167.

Okeke-Ogbuafor:2024:UOA

[OOG24] Nwamaka Okeke-Ogbuafor and Tim Gray. Unregulated open-access versus regulated openaccess fishing: Stakeholders perceptions in Sierra Leone. Fishes, 9 (7):282, July 15, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/7/282.

Okeke-Ogbuafor:2023:PSP

[OOGAS23] Nwamaka Okeke-Ogbuafor, Tim Gray, Kelechi Ani, and Selina Stead. Proposed solutions to the problems of the Lake Chad fisheries: Resilience lessons for Africa? Fishes, 8(2): 64, January 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/64.

Oliveira:2024:TBH

[OOTC24] Joana Oliveira, Aires Oliva-Teles, and Ana Couto. Tracking biomarkers for the health and welfare of aquaculture fish. Fishes, 9(7):289, July 19, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/7/289.

Ord:2019:ISP

[Ord19]  James Ord.  Ionic

stress prompts premature hatching of zebrafish (Danio rerio) embryos. Fishes, 4(1): 20, March 13, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/20.

Osorio:2023:PCB

[OSM23] Brendon James Osorio, Grzegorz Skrzypek, and Mark Meekan. Parasitic copepods as biochemical tracers of foraging patterns and dietary shifts in whale sharks (Rhincodon typus Smith, 1828). Fishes, 8 (5):261, May 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/261.

Pedescoll:2019:PPW

[PAMG19] Anna Pedescoll, Rafael Aguado, Carlos Marcos, and Gustavo González. Performance of a pool and weir fishway for Iberian cyprinids migration: a case study. Fishes, 4(3):45, August 16, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/45.

Paolucci:2023:FNF

[Pao23] Marina Paolucci. Fish nutrition and feed technology. Fishes, 8(3): 146, February 28, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/146.

Pena-Almaraz:2024:BCE

[PAVCCJVV24] Omar Alejandro Peña-Almaraz, Manuel Alejandro Vargas-Ceballos, Edilmar Cortés-Jacinto, and Fernando Vega-Villasante. Biochemical composition of eggs, larvae and tissues of Macrobrachium tenellum females fed diets with different lipid and protein levels. Fishes, 9 (4):145, April 21, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/145.

Petriki:2024:UGI

[PB24] Olga Petriki and Dimitra C. Bobori. Unraveling Greek inland competitive fishing: Historical insights, angler profiles, and motivations through limited data integration in recreational fishing research. Fishes, 9(7):278, July 13, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/278.

Pan:2022:CFP

[PBS^{+}22] Weicong Pan, Soottawat Benjakul, Chiara Sanmartin, Alessandra Guidi, Xiaoguo Ying, Lukai Ma, Xudong Weng, Jin Yu, and Shanggui Deng. Characterization of the flavor profile of bigeye tuna slices treated by cold plasma using E-Nose and GC-IMS. Fishes, 7 (1):13, January 08, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/13.

Pereira:2024:FPE

[PCH^{+}24] Luciana Alves Pereira, Leandro Castello, Eric Hallerman, Edson Rubens Ferreira Rodrigues, Carolina Rodrigues da Costa Doria, and Fabrice Duponchelle. Flood pulse effects on the growth of Pseudoplatystoma fasciatum in the Amazon Basin. Fishes, 9(6):223, June 12, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/223.

Pereira:2023:SEC

[PCO^{+}23] Luciana A. Pereira, Leandro Castello, Donald J. Orth, Fabrice Duponchelle, and Eric M. Hallerman. A synthesis of the ecology and conservation of Pseudoplatystoma catfishes in the neotropics. Fishes, 8(6):306, June 07, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/6/306.

Poisson:2024:PIH

[PEP24] François Poisson, Jim R. Ellis, and Sophy R. Mc-Cully Phillips. Preliminary insights on the habitat use and vertical movements of the pelagic stingray (Pteroplatytrygon violacea) in the Western Mediterranean Sea. Fishes, 9 (6):238, June 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/6/238.

Pimentel:2020:SLP

[PFM^{+}20] Marta S. Pimentel, Filipa Faleiro, Jorge Machado, Pedro Pousão-Ferreira, and Rui Rosa. Seabream larval physiology under ocean warming and acidification. Fishes, 5(1):1, December 20, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/1/1.

Puvanendran:2023:SLT

[PH23] Velmurugu Puvanendran and Øyvind J. Hansen. Short and longterm effects of early and late weaning on Atlantic cod, Gadus morhua. Fishes, 8(6):

312, June 13, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/312.

Peele:2023:EIT

[PHB^{+}23] Emily E. Peele, Charlie Huveneers, Culum Brown, Connor R. Gervais, and Kara E. Yopak. Effects of increased temperature on brain and sensory development in the Port Jackson shark (Heterodontus portusjacksoni). Fishes, 8(12): 611, December 17, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/611.

Pizzul:2022:ENT

[Piz22] Elisabetta Pizzul. Editorial: New trends in freshwater fishes. Fishes, 7(6):388, December 13, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/388.

Perez-Jimenez:2022:IFD

[PJPMMV^{+}22] Graciela M. Pérez-Jiménez, Emyr Saul Peña-Marín, Claudia I. Maytorena-Verdugo, Cesar Antonio Sepúlveda-Quiroz, Luis Daniel Jiménez-Martínez, Susana De la Rosa-García, Gloria Gertrudys Asencio-Alcudia, Rafael Martínez, Dariel Tovar-Ramírez, Mario A. Galaviz, Talhia Martínez-Burguete, Carlos A. Alvarez-González, and Carina Shianya Alvarez-Villagomez. Incorporation of fructooligosaccharides in diets influence growth performance, digestive enzyme activity, and expression of intestinal barrier function genes in tropical gar (Atractosteus tropicus) larvae. Fishes, 7(3):137, June 10, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/137.

Papai:2019:NCB

[PKC^{+}19] Nóra Pápai, Ferenc Kagan, György Csikós, Mónika Kosztelnik, Tibor Vellai, and Máté Varga. No correlation between endo- and exoskeletal regenerative capacities in teleost species. Fishes, 4(4): 51, October 14, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/51.

Przybyla-Kelly:2023:CPK

[PKSN23a] Kasia J. Przybyla-Kelly, Ashley M. Spoljaric, and Meredith B. Nevers.

Correction: Przybyla-Kelly et al. Round Goby Detection in Lakes Huron and Michigan — an Evaluation of eDNA and Fish Catches. Fishes 2023, 8, 41. Fishes, 8 (5):258, May 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/258.  See

[PKSN23b].

Przybyla-Kelly:2023:RGD

[PKSN23b] Kasia J. Przybyla-Kelly, Ashley M. Spoljaric, and Meredith B. Nevers. Round goby detection in Lakes Huron and Michigan — an evaluation of eDNA and fish catches. Fishes, 8(1): 41, January 06, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/41. See cor-

rection [PKSN23a].

Podhorec:2022:EHT

[PKV^{+}22] Peter Podhorec, Jindřiška Knowles, Jakub Vysloužil, Sergii Boryshpolets, Anatolii Sotnikov, Martina Holická, Jan Kouřil, and Borys Dzyuba. The effect of hormonal treatment on selected sperm quality parameters and sex steroids in tropical cyprinid bala shark Balantiocheilos melanopterus. Fishes, 7(3):122, May 30, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/122.

Patil:2023:CMP

[PKY^{+}23] Maheshkumar Prakash Patil, Jong-Oh Kim, Seung Hyun Yoo, Yong Bae Seo, Yu-Jin Lee, Jin-Koo Kim, Shin-Ichi Kitamura, and Gun-Do Kim. Complete mitogenome and phylogenetic analysis of a marine ray-finned fish, Alcichthys elongatus (Perciformes: Cottidae). Fishes, 8(10): 513, October 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/513.

Papadopoulos:2024:IWN

[PLC^{+}24] Dimitrios K. Papadopoulos, Athanasios Lattos, Ioanna Chatzigeorgiou, Aphrodite Tsaballa, Georgios K. Ntinas, and Ioannis A. Giantsis. The influence of water nitrate concentration combined with elevated temperature on rainbow trout Oncorhynchus mykiss in an experimental aquaponic setup. Fishes, 9(2): 74, February 13, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/74.

Peng:2023:HAG

[PLJ^{+}23] Zhilan Peng, Xiaomin Liu, Ming Jin, Yu Zhan, Xiaolin Zhang, Yongbo Bao, and Minhai Liu. Hypoxia activates HIF-1α and affects gene expression and transcriptional regulation of PHD in Tegillarca granosa. Fishes, 8(7):359, July 11, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/359.

Perez:2019:WPA

[PLV^{+}19] Evaristo Pérez, Fátima Linares, José Luis Rodríguez Villanueva, Antonio Vilar, Constantinos C. Mylonas, Ioannis Fakriadis, Maria Papadaki, Nikos Papandroulakis, Ioannis Papadakis, Rocío Robles, Christian Fauvel, Javier Roo, José Benito Peleteiro, Nuria Lluch, Gema Pazos, Belén Méndez, Irini Sigelaki, Castora Gómez, Montse Pérez, and Blanca Álvarez-Blázquez. Wreckfish (Polyprion americanus). new knowledge about reproduction, larval husbandry, and nutrition. Promise as a new species for aquaculture. Fishes, 4(1):14, February 25, 2019. CODEN ???? ISSN 2410-3888. URL

                                        https://www.mdpi.com/
                                        2410-3888/4/1/14.

Pan:2024:CTH

[PLY^{+}24] Tingshuang Pan, Tong Li, Min Yang, He Jiang, Jun Ling, and Qian Gao. Cardiac transcriptome and histology of the heart of the male Chinese mitten crab (Eriocheir sinensis) under high-temperature stress. Fishes, 9(3):92, February 28, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/92.

Peixoto:2022:TFF

[PMFBI22] Ualerson I. Peixoto, Adauto S. Mello-Filho, Bianca Bentes, and Victoria J. Isaac. Trawl fishing fleet operations used to illustrate the life cycle of the southern brown shrimp: Insights to management and sustainable fisheries. Fishes, 7 (3):141, June 17, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/141.

Pang:2022:CUL

[PNW^{+}22] Yangyang Pang, Chao Niu, Lifeng Wu, Yameng Song, Xiaozhe Song, Ao ya Shi, Xingliang Shi, Zong wen Wu, Boping Tang, Xiaozhen Yang, and Yongxu Cheng. Comprehensive utilization of land resources for photovoltaic power generation to culture Chinese mitten crab (Eriocheir sinensis): Growth performance, nutritional composition and tissue color. Fishes, 7(4):207, August 18, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/207.

Poisson:2024:CPL

[Poi24] François Poisson. Consumption of post-larval swordfish (Xiphias gladius) by dolphinfish (Coryphaena hippurus): New ecological insights into both species in the Tyrrhenian Sea. Fishes, 9(2):65, February 06, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/65.

Papadaki:2018:DAR

[PPAB^{+}18] Maria Papadaki, José Benito Peleteiro, Blanca Alvarez-Blázquez, José Luis Rodríguez Villanueva, Fatima Linares, Antonio Vilar, Evaristo Pérez Rial, Nuria Lluch, Ioannis Fakriadis, Irini Sigelaki, and Constantinos C. Mylonas. Description of the annual reproductive cycle of wreckfish Polyprion americanus in captivity. Fishes, 3(4):43, October 19, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/4/43.

Piper:2023:EAI

[PRPW23] Adam T. Piper, Paula J. Rosewarne, Charlotte Pike, and Rosalind M. Wright. The eel ascending: The influence of lateral slope, climbing substrate and flow rate on eel pass performance. Fishes, 8(12): 612, December 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/612.

Pasaribu:2018:ENT

[PSN18] Wesly Pasaribu, Sukenda Sukenda, and Sri Nuryati. The efficacy of Nile tilapia (Oreochromis niloticus) broodstock and larval immunization against Streptococcus agalactiae and Aeromonas hydrophila. Fishes, 3(1): 16, March 07, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/1/16.

Pontigo:2019:MEI

[PSO^{+}19] Juan Pablo Pontigo, Julia Saravia, Ricardo Oyarzún, Margarita P. González, Christopher Hawes, Francisco J. Morera, Jorge Pino, Simon Wadsworth, Jose Luis P. Muñoz, and Luis Vargas-Chacoff. Modulation of the expression of immune-related gene in Atlantic and Coho salmon during infestation with the sea lice Caligus rogercresseyi. Fishes, 4(3):42, July 25, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/42.

Politikos:2022:DVO

[PSP^{+}22] Dimitris V. Politikos, Nikolaos Sykiniotis, Georgios Petasis, Pavlos Dedousis, Alba Ordoñez, Rune Vabø, Aikaterini Anastasopoulou, Endre Moen, Chryssi Mytilineou, Arnt-Børre Salberg, Archontia Chatzispyrou, and Ketil Malde.

DeepOtolith v1.0: an

open-source AI platform for automating fish age reading from otolith or scale images. Fishes, 7 (3):121, May 29, 2022.

CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/121.

Pagliarulo:2018:GSS

[PSS^{+}18] Caterina Pagliarulo, Daniela Sateriale, Elisa Scioscia, Nunziatina De Tommasi, Roberta Colicchio, Chiara Pagliuca, Elena Scaglione, Japo Jussila, Jenny Makkonen, Paola Salvatore, and Marina Paolucci. Growth, survival and spore formation of the pathogenic aquatic oomycete Aphanomyces astaci and fungus Fusarium avenaceum are inhibited by Zanthoxylum rhoifolium bark extracts in vitro. Fishes, 3(1): 12, February 15, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/12.

Perri:2023:EMD

[PSW^{+}23] Edward Perri, Leslie Sturmer, Paul S. Wills, John Baldwin, and Susan Laramore. Effect of microalgal diets on sunray venus clam (Macrocallista nimbosa) production and fatty acid profile. Fishes, 8(2): 72, January 26, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/72.

Pradhan:2021:CCC

[PVY^{+}21] Pravata Kumar Pradhan, Dev Kumar Verma, Shrish Chandra Yadav, Atul Krishna Dev, Thangaraj Raja Swaminathan, Anutosh Paria, Rajendran Kooloth Valappil, and Neeraj Sood. Carps, Catla catla, Cirrhinus mrigala and Hypophthalmichthys molitrix are resistant to experimental infection with tilapia lake virus (TiLV). Fishes, 6(4):56, November 01, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/56.

Pham:2023:TEO

[PWCL23] Ca-Van Pham, Hui-Cheng Wang, Sheng-Hung Chen, and Jie-Min Lee. The threshold effect of overfishing on global fishery outputs: International evidence from a sustainable fishery perspective. Fishes, 8(2): 71, January 24, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/71.

Pompeu:2023:ISF

[PWH^{+}23] Paulo Santos Pompeu, Lídia Wouters, Heron Oliveira Hilário, Raquel Coelho Loures, Alexandre Peressin, Ivo Gavião Prado, Fábio Mineo Suzuki, and Daniel Cardoso Carvalho. Inadequate sampling frequency and imprecise taxonomic identification mask results in studies of migratory freshwater fish ichthyoplankton. Fishes, 8(10):518, October 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/518.

Pan:2024:EEH

[PXH^{+}24] Yucai Pan, Gefeng Xu, Tianqing Huang, Wei Gu, Enhui Liu, and Gaochao Wang. Effects of exogenous hormones on spawning performances, serum gonadotropin and sex steroid hormone in Manchurian trout (Brachymystax lenok) during sexual maturation. Fishes, 9(7):269, July 8, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/269.

Pan:2023:EAM

[PYJ^{+}23] Tingshuang Pan, Min Yang, He Jiang, Tong Li, Guoqing Duan, Jun Ling, and Qian Gao. Effect of Astragalus membranaceus on transcriptome and survival of hybrid yellow catfish (Pseudobagrus vachellii [male sign] × Tachysurus fulvidraco [female sign]) in response to Aeromonas hydrophila challenge. Fishes, 8(9): 454, September 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/454.

Palm:2017:TAI

[PYP17] Harry W. Palm, Irfan Yulianto, and Uwe Piatkowski. Trypanorhynch assemblages indicate ecological and phylogenetical attributes of their elasmobranch final hosts. Fishes, 2 (2):8, June 17, 2017. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/2/8.

Queller:2023:LLE

[QAC23] Philip S. Queller, Elena R. M. Adams, and Molly E. Cummings. Life-long experience with male mating tactics shapes spatial cognition and coercion evasion in female swordtails. Fishes, 8(11): 562, November 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/562.

Qu:2022:SAT

[QL22] Yunfeng Qu and Ruiyang Liu. A sustainable approach towards fisheries management: Incorporating the highseas fisheries issues into the BBNJ Agreement. Fishes, 7(6):389, December 14, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/389.

Quintanilla-Pineda:2023:PNP

[QPDGF^{+}23] Mario Quintanilla-Pineda, Jesús Díaz, Ana Gutiérrez-Falcon, Francisco C. Ibañez, and Florencio Marzo. Profiling a new postbiotic product for its application in fish aquaculture. Fishes, 8 (6):304, June 06, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/6/304.

Quigg:2023:FWC

[QWR^{+}23] Antonietta Quigg, R. J. David Wells, Jay R. Rooker, Ronald L. Hill, Larissa L. Kitchens, Michael A. Dance, David L. Moulton, Phillip J. Sanchez, and Beatrice Padovani Ferreira. Food web connectivity in a mangrove– seagrass–patch reef (MSP) seascape: Lessons from a tropical back-reef in Puerto Rico. Fishes, 8 (1):44, January 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/44.

Qiu:2024:ICC

[QWY^{+}24] Hengtong Qiu, Huan Wang, Xiaomin Yan, Lin Hu, Yonglin Huang, and Yanni Ye. The identification of a cell cycle regulation gene Cyclin E from Hong Kong oysters (Crassostrea hongkongensis) and its protein expression in response to salinity stress. Fishes, 9 (3):102, March 06, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/102.

Qin:2022:IIP

[QXAY22] Gaixiao Qin, Jin Xu, Xiaohui Ai, and Yibin Yang. Isolation, identification, and pathogenicity of Aeromonas veronii, the causal agent of hemorrhagic septicemia in channel catfish (Ictalurus punctatus) in China. Fishes, 7(6):394, December 17, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/394.

Qi:2024:EDS

[QXM^{+}24] Changle Qi, Qiyou Xu, Jianhua Ming, Fei Song, and Chuanpeng Zhou. Effect of dietary supplementation on the growth and immunity of fish and shellfish. Fishes, 9 (5):176, May 13, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/176.

Reading:2018:OEQ

[RAR^{+}18] Benjamin J. Reading, Linnea K. Andersen, Yong-Woon Ryu, Yuji Mushirobira, Takashi Todo, and Naoshi Hiramatsu. Oogenesis and egg quality in finfish: Yolk formation and other factors influencing female fertility. Fishes, 3(4):45, November 21, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/4/45.

Rosati:2023:SEH

[RCL^{+}23] Luigi Rosati, Ivana Caputo, Lillà Lionetti, Mayana Karoline Fontes, Camilo Dias Seabra Pereira, and Anna Capaldo. Side effects of human drug use: an overview of the consequences of eels’ exposure to cocaine. Fishes, 8 (3):166, March 17, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/166.

Romero:2023:HAC

[RCR^{+}23] Jaime Romero, Natalia Catalán, Carolina Ramírez, Claudio D. Miranda, Marcia Oliva, Héctor Flores, María Soledad Romero, and Rodrigo Rojas. High abundance of candidatus arthromitus in intestinal microbiota of Seriolella violacea (Palm Ruff) under reared conditions. Fishes, 8(2): 109, February 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/109.

Rodriguez-Dominguez:2024:FAD

[RDANPA^{+}24] Guillermo Rodríguez-Domínguez, Eugenio Alberto Aragón-Noriega, Jorge Payán-Alejo, Jaime Edzael Mendivil-Mendoza, Marcelo Vidal Curiel-Bernal, Wenceslao Valenzuela-Quiñonez, and Ricardo Urías-Sotomayor. The fractal approach to describe growth of farmed marine species: Using double and triple logistic models. Fishes, 9(3): 106, March 12, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/106.

Rusco:2023:ZBT

[RDE^{+}23] Giusy Rusco, Michele Di Iorio, Stefano Esposito, Pierpaolo Gibertoni, Emanuele Antenucci, Valentino Palombo, Alessandra Roncarati, and Nicolaia Iaffaldano. Zootechnical brown trout (Salmo trutta L. 1758) ovarian fluid fails to upregulate the swimming performances of native Mediterranean brown trout (Salmo cettii Rafinesque, 1810) sperm in the Biferno River. Fishes, 8 (4):190, March 31, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/190.

Richard:2020:MFW

[RDG^{+}20] Adandé Richard, Liady Mouhamadou Nourou Dine, Djidohokpin Gildas, Adjahouinou Dogbè Clément, Azon Mahuan Tobias Césaire, Micha Jean-Claude, and Fiogbe Didier Emile. Multispecies fresh water algae production for fish farming using rabbit manure. Fishes, 5(4):35, November 30, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/4/35.

Rusco:2021:CMB

[RDI^{+}21] Giusy Rusco, Michele Di Iorio, Roberta Iampietro, Alessandra Roncarati, Stefano Esposito, and Nicolaia Iaffaldano. Cryobank of Mediterranean brown trout semen: Evaluation of the use of frozen semen up to six hours postcollection. Fishes, 6(3): 26, August 02, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/26.

Ruban:2021:BCP

[RE21] Dmitry A. Ruban and Vladimir A. Ermolaev. Black caviar perturbs reflection of Russian geography: a research note of aquaculture-triggered place naming puzzle. Fishes, 6(2):13, April 06, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/2/13.

Ruiz-Garcia:2020:DSM

[RGABD20] David Ruiz-García, Kye Adams, Heidi Brown, and Andrew R. Davis. Determining stingray movement patterns in a wave-swept coastal zone using a blimp for continuous aerial video surveillance. Fishes, 5(4):

31, September 30, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/4/31.

Rodriguez-Garcia:2024:CSZ

[RGTPSCCC24] Carlos Rodríguez-García, Ana Toro-Podadera, Jesica Sarmiento-Carbajal, and Remedios Cabrera-Castro. Coexisting in the surf zone: Age and feeding habits of the spotted seabass (Dicentrarchus punctatus) and European seabass (Dicentrarchus labrax) on the Gulf of Cádiz beaches (Southwest Iberian Peninsula). Fishes, 9(5):173, May 9, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/173.

Rodriguez-Gonzalez:2019:PML

[RGVG19] Tania Rodríguez-González, Jesús Cerezo Valverde, and Benjamín García García. Performance of marine lecithin supplemented feeds for the common octopus (Octopus vulgaris) ongrowing: Changes in proximate composition and lipid classes’ profile. Fishes, 4(3):47, September 12, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/47.

Reinhardt:2019:THC

[RH19] Ulrich G. Reinhardt and Peter J. Hrodey. Trap happiness and catch bias in sea lamprey traps. Fishes, 4(2):34, June 12, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/2/34.

Rind:2023:EDC

[RHU^{+}23] Khalid Hussain Rind, Syed Sikandar Habib, Javed Ahmed Ujan, Francesco Fazio, Saira Naz, Aima Iram Batool, Mujeeb Ullah, Sobia Attaullah, Khayyam Khayyam, and Khalid Khan. The effects of different carbon sources on water quality, growth performance, hematology, immune, and antioxidant status in cultured Nile tilapia with biofloc technology. Fishes, 8(10):512, October 14, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/10/512.

Riepe:2024:TTL

[RHUJ24] Tawni B. Riepe, Zachary E.

Hooley-Underwood, and Megan Johnson. Thermal tolerance of larval flannelmouth sucker catostomus latipinnis acclimated to three temperatures. Fishes, 9 (5):181, May 16, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/181.

Rahman:2023:ISA

[RIF^{+}23] Afaf N. Abdel Rahman, Sameh H. Ismail, Moustafa M. S. Fouda, Abdelwahab A. Abdelwarith, Elsayed M. Younis, Samah S. Khalil, Mahmoud M. El-Saber, Ahmed E. Abdelhamid, Simon J. Davies, and Rowida E. Ibrahim. Impact of Streptococcus agalactiae challenge on immune response, antioxidant status and hepatorenal indices of Nile tilapia: The palliative role of chitosan white poplar nanocapsule. Fishes, 8(4):199, April 12, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/199.

Ruiz-Jarabo:2021:SPR

[RJFCJC^{+}21] Ignacio Ruiz-Jarabo, Miriam Fernández-Castro, Ismael Jerez-Cepa, Cristina Barragán-Méndez, Montse Pérez, Evaristo Pérez, Juan Gil, Jesús Canoura, Carlos Farias, Juan Miguel Mancera, and Ignacio Sobrino. Survival and physiological recovery after capture by hookline: the case study of the blackspot seabream (Pagellus bogaraveo). Fishes, 6(4):64, November 17, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/64.

Ruiz-Jarabo:2018:NRE

[RJMVC^{+}18] Ignacio Ruiz-Jarabo, Pura Márquez, Luis Vargas-Chacoff, Juan Antonio Martos-Sitcha, Salvador Cárdenas, and Juan Miguel Mancera. Narrowing the range of environmental salinities where juvenile meagre (Argyrosomus regius) can be cultured based on an osmoregulatory pilot study. Fishes, 3(4): 48, December 13, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/4/48.

Ren:2022:CFA

[RJR^{+}22] Xiaoliang Ren, Shulun Jiang, Long Ren, Yidong Liang, Di’an Fang, and Dongpo Xu. Changes in fish assemblage structure after pen culture removal in Gehu Lake, China. Fishes, 7(6): 382, December 09, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/382.

Ruiz-Jarabo:2019:ESA

[RJTVC^{+}19] Ignacio Ruiz-Jarabo, Ana Belén Tinoco, Luis Vargas-Chacoff, Juan Antonio Martos-Sitcha, Ana Rodríguez-Rúa, Salvador Cárdenas, and Juan Miguel Mancera. Environmental salinity affects growth and metabolism in fingerling meagre (Argyrosomus regius). Fishes, 4 (1):6, February 05, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/1/6.

Rabbane:2022:TEA

[RKHAMM22] Md. Golam Rabbane, Md. Alamgir Kabir, Md. Habibullah-Al-Mamun, and Md. Ghulam Mustafa. Toxic effects of arsenic in commercially important fish rohu carp, Labeo rohita of Bangladesh. Fishes, 7(5):217, October 25, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/5/217.

Robinson:2024:CTS

[RKL24] Wayne Robinson, John Koehn, and Mark Lintermans. Contemporary trends in the spatial extent of common riverine fish species in Australia’s Murray–Darling Basin. Fishes, 9(6):221, June 12, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/221.

Roubie:2024:TMD

[RKS^{+}24] Eleni Roubie, Sotirios Karavoltsos, Aikaterini Sakellari, Nikolaos Katsikatsos, Manos Dassenakis, and Persefoni Megalofonou. Trace metals distribution in tissues of 10 different shark species from the Eastern Mediterranean Sea. Fishes, 9(2):77, February 16, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/77.

Ramos:2023:WFP

[RLAE23] Jorge Ramos, Pedro G.

Lino, Jaime Aníbal, and Eduardo Esteves. (Un)wanted fish: Potential consumers’ acceptability of landings in the Portuguese case. Fishes, 8(6):324, June 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/324.

Rajalakshmi:2023:CED

[RLB^{+}23] Kaakarlu Shivakumar Vinanthi Rajalakshmi, Wen-Chao Liu, Balasubramanian Balamuralikrishnan, Arun Meyyazhagan, Govindharajan Sattanathan, Manikantan Pappuswamy, Kadanthottu Sebastian Joseph, Kuppusamy Alagesan Paari, and Jang-Won Lee. Cadmium as an endocrine disruptor that hinders the reproductive and developmental pathways in freshwater fish: a review. Fishes, 8(12): 589, November 30, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/589.

Rasul:2018:BMS

[RMA^{+}18] Md. Golam Rasul, Bhaskar Chandra Majumdar, Faria Afrin, Mohammad Abu Jafor Bapary, and A. K. M. Azad Shah. Biochemical, microbiological, and sensory properties of dried silver carp (Hypophthalmichthys molitrix) influenced by various drying methods. Fishes, 3(3):25, June 25, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/3/25.

Reyes-Mero:2022:YAR

[RMSPMC^{+}22] Byron Manuel Reyes-Mero, Ana María Santana-Piñeros, Leonela Griselda Muñoz-Chumo, Yanis Cruz-Quintana, and Enric Gisbert. Yolk absorption rate and mouth development in larvae of Dormitator latifrons (Perciformes: Eleotridae). Fishes, 7(6):375, December 07, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/375.

Rodriguez-Rey:2022:DDS

[RRG22] Marta Rodríguez-Rey and Gaël Grenouillet. Disentangling the drivers of the sampling bias of freshwater fish across Europe. Fishes, 7(6):383, December 10, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/383.

Reyes:2020:NGE

[RRM^{+}20] María Reyes, María Rodríguez, Juan Montes, Fernando G. Barroso, Dmitri Fabrikov, Elvira Morote, and María José Sánchez-Muros. Nutritional and growth effect of insect meal inclusion on seabass (Dicentrarchuss labrax) feeds.

Fishes, 5(2):16, June 01, 2020. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/2/16.

Romano:2017:WOC

[RSA17] Nicla Romano, Giuseppe Scapigliati, and Luigi Abelli. Water oxygen content affects distribution of T and B lymphocytes in lymphoid tissues of farmed sea bass (Dicentrarchus labrax). Fishes, 2(3):16, September 14, 2017. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/2/3/16.

Reverter:2018:BER

[RTBL^{+}18] Miriam Reverter, Nathalie Tapissier-Bontemps, David Lecchini, Bernard Banaigs, and Pierre Sasal. Biological and ecological roles of external fish mucus: a review. Fishes, 3 (4):41, October 09, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/4/41.

Ren:2023:CTA

[RWF^{+}23] Yuanhao Ren, Wei Wang, Yin Fu, Zhiqiang Liu, Ming Zhao, Likun Xu, Tianyong Zhan, Ting Huang, Minghao Luo, Wei Chen, Chunyan Ma, Fengying Zhang, Keji Jiang, and Lingbo Ma. Comparative transcriptome analysis identifies MAPK signaling pathway associated with regulating ovarian lipid metabolism during vitellogenesis in the mud crab, Scylla paramamosain. Fishes, 8(3):145, February 28, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/145.

Ruan:2024:EIA

[RXL^{+}24] Zhuohao Ruan, Ruilin Xie, Yifu Li, Yuanyuan Luo, Zufeng Weng, and Wensheng Liu. Effect of an Ipomoea aquatica floating raft on the water quality, antioxidant system, non-specific immune responses, and microbial diversity of Penaeus vannamei in an aquaculture system. Fishes, 9(5):175, May 12, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/175.

Stathopoulou:2022:AEP

[SAB^{+}22] Paraskevi Stathopoulou, Adamantia Asimaki, Panagiotis Berillis, Nikolaos Vlahos, Efi Levizou, Nikolaos Katsoulas, Ioannis T. Karapanagiotidis, Christos I. Rumbos, Christos G. Athanassiou, and Eleni Mente. Aqua-ento-ponics: Effect of insect meal on the development of sea bass, Dicentrarchus labrax, in co-culture with lettuce. Fishes, 7(6):397, December 18, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/397.

Shija:2023:EBP

[SAC23] Vicent Michael Shija, Kwaku Amoah, and Jia Cai. Effect of bacillus probiotics on the immunological responses of Nile tilapia (Oreochromis niloticus): a review. Fishes, 8(7): 366, July 13, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/366.

Sibiya:2023:BPG

[SAGG^{+}23] Ashokkumar Sibiya, Khalid A. Al-Ghanim, Marimuthu Govindarajan, Marcello Nicoletti, Nadezhda Sachivkina, and Baskaralingam Vaseeharan. Biochemical patterns and genotoxicity of the endocrine disruptor metformin in the freshwater fish Labeo rohita. Fishes, 8(7):

380, July 22, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/380.

Svendsen:2018:EHA

[SAHS18] Morten Bo Søndergaard Svendsen, Nikolaj Reducha Andersen, Per Juel Hansen, and John Fleng Steffensen. Effects of harmful algal blooms on fish: Insights from Prymnesium parvum. Fishes, 3(1):11, February 14, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/1/11.

Sharda:2018:STS

[SAL18] Sakshi Sharda, Emmanuelle Argenti, and Kay Lucek. On the status of threespine stickleback (Gasterosteus aculeatus Linnaeus 1758) in Lake Bracciano, Italy. Fishes, 3(1):17, March 15, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/1/17.

Saraiva:2019:WFN

[SAL19] João L. Saraiva and Pablo Arechavala-Lopez. Welfare of fish — no longer the elephant in the room. Fishes, 4 (3):39, July 03, 2019.

CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/3/39.

Saraiva:2019:GAW

[SALC^{+}19] João Luis Saraiva, Pablo Arechavala-Lopez, Maria Filipa Castanheira, Jenny Volstorf, and Billo Heinzpeter Studer. A global assessment of welfare in farmed fishes: The FishEthoBase. Fishes, 4(2):30, May 16, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/2/30.

Samarajeewa:2024:SPU

[Sam24] Upali Samarajeewa. Safety, processing, and utilization of fishery products. Fishes, 9(4):146, April 21, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/146.

Sorensen:2020:CSD

[SB20] Peter W. Sorensen and Przemyslaw G. Bajer. Case studies demonstrate that common carp can be sustainably reduced by exploiting source-sink dynamics in Midwestern lakes. Fishes, 5(4):36, December 04, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/4/36.

Strauch:2019:EOP

[SBB^{+}19] Sebastian M. Strauch, Judith Bahr, Björn Baßmann, Adrian A. Bischoff, Michael Oster, Berit Wasenitz, and Harry W. Palm. Effects of ortho-phosphate on growth performance, welfare and product quality of juvenile African catfish (Clarias gariepinus). Fishes, 4(1):3, January 23, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/3.

Svenning:2024:ECC

[SBG^{+}24] Martin A. Svenning, Eigil T. Bjørvik, Jane A. Godiksen, Johan Hammar, Jack Kohler, Reidar Borgstrøm, and Nigel G. Yoccoz. Expected climate change in the High Arctic — good or bad for Arctic charr? Fishes, 9(1): 8, December 23, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/1/8.

Schemmel:2023:HEH

[SBP23] Eva Schemmel and Nancy J. Brown-Peterson. Handling effects on histological identification of female reproductive status: Examples from tropical deepwater snappers. Fishes, 8(8):406, August 04, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/406.

Santos:2023:GPH

[SBS^{+}23] Fabio A. C. Santos, Felipe S. Batista, André S. Souza, Gustavo S. C. Julio, Gisele C. Favero, José F. V. Junior, Sílvio T. Costa, Carla C. Zeppenfeld, Nadia H. Bianchini, Berta M. Heinzmann, Bernardo Baldisserotto, and Ronald K. Luz. Growth performance and histomorphology of intestine, skin, gills and liver of juvenile Colossoma macropomum fed diets containing different levels of the essential oil of Nectandra grandiflora. Fishes, 8(10):509, October 11, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/509.

Senathipathi:2024:DBR

[SBS^{+}24] Deep Nithun Senathipathi, Soottawat Benjakul, Phutthipong Sukkapat, Matsapume Detcharoen, Gururaj Moorthy, and Jirakrit Saetang. DNA barcoding revealed mislabeling of imported seafood products in Thailand. Fishes, 9 (6):215, June 5, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/215.

Shrimpton:2022:FNT

[SBT22] J. Mark Shrimpton, Paige W. Breault, and Luc A. Turcotte. Fidelity to Natal tributary streams by Kokanee following introduction to a large oligotrophic reservoir. Fishes, 7(3):123, May 30, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/123.

Sanchez-Caballero:2024:ACV

[SCBSSMA24] Carlos Adrián Sánchez-Caballero, José Manuel Borges-Souza, Ricardo Javier Saldierna-Martínez, and Avigdor Abelson. Assessing the conservation value of artificial and natural reefs via ichthyoplankton spatiotemporal dynamics. Fishes, 9(5):166, May 4, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/166.

Sorensen:2023:HRJ

[SCCM23] Karl Sorensen, Steven R.

Craig, Avner Cnaani, and Ewen McLean. Hematological response of juvenile cobia to three anesthetics. Fishes, 8 (1):31, January 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/31.

Segarra:2023:IRA

[SCHT23] Sergi Segarra, Thanh Chau, Phuc Hoang, and Loc Tran. Immunoregulation and resistance to aquatic pathogens with dietary nucleotides in Pacific white shrimp, Litopenaeus vannamei. Fishes, 8(6):308, June 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/308.

Seibel:2022:MAE

[SCSR22] Henrike Seibel, Elvis Chikwati, Carsten Schulz, and Alexander Rebl. A multidisciplinary approach evaluating soybean meal-induced enteritis in rainbow trout Oncorhynchus mykiss. Fishes, 7(1):22, January 14, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/22.

Schakmann:2023:IBS

[SCSS23] Mathias Schakmann, Emil Aputsiaq Flindt Christensen, John Fleng Steffensen, and Morten Bo Søndergaard Svendsen. The influence of body size on behavioral thermal preference in Atlantic cod (Gadus morhua): Larger fish favor colder waters. Fishes, 8(12):596, December 02, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/596.

Stroe:2024:AGM

[SCT^{+}24] Desimira Maria Stroe, Mirela Cretu, Magdalena Tenciu, Floricel Maricel Dima, Neculai Patriche, George Tiganov, and Lorena Dediu. Age, growth, and mortality of Pontic shad, Alosa immaculata Bennett, 1835, in the Danube River, Romania. Fishes, 9(4): 128, April 2, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/128.

Seixas:2023:DTS

[SDA23] Manuel J. Seixas, Rodrigo R. Domingues, and Agostinho Antunes. Decoding the transcriptome of sharks, rays, and chimaeras: Insights into their physiology, morphology, evolution, and biomedical applications. Fishes, 8(5):271, May 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/271.

Saliba:2023:DAM

[SdSdOSSL23] Jaqueline Simões Saliba, Fabio Aremil Costa dos Santos, Eloísa de Oliveira Simões Saliba, and Ronald Kennedy Luz. Different animal metabolism markers for artemia nauplii in crude protein digestibility assay for Lophiosilurus alexandri larvae. Fishes, 8(2): 110, February 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/110.

Shakweer:2023:CIR

[SEA^{+}23] Medhat S. Shakweer, Gehad E. Elshopakey, Abdelwahab A. Abdelwarith, Elsayed M. Younis, Simon John Davies, and Samia Elbahnaswy. Comparison of immune response of Litopenaeus vannamei shrimp naturally infected with Vibrio species, and after being fed with florfenicol. Fishes, 8(3):148, March 02, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/148.

Sateriale:2023:BPB

[SFK^{+}23] Daniela Sateriale, Serena Facchiano, Katrin Kaldre, Giuseppina Forgione, Giuseppa Anna De Cristofaro, Caterina Pagliarulo, and Marina Paolucci. Benefits of polyphenol-based synbiotics in crustacean diet. Fishes, 8(5):255, May 10, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/255.

Sakihara:2017:MDD

[SFP17] Troy S. Sakihara, Atsuko Fukunaga, and Kimberly A. Peyton. Mugilids display distinct trait-mediated patterns with a reinvasion of Para grass Urochloa mutica in a tropical estuary. Fishes, 2(2):7, May 23, 2017. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/2/2/7.

Sanchez-Gonzalez:2020:VTL

[SGAC20] Jorge Rubén Sánchez-González, Amadeo Arbonés, and Frederic Casals. Variation over time of length–weight relationships and condition factors for four exotic fish species from a restored shallow lake in NE Iberian Peninsula. Fishes, 5(1):7, February 04, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/1/7.

Sanz-Gonzalez:2024:SOA

[SGANM^{+}24] Juan Carlos Sanz-González, Amalia Jurado-Mc Allister, Mercedes Navarro-Martínez, Rosa Martínez Álvarez-Castellanos, Ivan Felis-Enguix, Yassine Yazid, Yahya El-Mansouri, Fernando De Miquel-Moral, Hamid Errachdi, and Ana Juan-Licián. Sensing offshore aquaculture infrastructures for data-driven dynamic stress analysis. Fishes, 9 (2):61, January 31, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/61.

Sorensen:2021:GCC

[SGG^{+}21] Solveig Lysfjord Sørensen, Ateshm Ghirmay, Yangyang Gong, Dalia Dahle, Ghana Vasanth, Mette Sørensen, and Viswanath Kiron. Growth, chemical composition, histology and antioxidant genes of Atlantic Salmon (Salmo salar) fed whole or pre-processed Nannochloropsis oceanica and Tetraselmis sp. Fishes, 6(3):23, July 23, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/23.

Shamsi:2019:SBP

[Sha19] Shokoofeh Shamsi. Seafoodborne parasitic diseases: a “One-Health” approach is needed. Fishes, 4(1):9, February 09, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/9.

Shi:2023:EPT

[SHT^{+}23] Yongchuang Shi, Haibin Han, Fenghua Tang, Shengmao Zhang, Wei Fan, Heng Zhang, and Zuli Wu. Evaluation performance of three standardization models to estimate catchper-unit-effort: a case study on Pacific sardine (Sardinops sagax) in the Northwest Pacific Ocean. Fishes, 8(12): 606, December 11, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/606.

Sun:2023:BTF

[SHW^{+}23] Yan Sun, Yi Huang, Ying Wang, Yanqun Wang, Guiying Hao, Changwei Jiang, and Zhiqiu Huang. The bZIP transcription factor family orchestrates the molecular response to nitrite stress in the largemouth bass spleen. Fishes, 8(11): 540, November 01, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/11/540.

Subhan:2022:EUB

[SIZ^{+}22] Ujang Subhan, Iskandar, Zahidah, Camellia Panatarani, and I Made Joni. Effect of ultrafine bubbles on various stocking density of striped catfish larviculture in recirculating aquaculture system. Fishes, 7(4):190, July 29, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/190.

Su:2022:CAI

[SJY^{+}22] Xin Su, Da Ji, Junjie Yao, Yuanlong Zou, and Mengzhi Yan. Comparative analysis of intestinal characteristics of largemouth bass (Micropterus salmoides) and intestinal flora with different growth rates. Fishes, 7(2):65, April 15, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/2/65.

Shahraki:2023:DEA

[SKD^{+}23] Mojgan Zare Shahraki, Yazdan Keivany, Eisa Ebrahimi Dorche, Karen Blocksom, Andreas Bruder, Joseph Flotemersch, and Doru Bănăduc. Distribution and expansion of alien fish species in the Karun River Basin, Iran. Fishes, 8(11): 538, October 31, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/538.

Shi:2023:STV

[SKF^{+}23] Yongchuang Shi, Bo Kang, Wei Fan, Lingling Xu, Shengmao Zhang, Xuesen Cui, and Yang Dai. Spatio-temporal variations in the potential habitat distribution of Pacific sardine (Sardinops sagax) in the Northwest Pacific Ocean. Fishes, 8 (2):86, January 31, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/86.

Shin:2023:SRS

[SKK^{+}23] So Ryung Shin, Hyeon Jin Kim, Jae Won Kim, Dae-Hyeon Kwon, Junghwa Choi, Jung Jun Park, and Jung Sick Lee. Sex ratio, spawning period, and sexual group maturity of the largehead hairtail Trichiurus japonicus (Teleostei: Trichiuridae) in Korean waters. Fishes, 8(4): 194, April 07, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/194.

Sintori:2023:IAA

[SKL^{+}23] Alexandra Sintori, Vasilia Konstantidelli, Angelos Liontakis, Stamatis Mantziaris, and Irene Tzouramani. Is it all about profit? Greek fishers’ motives and objective profiles. Fishes, 8(10):527, October 22, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/10/527.

Samy-Kamal:2023:MET

[SKSL23] Mohamed Samy-Kamal, Tatiana Shulezhko, and Natalia Lisitcyna. Marine endangered and threatened species in Russia: a review of current conservation strategies and management legislative tools. Fishes, 8(8):399, August 02, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/399.

Samy-Kamal:2023:DMS

[SKT23] Mohamed Samy-Kamal and Célia M. Teixeira. Diagnosis and management of small-scale and data-limited fisheries. Fishes, 8(1):39, January 05, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/39.

Seo:2022:DID

[SLC^{+}22] Haneul Seo, Andre Ditya Maulana Lubis, Tae-Jin Choi, Tae-Sung Jung, Taek-Kyun Lee, and Sukchan Lee. Development of an immunoassay detection system for koi herpesvirus using recombinant singlechain variable fragments. Fishes, 7(6): 370, December 02, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/370.

Sun:2023:TDF

[SLDC23] Lufeng Sun, Xiuqi Li, Guancang Dong, and Xuri Cong. Taxonomic diversity of fish species in the lower reaches of the Yellow River. Fishes, 8(10):503, October 09, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/503.

Sung:2022:PTF

[SLL22] Wei-Ying Sung, Hsiao-Chien Lee, and Wen-Hong Liu. The path from traditional fisheries to ecotourism in Cimei Island. Fishes, 7(4): 200, August 10, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/200.

Sultana:2022:SAS

[SLS^{+}22] Rokeya Sultana, Qun Liu, Petra Schneider, Md. Abdullah Al-Mamun, Al Mamun, Md. Farhan Tazim, Mohammad Mojibul Hoque Mozumder, Mohammed Rashed Parvej, and Md. Mostafa Shamsuzzaman. Stock assessment of six Sciaenidae species in the Bay of Bengal, Bangladesh water using a Length-Based Bayesian Biomass (LBB) method. Fishes, 7(5):214, October 24, 2022. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/7/5/214.

Suhnel:2024:SDD

[SLSC^{+}24] Simone Sühnel, Francisco José Lagreze-Squella, Gabriel Nandi Corrêa, Jaqueline Araújo, Glauber de Souza, João Paulo Ramos Ferreira, Francisco Carlos da Silva, Carlos Henrique Araújo de Miranda Gomes, and Claudio Manoel Rodrigues de Melo. Stocking density and diet of two oyster (Crassostrea gasar and Crassostrea gigas) seeds in fluidized bed bottle nursery system. Fishes, 9(5):183, May 17, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/183.

Shan:2021:CTA

[SLY^{+}21] Binbin Shan, Yan Liu, Changping Yang, Yuan Li, Liangming Wang, and Dianrong Sun. Comparative transcriptome analysis of female and male finepatterned puffer: Identification of candidate genes associated with growth and sex differentiation. Fishes, 6(4): 79, December 12, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/79.

Shuai:2023:TPO

[SLYY23] Fangmin Shuai, Jie Li, Shunchao Yu, and Jian Yang. Temporal pattern of the occurrence of Japanese glass eels (Anguilla japonica) in the Pearl River Estuary. Fishes, 8(5):256, May 11, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/5/256.

Sievers:2024:MSC

[SMAR24] Katie T. Sievers, Eva C. McClure, Rene A. Abesamis, and Garry R. Russ. Multi-scale coral reef and seascape habitat variables combine to influence reef fish assemblages. Fishes, 9(4): 137, April 15, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/137.

Sikora:2022:DLH

[SMH^{+}22a] Logan W. Sikora, Joseph T. Mrnak, Rebecca Henningsen, Justin A. Van-DeHey, and Greg G. Sass. Demographic and life history characteristics of black bullheads Ameiurus melas in a north temperate USA lake. Fishes, 7(1):21, January 14, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/21.

Sumon:2022:EPA

[SMH^{+}22b] Md Afsar Ahmed Sumon, Mohammad Habibur Rahman Molla, Israa J. Hakeem, Foysal Ahammad, Ramzi H. Amran, Mamdoh T. Jamal, Mohamed Hosny Gabr, Md. Shafiqul Islam, Md. Tariqul Alam, Christopher L. Brown, Eun-Woo Lee, Mohammed Moulay, Amer H. Asseri, F A Dain Md Opo, Ahad Amer Alsaiari, and Md. Tawheed Hasan. Epigenetics and probiotics application toward the modulation of fish reproductive performance. Fishes, 7 (4):189, July 28, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/189.

Sayed:2024:ICP

[SMH^{+}24] Ramy K. A. Sayed, Doaa M. Mokhtar, Madeha Ahmed Hashim, Ahmed S. Aly, Giacomo Zaccone, Marco Albano, Alessio Alesci, and Nada Abdellah. Immune cell profiling in the ovarian stroma of a viviparous fish during the breeding season: a histological and immunohistochemical investigation. Fishes, 9(1): 10, December 25, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/1/10.

Saavedra:2023:TPP

[SML^{+}23a] Carlos Saavedra, Massimo Milan, Ricardo B. Leite, David Cordero, Tomaso Patarnello, M. Leonor Cancela, and Luca Bargelloni. Transcriptional profiling of populations in the clam Ruditapes decussatus suggests genetically determined differentiation in gene expression along parallel temperature gradients and between races of the Atlantic Ocean and West Mediterranean Sea. Fishes, 8(4):203, April 14, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/203.

Sun:2023:RAO

[SML^{+}23b]  Shuang-Shuang  Sun,

Shi-Wei Ma, Jun Li, Qin Zhang, and Guang-Zhou Zhou. Review on the antiviral organic agents against fish rhabdoviruses. Fishes, 8(1): 57, January 16, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/57.

Simeon:2018:IAS

[SMM^{+}18] Benaya Meitasari Simeon, Efin Muttaqin, Ulfah Mardhiah, Muhammad Ichsan, Dharmadi, Andhika Prima Prasetyo, Fahmi, and Irfan Yulianto. Increasing abundance of silky sharks in the Eastern Indian Ocean: Good news or a reason to be cautious? Fishes, 3(3): 29, July 18, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/29.

Sarais:2022:EIR

[SMO^{+}22] Fabio Sarais, Ruth Montero, Sven Ostermann, Alexander Rebl, Bernd Köllner, and Tom Goldammer. The early immune response of lymphoid and myeloid headkidney cells of rainbow trout (Oncorhynchus mykiss) stimulated with Aeromonas salmonicida. Fishes, 7(1):12, January 07, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/12.

Sardi:2023:SET

[SMO^{+}23] Adriana E. Sardi, José M.

Moreira, Lisa Omingo, Xavier Cousin, Marie-Laure Bégout, Manuel Manchado, and Nina Marn. Simulating the effects of temperature and food availability on true soles (Solea spp.) early-life history traits: a tool for understanding fish recruitment in future climate change scenarios. Fishes, 8(2): 68, January 22, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/68.

Suma:2023:BEG

[SNK^{+}23] Afrina Yeasmin Suma, Shishir Kumar Nandi, Zulhisyam Abdul Kari, Khang Wen Goh, Lee Seong Wei, Albaris B. Tahiluddin, Paul Seguin, Mikael Herault, Abdullah Al Mamun, Guillermo Téllez-Isaías, and Muhammad Anamul Kabir. Beneficial effects of graded levels of fish protein hydrolysate (FPH) on the growth performance, blood biochemistry, liver and intestinal health, economics efficiency, and disease resistance to Aeromonas hydrophila of Pabda (Ompok pabda) fingerling. Fishes, 8(3):147, March 02, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/147.

Sanchez-Nuno:2019:MPR

[SNSG^{+}19] Sergio Sánchez-Nuño, Sandra C. Silva, Pedro M. Guerreiro, Borja Ordóñez-Grande, Ignasi Sanahuja, Laura Fernández-Alacid, and Antoni Ibarz. Modulation of pituitary response by dietary lipids and throughout a temperature fluctuation challenge in gilthead sea bream. Fishes, 4(4): 55, November 22, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/55.

Santander-Neto:2023:AGT

[SNSVFL23] Jones Santander-Neto, Francisco Marcante Santana, Jonas Eloi Vasconcelos-Filho, and Rosângela Lessa. Age and growth of the threatened smalleye round ray, Urotrygon microphthalmum, Delsman, 1941, from Northeastern Brazil. Fishes, 8 (3):160, March 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/160.

Siamujompa:2023:IBP

[SNZ^{+}23] Mazuba Siamujompa, Kunda Ndashe, Frederick Chitonga Zulu, Chanda Chitala, Mwansa M. Songe, Katendi Changula, Ladslav Moonga, Emmanuel Shamulai Kabwali, Stephen Reichley, and Bernard Mudenda Hang’ombe. An investigation of bacterial pathogens associated with diseased Nile tilapia in small-scale cage culture farms on Lake Kariba, Siavonga, Zambia. Fishes, 8(9): 452, September 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/452.

Sorensen:2021:IBC

[Sor21] Peter W. Sorensen. Introduction to the biology and control of invasive fishes and a special issue on this topic. Fishes, 6(4):69, November 30, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/69.

Shelledy:2023:RST

[SOT^{+}23] Katharine N. Shelledy, Amy Y. Olsen, Alexander Tanz, Megan H. Williams, Jeff Christiansen, Heidi Wilken, Chris Van Damme, and Shawn Larson. Rockfish species trends in Puget Sound, Washington, USA, 2009–2023.

Fishes, 8(10):508, October 11, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/508.

Smith:2023:ALL

[SOW^{+}23] Cheyenne R. Smith, Christopher A. Ottinger, Heather L. Walsh, Patricia M. Mazik, and Vicki S. Blazer. Application of a lipopolysaccharide (LPS)-stimulated mitogenesis assay in smallmouth bass (Micropterus dolomieu) to augment wild fish health studies. Fishes, 8(3): 159, March 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/159.

Sanchez-Paz:2024:SEP

[SPEGMC24]  Arturo  Sánchez-Paz,

Trinidad Encinas-García, and Fernando Mendoza-Cano. In silico evaluation of the PCR performance of different tests for detection of WSSV. Fishes, 9(1):5, December 20, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/1/5.

Song:2024:EVP

[SPJ^{+}24] Jin Ah Song, Heung-Sik Park, Yun-Hwan Jung, Dong Mun Choi, Cheol Young Choi, and Dae-Won Lee. Effects of various photoperiods and specific wavelengths on retinal changes and oxidative stress in the conch Tegula rustica. Fishes, 9(6):226, June 13, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/226.

Sun:2024:TRS

[SPQ^{+}24] Qingsong Sun, Yuran Pang, Yuhan Qin, Ziting Dong, Yanling Ma, Yuan Zhao, Zhanning Zhang, Jinmei Liu, Binghui Mao, and Baishuang Yin. Transcriptomic revealed that selenium-rich Lactobacillus plantarum alleviated cadmium-induced immune responses in Bulatmai barbel Luciobarbus capito kidneys. Fishes, 9(6):230, June 16, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/230.

Seret:2023:FRH

[SQ23] Bernard Séret and Jean-Pascal Quod. First records of a hydrolagus species (Holocephali: Chimaeridae) from Reunion Island and Mayotte (Southwestern Indian Ocean). Fishes, 8(10):522, October 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/522.

Sanz-Ronda:2021:FUP

[SRBCGV^{+}21] Francisco Javier Sanz-Ronda, Francisco Javier Bravo-Córdoba, Ana García-Vega, Jorge Valbuena-Castro, Andrés Martínez de Azagra, and Juan Francisco Fuentes-Pérez. Fish upstream passage through gauging stations: Experiences with Iberian barbel in flat-V weirs. Fishes, 6(4):81, December 14, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/81.

Selvaraj:2023:SSF

[SRHCO23] John Josephraj Selvaraj, Leidy Viviana Rosero-Henao, and Maria Alejandra Cifuentes-Ossa. Small-scale fisheries in the Colombian Pacific: Understanding the impact of climate change on fishermen’s livelihoods. Fishes, 8(9): 453, September 09, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/453.

Sorensen:2019:BFP

[SRL^{+}19] Peter W. Sorensen, Mara C. P. Rue, Joseph M. Leese, Ratna Ghosal, and Hangkyo Lim. A blend of F prostaglandins functions as an attractive sex pheromone in silver carp. Fishes, 4 (2):27, April 09, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/2/27.

Shen:2023:IEA

[SS23] Huihui Shen and Liming Song. Implementing ecosystem approach to fisheries management in the Western and Central Pacific Fisheries Commission: Challenges and prospects. Fishes, 8 (4):198, April 12, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/198.

Silkin:2022:ERA

[SSK^{+}22] Yuriy A. Silkin, Mikhail Yu. Silkin, Sergey M. Korotkov, Elizaveta N. Silkina, and Alla Silkina. Erythrocyte respiratory activity of the mitochondrial complex of the black sea thornback ray (Raja clavata L.) under the influence of certain activators and inhibitors in vitro. Fishes, 7(6):

376, December 07, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/376.

Safari:2021:ESC

[SSSP21] Omid Safari, Mehrdad Sarkheil, Davar Shahsavani, and Marina Paolucci. Effects of single or combined administration of dietary synbiotic and sodium propionate on humoral immunity and oxidative defense, digestive enzymes and growth performances of African cichlid (Labidochromis lividus) challenged with Aeromonas hydrophila. Fishes, 6(4):63, November 15, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/63.

Silkin:2023:TPE

[SSSS23] Yuriy A. Silkin, Mikhail Yu.

Silkin, Elizaveta N. Silkina, and Alla Silkina. Thermal phenomena in erythrocytes of the Black Sea thornback ray (Raja clavata L.) and black scorpionfish (Scorpaena porcus L.) in vitro. Fishes, 8(2): 82, January 31, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/82.

Setiawati:2017:USS

[ST17] Martiwi Diah Setiawati and Tasuku Tanaka. Utilization of scatterplot smoothers to understand the environmental preference of bigeye tuna in the southern waters off Java–Bali: Satellite remote sensing approach. Fishes, 2(1): 2, February 09, 2017. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/2/1/2.

Simat:2019:VAS

[ŠT19] Vida Šimat and Željka Trumbić. Viability of Anisakis spp. larvae after direct exposure to different processing media and non-thermal processing in anchovy fillets. Fishes, 4(1):19, March 13, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/1/19.

Shi:2023:EAG

[STZ^{+}23] Xuehui Shi, Xinyi Tang, Yichao Zhang, Wenqi Wang, Siyong Qin, Qinghua Liu, and Jie Mei. The early allometric growth and osteological ontogeny of potbellied seahorse (Hippocampus abdominalis, L. 1827) under massscale captive breeding conditions in North China. Fishes, 8(12): 604, December 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/604.

Sava:2023:PPM

[SUL^{+}23] Alexandru Sava, Paul Uiuiu, Călin Laţiu, Daniel Cocan, George-Cătălin Muntean, Tudor Papuc, Andrada Ihuţ, Camelia Răducu, Anca Becze, Cristina Craioveanu, Camelia Munteanu, Radu Constantinescu, and Vioara Mireşan. PAHs, physicochemical and microbiological analyses of trout processed by traditional smoking, in different types of packaging. Fishes, 8(8):424, August 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/424.

Suski:2020:DCD

[Sus20] Cory D. Suski. Development of carbon dioxide barriers to deter invasive fishes: Insights and lessons learned from bigheaded carp. Fishes, 5 (3):25, August 13, 2020.

CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/3/25.

Smit:2023:MPM

[SVMLP23] Willem J. Smit, Maarten P. M. Vanhove, Ngonidzashe A. G. Moyo, and Wilmien J. Luus-Powell. Metazoan parasites of Mozambique tilapia (Oreochromis mossambicus) native to Lake Urema, Mozambique. Fishes, 8(5):273, May 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/5/273.

Shertzer:2022:MDS

[SWS22] Kyle W. Shertzer, Erik H. Williams, and Skyler R. Sagarese. Modeling discards in stock assessments: Red grouper Epinephelus morio in the U.S. Gulf of Mexico. Fishes, 7(1):7, December 28, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/7.

Sadler:2023:RHF

[SWUH23] Daniel E. Sadler, Phillip C. Watts, and Silva Uusi-Heikkilä. The riddle of how fisheries influence genetic diversity.

Fishes, 8(10):510, October 13, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/510.

Su:2022:LTC

[SXQ^{+}22] Li Su, Youwei Xu, Yongsong Qiu, Mingshuai Sun, Kui Zhang, and Zuozhi Chen. Long-term change of a fish-based index of biotic integrity for a semi-enclosed bay in the Beibu Gulf. Fishes, 7(3):124, May 30, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/124.

Sun:2024:RUC

[SYL^{+}24] Yueping Sun, Bikang Yuan, Ziqiang Li, Yong Liu, and Dean Zhao. Rethinking underwater crab detection via defogging and channel compensation. Fishes, 9(2): 60, January 30, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/60.

Shen:2021:EFD

[SZF^{+}21] Fengyuan Shen, Zonghang Zhang, Yiqiu Fu, Zhen Zhang, Xin Sun, Jianyu Dong, Xiayang Ding, Muyan Chen, and Xiumei Zhang. Effects of food deprivation duration on the behavior and metabolism of black rockfish (Sebastes schlegelii). Fishes, 6(4): 58, November 06, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/58.

Setyohadi:2021:UIS

[SZSW21] Daduk Setyohadi, Umi Zakiyah, Abu Bakar Sambah, and Adi Wijaya. Upwelling impact on Sardinella lemuru during the Indian Ocean dipole in the Bali Strait, Indonesia. Fishes, 6 (1):8, March 16, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/1/8.

Song:2023:BRS

[SZT^{+}23] Yifan Song, Shengmao Zhang, Fenghua Tang, Yongchuang Shi, Yumei Wu, Jianwen He, Yunyun Chen, and Lin Li. Behavior recognition of squid jigger based on deep learning. Fishes, 8(10):502, October 08, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/502.

Shi:2023:EFG

[SZZ^{+}23] Lian Shi, Junjie Zhang, Lingmin Zhao, Qi Li, Lixing Huang, Yingxue Qin, and Qingpi Yan. Effect of the flagellar gene fliL on the virulence of Pseudomonas plecoglossicida to hybrid grouper (Epinephelus fuscoguttatus [female sign] × E. lanceolatus [male sign]). Fishes, 8(8):397, August 01, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/8/397.

Taylor:2019:SBC

[Tay19] Brett M. Taylor. Standing out in a big crowd: High cultural and economic value of Naso unicornis in the insular Pacific. Fishes, 4 (3):40, July 03, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/40.

Turcotte:2023:POG

[TBPJ23] Lenora D. M. Turcotte, Julia C. Bradshaw, Mark P. Polinski, and Stewart C. Johnson. Piscine orthoreovirus genotype-1 (PRV-1) in wild Pacific salmon of British Columbia, Canada: 2011–2020. Fishes, 8(5):252, May 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/5/252.

Torres:2024:PPE

[TCB^{+}24] Susana Torres, Montserrat Compa, Antonio Box, Samuel Pinya, and Antoni Sureda. Presence and potential effects of microplastics in the digestive tract of two small species of shark from the Balearic Islands. Fishes, 9(2): 55, January 29, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/2/55.

Thomson:2021:HGS

[TCD^{+}21] Damian P. Thomson, Anna K. Cresswell, Christopher Doropoulos, Michael D. E. Haywood, Melanie Orr, and Andrew S. Hoey. Hidden giants: the story of Bolbometopon muricatum at Ningaloo Reef. Fishes, 6(4):73, December 06, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/4/73.

Tarasco:2019:AOA

[TCV^{+}19] Marco Tarasco, João Cardeira, Michael N. Viegas, Joana Caria, Gil Martins, Paulo J. Gavaia, M. Leonor Cancela, and Vincent Laizé.

Anti-osteogenic activity of cadmium in zebrafish. Fishes, 4(1): 11, February 15, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/11.

Tran:2022:CFW

[TDN^{+}22] Thi Ngoc Tran, Chien Thang Doan, Van Bon Nguyen, Anh Dzung Nguyen, and San-Lang Wang. Conversion of fishery waste to proteases by Streptomyces speibonae and their application in antioxidant preparation. Fishes, 7(3):140, June 14, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/140.

Tanaka:2022:BCO

[THS^{+}22] Teruyoshi Tanaka, Tomoki Honryo, Yoshifumi Sawada, Daniel Margulies, Vernon Scholey, Jeanne Wexler, Maria Stein, Amal Biswas, and Kenji Takii. Biochemical changes occurring in yellowfin tuna eggs during embryonic development. Fishes, 7(2):62, April 11, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/2/62.

Tefal:2023:NOR

[TJTV^{+}23] Eslam Tefal, Ignacio Jauralde, Ana Tomás-Vidal, Silvia Martínez-Llorens, David S. Peñaranda, and Miguel Jover-Cerdá. New organic raw materials for gilthead seabream (Sparus aurata) feeding and the effects on growth, nutritive parameters, digestibility, and histology. Fishes, 8 (6):330, June 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/6/330.

Tang:2022:TSA

[TJW^{+}22] Xuemei Tang, Shulun Jiang, Henglin Wang, Yanfeng Zhou, Fei Peng, Xizhao Zhang, Yifan Zhou, Shiyue Guo, and Yang You. Transcriptome sequencing analysis reveals dynamic changes in major biological functions during the early development of clearhead icefish, Protosalanx chinensis. Fishes, 7(3):115, May 21, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/115.

Tsaousi:2024:FRR

[TK24] Nefeli Tsaousi and Stefanos Kalogirou. First record of the red cornetfish Fistularia petimba Lacepède, 1803 from Amorgos Island (Central Aegean Sea; Greece) and a review of its current distribution in the Mediterranean Sea. Fishes, 9(6):237, June 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/237.

Tahar:2018:FWQ

[TKF^{+}18] Alexandre Tahar, Alan Kennedy, Richard D. Fitzgerald, Eoghan Clifford, and Neil Rowan. Full water quality monitoring of a traditional flow-through rainbow trout farm. Fishes, 3 (3):28, July 17, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/3/28.

Torrecillas:2019:SMF

[TMD^{+}19] Silvia Torrecillas, Daniel Montero, David Domínguez, Lidia Robaina, and Marisol Izquierdo. Skin mucus fatty acid composition of gilthead sea bream (Sparus aurata): a descriptive study in fish fed low and high fish meal diets. Fishes, 4(1): 15, February 27, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/1/15.

Trifonova:2018:CAB

[TMM^{+}18] Oxana P. Trifonova, Dmitry L. Maslov, Anton N. Mikhailov, Konstantin V. Zolotarev, Kirill V. Nakhod, Valeriya I. Nakhod, Nataliya F. Belyaeva, Marina V. Mikhailova, Petr G. Lokhov, and Alexander I. Archakov. Comparative analysis of the blood plasma metabolome of negligible, gradual and rapidly ageing fishes. Fishes, 3(4):46, December 04, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/4/46.

Tserkova:2023:BPB

[TMPP23] Feriha M. Tserkova, Vesselina V. Mihneva, and Elitsa P. Petrova-Pavlova. Biological parameters and biomass and abundance indices of two demersal species, turbot (Scophthalmus maximus) and thornback ray (Raja clavata), estimated by a trawl survey in Western Black Sea. Fishes, 8(8):400, August 02, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/400.

Tweedley:2023:SFM

[TOB^{+}23] James R. Tweedley, Clara Obregón, Sarah J. Beukes, Neil R. Loneragan, and Michael Hughes. Selecting from the fisheries managers’ tool-box: Recreational fishers’ views of stock enhancement and other management options. Fishes, 8(9): 460, September 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/460.

Tsotsios:2024:EFE

[TPA^{+}24] Dimitrios Tsotsios, Dimitrios K. Papadopoulos, Maria V. Alvanou, Ioannis Georgoulis, Athanasios Lattos, Basile Michaelidis, Konstantinos Feidantsis, Ioannis A. Giantsis, and John A. Theodorou. Environmentally friendly and efficient methods for mitigating the density of ascidian fouling in Mediterranean mussel farming. Fishes, 9 (4):135, April 14, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/135.

Teixeira:2023:DTS

[TPC^{+}23] Cláudia Teixeira, Rita Pedrosa, Carolina Castro, Rui Magalhães, Elisabete Matos, Aires Oliva-Teles, Helena Peres, and Amalia Pérez-Jiménez. Dietary tryptophan supplementation implications on performance, plasma metabolites, and amino acid catabolism enzymes in meagre (Argyrosomus regius). Fishes, 8(3):141, February 28, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/3/141.

Totoiu:2023:ETS

[ŢPN^{+}23] Aurelia Ţoţoiu, Neculai Patriche, Victor Niţă, Elena Sîrbu, Floricel Maricel Dima, Magda Ioana Nenciu, and Veta Nistor. Epidemiology of turbot (Scophthalmus maeoticus) bacterial contamination, a fishery limiting factor on the Romanian Black Sea. Fishes, 8(8): 418, August 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/8/418.

Tringali:2023:RSD

[Tri23] Michael D. Tringali. Reproductive success dynamics could limit precision in close-kin mark–recapture abundance estimation for Atlantic goliath grouper (Epinephelus itajara). Fishes, 8(5):254, May 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/254.

Toledo:2023:RBJ

[TRM^{+}23] Cesar Toledo, Eduardo Rubilar, Lorena Marchant, Jessica Dörner, Lorenzo Márquez, Víctor Martínez, and Patricio Dantagnan. Relationship between jaw malformations and long-chain PUFA’s in Seriola lalandi larvae during the spawning season at a commercial hatchery. Fishes, 8(4): 200, April 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/200.

Tseng:2023:LRP

[TTT23] Huan-Sheng Tseng, Hsin-Hua Tsai, and Po-Hsing Tseng. The labour rights protection of migrant fishing workers in Taiwan: Case study of Nan-Fang-Ao fishing harbor. Fishes, 8(2): 73, January 26, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/73.

Takase:2018:CIG

[TU18] Mai Takase and Hideki Ushio. Changes in intestinal gene expression of zebrafish (Danio rerio) related to sterol uptake and excretion upon β-sitosterol administration. Fishes, 3(1):1, January 04, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/1.

Thach:2021:TEO

[TVL21] Keo Sa Rate Thach, Hong Tu Vo, and Ji-Yong Lee. Technical efficiency and output losses in shrimp farming: a case in Mekong Delta, Vietnam. Fishes, 6(4): 59, November 11, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/4/59.

Uribe:2019:SAO

[UDG^{+}19] Mari Carmen Uribe, Gabino De la Rosa Cruz, Adriana García Alarcón, Juan Carlos Campuzano Caballero, and María Guadalupe Guzmán Bárcenas. Structures associated with oogenesis and embryonic development during intraovarian gestation in viviparous teleosts (Poeciliidae). Fishes, 4(2): 35, June 19, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/2/35.

Ulman:2024:IPC

[UGH^{+}24] Aylin Ulman, Cat A.

Gordon, Ali R. Hood, Melanie Warren, and Daniel Pauly. Identifying potential critical angel shark areas in Türkiye, Eastern Mediterranean based on new records of Squatina spp. identified through fisher engagement. Fishes, 9 (7):270, July 8, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/270.

Unger:2019:RST

[UH19] Shem Unger and Caleb Hickman. Report on the short-term scavenging of decomposing native and non-native trout in Appalachian streams. Fishes, 4(1): 17, March 01, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/1/17.

Uyisenga:2023:ERF

[ULR^{+}23] Adolphe Uyisenga, Hualiang Liang, Mingchun Ren, Dongyu Huang, Chunyu Xue, Heng Yin, and Haifeng Mi. The effects of replacing fish meal with enzymatic soybean meal on the growth performance, wholebody composition, and health of juvenile Gibel carp (Carassius auratus gibelio). Fishes, 8(8): 423, August 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/423.

Urias-Sotomayor:2022:SRA

[USRDFO^{+}22] Ricardo Urías-Sotomayor, Guillermo Rodríguez-Domínguez, José Adán Félix-Ortiz, Gilberto G. Ortega-Lizárraga, Horacio A. Muñoz-Rubí, and Eugenio Alberto Aragón-Noriega. Stock reduction analysis of bigeye croaker micropogonias megalops in the Upper Gulf of California, Mexico. Fishes, 7(1): 15, January 08, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/15.

Vercelli:2023:TRI

[VAT^{+}23] Cristina Vercelli, Michela Amadori, Massimiliano Tursi, Graziana Gambino, Paolo Pastorino, Marino Prearo, Ugo Ala, Raffaella Barbero, and Giovanni Re. TRPV1 receptor identification in rainbow trout (Oncorhynchus mykiss) and evaluation of the effects produced by Ocimum basilicum super critical fluid extract. Fishes, 8 (1):38, January 04, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/38.

Vaibarova:2024:SVF

[VČ24] Věra Vaibarová and Alois Čížek. Supposed virulence factors of flavobacterium psychrophilum: a review. Fishes, 9(5):163, April 30, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/163.

Valero:2018:IEI

[VCC^{+}18] Yulema Valero, Alberto Cuesta, Matteo Cammarata, María Angeles Esteban, and Elena Chaves-Pozo. Immuneendocrine interactions in the fish gonad during infection: an open door to vertical transmission. Fishes, 3(2): 24, June 13, 2018. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/2/24.

Vijayakumar:2020:ICD

[VCL20] Parameswaran Vijayakumar, M. Leonor Cancela, and Vincent Laizé. Isolation, culture, and differentiation of blastema cells from the regenerating caudal fin of zebrafish. Fishes, 5(1): 6, January 30, 2020. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/1/6.

Verhille:2023:ESR

[VF23] Christine E. Verhille and Anthony P. Farrell. Endurance swimming is related to summer lake survival of rainbow trout in a warm lake with avian piscivores. Fishes, 8(4):213, April 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/213.

Vu:2020:OLF

[VH20] Ngoc-Ut Vu and Truong-Giang Huynh. Optimized live feed regime significantly improves growth performance and survival rate for early life history stages of pangasius catfish (Pangasianodon hypophthalmus). Fishes, 5(3):20, June 28, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/3/20.

Vaihola:2023:EFD

[VK23] Siosaia Vaihola and Stuart Kininmonth. Environmental factors determine tuna fishing vessels’ behavior in Tonga. Fishes, 8(12):602, December 07, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/602.

Vuorinen:2024:TDM

[VKP^{+}24] Pekka J. Vuorinen, Reijo Käkelä, Tapani Pakarinen, Petri Heinimaa, Tiina Ritvanen, Soili Nikonen, Mervi Rokka, and Marja Keinänen. Thiamine deficiency M74 developed in salmon (Salmo salar) stocks in two Baltic Sea areas after the hatching of large year-classes of two clupeid species — detected by fatty acid signature analysis. Fishes, 9(2): 58, January 30, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/58.

vonKrogh:2019:VEB

[vKRNL^{+}19] Kristine von Krogh, Erik Ropstad, Rasoul Nourizadeh-Lillabadi, Trude Marie Haug, and Finn-Arne Weltzien. In vitro effects of bisphenol A and tetrabromobisphenol A on cell viability and reproductionrelated gene expression in pituitaries from sexually maturing Atlantic cod (Gadus morhua L.). Fishes, 4(3):48, September 17, 2019. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/48.

Velazquez-Landa:2023:ZSB

[VLCCA^{+}23] Xiomara Velázquez-Landa, Porfirio Carrillo, Genaro A. Coria-Avila, Deissy Herrera-Covarrubias, Luis I. García, María Rebeca Toledo-Cárdenas, María Elena Hernández-Aguilar, and Jorge Manzo. Zebrafish sexual behavior in plain and enriched environments: Parameters in the valproate model of autism. Fishes, 8(3): 156, March 05, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/156.

Valdez:2019:ASC

[VM19] Jose W. Valdez and Kapil Mandrekar. Assessing the species in the CARES Preservation Program and the role of aquarium hobbyists in freshwater fish conservation. Fishes, 4(4): 49, September 29, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/4/49.

Valic:2024:MAT

[VMBT24] Damir Valić, Matej Kristan Mirković, Višnja Besendorfer, and Emin Teskeredžić. Molecular analysis of two endemic Squalius species: Evidence for intergeneric introgression among Cyprinids and conservation issues. Fishes, 9(1):4, December 20, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/1/4.

Vallecillos:2022:DFM

[VMDV^{+}22] Antonio Vallecillos, Emilio María-Dolores, Javier Villa, Francisco Miguel Rueda, José Carrillo, Guillermo Ramis, Mohamed Soula, Juan Manuel Afonso, and Eva Armero. Development of the first microsatellite multiplex PCR panel for meagre (Argyrosomus regius), a commercial aquaculture species. Fishes, 7 (3):117, May 24, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/117.

Voorhees:2023:ECI

[VMS^{+}23] Jill M. Voorhees, Elizabeth R. J. M. Mamer, Daniel J. Schill, Mitchel Adams, Carlos Martinez, and Michael E. Barnes. 17β-estradiol can induce sex reversal in brown trout. Fishes, 8(2):103, February 09, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/103.

Vaz:2022:EIM

[VPP^{+}22] Mariana Vaz, Damiana Pires, Pedro Pires, Marco Simões, Ana Pombo, Paulo Santos, Beatriz do Carmo, Ricardo Passos, Janina Z. Costa, Kim D. Thompson, and Teresa Baptista. Early immune modulation in European seabass (Dicentrarchus labrax) juveniles in response to Betanodavirus infection. Fishes, 7 (2):63, April 11, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/2/63.

Vieira:2019:SVC

[VPPF^{+}19] Manuel Vieira, Beatriz P. Pereira, Pedro Pousão-Ferreira, Paulo J. Fonseca, and M. Clara P. Amorim.

Seasonal variation of captive meagre acoustic signalling: a manual and automatic recognition approach. Fishes, 4(2):28, April 18, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/2/28.

Vijayaram:2024:UAA

[VRG^{+}24] Srirengaraj Vijayaram, Einar Ringø, Hamed Ghafarifarsani, Seyed Hossein Hoseinifar, Saman Ahani, and Chi-Chung Chou. Use of algae in aquaculture: a review. Fishes, 9(2):63, February 01, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/63.

Vetorelli:2024:IAR

[VRKV24] Michelle Pinheiro Vetorelli, Laurindo André Rodrigues, Janaina Mitsue Kimpara, and Wagner C. Valenti. Intensification of Amazon River prawn hatchery. Fishes, 9(3):82, February 22, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/82.

Vehanen:2023:PIC

[VSH23] Teppo Vehanen, Tapio Sutela, and Ari Huusko. Potential impact of climate change on salmonid smolt ecology. Fishes, 8(7):382, July 24, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/382.

vonSiebenthal:2018:TOU

[vSRB^{+}18] Elena Wernicke von Siebenthal, Kristina Rehberger, Christyn Bailey, Albert Ros, Elio L. Herzog, and Helmut Segner. Trade-offs underwater: Physiological plasticity of rainbow trout (Oncorhynchus mykiss) confronted by multiple stressors. Fishes, 3(4): 49, December 16, 2018. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/3/4/49.

Vella:2023:PAS

[VV23] Noel Vella and Adriana Vella. Phylogeographic analyses of the shortfin mako, Isurus oxyrinchus Rafinesque, 1810 (Chondrichthyes: Lamniformes) from the Central Mediterranean Sea, a critically endangered species in the region. Fishes, 8(10): 520, October 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/520.

Wuertz:2023:TPC

[WBK^{+}23] Sven Wuertz, Filipa Beça, Eva Kreuz, Konrad M. Wanka, Rita Azeredo, Marina Machado, and Benjamin Costas. Two probiotic candidates of the genus Psychrobacter modulate the immune response and disease resistance after experimental infection in turbot (Scophthalmus maximus, Linnaeus 1758). Fishes, 8(3): 144, February 28, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/144.

Wu:2024:ICP

[WCX^{+}24] Xingbing Wu, Baolin Cheng, Mingyang Xue, Nan Jiang, Xuemei Li, Xiaona Hu, Xiaoli Li, Tingbing Zhu, Yongjiu Zhu, and Yong Zhou. Isolation, characterization, and pathogenicity of an Aeromonas veronii strain causing disease in Rhinogobio ventralis. Fishes, 9 (5):188, May 18, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/188.

Wu:2024:ECC

[WCY^{+}24] Yijie Wu, Liting Chen, Xin Yan, Jun Xiao, Zhirui Ma, Zhanyang Tang, Zhongbao Guo, Liping Li, Guixiang Tong, Honglian Tan, Fuyan Chen, Xinxian Wei, Ting Huang, and Yongju Luo. The effect of copper–cadmium co-exposure and hormone remediation on the ovarian transcriptome of Nile tilapia (Oreochromis niloticus). Fishes, 9(2):67, February 08, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/2/67.

Wu:2023:CPC

[WDL^{+}23] Yingru Wu, Qi Du, Yueqin Liao, Shanshan Shui, Jie Pang, Soottawat Benjakul, and Bin Zhang. Changes in physicochemical characteristics and volatile flavor compounds of brinepreserved ready-to-eat shrimp (Solenocera crassicornis) during chilled storage. Fishes, 8(7): 372, July 19, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/372.

Weber:2023:EII

[Web23] Gregory M. Weber.

Effects of IGF1 and IGF2 on in vitro ovarian follicle maturation in rainbow trout, Oncorhynchus mykiss. Fishes, 8(7):367, July 14, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/367.

Wei:2023:GSP

[WFL^{+}23] Xuelian Wei, Zeqin Fu, Jiji Li, Baoying Guo, and Yingying Ye. Genetic structure and phylogeography of commercial Mytilus unguiculatus in China based on mitochondrial COI and Cytb sequences. Fishes, 8(2): 89, February 02, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/89.

Wu:2023:IIS

[WFX^{+}23] Yeying Wu, Yalan Feng, Mingyang Xue, Zidong Xiao, Lijuan Jin, Ren Gao, Yahong Chen, Tianwang Liang, and Yong Zhou. Isolation and identification of Staphylococcus saprophyticus from diseased hybrid sturgeon. Fishes, 8(5):250, May 09, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/5/250.

Wu:2023:FHS

[WFZ^{+}23] Di Wu, Hao Feng, Ying Zou, Juan Xiao, Pengfei Zhang, Yuxiang Ji, Sovan Lek, Zhiqiang Guo, and Qiongyao Fu. Feeding habit-specific heavy metal bioaccumulation and health risk assessment of fish in a tropical reservoir in Southern China. Fishes, 8 (4):211, April 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/4/211.

Wang:2024:EDS

[WFZ^{+}24a] Jing Wang, Dan Fan, Ran Zhao, Tongyan Lu, Shaowu Li, and Di Wang. Effects of dietary supplementation with endogenous probiotics Bacillus subtilis on growth performance, immune response and intestinal histomorphology of juvenile rainbow trout (Oncorhynchus mykiss). Fishes, 9(6): 229, June 15, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/6/229.

Wu:2024:AFS

[WFZ^{+}24b] Jingjing Wu, Jinghua Fu, Dingkang Zhou, Jiasen Huang, and Minjun Xu. Assessment of fish species in Wanlv Lake, the largest drinking water source in South China, by environmental DNA metabarcoding technology. Fishes, 9(3): 86, February 24, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/86.

Wagner:2019:NBM

[WGRM^{+}19] Liane Wagner, Pedro Gómez-Requeni, Ali A. Moazzami, Torbjörn Lundh, Aleksandar Vidakovic, Markus Langeland, Anders Kiessling, and Jana Pickova. ^{1}H NMR-based metabolomics and lipid analyses revealed the effect of dietary replacement of microbial extracts or mussel meal with fish meal to Arctic charr (Salvelinus alpinus). Fishes, 4(3): 46, September 10, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/46.

Wang:2023:ECA

[WGW^{+}23] Jiaqi Wang, Chunxia Gao, Feng Wu, Libin Dai, Qiuyun Ma, and Siquan Tian. Environmental characteristics associated with the presence of the pelagic stingray (Pteroplatytrygon violacea) in the Pacific high sea. Fishes, 8 (1):46, January 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/1/46.

Wang:2024:CGP

[WHL^{+}24] Ziyu Wang, Peiying Huang, Xiaoyong Li, Jianmin Pei, Wenzhen Liu, Jiahao Hou, Linjie Li, Hongxiang Fan, Liugen Zeng, and Daxian Zhao. Comparison of growth performance, nutritional composition, and muscle transcriptome between two cultured varieties of the Chinese mitten crab (Eriocheir sinensis). Fishes, 9(4):132, April 9, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/132.

Wei:2021:EEF

[WHM21] Xinyi Wei, Qiuguang Hu, and Jintao Ma. Evaluation on the effect of fishery insurance policy: Evidence based on text mining. Fishes, 6(3):41, September 13, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/3/41.

Wang:2023:OPA

[WHX^{+}23] Shuo Wang, Yingxia He, Feng Xi, Ying Liang, and Shaowei Zhai. Oligomeric proanthocyanidins alleviate the detrimental effects of dietary histamine on intestinal health of juvenile American eels (Anguilla rostrata). Fishes, 8(8):413, August 12, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/413.

Wei:2024:COD

[WHY^{+}24] Shuwei Wei, Zhong Hua, Yanping Yang, Fengjiao Ma, Wei Han, Wei Zhang, Congping Ying, Yanmin Deng, and Kai Liu. Characterization of the ovarian development and associated factors during the breeding migration of Coilia nasus in the Yangtze River. Fishes, 9(3): 90, February 28, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/90.

Wahltinez:2023:CSS

[WKB^{+}23] Sarah J. Wahltinez, Kevin J. Kroll, Donald C. Behringer, Jill E.

Arnold, Brent Whitaker, Alisa L. Newton, Kristina Edmiston, Ian Hewson, and Nicole I. Stacy. Common sea star (Asterias rubens) coelomic fluid changes in response to short-term exposure to environmental stressors. Fishes, 8(1):51, January 12, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/51.

Wang:2023:ICG

[WLC^{+}23] Yan Wang, Zhihong Liu, Xi Chen, Liqing Zhou, Xiujun Sun, Tao Yu, Xiaomei Wang, Yanxin Zheng, and Biao Wu. Identification and characterization of GYS and GSK3β provides insights into the regulation of glycogen synthesis in Jinjiang oyster Crassostrea ariakensis. Fishes, 8(2): 65, January 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/65.

Wang:2022:EDP

[WLL^{+}22] Ronghua Wang, Chaobo Lei, Zhenyu Li, Yanju Lei, Congqiang Luo, Liye Shao, Chunhong Huang, and Pinhong Yang. Effects of a diet of Phragmites australis instead of Triticum aestivum L. on immune performance and liver tissue structure of Ctenopharyngodon idellus. Fishes, 7(6):378, December 08, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/378.

Wang:2023:MBF

[WLL^{+}23] Jilong Wang, Peilun Li, Wei Liu, Wanqiao Lu, and Fujiang Tang. The migratory biology and feeding habits of downstreammigrating juvenile chum salmon Oncorhynchus keta in the Amur River of Northeast China. Fishes, 8(9):458, September 14, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/458.

Wang:2024:ETS

[WLL^{+}24] Lei Wang, Xiaowen Long, Yaopeng Li, Yong Zhang, Weihong Sun, and Xugan Wu. Effects of three sources of astaxanthin on the growth, coloration, and antioxidant capacity of rainbow trout (Oncorhynchus mykiss) during longterm feeding. Fishes, 9 (5):174, May 10, 2024.

CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/5/174.

Wiryawan:2020:CPU

[WLM^{+}20] Budy Wiryawan, Neil Loneragan, Ulfah Mardhiah, Sonja Kleinertz, Prihatin Ika Wahyuningrum, Jessica Pingkan, Wildan, Putra Satria Timur, Deirdre Duggan, and Irfan Yulianto. Catch per unit effort dynamic of yellowfin tuna related to sea surface temperature and chlorophyll in Southern Indonesia. Fishes, 5(3): 28, August 31, 2020. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/5/3/28.

Wu:2023:TAR

[WLN^{+}23] Ren-Xie Wu, Yan-Shan Liang, Su-Fang Niu, Jing Zhang, Bao-Gui Tang, and Zhen-Bang Liang. Transcriptomic analysis reveals circadian rhythm homeostasis in pearl gentian grouper under acute hypoxia. Fishes, 8(7):358, July 10, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/358.

Wu:2023:GRM

[WLT23] Chien-Yu Wu, Tsung-Han Lee, and Deng-Yu Tseng. Glucocorticoid receptor mediates cortisol regulation of glycogen metabolism in gills of the euryhaline tilapia (Oreochromis mossambicus). Fishes, 8(5):267, May 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/267.

Wang:2023:EFC

[WLW^{+}23] Zhanzhan Wang, Shuling Liao, Jun Wang, Yun Wang, Zhong Huang, Wei Yu, Xiaolin Huang, Heizhao Lin, Maoyan Luo, Zhenyan Cheng, and Chuanpeng Zhou. Effects of fermented cottonseed meal substitution for fish meal on intestinal enzymatic activity, inflammatory and physical-barrier-related gene expression, and intestinal microflora of juvenile golden pompano (Trachinotus ovatus). Fishes, 8(9):466, September 18, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/466.

Wang:2024:EGI

[WLW^{+}24] Chang’an Wang, Fangyuan Li, Di Wang, Shaoxia Lu, Shicheng Han, Wei Gu, Haibo Jiang, Zhuang Li, and Hongbai Liu. Enhancing growth and intestinal health in triploid rainbow trout fed a low-fish-meal diet through supplementation with Clostridium butyricum. Fishes, 9 (5):178, May 14, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/178.

Wang:2022:PGD

[WLZ^{+}22a] Shihui Wang, Liang Luo, Rui Zhang, Kun Guo, Wei Xu, and Zhigang Zhao. Population genetic diversity and differentiation of mitten crab, genus Eriocheir, based on microsatellite markers. Fishes, 7(4): 182, July 22, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/182.

Wu:2022:EWT

[WLZ^{+}22b] Xingbing Wu, Xuemei Li, Yongjiu Zhu, Jinling Gong, Tingbing Zhu, Jiajia Ni, and Deguo Yang. Effects of water temperature on the growth, antioxidant capacity, and gut microbiota of Percocypris pingi juveniles.

Fishes, 7(6):374, December 06, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/374.

Wang:2023:DFD

[WLZ^{+}23] Zhen Wang, Haolu Liu, Guangyue Zhang, Xiao Yang, Lingmei Wen, and Wei Zhao. Diseased fish detection in the underwater environment using an improved YOLOV5 network for intensive aquaculture. Fishes, 8 (3):169, March 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/169.

Wang:2022:ACE

[WM22] Peiwen Wang and Isabel Mendes. Assessment of changes in environmental factors affecting aquaculture production and fisherfolk incomes in China between 2010 and 2020. Fishes, 7(4): 192, August 05, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/192.

Wei:2024:EFT

[WMD^{+}24] Yuchong Wei, Aixing Ma, Ya Deng, Minxiong Cao, Ying Hu, Zhaoyi Cheng, and Lijun Zhu. Effects of flow turbulence on the entire development process of drifting fish eggs. Fishes, 9(3):88, February 27, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/88.

Wang:2021:EDT

[WML^{+}21] Bin Wang, Guoxin Ma, Yong Liu, Yafei Wang, Xiaoxue Du, Qiang Shi, and Hanping Mao. Effects of different temperatures on the antibacterial, immune and growth performance of crucian carp epidermal mucus. Fishes, 6(4): 66, November 22, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/66.

Wang:2022:DMP

[WMZ^{+}22] Bin Wang, Hanping Mao, Jian Zhao, Yong Liu, Yafei Wang, and Xiaoxue Du. Designing a multi-parameter method to assess the adaptation period of crucian carp under stress conditions of the bionic robot fish. Fishes, 7(4): 198, August 08, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/198.

Wikumpriya:2023:MCG

[WPLK23] Gunasekara Chathura Wikumpriya, Madhuranga Walawedurage Srinith Prabhatha, Jiye Lee, and Chan-Hee Kim. Molecular cloning and gene expression of type I suppressors of cytokine signaling 6 and 7 (SOCS6 and SOCS7) in whiteleg shrimp (Litopenaeus vannamei). Fishes, 8(8):416, August 14, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/8/416.

Watz:2023:ITD

[WSA^{+}23] Johan Watz, Joel Schill, Louis Addo, John J. Piccolo, and Mahboobeh Hajiesmaeili. Increased temperature and discharge influence overwinter growth and survival of juvenile salmonids in a hydropeaking river: Simulating effects of climate change using individual-based modelling. Fishes, 8(6):323, June 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/323.

Wang:2023:ESP

[WSD^{+}23] Wenqian Wang, Shengyan Su, Ping Dong, Wenrong Feng, Jianlin Li, Chengfeng Zhang, and Yongkai Tang. Effects of seasonal photoperiod on growth, lipid metabolism, and antioxidant response in the Huanghe carp (Cyprinus carpio haematopterus). Fishes, 8(12):595, December 01, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/595.

Wenzel:2021:EDP

[WSE^{+}21] Lisa Carolina Wenzel, Sebastian Marcus Strauch, Ep Eding, Francisco Xose Presas-Basalo, Berit Wasenitz, and Harry Wilhelm Palm. Effects of dissolved potassium on growth performance, body composition, and welfare of juvenile African catfish (Clarias gariepinus). Fishes, 6(2):11, March 28, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/2/11.

Wylie:2019:ISF

[WSI^{+}19] Matthew J. Wylie, Alvin N. Setiawan, Glen W. Irvine, Abigail Elizur, Yonathan Zohar, Jane E. Symonds, and P. Mark Lokman. Induced spawning of F1 wreckfish (Hāpuku) Polyprion oxygeneios using a synthetic agonist of gonadotropinreleasing hormone. Fishes, 4(3):41, July 08, 2019. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/4/3/41.

Wan:2023:UEM

[WSL^{+}23] Rong Wan, Pengbo Song, Zengguang Li, Xiangyu Long, Dong Wang, and Lu Zhai. Use of ensemble model for modeling the larval fish habitats of different ecological guilds in the Yangtze Estuary. Fishes, 8(4):209, April 17, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/209.

White:2024:DNF

[WSON24] William T. White, Andrew L. Stewart, Helen L. O’Neill, and Gavin J. P. Naylor. Dichichthyidae, a new family of deepwater sharks (Carcharhiniformes) from the Indo– West Pacific, with description of a new species. Fishes, 9(4): 121, March 28, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/121.

Wylie:2019:TCD

[WSS^{+}19] Matthew J. Wylie, Jane E. Symonds, Alvin N. Setiawan, Glen W. Irvine, Hui Liu, Abigail Elizur, and P. Mark Lokman. Transcriptomic changes during previtellogenic and vitellogenic stages of ovarian development in wreckfish (Hāpuku), Polyprion oxygeneios (Perciformes). Fishes, 4(1):16, February 28, 2019. CODEN ???? ISSN 2410-3888. URL

                                        https://www.mdpi.com/
                                        2410-3888/4/1/16.

Wan:2023:ELP

[WSZ^{+}23] Shuai Wan, Zhaosheng Sun, Chang Zhang, Tingshuang Pan, Shuya Yuan, Yuxi Chen, Jun Zou, and Qian Gao. Effects of LPS, poly (I:C) and Edwardsiella tarda on the expression patterns of IL-17 family members and their receptors in spotted sea bass (Lateolabrax maculatus). Fishes, 8(8): 405, August 04, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/8/405.

Wang:2022:OMD

[WTC^{+}22] Gongpei Wang, Qindong Tang, Zhi Chen, Dingli Guo, Lei Zhou, Han Lai, and Guifeng Li. Otolith microchemistry and demographic history provide new insight into the migratory behavior and heterogeneous genetic divergence of Coilia grayii in the Pearl River. Fishes, 7 (1):23, January 17, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/23.

Wildhaber:2023:RGC

[WWD^{+}23] Mark L. Wildhaber, Benjamin M. West, Karlie K. Ditter, Adrian P. Moore, and Alex S. Peterson. A review of grass carp and related species literature on diet, behavior, toxicology, and physiology focused on informing development of controls for invasive grass carp populations in North America. Fishes, 8(11):547, November 10, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/547.

Woodward:2019:EIA

[WWW19] Melanie A. Woodward, Lucy A. Winder, and Penelope J. Watt. Enrichment increases aggression in zebrafish. Fishes, 4(1):22, March 19, 2019. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/4/1/22.

Wei:2024:MEM

[WWW^{+}24] Qingcong Wei, Dan Wang, Kaijin Wei, Bin Xu, and Jin Xu. The mechanism of Elizabethkingia miricola infection of the black spotted frog as revealed by multi-omics analysis. Fishes, 9(3):91, February 28, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/91.

Wu:2023:CBE

[WWY^{+}23] Jinming Wu, Jinping Wu, Huan Ye, Wei Xiong, Wanmin Qu, Xiaoqian Leng, and Hao Du. Correlations between environmental factors and the distribution of juvenile Hucho bleekeri in the Taibai River, Shaanxi, China. Fishes, 8(7):379, July 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/379.

Wu:2024:TOS

[WWZ^{+}24] Song Wu, Peng Wang, Ying Zhang, Li Huang, Qirui Hao, Lei Gao, Dongli Qin, and Xiaoli Huang. Toxicity, oxidative stress, and tissue distribution of butachlor in the juvenile Chinese mitten crab (Eriocheir sinensis). Fishes, 9 (5):177, May 13, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/5/177.

Wang:2023:EIH

[WXL^{+}23] Xinxin Wang, Junyi Xie, Yan Luo, Xiao Wang, Gaobo Guo, and Xinxing You. Experimental investigation of the hydrodynamic characteristics of longline aquaculture facilities under current and wave conditions. Fishes, 8(4): 204, April 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/204.

Wang:2023:TAR

[WYG^{+}23] Tong Wang, Yang Yang, Shirui Gong, Xi Wu, Leilei Zeng, Yuhao Tao, Chaoyue Zhong, Leling Song, and Xiaochun Liu. Transcriptome analysis reveals differences in gene expression in the muscle of the brown-marbled grouper (Epinephelus fuscoguttatus) with different growth rates. Fishes, 8 (6):309, June 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/6/309.

Wang:2023:RTL

[WYL23] Jiaming Wang, Xiangbo Yin, and Guodong Li. A real-time lightweight detection algorithm for deck crew and the use of fishing nets based on improved YOLOv5s network. Fishes, 8(7):376, July 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/7/376.

Wang:2023:ECV

[WZG^{+}23] Yang Wang, Fan Zhang, Zhe Geng, Yuying Zhang, Jiangfeng Zhu, and Xiaojie Dai. Effects of climate variability on two commercial tuna species abundance in the Indian Ocean. Fishes, 8(2): 99, February 07, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/99.

Wang:2024:EDF

[WZL^{+}24] Hongquan Wang, Zheming Zhang, Feilong Li, Liang Hu, Tiaoyi Xiao, Yurong Zhao, and Mengxi Yang. The effects of dietary fermented soybean residue on the growth, antioxidant capacity, digestive enzyme activities, and microbial compositions of the intestine in Furong crucian carp (Furong carp[female-sign] × red crucian carp[male sign]). Fishes, 9(4):138, April 16, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/138.

Wang:2024:IEP

[WZX^{+}24] Ping Wang, Ziling Zhu, Qinglai Xu, Yangfen Xing, Mingyue Zhang, and Jishu Zhou. Inhibitory effect of prickly ash (Zanthoxylum bungeanum) seed kernel oil on lipid metabolism of grass carp (Ctenopharyngodon idellus) in high-fat diet. Fishes, 9(4):123, March 29, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/123.

Xu:2023:EAM

[XCW^{+}23] Jing Xu, Gangfu Chen, Min Wu, Qihui Yang, and Huatao Li. The extract of Astragalus membranaceus inhibits lipid oxidation in fish feed and enhances growth performance and antioxidant capacity in Jian carp (Cyprinus carpio var.

Jian). Fishes, 8(12): 594, December 01, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/12/594.

Xiao:2022:MCD

[XHC^{+}22] Yangbo Xiao, Rong Huang, Shenping Cao, Dafang Zhao, Zhuangwen Mao, Chuchu Xiao, Zhehua Xu, Xiaomei Zhou, Xinran Zhang, Yu Zhang, Jianzhou Tang, Junyan Jin, Yaoguo Li, Jun Zou, and Zhen Liu. Molecular characterization and dietary regulation of glutaminase 1 (gls1) in triploid crucian carp (Carassius auratus). Fishes, 7(6): 377, December 07, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/6/377.

Xue:2022:MME

[XJC^{+}22] Junren Xue, Tao Jiang, Xiubao Chen, Hongbo Liu, and Jian Yang. Multi-mineral element profiles in genuine and “Bathing” cultured Chinese mitten crabs (Eriocheir sinensis) in Yangcheng Lake, China. Fishes, 7 (1):11, January 07, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/11.

Xie:2024:DLB

[XLC^{+}24] Mingyang Xie, Bin Liu, Xinjun Chen, Wei Yu, and Jintao Wang. Deep learning-based fishing ground prediction using asymmetric spatiotemporal scales: a case study of Ommastrephes bartramii. Fishes, 9(2): 64, February 04, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/64.

Xu:2023:RLR

[XLP23] Chang Xu, Yang Liu, and Zhaobin Pei. Research on legal risk identification, causes and remedies for prevention and control in China’s aquaculture industry. Fishes, 8(11):537, October 29, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/537.

Xu:2022:EDE

[XLX^{+}22] Yan Xu, Yiqun Li, Mingyang Xue, Zidong Xiao, Yuding Fan, Lingbing Zeng, and Yong Zhou. Effects of dietary Enterococcus faecalis YFI-G720 on the growth, immunity, serum biochemical, intestinal morphology, intestinal microbiota, and disease resistance of crucian carp (Carassius auratus). Fishes, 7(1): 18, January 12, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/1/18.

Xenarios:2018:IRP

[XQLA18] Stefanos Xenarios, Henrique Queiroga, Ana I. Lillebø, and Ana Aleixo. Introducing a regulatory policy framework of bait fishing in European coastal lagoons: the case of Ria de Aveiro in Portugal. Fishes, 3 (1):2, January 05, 2018. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/3/1/2.

Xu:2023:CDI

[XRX^{+}23] Hongxiang Xu, Jindong Ren, Xiaojun Xu, Bao Lou, and Demin Zhang. The composition and diversity of intestinal microbes at different growth stages of giant freshwater prawns (Macrobrachium rosenbergii). Fishes, 8(9): 473, September 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/9/473.

Xu:2023:AFH

[XSZ^{+}23] Siqing Xu, Peng Sun, Chi Zhang, Jianchao Li, Xiaoyu Xi, Shuyang Ma, Wenchao Zhang, and Yongjun Tian. Age and feeding habits of caml grenadier Macrourus caml in Cosmonauts Sea. Fishes, 8(1):56, January 16, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/56.

Xiong:2024:PTU

[XW24] Minsi Xiong and Zuli Wu. Practice of territorial use rights in fisheries in coastal fishery management in China: a case study of the Island Reefs Fishery Lease Policy from Shengsi County in Zhejiang Province. Fishes, 9(7):292, July 22, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/292.

Xu:2024:EDF

[XWP^{+}24] Guangming Xu, Huijie Wei, Di Peng, Lang Zhang, Xing Lu, Qing Li, Lixue Dong, Juan Tian, Hua Wen, Guiying Wang, and Ming Jiang. Effects of dietary fish meal replaced by cottonseed protein concentrate on growth performance, antioxidant capacity, and liver and intestinal health of juvenile hybrid culter. Fishes, 9(4):127, March 31, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/127.

Xing:2023:CCM

[XWR^{+}23] Lili Xing, Lingyu Wang, Femke Roos, Michelle Lee, and Gregory A. Wray. CRISPR/Cas9mediated disruption of Endo16 cis-regulatory elements in sea urchin embryos. Fishes, 8(2): 118, February 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/2/118.

Xu:2024:HEG

[XWW^{+}24] Wencheng Xu, Yanzhe Wang, Guodong Wang, Lili Zhang, Guiling Zhang, Zhipeng Huo, and Hui Ge. Heritability estimates for growth traits and correlation analysis between weight and metamorphosis rate in the bullfrog Rana (Aquarana) catesbeiana. Fishes, 9 (3):105, March 09, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/105.

Xue:2022:IIC

[XXL^{+}22] Mingyang Xue, Zidong Xiao, Yiqun Li, Nan Jiang, Wenzhi Liu, Yan Meng, Yuding Fan, Lingbing Zeng, and Yong Zhou. Isolation, identification and characteristics of Aeromonas caviae from diseased largemouth bass (Micropterus salmoides). Fishes, 7(3):119, May 28, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/119.

Xu:2024:IBC

[XXL^{+}24] Yiqian Xu, Junren Xue, Hongbo Liu, Tao Jiang, Xiubao Chen, and Jian Yang. Identification of “Bathed” Chinese mitten crabs (Eriocheir sinensis) using geometric morphological analysis of the carapace. Fishes, 9(1):6, December 21, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/1/6.

Xu:2022:ERD

[XYC^{+}22] Qian Xu, Zheng Yang, Siyu Chen, Wenjuan Zhu, Siyuan Xiao, Jing Liu, Hongquan Wang, and Shile Lan. Effects of replacing dietary fish meal by soybean meal co-fermented using Bacillus subtilis and Enterococcus faecium on serum antioxidant indices and gut microbiota of crucian carp Carassius auratus. Fishes, 7 (2):54, April 25, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/2/54. See cor-

rection [XYC^{+}23].

Xu:2023:CXA

[XYC^{+}23] Qian Xu, Zheng Yang, Siyu Chen, Wenjuan Zhu, Siyuan Xiao, Jing Liu, Hongquan Wang, and Shile Lan. Correction: Xu et al. Effects of Replacing Dietary Fish Meal by Soybean Meal Co-Fermented Using Bacillus subtilis and Enterococcus faecium on Serum Antioxidant Indices and Gut Microbiota of Crucian Carp Carassius auratus. Fishes 2022, 7, 54. Fishes, 8(12):601, December 07, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/601.  See

[XYC^{+}22].

Xiu:2024:EIP

[XYF^{+}24] Yunji Xiu, Jingyuan Yi, Ruixin Feng, Jiaxue Song, Yunfei Pang, Peng Liu, and Shun Zhou. Evaluation of immune protection of a bivalent inactivated vaccine against Aeromonas salmonicida and Vibrio vulnificus in turbot. Fishes, 9(4):131, April 9, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/131.

Xu:2023:EDL

[XYL^{+}23] Jiaming Xu, Zhoulin Yu, Guangye Liu, Sijie Li, Guoyong Zhou, Hanhua Wang, Yewei Dong, Cuihong You, Weidong Bai, Meng Zhou, Yanhua Huang, and Xiaohong Tan. Effects of dietary Lentinus edodes fermentation supplementation on digestive enzyme activity, antioxidant capacity and morphology of the liver and intestine in largemouth bass (Micropterus salmoides) fed high plant protein diets. Fishes, 8(6):329, June 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/329.

Xu:2023:UST

[XYT23] Ran Xu, Xiaoming Yang, and Siquan Tian. Use of space-time cube model and spatiotemporal hot spot analyses in fisheries — a case study of tuna purse seine. Fishes, 8(10): 525, October 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/10/525.

Xu:2024:FCP

[XZD^{+}24] Xiandong Xu, Yanping Zhang, Liyun Ding, Jiangfeng Huang, Zhiyong Zhou, and Wenjing Chen. Farmed Chinese perch (Siniperca chuatsi) coinfected with parasites and oomycete pathogens. Fishes, 9 (3):97, March 04, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/97.

Xuan:2024:DCD

[XZZ^{+}24] Wendan Xuan, Hongliang Zhang, Haobo Zhang, Tian Wu, Yongdong Zhou, and Wenbin Zhu. Distribution characteristics and driving factors of Collichthys lucidus species in offshore waters of Zhejiang Province, China. Fishes, 9(3): 83, February 23, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/83.

Yildirim-Aksoy:2023:FBS

[YAEAB23] Mediha Yıldırım-Aksoy, Rashida Eljack, Janset Aksoy, and Benjamin H. Beck. Frass from black soldier fly larvae, Hermetia illucens, as a possible functional dietary ingredient in channel catfish feed. Fishes, 8(11):542, November 02, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/542.

Young:2021:ASL

[YAS21] Bradley Young, B. J.

Allaire, and Stephen Smith. Achieving sea lamprey control in Lake Champlain. Fishes, 6 (1):2, January 26, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/1/2.

Yi:2022:SAM

[YBL^{+}22] Hong Yi, Qinglu Bai, Ying Li, Honglei Zhan, Yujia Liu, Bingnan Liu, and Jihui Wang. Sporosarcina aquimarina MS4 regulates the digestive enzyme activities, body wall nutrients, gut microbiota, and metabolites of Apostichopus japonicus. Fishes, 7(3):134, June 07, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/134.

Yu:2023:BAU

[YCL^{+}23] Young-Bin Yu, Jae-Ho Choi, Ju-Hyeong Lee, A-Hyun Jo, Sung Won Han, Song-Hun Han, Hee Jae Choi, Cheol Young Choi, Ju-Chan Kang, EunYoung Min, and Jun-Hwan Kim. Biofloc application using aquaponics and vertical aquaculture technology in aquaculture: Review. Fishes, 8(11):543, November 04, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/543.

Youn:2023:AGC

[YCR^{+}23] Byoung-Il Youn, Dong-Hyuk Choi, Tae-Hyoung Roh, Seung-Hwan Lee, Kyeong-Ho Han, Dae-Hyeon Kwon, and Maeng-Jin Kim. Age and growth characteristics of Okamejei kenojei in the West Sea of South Korea according to coronal vertebral microstructure. Fishes, 8(4):197, April 10, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/4/197.

Young:2020:LHP

[YE20] Alan M. Young and James A. Elliott. Life history and population dynamics of green crabs (Carcinus maenas). Fishes, 5(1):4, December 31, 2020. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/1/4.

Yang:2023:HDD

[YGD^{+}23] Xuanyi Yang, Xinwei Guo, Xiaohui Dong, Qihui Yang, Hongyu Liu, Shuang Zhang, Beiping Tan, and Shuyan Chi. How do different dietary carbohydrate/lipid ratios influence intestinal morphology and glycolipid metabolism capacity in hybrid grouper (Epinephelus fuscoguttatus [female sign] × E. lanceolatus [male sign]). Fishes, 8(9): 467, September 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/467.

Ytteborg:2020:MTV

[YHH^{+}20] Elisabeth Ytteborg, Øyvind Johannes Hansen, Vibeke Høst, Sergey Afanasyev, Ireen Vieweg, Jasmine Nahrgang, and Aleksei Krasnov. Morphology, transcriptomics and in vitro model of skin from polar cod (Boreogadus saida) and Atlantic cod (Gadus morhua). Fishes, 5(4): 34, November 04, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/4/34.

Yuan:2023:CTP

[YHZ^{+}23] Hang Yuan, Naijie Hu, Yudong Zheng, Cuihong Hou, Beiping Tan, Lili Shi, and Shuang Zhang. A comparison of three protein sources used in medium-sized Litopenaeus vannamei: Effects on growth, immunity, intestinal digestive enzyme activity, and microbiota structure. Fishes, 8(9):449, September 06, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/449.

Yu:2023:HIB

[YJK23] Tae-Sik Yu, Chang Woo Ji, and Ihn-Sil Kwak. Hierarchical interaction between food diversity and competition in brackish fish species in South Korea. Fishes, 8 (6):313, June 13, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/6/313.

Yang:2024:CMP

[YJZ^{+}24] Zhigang Yang, Qingqing Jiang, Wuxiao Zhang, Silei Xia, Hongyan Tian, Fei Liu, Wenping Yang, Yebing Yu, Yanmin Wu, Yongzhi Zhu, Zhiqiang Xu, Zemao Gu, Aimin Wang, and Aqin Chen. Comparison of morphometric parameters, nutritional composition, and textural properties of seven crustaceans species. Fishes, 9(4): 141, April 17, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/141.

Yang:2024:CSB

[YLH^{+}24] Yukai Yang, Heizhao Lin, Xiaolin Huang, Hongbiao Dong, and Yafei Duan. Changes in the serum biochemical indices and intestinal microbial community of rabbitfish (Siganus oramin) in three different habitats. Fishes, 9 (3):96, March 04, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/96.

Yang:2022:FVC

[YLL22] Tsung-Yu Yang, Hsing-Chun Lin, and Wen-Hong Liu. The fishery value chain analysis in Taiwan. Fishes, 7 (3):114, May 16, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/114.

Yang:2023:EIA

[YLW^{+}23a] Jicheng Yang, Yaoyao Lin, Zhaohui Wei, Zhenbing Wu, Qianqian Zhang, Jingwen Hao, Shuyi Wang, and Aihua Li. Edwardsiella ictaluri almost completely occupies the gut microbiota of fish suffering from enteric septicemia of catfish (Esc). Fishes, 8(1): 30, January 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/30.

Yang:2023:EUP

[YLW^{+}23b] Yingming Yang, Wenlong Li, Run Wang, Dan Xu, Yadong Chen, Zhongkai Cui, and Songlin Chen. Effects of Ulva prolifera degradation on growth performance and antioxidant capacity of Japanese flounder (Paralichthys olivaceus) family. Fishes, 8(12):598, December 05, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/598.

Yu:2022:CGC

[YLX^{+}22] Hailuo Yu, Yi Liu, Tiaoyi Xiao, Hongquan Wang, and Baohong Xu. Cloning of grass carp chemokine XC receptor 1 (XCR1) gene and evaluation of its expression in various organs after GCRV infection. Fishes, 7(3):130, May 31, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/130.

Yang:2021:IIR

[YMD^{+}21] Yanping Yang, Fengjiao Ma, Juanjuan Dong, Lianxing Li, Ping Ren, Yuning Zhang, Yatao Wu, Yinping Wang, Kai Liu, and Fang Zhang. The innate immune response to infection by Polyascus gregaria in the male Chinese mitten crab (Eriocheir sinensis), revealed by proteomic analysis. Fishes, 6(4): 57, November 04, 2021. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/4/57.

Yuan:2023:MCE

[YSG^{+}23] Shuya Yuan, Zhaosheng Sun, Qian Gao, Zhen Li, Zhitao Qi, Sidi Zheng, and Danjie Liu. Molecular characterization and expression analysis of NLRC3-like, ASC, and Caspase1 in spotted sea bass (Lateolabrax maculatus). Fishes, 8(7): 378, July 21, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/378.

Yu:2023:IMS

[YSY^{+}23] Hyun-Ji Yu, Young Il Seo, Jae-Hyeong Yang, Jeong-Ik Baek, and Seong hun Kim. Investigating the mesh size selectivity of olive flounder (Paralichthys olivaceus) gillnets for fisheries resource management in the East Sea. Fishes, 8(11):560, November 20, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/560.

Yick:2021:EIC

[YWDP21] Jonah L. Yick, Chris Wisniewski, John Diggle, and Jawahar G. Patil. Eradication of the invasive common carp, Cyprinus carpio from a large lake: Lessons and insights from the Tasmanian experience. Fishes, 6(1):6, February 23, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/6/1/6.

Yu:2024:HSS

[YWL^{+}24] Jun Yu, Lihong Wen, Siyuan Liu, Heng Zhang, and Zhou Fang. Habitat suitability of the squid Sthenoteuthis oualaniensis in Northern Indian Ocean based on different weights. Fishes, 9 (3):107, March 15, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/3/107.

Yin:2024:DDI

[YXS^{+}24] Zifeng Yin, Lingzhan Xue, Peng Song, Muhammad Jawad, Cong Xu, Weishao Bu, and Mingyou Li. Dazl and dnd identify both embryonic and gonadal germ cells in Chinese hook snout carp (Opsariichthys bidens). Fishes, 9(6):214, June 5, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/6/214.

Yu:2024:MTC

[YYH^{+}24] Linghui Yu, Weijun Yin, Senrong Han, Tanjun Zhao, Zhenlin Hao, Donghong Yin, Yaoyao Zhan, and Yaqing Chang. Morphological trait correlations, gonadal development characteristics and pleopod nutrient compositions of the whelk Volutharpa perryi perryi. Fishes, 9(2):72, February 11, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/72.

Yang:2023:ENT

[YYW^{+}23] Shenglong Yang, Linlin Yu, Fei Wang, Tianfei Chen, Yingjie Fei, Shengmao Zhang, and Wei Fan. The environmental niche of the tuna purse seine fleet in the Western and Central Pacific Ocean based on different fisheries data. Fishes, 8(2):78, January 29, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/78.

Yazici:2024:ARM

[YZH^{+}24] Metin Yazici, Fatemeh Zavvar, Seyed Hossein Hoseinifar, Shiva Nedaei, and Hien Van Doan. Administration of red macroalgae (Galaxaura oblongata) in the diet of the rainbow trout (Oncorhynchus mykiss) improved immunity and hepatic gene expression. Fishes, 9(2):48, January 26, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/2/48.

Yang:2023:CGA

[YZL^{+}23] Yuanhao Yang, Yameng Zhang, Fenggang Li, Sien Wen, Lvzhou Wang, Guozhu Lan, Jishu Zhou, Yang Li, and Zilin Zhu. Comparative genomic analysis unveils potential factors contributing to the endangerment of Silurus lanzhouensis. Fishes, 8(12):613, December 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/613.

Zaccone:2022:INI

[Zac22] Giacomo Zaccone. Immunity and neuroimmune interactions at the mucosal barriers in fish. Fishes, 7(6):381, December 08, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/381.

Zaccone:2023:LAA

[ZAM^{+}23] Giacomo Zaccone, Alessio Alesci, Doaa M. Mokhtar, Marialuisa Aragona, Maria Cristina Guerrera, Gioele Capillo, Marco Albano, Jorge de Oliveira Fernandes, Viswanath Kiron, Ramy K. A. Sayed, Marwa M. Hussein, Patrizia Lo Cascio, Michal Kuciel, Krystyna Zuwala, Antonino Germanà, Jose Manuel Icardo, and Eugenia Rita Lauriano. Localization of acetylcholine, alpha 7-NAChR and the antimicrobial peptide piscidin 1 in the macrophages of fish gut: Evidence for a cholinergic system, diverse macrophage populations and polarization of immune responses. Fishes, 8(1): 43, January 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/43.

Zhong:2023:TAF

[ZCT^{+}23] Zhenxiao Zhong, Guozhu Chen, Haihui Tu, Xinyi Yao, Xin Peng, Xuan Lan, Qiongying Tang, Shaokui Yi, Zhenglong Xia, Miaoying Cai, and Guoliang Yang. Transcriptomic analysis and functional gene expression in different stages of gonadal development of Macrobrachium rosenbergii. Fishes, 8(2): 94, February 05, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/94.

Zhou:2022:EST

[ZCX^{+}22] Xingxing Zhou, Zuozhi Chen, Pengli Xiong, Yancong Cai, Jie Li, Peng Zhang, Jun Zhang, Miao Li, and Jiangtao Fan. Exploring the spatial and temporal distribution of frigate tuna (Auxis thazard) habitat in the South China Sea in spring and summer during 2015–2019 using fishery and remote sensing data. Fishes, 7(5): 218, October 25, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/5/218.

Zhou:2023:GDP

[ZGY^{+}23] Lin Zhou, Jiancao Gao, Yanping Yang, Zhijuan Nie, Kai Liu, and Gangchun Xu. Genetic diversity and population structure analysis of Chinese mitten crab (Eriocheir sinensis) in the Yangtze and Liaohe Rivers. Fishes, 8 (5):253, May 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/253.

Zhang:2022:MCH

[ZHF^{+}22a] Jianlu Zhang, Jiqin Huang, Cheng Fang, Wanchun Li, Hu Zhao, Fei Kong, Han Zhang, Hongxing Zhang, and Qijun Wang. Molecular cloning of heat shock protein 60 (Sp HSP60) from Schizothorax prenanti and the gene expressions of four Sp HSPs during lipopolysaccharide (LPS) infection. Fishes, 7(3):139, June 13, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/139.

Zhao:2022:PSR

[ZHF^{+}22b] Hu Zhao, Jiqin Huang, Cheng Fang, Hongying Ma, Han Zhang, Jie Deng, Wei Jiang, Fei Kong, Hongxing Zhang, Hong Liu, and Qijun Wang. Predicting sexrelated transcripts in the Chinese giant salamander (Andrias davidianus): a transcriptomics study, selection gender for preservation, breeding and reintroduction. Fishes, 7(6):399, December 19, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/6/399.

Zuluaga-Hernandez:2023:NCI

[ZHHO23] Christian David Zuluaga-Hernández, Carlos A. Hincapié, and Marisol Osorio. Non-conventional ingredients for tilapia (Oreochromis spp.) feed: a systematic review. Fishes, 8(11):556, November 18, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/556.

Zhao:2023:HLG

[ZHQ^{+}23] Lingmin Zhao, Lixing Huang, Yingxue Qin, Dou Yang, Jiaonan Zhang, Jiaolin Zhang, and Qingpi Yan. How the luxR gene affects the pathogenicity of Pseudomonas plecoglossicida and the immune response of Epinephelus coioides. Fishes, 8(10): 507, October 11, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/507.

Zhu:2024:COL

[ZHX^{+}24] Jian Zhu, Nanjun Hu, Yao Xiao, Xiaohong Lai, Lingjiao Wang, and Yufeng Song. Characterization of ovarian lipid composition in the largemouth bronze gudgeon (Coreius guichenoti) at different development stages. Fishes, 9(7): 291, July 22, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/7/291.

Zhu:2022:IOD

[ZJC^{+}22a] Yahua Zhu, Tao Jiang, Xiubao Chen, Hongbo Liu, Quinton Phelps, and Jian Yang. Interotolith differences in strontium markings: a case study on the juvenile crucian carp Carassius carassius (Linnaeus, 1758). Fishes, 7(3):112, May 15, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/3/112.

Zhu:2022:PSA

[ZJC^{+}22b] Yahua Zhu, Tao Jiang, Xiubao Chen, Hongbo Liu, Quinton Phelps, and Jian Yang. A pilot study assessing a concentration of 100 mg/L alizarin complexone (ALC) to mark calcified structures in Hypophthalmichthys molitrix. Fishes, 7(2):66, April 16, 2022. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/2/66.

Zhang:2022:SDR

[ZJJ^{+}22] Zhicong Zhang, Fengyu Ji, Shouwen Jiang, Zhichao Wu, and Qianghua Xu. Scale developmentrelated genes identified by transcriptome analysis. Fishes, 7(2):64, April 12, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/2/64.

Zhou:2023:TBA

[ZJM^{+}23] Tangjian Zhou, Chaofeng Jia, Qian Meng, Dafeng Xu, Zhiwei Zhang, Fei Zhu, Yonglei Zhao, Ruijian Sun, Yunxia Yang, and Shuyin Chen. Transcriptome-based analysis of the liver response mechanism of black porgy (Acanthopagrus schlegelii) to stocking density. Fishes, 8 (7):356, July 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/356.

Zhang:2023:EAC

[ZLF23] Ying Zhang, Meng-Fei Li, and Xiao-Han Fang. Efficiency analysis of China deep-sea cage aquaculture based on the SBM–Malmquist model. Fishes, 8(10): 529, October 23, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/529.

Zhang:2023:GDG

[ZLL^{+}23] Zhe Zhang, Chengkuan Lu, Kebing Lin, Weiwei You, and Zhangwu Yang. Genetic diversity and genetic structure among four selected strains of whiteleg shrimp (Litopenaeus vannamei) using SSR markers. Fishes, 8(11): 544, November 06, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/544.

Zhang:2022:ICD

[ZLSB22] Dunhua Zhang, Miles D. Lange, Craig A. Shoemaker, and Benjamin H. Beck. Identification and characterization of differentially expressed IgM transcripts of channel catfish vaccinated with antigens of virulent Aeromonas hydrophila. Fishes, 7(1): 24, January 19, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/1/24.

Zhan:2023:GWI

[ZLW^{+}23] Feifei Zhan, Liqun Liang, Shuangyi Wang, Honjung Liew, Yumei Chang, and Limin Zhang. Genomewide identification, phylogenetic analysis and expression pattern profiling of the aquaporin family genes in Leuciscus waleckii. Fishes, 8(2):107, February 11, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/2/107.

Zheng:2023:CAF

[ZLX^{+}23] Yueping Zheng, Jiehao Liu, Jianan Xu, Houyong Fan, Youji Wang, Ping Zhuang, and Menghong Hu. Comparison of artificial feed and natural food by the growth and blood biochemistry in Chinese sturgeon Acipenser sinensis. Fishes, 8(1):45, January 10, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/45.

Zhang:2022:MFT

[ZLZ^{+}22] Haoran Zhang, Kun Li, Fayang Zhang, Jikui Wu, and Junling Zhang. The miR-200 family targeting amh affects the gonadal development of Japanese flounder. Fishes, 7(3):129, May 31, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/3/129.

Zhai:2024:EWF

[ZLZ^{+}24] Cunhua Zhai, Yutao Li, Guanyu Zhu, Wenjie Peng, Qiuxu E, Ying Zhang, and Bo Ma.

The effects of water flow speed on swimming capacity and energy metabolism in adult Amur grayling (Thymallus grubii). Fishes, 9 (7):272, July 10, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/272.

Zhang:2023:MCR

[ZLZL23] Zewen Zhang, Xiaohui Li, Guiwei Zou, and Hongwei Liang. Molecular characterization and response of silver carp (Hypophthalmichthys molitrix) GLUT1 under hypoxia stress. Fishes, 8(8): 425, August 20, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/8/425.

Zaccone:2024:PPI

[ZMA^{+}24] Giacomo Zaccone, Doaa M. Mokhtar, Alessio Alesci, Gioele Capillo, Marco Albano, Manal T. Hussein, Marialuisa Aragona, Antonino Germanà, Eugenia R. Lauriano, and Ramy K. A. Sayed. From proliferation to protection: Immunohistochemical profiling of cardiomyocytes and immune cells in molly fish hearts. Fishes, 9(7): 283, July 16, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/283.

Zaragoza:2020:SFS

[ZMLFS^{+}20] Patricia Zaragozá, Silvia Martínez-Llorens, Isabel Fernández-Segovia, José-Luis Vivancos, Ana Tomas-Vidal, Ana Fuentes, José Vicente Ros-Lis, Ramón Martínez-Máñez, and José Manuel Barat. Study of fishmeal substitution on growth performance and shelf-life of giltheadsea bream (Sparusaurata). Fishes, 5(2):15, May 15, 2020. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/5/2/15.

Zhang:2023:MAP

[ZQL^{+}23] Yu Zhang, Lu Qi, Fengping Li, Yi Yang, Zhifeng Gu, Chunsheng Liu, Qi Li, and Aimin Wang. Mitogenomic analysis of Pterioidea (Bivalvia: Pteriomorphia): Insights into the evolution of the gene rearrangements. Fishes, 8(10):528, October 23, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/10/528.

Zhang:2023:GGC

[ZQLW23] Yichao Zhang, Siyong Qin, Qinghua Liu, and Wenqi Wang. Gonad and germ cell development and maturation characteristics of the pot-bellied seahorse (Hippocampus abdominalis) under captive breeding conditions in Northern China. Fishes, 8(11):551, November 15, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/11/551.

Zielinski:2021:NSD

[ZS21] Daniel Patrick Zielinski and Peter W. Sorensen. Numeric simulation demonstrates that the upstream movement of invasive bigheaded carp can be blocked at sets of Mississippi River locksand-dams using a combination of optimized spillway gate operations, lock deterrents, and carp removal. Fishes, 6(2): 10, March 26, 2021. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/2/10.

Zhang:2023:AIA

[ZSH^{+}23] Yadong Zhang, Abubakar Shitu, Shengyu Hang, Zhangying Ye, Wen Xu, Hangfang Zhao, Jian Zhao, and Songming Zhu. Assessing the impacts of aquaculture soundscapes on the growth, physiology and behavior of Micropterus salmoides. Fishes, 8 (7):377, July 21, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/7/377.

Zhou:2023:NTA

[ZSL^{+}23] Ziyu Zhou, Gu Shan, Xin Li, Wenxuan Guo, Kexin Ma, Yueyao Yang, Yifan Li, Yunbang Zhang, He Zhou, and Xiaojuan Cao. De novo transcriptome analysis of the early hybrid triploid loach (Misgurnus anguillicaudatus) provides novel insights into fertility mechanism. Fishes, 8(2):70, January 23, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/2/70.

Zhang:2024:MCC

[ZSL^{+}24] Yang Zhang, Jiayuan Shi, Yuntao Lu, Qing Luo, Pengfei Chu, Rong Huang, Kunci Chen, Jian Zhao, Yaping Wang, and Mi Ou. Molecular cloning and characterization of scavenger receptor class B type 1 in grass carp (Ctenopharyngodon idellus) and its expression profile following grass carp reovirus challenge. Fishes, 9(7):276, July 12, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/7/276.

Zhang:2021:SIS

[ZSQ^{+}21] Xi Zhang, Zhiyuan Shen, Tianpeng Qi, Rujuan Xi, Xiao Liang, Li Li, Rong Tang, and Dapeng Li. Slight increases in salinity improve muscle quality of grass carp (Ctenopharyngodon idellus). Fishes, 6(1):7, February 24, 2021. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/6/1/7.

Zhang:2023:MCE

[ZSZ^{+}23] Lei Zhang, Peng Shi, Lin Zhang, Yajun Wang, Jilin Xu, Xiaojun Yan, and Kai Liao. Molecular characterization and expression response of Ghrelin, GLP-1 and PYY to fasting, dietary lipid, and fatty acids in silver pomfret (Pampus argenteus). Fishes, 8 (3):170, March 22, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/3/170.

Zemah-Shamir:2022:SRB

[ZSZSS^{+}22] Ziv Zemah-Shamir, Shiri Zemah-Shamir, Aviad Scheinin, Dan Tchernov, Teddy Lazebnik, and Gideon Gal. A systematic review of the behavioural changes and physiological adjustments of elasmobranchs and teleost’s to ocean acidification with a focus on sharks. Fishes, 7(2):56, April 28, 2022. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/2/56.

Zhou:2024:MEI

[ZTWK24] Jinxin Zhou, Teng Tu, Huajin Wang, and Daisuke Kitazawa. Modeling environmental impacts of intensive shrimp aquaculture: a threedimensional hydrodynamic ecosystem approach. Fishes, 9(4): 126, March 31, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/4/126.

Ziebell:2023:HUT

[ZVRH23] Ann-Christin Ziebell, Maite L. Vogel, Niklas Kjell Ratajczak, and Bert W. Hoeksema. Habitat use of two coral-associated cryptobenthic gobiid fishes (family: Gobiidae) in the Southern Caribbean. Fishes, 8 (10):531, October 23, 2023. CODEN ???? ISSN 2410-3888. URL

https://www.mdpi.com/
2410-3888/8/10/531.

Zhou:2023:TTE

[ZWD^{+}23] Zhibin Zhou, Lubo Wang, Mingqin Dai, Qingyan Gao, Peng Wang, Lili Zhao, Yanlu Li, Rui Xi, Mingchao Pan, Qiang Ma, Houguo Xu, Mengqing Liang, and Yuliang Wei. Three types of Enteromorpha prolifera bio-products based on different processing procedures as feed additives in the diets of Pacific white shrimp (Litopenaeus vannamei). Fishes, 8(12):587, November 29, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/12/587.

Zheng:2023:ESD

[ZWP^{+}23] Jishu Zheng, Zhengxi Wang, Decheng Pu, Peiyuan Li, Xiuli Wei, Mai Li, Dongsheng Li, Lihong Gao, and Xuliang Zhai. Effects of stocking density on intestinal health of juvenile Micropterus salmoides in industrial aquaponics. Fishes, 8(11):555, November 17, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/555.

Zhang:2022:EDC

[ZWZ22] Mingliang Zhang, Xinyi Wu, and Shaowei Zhai. Effect of dietary compound acidifiers supplementation on growth performance, serum biochemical parameters, and body composition of juvenile American eel (Anguilla rostrata). Fishes, 7(4):203, August 12, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/4/203.

Zhang:2024:MCE

[ZWZ^{+}24] Xiaotian Zhang, Yuxia Wu, Yang Zhang, Jin Zhang, Kunci Chen, Haiyang Liu, Qing Luo, Shuzhan Fei, Jian Zhao, and Mi Ou. Molecular characteristics, expression patterns, and response of insulin-like growth factors gene induced by sex steroid hormones in blotched snakehead (Channa maculata). Fishes, 9(4):120, March 28, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/4/120.

Zhou:2023:DSG

[ZXH^{+}23] Yang Zhou, Li Xu, Zhongtang He, Weijie Cui, Qun Lu, Jianguang Qin, Shengqi Su, and Tao He. Discrimination of Schizothorax grahami (Regan, 1904) stocks based on otolith morphology. Fishes, 8(10): 504, October 10, 2023. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/8/10/504.

Zhang:2024:PFG

[ZXY^{+}24] Tianjiao Zhang, Jia Xin, Wei Yu, Hongchun Yuan, Liming Song, and Zhuo Yang. Predicting the fishery ground of jumbo flying squid (Dosidicus gigas) off Peru by extracting features of the ocean environment. Fishes, 9(3): 81, February 21, 2024. CODEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/81.

Zhang:2023:EDF

[ZYG^{+}23] Qin Zhang, Qiuyue Yang, Mengjie Guo, Fanghui Li, Meilan Qin, Yi Xie, Jian Xu, Yongqiang Liu, and Tong Tong. The effects of dietary fermented soybean meal supplementation on the growth, antioxidation, immunity, and mTOR signaling pathway of juvenile Coho salmon (Oncorhynchus kisutch). Fishes, 8(9):448, September 05, 2023. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/9/448.

Zhong:2024:FSB

[ZYH^{+}24] Botao Zhong, Huaihua Yu, Shengming Han, Weiwei Song, Zhiming Ren, Chunlin Wang, and Changkao Mu. Functional study on the BMP signaling pathway in the molting of Scylla paramamosain. Fishes, 9 (7):263, July 4, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/263.

Zhong:2024:ERI

[ZYT^{+}24] Hua Zhong, Xinyi Yao, Haihui Tu, Zhenglong Xia, Miaoying Cai, Qiang Sheng, Shaokui Yi, Guoliang Yang, and Qiongying Tang. Effects of RNA interference with acetyl-CoA carboxylase gene on expression of fatty acid metabolism-related genes in Macrobrachium rosenbergii under cold stress. Fishes, 9(5): 170, May 8, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/5/170.

Zhao:2024:MMG

[ZZ24] Yun Zhao and Junming Zhou. Molecular mechanisms of growth differences in Gymnocypris przewalskii and Gymnocypris eckloni through a comparative transcriptome perspective. Fishes, 9(3):89, February 28, 2024. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/9/3/89.

Zhou:2022:MCN

[ZZC^{+}22] Xiaomei Zhou, Dafang Zhao, Yuan Chen, Yangbo Xiao, Zhuangwen Mao, Shenping Cao, Fufa Qu, Yutong Li, Junyan Jin, Zhen Liu, Jianzhong Li, and Zhimin He. Molecular characterization and nutrition regulation of the glutamine synthetase gene in triploid crucian carp. Fishes, 7(4):196, August 08, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

                                        //www.mdpi.com/2410-
                                        3888/7/4/196.

Zhang:2023:DDM

[ZZL^{+}23] Kexin Zhang, Weiren Zhang, Ronghua Li, Junkai Lu, Qingwei Chen, Haojie Hu, Fei Yin, Changkao Mu, Weiwei Song, and Chunlin Wang. Dynamic distribution of Mesanophrys sp. and tissue enzyme activities in experimentally infected mud crab Scylla paramamosain. Fishes, 8(5):249, May 08, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/5/249.

Zhang:2023:TAG

[ZZT^{+}23] Yifan Zhang, Dandan Zhang, Ying Tian, Junxia Mao, Yang Liu, and Zhenlin Hao. Transcriptome analysis of gill tissues from Neptunea cumingii in different seasons. Fishes, 8(11): 549, November 11, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/549.

Zhou:2022:RHT

[ZZW^{+}22] Yan Zhou, Yanjie Zhang, Shang Wei, Wei Li, Wenhao Li, Zhichao Wu, Shouwen Jiang, Ying Lu, Qianghua Xu, and Liangbiao Chen. Reduced hypoxia tolerance and altered gill morphology at elevated temperatures may limit the survival of tilapia (GIFT, Oreochromis niloticus) under global warming. Fishes, 7(5):216, October 24, 2022. CO-DEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/7/5/216.

Zhang:2024:CSC

[ZZW^{+}24] Yanping Zhang, Haixin Zhang, Zijun Wu, Mingguang Zhao, and Guangpeng Feng. Community structure characteristics and changes in fish species at Poyang Lake after the Yangtze River fishing ban. Fishes, 9 (7):281, July 15, 2024. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/9/7/281.

Zhou:2023:TCL

[ZZY^{+}23] Fan Zhou, Xuelin Zhang, Gaohua Yao, Xiaoming Chen, Ming Qi, Qin Zhou, Ningyu Zhu, Qinghui Meng, Yu Zhang, and Xueyan Ding. Transcriptomic comparison of liver tissue across different largemouth bass (Micropterus salmoides) strains. Fishes, 8(11): 558, November 19, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/11/558.

Zuo:2023:EUF

[ZZZ^{+}23] Chenxia Zuo, Tingting Zhang, Chenchen Zhang, Daotan Zhao, Yi Zhu, Xiaojie Ma, Haiyan Wang, Peizhen Ma, and Zhen Zhang. Evaluating the utility of five gene fragments for genetic diversity analyses of Mytella strigata populations. Fishes, 8 (1):34, January 03, 2023. CODEN ???? ISSN 2410-3888. URL https:

//www.mdpi.com/2410-
3888/8/1/34.